### v1

# Model-oriented Graph Distances via Partially Ordered Sets

Abstract
A well-defined distance on the parameter space is key to evaluating estimators, ensuring consistency, and building confidence sets. While there are typically standard distances to adopt in a continuous space, this is not the case for combinatorial parameters such as graphs that represent statistical models. Existing proposals like the structural Hamming distance are defined on the graphs rather than the models they represent and can hence lead to undesirable behaviors. We propose a model-oriented framework for defining the distance between graphs that is applicable across many different graph classes. Our approach treats each graph as a statistical model and organizes the graphs in a partially ordered set based on model inclusion. This induces a neighborhood structure, from which we define the model-oriented distance as the length of a shortest path through neighbors, yielding a metric in the space of graphs. We apply this framework to both probabilistic graphical models (e.g., undirected graphs and completed partially directed acyclic graphs) and causal graphical models (e.g., directed acyclic graphs and maximally oriented partially directed acyclic graphs). We analyze the theoretical and empirical behaviors of model-oriented distances. Algorithmic tools are also developed for computing and bounding these distances.
structural Hamming distance,
keywords:
[class=MSC]
keywords:
, , and
t1All authors contributed equally to this work.
1 Introduction
Graphs provide an intuitive framework for encoding and reasoning about the dependency structure among a set of variables. In statistics, they have become popular in two closely related but different contexts: probabilistic modeling and causal modeling. For the former context, graphs typically encode conditional independence relations among a set of (factual) random variables (Lauritzen, 1996); for certain scenarios, they can also encode generalized versions of conditional independence and other constraints (Evans, 2016; Richardson et al., 2023) — we refer to these as probabilistic graphical models. Among other applications, these graphs lead to sound and efficient algorithms for probabilistic reasoning (Lauritzen and Spiegelhalter, 1988; Lunn et al., 2000). For the latter context, a graph models a causal system by encoding the relations of its building blocks, namely counterfactual (as well as factual) random variables; as such, the graph not only describes the system’s behavior as things naturally occur, but also its behavior under interventions (Wright, 1934; Spirtes, Glymour and Scheines, 2000; Pearl, 2009) — we use the term causal graphical models to refer to graphs like this. These graphs are used to determine if a causal effect can be identified from observational data (Huang and Valtorta, 2006; Shpitser and Pearl, 2006), and if so, to also inform the best way to estimate the effect (Rotnitzky and Smucler, 2020; Henckel, Perković and Maathuis, 2022; Guo, Perković and Rotnitzky, 2023). There is also a large literature on learning probabilistic or causal graphs from data, a problem known as structure learning or causal discovery; the reader is referred to Drton and Maathuis (2017); Glymour, Zhang and Spirtes (2019) for two recent surveys.
In whichever context these graphs are used, a key unresolved issue is how to define a suitable distance between them. Having a distance is essential for several tasks, including benchmarking causal discovery algorithms (Scutari, Graafland and Gutiérrez, 2019; Rios, Moffa and Kuipers, 2023), characterizing their statistical performance (Jin, Zhang and Zhang, 2012; Ke, Jin and Fan, 2012; Han et al., 2016) or finding a consensus graph among a collection of graphs through a Frechét mean (Ferguson and Meyer, 2023; Wang, Kolar and Drton, 2025). Other potential applications include constructing ball-like confidence regions for a graph and and conducting sensitivity analysis with respect to potential mistakes in the causal graph. However, defining a suitable distance is challenging for two reasons. First, there are many types of graphs, such as undirected graphs (UGs), directed acyclic graphs (DAGs), completed partially directed acyclic graphs (CPDAGs), and maximally oriented partially directed acyclic graphs (MPDAGs), each with its own edge types and rules on how edges may be configured. Second, each type of graphs rely on a set of semantics (sometimes called Markov properties) to be interpreted as probabilistic or causal models, and a good distance between two graphs should reflect how close they are in terms of the models they represent. That is, we seek a semantic rather than structural definition of the distance. Despite these challenges, several distances have been proposed in the literature (Tsamardinos, Brown and Aliferis, 2006; Peters and Bühlmann, 2015; Viinikka, Eggeling and Koivisto, 2018), which we divide into two categories — model-agnostic and task-oriented distances.
Arguably, the most popular distance is the structural Hamming distance (SHD) (Tsamardinos, Brown and Aliferis, 2006). It counts the number of edges that differ between two graphs with various proposals for how to weigh different types of edge mismatches (Perrier, Imoto and Miyano, 2008). Irrespective of the concrete weighing the SHD is computationally inexpensive, flexible with respect to the graph class, and guaranteed to define a metric. It is, however, model-agnostic — it disregards the specific models the graphs represent and is solely defined on the graphs themselves. This results in undesirable behaviors. First, for a pair of graphs, the SHD provides a fixed answer regardless of how the graphs are interpreted, whether probabilistically or causally. Second, the SHD does not depend on the set of graphs in consideration — it is agnostic to constraints posed by the graph type and domain specific considerations (e.g., one may wish to consider acyclic graphs). Finally, by construction the SHD treats each edge individually. This ignores that not all edge configurations may give a valid graph in the graph class under consideration. We illustrate these issues with an example.
Example 1.
Consider the four CPDAGs in Fig.˜1(a-d). For two of the more prominent types of SHD (see Section˜4.3), we have $\mathrm{SHD}(\mathcal{G}_{1},\mathcal{G}_{2})=n+1$, respectively $n+2$. Similarly, $\mathrm{SHD}(\mathcal{G}_{1},\mathcal{G}_{3})=2\text{ or }4$, and for even $n$, $\mathrm{SHD}(\mathcal{G}_{1},\mathcal{G}_{4})=n/2+1\text{ or }n+2$. However, in the context of probabilistic models, $\mathcal{G}_{2}$ sits “just between” $\mathcal{G}_{1}$ and $\mathcal{G}_{3}$: $\mathcal{G}_{2}$ encodes fewer conditional dependencies than $\mathcal{G}_{1}$ and no graph encodes an intermediary collection of dependencies. The same is true for $\mathcal{G}_{3}$ with respect to $\mathcal{G}_{2}$. Logically, $\mathcal{G}_{3}$ should be closer to $\mathcal{G}_{2}$ than $\mathcal{G}_{1}$. The SHD overestimates the semantic distance between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, as it wrongly assumes that undirecting any edge in $\mathcal{G}_{2}$ leads to a valid CPDAG that is “in between” $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. Similarly, since $\mathcal{G}_{4}$ removes far more dependencies from $\mathcal{G}_{1}$ than $\mathcal{G}_{2}$ does, we expect $\mathcal{G}_{1}$ to be closer to $\mathcal{G}_{2}$ than to $\mathcal{G}_{4}$. Our proposed model-oriented distance reflects the semantics of the graphs, and yields $d(\mathcal{G}_{1},\mathcal{G}_{2})=1$, $d(\mathcal{G}_{1},\mathcal{G}_{3})=2$, and $d(\mathcal{G}_{1},\mathcal{G}_{4})=n/2+1$. As a further illustration, we set $n=10$, and generate $100$ datasets of $1000$ samples each from a linear Gaussian structural equation model from graph $\mathcal{G}_{1}$, with unit edge weight and noise variance. We can fit a given graph $\mathcal{G}$ to the data and compute the Bayesian Information Criterion (BIC), which balances model fit and complexity. As shown in Fig.˜1(e-f), our proposed model-oriented distance better reflects fit to data: further away graphs tend to yield worse BIC scores. This indicates that using SHD to benchmark score-based causal discovery algorithms that rely on BIC, such as the well-known greedy equivalence search algorithm (Chickering, 2002), may be misleading.
To address the limitation of model-agnostic distances, Peters and Bühlmann (2015) proposed an alternative, task-oriented approach, based on the idea that the distance between two graphs should be measured based on the specific task they serve. In the context of causal inference, the central task is to identify a causal effect of interest, i.e., to express the effect as a functional of the observed distribution that holds under every causal system that is compatible with a graph. Peters and Bühlmann (2015) introduced the structural intervention distance (SID) between two DAGs, which counts how many identifying formulae under one graph would be invalid if the truth lies within the other graph. This approach was also adapted and generalized by Henckel, Würtzen and Weichwald (2024) who establish that the SID is merely one of a class of closely related distances they call adjustment identification distances. The SID is, however, not a metric: it can be zero between two different graphs, is not symmetric and does not obey the triangle inequality. It is possible to define a symmetrized version of SID but we show in Section˜H.1 that this version still violates the triangle inequality. In general task-oriented distance are not constructed to be metrics, limiting their usefulness.
(a) $\mathcal{G}_{1}$
(c) $\mathcal{G}_{3}$
(b) $\mathcal{G}_{2}$
(d) $\mathcal{G}_{4}$
(e) SHD distance vs BIC score
(f) Proposed model-oriented distance vs BIC score
Figure 1: (a–f) Four CPDAGs in Example˜1 and their associated BIC scores for synthetically generated data.
1.1 Our contributions
We introduce a model-oriented framework that defines distance between graphs through the models they represent. Given a class of graphs of the same type (e.g., the collection of CPDAGs over five vertices), our approach begins by interpreting a probabilistic or causal graph as a statistical model, namely a set of distributions over a common collection of random variables — factual or counterfactual — that obey the constraints encoded by the graph. This viewpoint naturally leads to the organization of graphs into a partially ordered set (poset), where the ordering is defined by model inclusion: for two graphs $\mathcal{G}$ and $\mathcal{G}^{\prime}$, we say $\mathcal{G}\preceq\mathcal{G}^{\prime}$ if and only if the model represented by $\mathcal{G}$ is contained by the model represented by $\mathcal{G}^{\prime}$. Posets are equipped with a natural neighborhood notion called covering elements. Exploiting this neighborhood notion, we define our distance as the length of a shortest path between two graphs when moving only through neighbors. By construction, our distance is a metric and applicable to many classes of graphs of varying type.
We study in detail the model-oriented distance for four graph types, listed in order of increasing complexity of their associated posets: probabilistic UGs, causal DAGs, probabilistic CPDAGs, and causal MPDAGs. For each, we study algorithms for computing or bounding the model-oriented distance. Here, a key ingredient is a graphical characterization of model containment, and whenever possible, of neighborhood. When the latter proves difficult, we study the neighborhood for a restricted class of graphs where the characterization is tractable. Further, we show that poset-theoretic properties, such as gradedness and semimodularity, can facilitate computing the distances. We also show that, for CPDAGs and MPDAGs, the model-oriented distance can differ substantially from the structural Hamming distance. For CPDAGs and MPDAGs restricted to polytrees, we also prove that the model-oriented distance admits several simplifications that make it easier to compute and interpret.
By establishing that many statistical graph classes are naturally equipped with partial orderings that can be studied using theoretical tools from order theory, we open a new avenue for investigating graphical models and their properties. Partial orders, for example, have been shown to be natural and useful structures for model selection over combinatorial objects, including graphs (Taeb, Bühlmann and Chandrasekaran, 2023). Similarly, neighborhood notions are central to score-based causal discovery methods where they used to decide how to explore the space of all graphs in a class while searching for a score-minimizing graph (Chickering, 2002).
2 Background
2.1 Graphical preliminaries
We consider graphs with nodes representing random variables and edges between them representing probabilistic or causal dependencies between these variables. Basic graphical concepts are briefly described in below.
2.1.1 Node, edge, subgraph, cycle, PDAG
We write a graph over vertex set $\mathrm{V}$ with $k\geq 1$ types of edges (e.g., undirected ‘-’ and directed $\rightarrow$’) as $\mathcal{G}=(\mathrm{V},\mathrm{E}^{1},\dots,\mathrm{E}^{k})$. Depending on the edge type $i$, edge set $\mathrm{E}^{i}$ is either a set of unordered tuples ($(1,2)=(2,1)$) or a set of ordered tuples ($(1,2)\neq(2,1)$). In either case, we forbid self-loops and let $\mathrm{E}^{i}\subseteq(\mathrm{V}\times\mathrm{V})\setminus\cup_{i\in\mathrm{V}}\{(i,i)\}$ for $i=1,\dots,k$. A graph $\mathcal{G}_{1}=(\mathrm{V},\mathrm{E}_{1}^{1},\dots,\mathrm{E}_{1}^{k})$ is said to be a subgraph of $\mathcal{G}_{2}=(\mathrm{V},\mathrm{E}_{2}^{1},\dots,\mathrm{E}_{2}^{k})$ if $\mathrm{E}_{1}^{i}\subseteq\mathrm{E}_{2}^{i}$ for each edge type $i$. Given a graph $\mathcal{G}=(\mathrm{V},\mathrm{E}^{1},\dots,\mathrm{E}^{k})$ and a subset of vertices $\mathrm{V}^{\prime}\subseteq\mathrm{V}$, the induced graph of $\mathrm{V}^{\prime}$ is $\mathcal{G}_{\mathrm{V}^{\prime}}:=(\mathrm{V}^{\prime},{\mathrm{E}^{1}}^{\prime},\dots,{\mathrm{E}^{k}}^{\prime})$ where ${\mathrm{E}^{i}}^{\prime}:=\mathrm{E}^{i}\cap(\mathrm{V}^{\prime}\times\mathrm{V}^{\prime})$ for $i=1,\dots,k$. A graph is simple if between every two vertices there is at most one edge (of any type). Throughout, we only consider simple graphs. A cycle is a sequence of vertices $v_{1},v_{2},\dots,v_{k}$ ($k\geq 3$) such that there is an edge between every two adjacent vertices, $v_{k}=v_{1}$ and $v_{1},\dots,v_{k-1}$ are all distinct. A directed cycle is a cycle of the form $v_{1}\rightarrow v_{2}\rightarrow\dots\rightarrow v_{k}$. In contrast, a partially directed cycle is a cycle where the edges can be directed or undirected such that (i) at least one edge is directed and (ii) all the directed edges in the cycle point to the same direction. A graph is chordal if every cycle of length 4 or more has a chord — an edge connecting two non-consecutive vertices. A partially directed acyclic graph (PDAG) is a simple graph with two types of edges, directed and undirected, such that there is no directed cycle.
2.1.2 Undirected graphs (UGs)
An undirected graph is a graph consisting of undirected edges. Suppose a random vector $X=(X_{v}:v\in\mathrm{V})$ is indexed by vertices of the undirected graph $\mathcal{G}=(\mathrm{V},\mathrm{E})$. An undirected graph $\mathcal{G}$ defines a probabilistic model for the distribution of $X$ through a set of semantics known as the Markov properties of UGs. In the literature, there are four notions of such semantics: factorization, as well as, global, local and pairwise Markov properties, which are not equivalent in general; see Lauritzen (1996, §3.2.1). The exact choice is not essential for our approach. Yet, for simplicity, in Section˜2.2 we will assume regularity conditions so that these four notions are equivalent.
2.1.3 Directed acyclic graphs (DAGs)
A directed acyclic graph is a PDAG with only directed edges. Depending on the semantics chosen, a DAG can be interpreted either as a probabilistic model or a causal model. For the purpose of this paper, we only consider DAGs as causal models since distinct DAGs (over the same set of vertices) may encode the same probabilistic model — these DAGs are called Markov equivalent — and we require the map from graph to model to be injective. For probabilistic modeling, we instead use a special type of graphs to represent the Markov equivalence classes of DAGs, which is described next.
2.1.4 Completed partially directed acyclic graphs (CPDAGs)
A CPDAG, also known as an essential graph, is a partially directed acyclic graph (PDAG) that represents a set of DAGs that are Markov equivalent. By a classical result (Verma and Pearl, 1990; Andersson, Madigan and Perlman, 1997), two DAGs are Markov equivalent if they share (1) the same skeletons (e.g., edges ignoring the directions) and (2) the same set of v-structures (i.e., unshielded colliders). We have a v-structure at a vertex $v_{2}$ if there are two edges, e.g., $v_{1}\rightarrow v_{2}$ and $v_{2}\leftarrow v_{3}$, pointing to $v_{2}$, but there is no edge between $v_{1}$ and $v_{3}$. Given a class of Markov equivalent DAGs, the CPDAG $\mathcal{G}$ representing the class can be obtained as follows. First, let $\mathcal{G}$ have the same skeleton as any (and hence every) DAG in the class. Then, for each edge in the skeleton, orient the edge accordingly if the corresponding edge points to the same direction in every DAG in the equivalence class; otherwise, let the edge be undirected. In the rest of this paper, we only consider a CPDAG $\mathcal{G}$ as the probabilistic model associated with any (and consequently every) DAG in the corresponding Markov equivalent class. It is worth mentioning that CPDAGs, or a larger type of graphs introduced below, can also be used to represent causal models. Not every PDAG is a valid CPDAG; see Section˜B.1 for details.
2.1.5 Maximally oriented partially directed acyclic graphs (MPDAGs)
A CPDAG $v_{1}-v_{2}-v_{3}$ represents three DAGs, namely $v_{1}\rightarrow v_{2}\rightarrow v_{3}$, $v_{1}\leftarrow v_{2}\leftarrow v_{3}$ and $v_{1}\leftarrow v_{2}\rightarrow v_{3}$, which are equivalent probabilistic models but distinct causal models. In the causal context, suppose we also have the background knowledge that $v_{2}$ temporally precedes $v_{3}$. Then, we can narrow down to $v_{1}\rightarrow v_{2}\rightarrow v_{3}$ or $v_{1}\leftarrow v_{2}\rightarrow v_{3}$, which together can be represented by a PDAG $v_{1}-v_{2}\rightarrow v_{3}$. PDAGs like this are MPDAGs, which represent the conjunction of several distinct causal models that are observationally Markov equivalent. They generalize CPDAGs by incorporating additional background (or experimental) knowledge on the orientation of certain edges (Hauser and Bühlmann, 2012; Perkovic, 2020; Hauser and Bühlmann, 2015). In the context of causal modeling, CPDAGs are special cases of MPDAGs (without any additional background knowledge). Therefore, in the rest of this paper, we will only focus on the probabilistic models represented by CPDAGs and the causal models represented by MPDAGs. Not every PDAG is valid MPDAG; see Section˜B.2 for structural constraints that an MPDAG must obey.
Figure 2: (a) A CPDAG representing three Markov equivalent DAGs. A subset of two DAGs can be represented by an MPDAG. (b) Structures that are forbidden from a valid MPDAG.
2.2 Graphs as statistical models
For the four types of graphs we introduced, we describe the models they represent. To ensure precision in deriving the corresponding model-oriented distances we do so rigorously, but the reader may set aside the technical details on first reading.
2.2.1 Graphs as probabilistic models
UGs and CPDAGs represent probabilistic models for a random vector $(X_{v}:v\in\mathrm{V})$. Suppose $X$ takes value in a product measurable space $\mathcal{X}=\prod_{v\in\mathrm{V}}\mathcal{X}_{v}$. Let $\mu=\prod_{v\in V}\mu_{v}$ be a product measure over $\mathcal{X}$. Let $P$ be the law of $X$. We use the nonparametric model
$\mathcal{M}^{\mathrm{o}}:=\left\{P:\text{$P$ is absolutely continuous with respect to $\mu$}\right\}$
as our ground set (here $\mathrm{o}$’ stands for observational) so all the models are subsets of $\mathcal{M}^{\mathrm{o}}$. For $P\in\mathcal{M}^{\mathrm{o}}$, we use $p:=\mathop{}\!\mathrm{d}P/\mathop{}\!\mathrm{d}\mu$ to denote its density.
For UGs, as explained in Section˜2.1.2, we pose additional regularity condition on the ground set such that four Markov properties all lead to the same model definition. Let the ground set be
$\mathcal{M}^{\mathrm{o}}_{+}:=\left\{P\in\mathcal{M}^{\mathrm{o}}:p>0\text{ $\mu$-a.e.}\right\}.$
Let $\mathcal{C}(\mathcal{G})$ be the set of cliques of $\mathcal{G}$. An undirected graph $\mathcal{G}$ represents the model
$\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}(\mathcal{G}):=\left\{P\in\mathcal{M}^{\mathrm{o}}_{+}:p(x)=\prod_{\mathrm{C}\in\mathcal{\mathrm{C}}(\mathcal{G})}\varphi_{\mathrm{C}}(x_{\mathrm{C}})\text{ for functions $\varphi_{\mathrm{C}}$}\right\},$
i.e., densities that factorize in terms of the cliques of $\mathcal{G}$. It can be shown that the model can also be written in terms of UG’s global Markov property (Lauritzen, 1996, §3.2.1) as
[…]
For a CPDAG $\mathcal{G}$, it represents the probabilistic model consisting of densities that factorize according any (and hence every) DAG $\mathcal{D}$ represented by $\mathcal{G}$ (i.e., $\mathcal{D}$ is in the Markov equivalence class represented by $\mathcal{G}$), given by
$\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}}(\mathcal{G}):=\left\{P\in\mathcal{M}^{\mathrm{o}}:\,p(x)=\prod_{v\in{V}}p_{v}(x_{v}|x_{{\text{pa}_{\mathcal{D}}(v)}})\text{ for some DAG $\mathcal{D}$ represented by $\mathcal{G}$}\right\},$
where $p_{v}$ is a conditional density function. This model can be equivalently described in terms of the global or local Markov property of any DAG represented by $\mathcal{G}$; this holds without requiring positivity, so we choose $\mathcal{M}^{\mathrm{o}}$ as the ground set above. The model is equivalent to the following global Markov property (Lauritzen, 1996, Theorem 3.27):
[…]
For separation and d-separation, the reader is referred to Lauritzen (1996, §3.2.1-2).
2.2.2 Graphs as causal models
We follow Zhao (2025) in defining causal models associated with DAGs and MPDAGs. Consider a causal system that generates observations of a random vector $X=(X_{v}:v\in\mathrm{V})$, which takes value in a product measurable space $\mathcal{X}\equiv\mathcal{X}_{\mathrm{V}}$. Here we use notation $\mathcal{X}_{I}:=\prod_{v\in\mathrm{I}}\mathcal{X}_{v}$ for every $\mathrm{I}\subseteq\mathrm{V}$. A causal model is a set of joint distributions of counterfactual variables as well as factual (i.e., observed) variables, both of which are part of the potential outcome schedule:
$X(\cdot):=\left(X_{v}(x_{I}):v\in V,I\subseteq V,x_{I}\in\mathcal{X}_{I}\right),$
where $X_{v}(x_{I})$ is the potential outcome of variable $v$ under the intervention that imposes value $x_{I}$ on $X_{I}$. In the special case when $I=\emptyset$, the margin $(X_{v}(\emptyset):v\in V)$ consists of all the factual variables, whose law is called the observed distribution. The potential outcome schedule is a stochastic process that takes value in $\mathcal{X}^{\prod_{I\subseteq V}\mathcal{X}_{I}}$, which can be equipped with a suitable $\sigma$-algebra to form a measurable space. We use $P$ to denote the law of $X(\cdot)$ (i.e., a probability measure over the measurable space). Similar to the definition of $\mathcal{M}^{\mathrm{o}}$, we consider the ground causal model to be
$\mathcal{M}^{\mathrm{c}}:=\big\{P:P\left(X(x_{I},x_{I^{\prime}})=X(x_{I})\mid X_{I^{\prime}}(x_{I})=x_{I^{\prime}}\right)=1,\\ \text{ for every disjoint $I,I^{\prime}\subseteq V$ and every $x\in\mathcal{X}$}\big\},$ (3)
where $\mathrm{c}$’ stands for causal. In words, $\mathcal{M}^{\mathrm{c}}$ is the set of distributions of $X(\cdot)$ that maintain consistency through recursive substitution (see, e.g., Shpitser, Richardson and Robins, 2022, §41.2): in a world that sets $X_{I}$ to be $x_{I}$, if we observe $X_{I^{\prime}}$ to be $x_{I}^{\prime}$, then the observed $X$ equals $X(x_{I},x_{I^{\prime}})$. Any causal model that we discuss next will be a subset of $\mathcal{M}^{\mathrm{c}}$.
A DAG $\mathcal{D}$ represents the causal model
$\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}(\mathcal{D}):=\left\{P\in\mathcal{M}^{\mathrm{c}}:\text{$P$ is causal Markov with respect to $\mathcal{D}$}\right\}.$ (4)
Let $\operatorname{Pa}_{\mathcal{D}}(v)$ denote the parent set of $v$ in $\mathcal{D}$. We say $P$ is causal Markov with respect to $\mathcal{D}$ if it satisfies the following two properties:
  1. 1.
    For every $v\in\mathrm{V}$, $\mathrm{I}\subseteq\mathrm{V}$ and $x\in\mathcal{X}$, $X_{v}(x_{\mathrm{I}})=X_{v}(x_{\mathrm{I}\cap\operatorname{Pa}_{\mathcal{D}}(v)})$ holds with probability 1;
  2. 2.
    For every $x\in\mathcal{X}$, the one-step-ahead counterfactuals $\left\{X_{v}(x_{\operatorname{Pa}_{\mathcal{D}}(v)}):v\in V\right\}$ are mutually independent.
In conjunction with Eq.˜3, these two properties imply that under any intervention (or no intervention), $(X_{v}(x):v\in\mathrm{V})$ can be constructed from a set of independent basic building blocks, the one-step-ahead counterfactuals under that intervention, through recursive substitution. Note that the second property does not impose cross-world independence of the counterfactuals; see also Robins (1986); Richardson and Robins (2013).
An MPDAG $\mathcal{G}$ represents the set of Markov equivalent DAGs that also obey any extra orientation encoded by $\mathcal{G}$. A DAG $\mathcal{D}$ is represented by $\mathcal{G}$ if $\mathcal{D}$ can be obtained by orienting the undirected edges in $\mathcal{G}$ without introducing any directed cycles or v-structures. The causal model represented by MPDAG $\mathcal{G}$ is simply defined in terms of a disjunction:
$\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}(\mathcal{G})=\bigcup\big\{\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}(\mathcal{D}):\text{$\mathcal{D}$ is a DAG represented by $\mathcal{G}$}\big\}.$
3 A general recipe for model-oriented distance using partially ordered sets
Let $\mathfrak{G}$ be a finite collection of graphs over a common set of vertices $\mathrm{V}$. Let $\mathcal{M}$ be an associated model map (i.e., semantics) that assigns a model to each graph in the class. Given $(\mathfrak{G},\mathcal{M})$, our goal is to define an $\mathcal{M}$-oriented distance $d(\mathcal{G}_{s},\mathcal{G}_{t})$ for every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$. For example, for $\mathrm{V}=\{1,2,3,4,5\}$, we can consider defining distances for $(\mathfrak{G}_{1},\mathcal{M}^{\mathrm{o}}_{\textsc{ug}})$ and $(\mathfrak{G}_{2},\mathcal{M}^{\mathrm{o}}_{\textsc{ug}})$ with
[…]
In Section 3.1, we use the model map $\mathcal{M}$ to define our model-oriented distance between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$. In Section 3.2, we provide an equivalent but more constructive characterization of this distance using the poset underlying $\mathfrak{G}$. Finally, in Section 3.3, we describe how the poset structure may enable faster computation of the model-oriented distance.
3.1 Model-oriented distance
Our approach for defining a model-oriented distance is based on the following notion of neighborhood.
Definition 1 (Neighbor).
For $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ and $\mathcal{G}_{1}\neq\mathcal{G}_{2}$, we call $\mathcal{G}_{1}$ a neighbor of $\mathcal{G}_{2}$ if either $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$ or $\mathcal{M}(\mathcal{G}_{2})\subseteq\mathcal{M}(\mathcal{G}_{1})$ holds, and there does not exist a third, different graph $\mathcal{G}_{3}\in\mathfrak{G}$ such that $\mathcal{M}(\mathcal{G}_{3})$ is nested between $\mathcal{M}(\mathcal{G}_{1})$ and $\mathcal{M}(\mathcal{G}_{2})$.
By construction, neighborhood is a symmetric relation. Intuitively, $\mathcal{G}_{2}$ is a neighbor of $\mathcal{G}_{1}$ if it represents the simplest refinement or coarsening of $\mathcal{G}_{1}$’s model, with no graph lying strictly between them in the model hierarchy. The model-oriented distance between two graphs counts the minimal number of such basic refinement or coarsening operations to transform one model into the other, which we define as follows.
Definition 2 (Model-oriented distance).
Given $(\mathfrak{G},\mathcal{M})$, for $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$, the model-oriented distance between $(\mathcal{G}_{s},\mathcal{G}_{t})$ is the minimal number of steps through neighbors to move from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$.
We would like our distance to be a metric, i.e., a distance function $d:\mathfrak{G}\times\mathfrak{G}\to[0,\infty)$ that satisfies the following conditions:
  1. 1.
    Symmetry: $d(\mathcal{G}_{s},\mathcal{G}_{t})=d(\mathcal{G}_{t},\mathcal{G}_{s})$,
  2. 2.
    Anti-symmetry: $d(\mathcal{G}_{s},\mathcal{G}_{t})=0$ if and only if $\mathcal{G}_{s}=\mathcal{G}_{t}$,
  3. 3.
    Triangle inequality: $d(\mathcal{G}_{s},\mathcal{G}_{t})\leq d(\mathcal{G}_{s},\mathcal{G})+d(\mathcal{G}_{t},\mathcal{G})$ for any $\mathcal{G}\in\mathfrak{G}$.
The significance of the symmetry and anti-symmetry properties is straightforward. We illustrate the importance of triangle inequality from both statistical and algorithmic perspectives. From a statistical perspective, when proving consistency or convergence of an estimator, one often introduces an intermediate “ideal" object (e.g., the population version of the estimator) to bridge the gap between the data and the truth. The triangle inequality ensures that the distance between an estimated graph and the true graph can be controlled via the intermediate graph. From an algorithmic perspective, the triangle inequality helps in enumerating graphs within a small distance of a reference graph. For example, when the distance takes integer values, one can use the triangle inequality to enumerate all graphs within distance two of a reference graph by first finding all graphs at distance one, and then enumerating all graphs that are distance one from those. Without the triangle inequality, this decomposition is no longer possible, making enumeration much more expensive.
For the model-oriented distance to be a metric, we require the following two conditions.
Condition 1 (Injectivity).
For any $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$ and $\mathcal{G}_{s}\neq\mathcal{G}_{t}$, we have $\mathcal{M}(\mathcal{G}_{s})\neq\mathcal{M}(\mathcal{G}_{t})$.
Condition 2 (Connectedness).
For every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$, there exists a finite sequence of graphs $\mathcal{G}_{s}=\mathcal{G}_{1},\dots,\mathcal{G}_{k}=\mathcal{G}_{t}$ with $k\geq 1$ such that every $\mathcal{G}_{i},\mathcal{G}_{i+1}$ are neighbors.
Theorem 1 (Metric).
The model-oriented distance in Definition˜2 is a metric if and only if $(\mathfrak{G},\mathcal{M})$ satisfies Conditions˜1 and 2.
Proof.
Let the model-oriented distance be denoted by $d$. Under Condition˜1, function $d$ can be viewed as the shortest-path distance, which is a metric, defined on an undirected graph with vertex set $\mathfrak{G}$ and edges incident to neighbors. Condition˜2 ensures that the graph is connected so $d$ is finite. For the reverse direction, if Condition˜1 fails then there exists $\mathcal{G}_{s}\neq\mathcal{G}_{t}$ with $d(\mathcal{G}_{s},\mathcal{G}_{t})=0$. If Condition˜2 fails, then there exist $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathcal{G}$ with $d(\mathcal{G}_{1},\mathcal{G}_{2})$ undefined. ∎
For natural choices of $(\mathfrak{G},\mathcal{M})$, we argue that Conditions˜1 and 2 are easily satisfied. In fact, in Section˜A.1, we show that all the four models $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}},\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}},\mathcal{M}^{\mathrm{c}}_{\textsc{dag}},\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}$ discussed in this paper satisfy Condition˜1. Typically, Condition˜2 can be ensured by the checking either of the two following stronger condition.
Condition 3 (Least element).
There exists a graph $\mathcal{G}_{\hat{0}}\in\mathfrak{G}$ such that $\mathcal{M}(\mathcal{G}_{\hat{0}})\subseteq\mathcal{M}(\mathcal{G})$ for every $\mathcal{G}\in\mathfrak{G}$.
When it exists, $\mathcal{G}_{\hat{0}}$ is the unique least element in $\mathfrak{G}$ and is usually taken to be the empty graph, which represents the smallest model. This condition has the following dual.
Condition 3 (Greatest element).
There exists a graph $\mathcal{G}_{\hat{1}}\in\mathfrak{G}$ such that $\mathcal{M}(\mathcal{G}_{\hat{1}})\supseteq\mathcal{M}(\mathcal{G})$ for every $\mathcal{G}\in\mathfrak{G}$.
Defining distance through the minimal number of basic operations to transform one object to the other is common in other combinatorial settings. In the context of rankings or permutations, the Kendall’s tau distance (Kendall, 1938) counts the smallest number of adjacent swaps needed to align two rankings. In clustering, the variation of information (Meila, 2005) counts the smallest weighted number of merges or splits needed to align two partitions. In both cases, the distance reflects how many basic, interpretable operations are required to transform one model into another. Here, we use the notion of neighbor to provide a parsimonious and model-oriented definition of what the basic operations are. In what follows, we use posets to formalize our approach.
3.2 Formulation of model-oriented distance via posets
Definition 2 provides a succinct yet principled strategy for defining a model-oriented distance across a wide range of graph classes. However, since the definition is implicit, it is not immediately clear how to compute the distance for any specific class. We now provide a constructive characterization of the model-oriented distance which we both use to study under what conditions it simplifies (see Sections 3.3 and 4) and to develop algorithmic tools for computing it (see Section 5).
Definition 3 (Poset, cover, connectedness, least element, comparability).
A poset (partially ordered set) $\mathcal{L}:=(L,\preceq)$ is a collection $L$ of elements and a relation $\preceq$ that is reflexive ($x\preceq x,~\forall x\in L$), transitive ($x\preceq y,y\preceq z\Rightarrow x\preceq z,~\forall x,y,z\in L$), and anti-symmetric ($x\preceq y,y\preceq x\Rightarrow x=y,~\forall x,y\in L$). We use notation $x\prec y$ if $x\preceq y$ and $x\neq y$. An element $y\in L$ covers $x\in L$, denoted as $x\lessdot y$, if $x\prec y$ and there is no $z\in L$ with $x\prec z\prec y$. The poset $\mathcal{L}$ is connected if for every $x,y\in L$, there exists a finite sequence of $k\in\mathbb{N}_{+}$ elements $x=x_{1},x_{2},\dots,x_{k}=y$ such that either $x_{i+1}\gtrdot x_{i}$ or $x_{i+1}\lessdot x_{i}$ for every $i=1,2,\dots,k-1$. An element $x\in L$ is called the least (greatest) element if $x\preceq y$ ($x\succeq y$) for all $y\in L$. Elements $x,y\in\mathcal{L}$ are said to be comparable if either $x\preceq y$ or $y\succeq x$; otherwise, they are incomparable.
Definition 4 (Model-oriented poset).
Given a graph class $\mathfrak{G}$ and an associated model map $\mathcal{M}$ that satisfy Condition˜1, the corresponding model-oriented poset is $\mathcal{L}:=(\mathfrak{G},\preceq)$ with $\preceq$’ given by model containment: for $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, we say $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ if $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$.
Reflexivity and transitivity of $\preceq$’ hold by definition, while the anti-symmetry follows from the injectivity of $\mathcal{M}(\cdot)$ (Condition˜1). The poset $\mathcal{L}$ offers a natural way to organize the graphs $\mathcal{G}\in\mathfrak{G}$ based on the models they represent. The neighbors in Definition˜1 are exactly the covering relations in the poset: $\mathcal{G}_{2}$ is a neighbor of $\mathcal{G}_{1}$ if either $\mathcal{G}_{2}$ covers $\mathcal{G}_{1}$ or $\mathcal{G}_{1}$ covers $\mathcal{G}_{2}$. Note that Condition˜2 is equivalent to requiring that the poset $\mathcal{L}$ is connected, which is typically ensured by the existence of a least element $\mathcal{G}_{\hat{0}}$ (usually the empty graph) in $\mathfrak{G}$.
Figure 3: Hasse diagrams for different statistical graphs over vertex set $\{1,2,3\}$: each box represents a graph, and each dashed arrow represents a covering relation, i.e., we draw $\mathcal{G}_{1}\dashrightarrow\mathcal{G}_{2}$ pointing upwards if $\mathcal{G}_{2}$ covers $\mathcal{G}_{1}$. We can determine whether $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ by checking whether there is a directed path from $\mathcal{G}_{1}$ to $\mathcal{G}_{2}$. In (b) and (d), only part of the poset is shown for simplicity. Each poset has a least element, but not all have a greatest element. The poset in panel (d) is not graded when the graphs have more than three vertices (see Proposition˜10). This means there is no consistent way to assign ranks to the graphs.
For example, consider $\mathfrak{G}$ to be all the UGs over the vertex set $\{1,2,3\}$ and let $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}$ be the model map. In this case, the partial order $\preceq$’ can be characterized graphically: for $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, one can show that $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ if and only if $\mathcal{G}_{1}$ is a subgraph of $\mathcal{G}_{2}$.
Further, $\mathcal{G}_{2}$ covers $\mathcal{G}_{1}$ if and only if $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ and $\mathcal{G}_{2}$ has exactly one more edge than $\mathcal{G}_{1}$. The least element $\mathcal{G}_{\hat{0}}$ in the poset is the empty graph. This poset can be visualized with a Hasse diagram shown in Fig.˜3(a): each box is a a graph; if $\mathcal{G}_{2}$ covers $\mathcal{G}_{1}$, we then place $\mathcal{G}_{2}$ above $\mathcal{G}_{1}$ and draw an upward arrow $\mathcal{G}_{1}\dashrightarrow\mathcal{G}_{2}$. Partial order can be read off from the diagram: $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ if there exists a directed path (can be of length zero) from $\mathcal{G}_{1}$ to $\mathcal{G}_{2}$. Hence, the poset $\mathcal{L}=(\mathfrak{G},\preceq)$ can be compactly represented by its Hasse diagram, which is directed acyclic graph over $\mathfrak{G}$. Finally, it is not hard to see that our model-oriented distance, denoted hereafter as $d_{\mathcal{L}}$ to highlight its dependence on $\mathcal{L}$ and hence $(\mathfrak{G},\mathcal{M})$, is given by
[…]
where the path can contain dashed arrows in both directions.
We note that the ‘shortest path length’ is a standard way to measure distance in posets (Monjardet, 1981; Trotter, 1992; Brualdi, Graves and Lawrence, 1995), although its use has primarily been confined to the order theory literature.
3.3 Poset structures that facilitate the computation of model-oriented distance
Having established that the model-oriented distance corresponds to the shortest-path distance in the Hasse diagram of the model-oriented poset, we can compute it by using using the well-known $A^{*}$ algorithm. In general, however, this can be expensive, particularly because it requires repeatedly identifying all neighbors of a graph while traversing many candidate paths in the Hasse diagram. Often, however, we can exploit additional structure in the poset to simplify computation — either by deriving a closed-form expression for the distance (eliminating the need for $A^{*}$), by streamlining the neighborhood search, or by reducing the number of paths that must be considered. In the following, we study several such structural properties in general, as we will rely on them when analyzing the model-oriented distances corresponding to probabilistic UG and CPDAG models, as well as causal DAGs and MPDAG models (see Section 4). We also examine the conditions under which the model-oriented distance reduces to an edge-by-edge distance similarly to the SHD.
3.3.1 Graded poset
A connected poset $\mathcal{L}$ is graded if there exists a function $\operatorname{\mathrm{rank}}:\mathfrak{G}\rightarrow\mathbb{N}$ such that $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$ implies $\operatorname{\mathrm{rank}}(\mathcal{G}_{2})=\operatorname{\mathrm{rank}}(\mathcal{G}_{1})+1$ and $\operatorname{\mathrm{rank}}(\mathcal{G})=0$ for a minimal $\mathcal{G}\in\mathfrak{G}$. It can be shown that if it exists, the rank function is unique. A graph $\mathcal{G}^{\prime}\in\mathfrak{G}$ is called a minimal element if there exists no $\mathcal{G}\in\mathfrak{G}$ such that $\mathcal{G}\prec\mathcal{G}^{\prime}$ ($\mathcal{G}\succ\mathcal{G}^{\prime}$).
If the model-oriented poset is graded, we can exploit this property to simplify computation in two ways. First, we obtain the following closed form characterization of the model-oriented distance between graphs comparable within the poset.
Proposition 1.
Suppose $\mathcal{L}=(\mathfrak{G},\preceq)$ is a graded poset with rank function $\mathrm{rank}(\cdot)$. For any $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$ with $\mathcal{G}_{s}\succeq\mathcal{G}_{t}$, every downward path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ has length $\mathrm{rank}(\mathcal{G}_{s})-\mathrm{rank}(\mathcal{G}_{t})$, which also equals $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
This is a well-known result that follows from the property that in a graded poset every maximal chain that is bounded between two comparable elements are of the same length. A subset of $\mathfrak{G}$ is a chain (i.e., totally ordered set) if every pair of elements are comparable. A chain is maximal if it is not contained by a larger chain. For completeness, we provide a short proof in Section˜C.1.
Second, gradedness makes it significantly easier to enumerate all neighbors of a given graph $\mathcal{G}$. Recall that $\mathcal{G}^{\prime}$ is a neighbor of $\mathcal{G}$ in $\mathcal{L}$ if either $\mathcal{G}^{\prime}\preceq\mathcal{G}$ and there is no distinct graph $\tilde{\mathcal{G}}$ with $\mathcal{G}^{\prime}\preceq\tilde{\mathcal{G}}\preceq\mathcal{G}$, or $\mathcal{G}\preceq\mathcal{G}^{\prime}$ and there is no distinct graph $\tilde{\mathcal{G}}$ with $\mathcal{G}\preceq\tilde{\mathcal{G}}\preceq\mathcal{G}^{\prime}$. In general, enumerating all neighbors can be cumbersome, since one must check for the absence of these intermediate elements. If $\mathcal{L}$ is graded, this process becomes much simpler since two graphs can only be neighbors if their ranks differ by exactly one. Combined with the requirement that one graph is either above or below the other in the poset, this condition becomes both necessary and sufficient. This characterization is particularly useful when the rank function has a simple form. For instance, we will show in Section 4.1 that the poset underlying undirected graphs (UGs) is graded with the rank of a graph equal to its number of edges. All neighbors of a UG $\mathcal{G}$ are therefore UGs $\mathcal{G}^{\prime}$ that differ by exactly one edge and are comparable to $\mathcal{G}$. In a non-graded poset, neighboring graphs may differ more substantially; for example, we show in Section 4.4 that for causal MPDAGs, a neighboring graph may have many more edges.
3.3.2 Down-up and up-down distance, semi-lattice, semimodularity
To reduce the number of paths we need to explore, it is natural to consider the shortest path among a restricted set of paths between two graphs in the Hasse diagram. Specifically, for $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$, we consider the set of down-up and up-down paths between them, denoted by $\mathrm{P}_{\downarrow\uparrow}{(\mathcal{G}_{s},\mathcal{G}_{t})}$ and $\mathrm{P}_{\uparrow\downarrow}{(\mathcal{G}_{s},\mathcal{G}_{t})}$ respectively. A path is down-up if it takes the form of
$\mathcal{G}_{s}=x_{0}\gtrdot\cdots\gtrdot x_{i^{\ast}}\lessdot\cdots\lessdot x_{d}=\mathcal{G}_{t},$
where $i^{\ast}\in\{0,\dots,d\}$ is the inflection point; chains ($i^{\ast}=0$ or $i^{\ast}=d$) are special cases of down-up paths. Conversely, an up-down path takes the form of
$\mathcal{G}_{s}=x_{0}\lessdot\cdots\lessdot x_{i^{\ast}}\gtrdot\cdots\gtrdot x_{d}=\mathcal{G}_{t}.$
The next result characterizes when down-up (up-down) paths connect every pair of graphs.
Proposition 2.
The set of down-up paths $\mathrm{P}_{\downarrow\uparrow}{(\mathcal{G}_{1},\mathcal{G}_{2})}$ is non-empty for all $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ if and only if $\mathcal{L}$ has a least element (Condition˜3). Similarly, the set of up-down paths $\mathrm{P}_{\uparrow\downarrow}{(\mathcal{G}_{1},\mathcal{G}_{2})}$ is non-empty for all $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ if and only if $\mathcal{L}$ has a greatest element (Condition˜3).
Proof.
It is clear that the existence of a least element guarantees a down-up path between any pair of graphs in $\mathfrak{G}$, since all elements share a common lower bound. Conversely, suppose $\mathcal{L}$ does not have a least element. Then there must exist at least two minimal elements $\mathcal{G}_{1},\mathcal{G}_{2}$ ($\mathcal{G}_{1}\neq\mathcal{G}_{2}$), which share no common lower bound. Hence, no down-up path can exist between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. The proof for the up-down paths follows similarly. ∎
The posets of all graph types (UGs, DAGs, CPDAGs, and MPDAGs) we consider in this paper have the empty graph as the least element. Further, except for DAGs, all of them also have a greatest element. See Table˜1 for a summary.
The down-up and up-down paths lead to the definition of down-up and up-down distances.
Definition 5 (Down-up and up-down distance).
Under Condition˜3, the down-up distance is defined as $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t}):=\min\{\mathrm{length}(p):p\in\mathrm{P}_{\downarrow\uparrow}{(\mathcal{G}_{s},\mathcal{G}_{t})}\}$ for every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$. Under Condition˜3, the up-down distance is defined as $d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t}):=\min\{\mathrm{length}(p):p\in\mathrm{P}_{\uparrow\downarrow}{(\mathcal{G}_{s},\mathcal{G}_{t})}\}$ for every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$.
Unlike $d_{\mathcal{L}}$, the down-up and up-down distances may not be metrics as they are not guaranteed to obey the triangle inequality. Nevertheless, by definition, they are upper bounds on $d_{\mathcal{L}}$ and are generally easier to compute. In particular, when $\mathcal{L}$ is graded, the down-up or up-down distances take a closed form.
Proposition 3.
Suppose $\mathcal{L}=(\mathfrak{G},\preceq)$ is graded with rank function $\mathrm{rank}(\cdot)$. Then, under Condition˜3, we have
$d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})=\operatorname{\mathrm{rank}}(\mathcal{G}_{s})+\operatorname{\mathrm{rank}}(\mathcal{G}_{t})-2\max_{\mathcal{G}\preceq\mathcal{G}_{s},\mathcal{G}_{t}}\operatorname{\mathrm{rank}}(\mathcal{G}),\quad\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}.$
Similarly, under Condition˜3, we have
$d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t})=2\min_{\mathcal{G}\succeq\mathcal{G}_{s},\mathcal{G}_{t}}\operatorname{\mathrm{rank}}(\mathcal{G})-\operatorname{\mathrm{rank}}(\mathcal{G}_{s})-\operatorname{\mathrm{rank}}(\mathcal{G}_{t}),\quad\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}.$
Proof.
Any down-up path between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ moves downward from $\mathcal{G}_{s}$, and then upward from an inflection point $\mathcal{G}$ to $\mathcal{G}_{t}$. Since $\mathcal{G}\preceq\mathcal{G}_{s}$ and $\mathcal{G}\preceq\mathcal{G}_{t}$, Proposition 1 implies that the length of the path is $\mathrm{rank}(\mathcal{G}_{s})+\mathrm{rank}(\mathcal{G}_{t})-2\mathrm{rank}(\mathcal{G})$. Therefore, the shortest down-up path corresponds to an inflection point $\mathcal{G}$ that maximizes $\mathrm{rank}(\mathcal{G})$ over all $\mathcal{G}\preceq\mathcal{G}_{s},\mathcal{G}\preceq\mathcal{G}_{t}$. The proof for the up-down distance follows similarly. ∎
The following additional simplification requires notions of meet and join semi-lattice structures. In particular, graph $\mathcal{G}_{3}\in\mathfrak{G}$ is the least upper bound (or join) of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, denoted as $\mathcal{G}_{1}\vee\mathcal{G}_{2}$, if $\mathcal{G}_{3}\succeq\mathcal{G}_{1},\mathcal{G}_{2}$ and $\mathcal{G}_{3}\preceq\mathcal{G}^{\prime}$ for every $\mathcal{G}^{\prime}$ satisfying $\mathcal{G}^{\prime}\succeq\mathcal{G}_{1},\mathcal{G}_{2}$. The greatest lower bound (or meet) of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ is defined similarly and is denoted as $\mathcal{G}_{1}\wedge\mathcal{G}_{2}$. Poset $\mathcal{L}$ is a join (meet) semi-lattice if $\mathcal{G}\vee\mathcal{G}^{\prime}$ ($\mathcal{G}\wedge\mathcal{G}^{\prime}$) exists for every $\mathcal{G},\mathcal{G}^{\prime}\in\mathfrak{G}$.
Corollary 1.
Suppose $\mathcal{L}=(\mathfrak{G},\preceq)$ is graded with rank function $\mathrm{rank}(\cdot)$. Under Condition˜3, if $\mathcal{L}$ is also a meet semi-lattice, we have
$d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})=\operatorname{\mathrm{rank}}(\mathcal{G}_{s})+\operatorname{\mathrm{rank}}(\mathcal{G}_{t})-2\operatorname{\mathrm{rank}}(\mathcal{G}_{s}\wedge\mathcal{G}_{t}),\quad\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}.$
Similarly, under Condition˜3, if $\mathcal{L}$ is also a join semi-lattice, we have
$d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t})=2\operatorname{\mathrm{rank}}(\mathcal{G}_{s}\vee\mathcal{G}_{t})-\operatorname{\mathrm{rank}}(\mathcal{G}_{s})-\operatorname{\mathrm{rank}}(\mathcal{G}_{t}),\quad\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}.$
Proof.
The result follows from the fact that $\mathcal{G}_{s}\wedge\mathcal{G}_{t}$ ($\mathcal{G}_{s}\vee\mathcal{G}_{t}$) has the maximum (minimum) rank among common lower (upper) bounds of $\mathcal{G}_{s},\mathcal{G}_{t}$. ∎
As described earlier, down-up or up-down distance is an upper bound on the model-oriented distance. The next key result characterizes the poset structure that makes them equal.
[…]
Figure 4: Upper and lower semimodularity of a poset. Between any $\mathcal{G}_{s},\mathcal{G}_{t}$ in a lower semimodular poset, a shortest path can always be transformed into a down-up path of the same length.
Definition 6 (Semimodularity; see Fig.˜4).
[…]
Theorem 2.
Given a graph class $\mathfrak{G}$ and a model map $\mathcal{M}$ that satisfy Condition˜1, let $\mathcal{L}=(\mathfrak{G},\preceq)$ be the model-oriented poset. Under Condition˜3 (least element), $\mathcal{L}$ is lower semimodular if and only if $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})=d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ holds for every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$. Similarly, under Condition˜3 (greatest element), $\mathcal{L}$ is upper semimodular if and only if $d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t})=d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ holds for every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$.
We prove Theorem˜2 in Appendix˜D. Fig.˜4 illustrates the basic idea: when $\mathcal{L}$ is semimodular, any shortest path can be transformed into a down-up or up-down path with the same length. We remark that, to our best knowledge, our characterization weakens the conditions for existing results in the literature and is well-suited for studying model-oriented graph distances. For example, Monjardet (1981, Theorem 8) additionally requires the poset to be graded, while Foldes and Radeleczki (2021, Proposition 2.5) instead requires the poset to be a semi-lattice. However, the model-oriented posets for graphs may not satisfy these conditions. In fact, in Section˜4.3 we will use Theorem˜2 to study polytree CPDAGs, where these earlier results are inapplicable.
3.3.3 Relation to edge-by-edge distances
Recall that SHDs measure the distance between two graphs by counting how many edges are mismatched. They are therefore by construction edge-by-edge style distance, in the sense that they can be computed considering each edge at a time. Different types of mismatches are often assigned different weights (Perrier, Imoto and Miyano, 2008). For PDAGs, for example, there are two common variants of SHD (Tsamardinos, Brown and Aliferis, 2006) defined as follows. Consider encoding a PDAG $\mathcal{G}$ over $\mathrm{V}$ with adjacency matrix
$A_{i,j}:=\begin{cases}1,\quad&\text{$i-j$ or $i\rightarrow j$}\\ 0,\quad&\text{otherwise}\end{cases},\quad i,j\in\mathrm{V}.$
Definition 7 (Structural Hamming distance).
Let PDAGs $\mathcal{G}_{s},\mathcal{G}_{t}$ over a vertex set be encoded by adjacency matrix $A^{s},A^{t}$ respectively. The two variants of SHD are given by
$\begin{aligned} \operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})&:=\sum_{i<j}\mathbb{I}\left[A^{s}_{i,j}\neq A^{t}_{i,j}\text{ or }A^{s}_{j,i}\neq A^{t}_{j,i}\right],\\ \operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})&:=\sum_{i<j}\mathbb{I}\left[A^{s}_{i,j}\neq A^{t}_{i,j}\right]+\mathbb{I}\left[A^{s}_{j,i}\neq A^{t}_{j,i}\right].\end{aligned}$
Here, $\mathbb{I}[\cdot]$ is the indicator function. In the following corollary to Theorem˜2, we study for general graph-types when the model-oriented distance also behaves edge-by-edge, and discuss the implications for certain PDAGs considered in this work.
Lemma 1.
Suppose we consider graphs $\mathfrak{G}$ with $k$ edge types (see Section 2.1.1) that define a poset $\mathcal{L}$, in which they are ordered based on subgraph containment; that is, $\mathcal{G}_{1}=(V,E^{1}_{1}\cup\dots\cup E_{1}^{k})\preceq\mathcal{G}_{2}=(V,E^{1}_{2}\cup\dots\cup E^{k}_{2})$ if $E^{i}_{1}\subseteq E^{i}_{2}$ for all $i$. Further, suppose $\mathcal{L}$ is graded with $\mathrm{rank}(\mathcal{G})=\sum_{i=1}^{k}|E_{i}|$ for any $\mathcal{G}\in\mathfrak{G}$, and that it is lower semimodular. Then, $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=\sum_{i=1}^{k}|E^{1}_{i}\triangle E^{2}_{i}|$, where $\triangle$ denotes the symmetric set difference. If the graph class contains only undirected edges, this is equal to $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})$. If the class contains only directed edges, this is equal to $\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})$.
Proof.
Since the poset is lower semimodular, we have from Theorem 2 that $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$. By construction, $\mathcal{L}$ is a meet semi-lattice with $\mathcal{G}_{1}\wedge\mathcal{G}_{2}=(V,E_{1}^{1}\cap E_{2}^{1}\cup\dots\cup E_{1}^{k}\cap E_{2}^{k})$. Appealing to Corollary 1, we then have that: $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=\mathrm{rank}(\mathcal{G}_{s})+\mathrm{rank}(\mathcal{G}_{t})-2\mathrm{rank}(\mathcal{G}_{s}\wedge\mathcal{G}_{t})=\sum_{i=1}^{k}|E^{1}_{i}\triangle E^{2}_{i}|$. ∎
Effectively, this shows that the model-oriented distance is an edge-by-edge distance whenever the partial ordering can be determined by examining one edge at a time (subgraph containment) and the Hasse diagram is sufficiently dense (lower semimodularity). We will see that this typically holds for graph classes that impose few constraints, such as UGs and DAGs, but not for CPDAGs or MPDAGs. Which version of the SHD it coincides with then depends on whether the edges are directed or undirected. In particular, as we will see in the next section, for UGs with model map $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}$, the model-oriented distance is equivalent to $\operatorname{\mathrm{SHD}}_{1}$. However, the model-oriented distance diverges from $\operatorname{\mathrm{SHD}}_{1}$ for all other graph types considered in this paper.
4 Model-oriented distance $d_{\mathcal{L}}$ in probabilistic and causal graphs
Next, we apply the poset framework described in Section˜3 to derive the model-oriented distance for UGs, DAGs, CPDAGs, and MPDAGs, presented in order of increasing complexity in their associated poset. Table˜1 highlights the different attributes for each associated poset, with the class of UGs being the most structured, and the class of MPDAGs being the least structured. We will formally show these properties in this section. For each graph class, we first study the corresponding model-oriented poset, then use it to derive the model-oriented distance, and finally examine simpler characterizations of the model-oriented distance, for example in terms of the SHD, or when this is not possible approximations to the model-oriented distance.
Table 1: Structural attributes of the underlying model-oriented poset for different graph classes
Graph class Model Least element Greatest element Graded Semi-lattice Semimodular
UG prob. empty graph full graph join + meet lower + upper
DAG causal empty graph meet lower
CPDAG prob. empty graph full graph
Polytree CPDAG prob. empty graph lower
MPDAG causal empty graph full graph
Polytree MPDAG causal empty graph
4.1 Probabilistic UGs
As a warm-up, consider $\mathfrak{G}$ the class of UGs over a finite vertex set $\mathrm{V}$ representing probabilistic models in the sense of $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}$ given by Eq.˜1. The next result states that the model inclusion becomes graph inclusion, which we prove in Section˜A.2.
Proposition 4.
For UGs $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, we have $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}(\mathcal{G}_{1})\subseteq\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}(\mathcal{G}_{2})$ if and only if $\mathcal{G}_{1}$ is a subgraph of $\mathcal{G}_{2}$.
Because graph containment in $\mathfrak{G}$ is isomorphic to set containment in the power set of $\{(i,j)\in\mathrm{V}\times\mathrm{V}:i<j\}$, we have a model-oriented poset $\mathcal{L}$ that is highly structured. In fact, $\mathcal{L}$ is known as the Boolean lattice, which gives the next result. Fig.˜3(c) shows its Hasse diagram when $|\mathrm{V}|=3$.
Proposition 5.
Let $\mathcal{L}$ be the model-oriented poset associated with $(\mathfrak{G},\mathcal{M}^{\mathrm{o}}_{\textsc{ug}})$, where $\mathfrak{G}$ is the class of UGs over a finite vertex set $\mathrm{V}$. We have the following results.
  1. 1.
    $\mathcal{G}_{\hat{0}}$ is the empty graph; $\mathcal{G}_{\hat{1}}$ is the full graph.
  2. 2.
    $\mathcal{L}$ is graded with $\operatorname{\mathrm{rank}}(\cdot)$ equal to the number of edges.
  3. 3.
    For $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, we have $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$ if and only if $\mathcal{G}_{1}$ is a subgraph of $\mathcal{G}_{2}$ with exactly one fewer edge.
  4. 4.
    $\mathcal{L}$ is a lattice (i.e., both a meet and a join semi-lattice).
  5. 5.
    $\mathcal{L}$ is both lower and upper semimodular.
From these properties, we obtain a closed form for the model-oriented distance.
Corollary 2.
The model-oriented distance between two UGs is given by
$d_{\mathcal{L}}(\mathcal{G}_{1},\mathcal{G}_{2})=|\mathrm{E}_{1}\triangle\mathrm{E}_{2}|=\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{1},\mathcal{G}_{2}),\quad\mathcal{G}_{1}=(\mathrm{V},\mathrm{E}_{1}),\,\mathcal{G}_{2}=(\mathrm{V},\mathrm{E}_{2})\in\mathfrak{G}.$
Proof.
It follows from Proposition˜5 and Lemma˜1. ∎
4.2 Causal DAGs
Consider $\mathfrak{G}$ the class of DAGs over a finite vertex set representing causal models in the sense of $\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}$ given by Eq.˜4. The partial order is given by subgraph containment, stated as follows.
Proposition 6.
For DAGs $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, we have $\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}(\mathcal{G}_{1})\subseteq\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}(\mathcal{G}_{2})$ if and only if $\mathcal{G}_{1}$ is a subgraph of $\mathcal{G}_{2}$.
We prove Proposition˜6 in Section˜A.2. Using this result, the model-oriented poset can be characterized as follows; see Section˜E.1 for its proof. Fig.˜3(b) shows the Hasse diagram of this poset when the vertex set is of size 3.
Proposition 7.
Let $\mathcal{L}$ be the model-oriented poset associated with $(\mathfrak{G},\mathcal{M}^{\mathrm{c}}_{\textsc{dag}})$, where $\mathfrak{G}$ is the class of DAGs over a finite vertex set $\mathrm{V}$. We have the following results.
  1. 1.
    $\mathcal{G}_{\hat{0}}$ is the empty graph. There is no $\mathcal{G}_{\hat{1}}$.
  2. 2.
    $\mathcal{L}$ is graded with $\operatorname{\mathrm{rank}}(\cdot)$ equal to the number of edges.
  3. 3.
    For $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, we have $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$ if and only if $\mathcal{G}_{1}$ is a subgraph of $\mathcal{G}_{2}$ with exactly one fewer edge.
  4. 4.
    $\mathcal{L}$ is a meet semi-lattice.
  5. 5.
    $\mathcal{L}$ is lower semimodular.
From Propositions˜7 and 1, we obtain the model-oriented distance as follows.
Corollary 3.
The model-oriented distance between two DAGs is given by
$d_{\mathcal{L}}(\mathcal{G}_{1},\mathcal{G}_{2})=|\mathrm{E}_{1}\triangle\mathrm{E}_{2}|=\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{1},\mathcal{G}_{2}),\quad\mathcal{G}_{1}=(\mathrm{V},\mathrm{E}_{1}),\,\mathcal{G}_{2}=(\mathrm{V},\mathrm{E}_{2})\in\mathfrak{G}.$
4.3 Probabilistic CPDAGs
Let $\mathfrak{G}$ be the class of CPDAGs over a finite vertex set $\mathrm{V}$, which represent probabilistic models in the sense of $\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}}$ given by Eq.˜2. We use $\mathcal{G}]$ to denote the set of Markov equivalent DAGs represented by a CPDAG $\mathcal{G}$. The model containment is characterized by a basic result due to Chickering (2002), restated as follows.
Lemma 2 (Chickering, 2002, Theorem 4).
For CPDAGs $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$, the following are equivalent:
  1. 1.
    $\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}}(\mathcal{G}_{s})\subseteq\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}}(\mathcal{G}_{t})$.
  2. 2.
    There exist DAGs $\mathcal{D}_{s}\in[\mathcal{G}_{s}]$ and $\mathcal{D}_{t}\in[\mathcal{G}_{t}]$ such that the set of d-separations in $\mathcal{D}_{t}$ are also present in $\mathcal{D}_{s}$.
  3. 3.
    There exist DAGs $\mathcal{D}_{s}\in[\mathcal{G}_{s}]$ and $\mathcal{D}_{t}\in[\mathcal{G}_{t}]$ such that there is a sequence of DAGs $\mathcal{D}_{s}=:\mathcal{D}_{0},\mathcal{D}_{1},\dots,\mathcal{D}_{k}:=\mathcal{D}_{t}$ $(k\geq 0)$, where each $\mathcal{D}_{i}$ is obtained from $\mathcal{D}_{i-1}$ by one of the two operations: (i) reversing a covered edge, or (ii) adding an edge. For operation (i), $\mathcal{D}_{i-1}$ and $\mathcal{D}_{i}$ are Markov equivalent; for (ii), the d-separations in $\mathcal{D}_{i-1}$ contain those in $\mathcal{D}_{i}$.
Here, an edge $u\rightarrow v$ is covered in DAG $\mathcal{D}$ if $\operatorname{Pa}_{\mathcal{D}}(v)=\operatorname{Pa}_{\mathcal{D}}(u)\cup\{u\}$. Using the result above, we can derive the following properties of CPDAG’s model-oriented poset. Fig.˜3(a) shows the Hasse diagram of this poset when $|\mathrm{V}|=3$.
Proposition 8.
Let $\mathcal{L}$ be the model-oriented poset associated with $(\mathfrak{G},\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}})$, where $\mathfrak{G}$ is the class of CPDAGs over a finite vertex set $\mathrm{V}$. We have the following results.
  1. 1.
    $\mathcal{G}_{\hat{0}}$ is the empty graph; $\mathcal{G}_{\hat{1}}$ is the full graph.
  2. 2.
    For $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, we have $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$ if and only if there exist DAGs $\mathcal{D}_{1}\in[\mathcal{G}_{1}]$, $\mathcal{D}_{2}\in[\mathcal{G}_{2}]$ such that $\mathcal{D}_{1}$ is a subgraph of $\mathcal{D}_{2}$ with exactly one fewer edge.
  3. 3.
    $\mathcal{L}$ is graded with $\operatorname{\mathrm{rank}}(\cdot)$ equal to the number of edges.
Additionally, $\mathcal{L}$ is neither a (join or meet) semi-lattice nor (upper or lower) semimodular in general. For proofs of these results, see Section˜E.2.
4.3.1 Upper and lower bounds on $d_{\mathcal{L}}$
The lack of semi-lattice and semimodularity structures makes computing our model-oriented distance between CPDAGs much harder than UGs and DAGs. First, Lemma˜1 is no longer applicable and indeed, as we demonstrate in Example˜1, the SHD behaves very differently from our distance. Further, by Theorem˜2, because the poset is not semimodular, up-down and down-up distances can only provide upper bounds on the model-oriented distance. Indeed, consider the example in Fig.˜5, where the shortest path, taking a ‘zigzag’ shape, is of length 4, while both the shortest down-up and up-down paths have length 8.
Figure 5: The model-oriented distance between probabilistic CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ is 4, as shown by the shortest path in red. Meanwhile, the up-down distance is 8 (blue path) and the down-up distance is also 8 (orange path).
Due to these difficulties, in Section˜5, we propose using the $\text{A}^{\ast}$ algorithm to compute $d_{\mathcal{L}}$. As an informed search algorithm, $\text{A}^{\ast}$ relies on good, custom-made lower and upper bounds on the distance to speed up the search: the lower bound, often called the heuristic function in the literature, allows one to disregard some paths without discarding a potentially optimal path; the upper bound let one prune all the paths that are longer than the current best solution. In this section, we investigate lower and upper bounds on $d_{\mathcal{L}}$ between CPDAGs.
To formulate the lower bound we need some notation. For a PDAG $\mathcal{G}$, we use $\operatorname{sk}(\mathcal{G})$ to denote its skeleton, which is the set of edges after undirecting all the edges in $\mathcal{G}$. We also use $\operatorname{v-str}(\mathcal{G})$ to denote the set of triplets $(a,b,c)$ that form a v-structure at $b$ in $\mathcal{G}$ (see Section˜2.1.4); here we do not distinguish between $(a,b,c)$ and $(c,b,a)$. Let us fix two CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ over a vertex set $\mathrm{V}$. Let us fix two CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ over a vertex set $\mathrm{V}$. Define the set
$\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t}):=\Big\{(a,b,c)\subseteq\mathrm{V}\times\mathrm{V}\times\mathrm{V}:(a,b),(b,c)\in\operatorname{sk}(\mathcal{G}_{s})\cap\operatorname{sk}(\mathcal{G}_{t}),\\ (a,c)\not\in\operatorname{sk}(\mathcal{G}_{s})\cup\operatorname{sk}(\mathcal{G}_{t}),\text{ and }(a,b,c)\in\operatorname{v-str}(\mathcal{G}_{s})\triangle\operatorname{v-str}(\mathcal{G}_{t})\Big\},$
which consists of v-structures that appear in only one of the two graphs. The set above can always be partitioned as
$\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t})=\mathrm{S}_{1}\cup\dots\cup\mathrm{S}_{L},\quad L\geq 1,$
where two triplets $(a,b,c)$ and $(d,e,f)$ are placed in the same partition if they have any vertex in common (e.g., $a=d$ or $a=e$).
For a pair of vertices $(i,j)$, we define operations $\operatorname{OP}_{(i,j),\text{CE}}$ (cover edge) and $\operatorname{OP}_{(i,j),\text{ER}}$ (edge removal) that map a subset of $\mathrm{S}$ to a subset of $\mathrm{S}$, given by
$\begin{array}[]{ll}\operatorname{OP}_{(i,j),\text{CE}}(\mathrm{T})&:=\mathrm{T}\setminus\{(i,k,j):k\in\mathrm{V}\},\\ \operatorname{OP}_{(i,j),\text{ER}}(\mathrm{T})&:=\mathrm{T}\setminus\{(i,j,k),(j,i,k):k\in\mathrm{V}\}\end{array},\quad\mathrm{T}\subseteq\mathrm{S}.$
Here, $\operatorname{OP}_{(i,j),\text{CE}}$ removes v-structures of the form $i\rightarrow k\leftarrow j$ by adding a covering edge (see Lemma˜2) between $i$ and $j$; $\operatorname{OP}_{(i,j),\text{ER}}$ removes v-structures of the form $i\rightarrow j\leftarrow k$ or $j\rightarrow i\leftarrow k$ by removing the edge between $i$ and $j$. For each partition $\mathrm{S}_{l}$, consider the minimum number of operations taken to eliminate all the v-structures in the partition, namely
$m_{\text{op}}(\mathrm{S}_{l}):=\min\left\{m\in\mathbb{N}:\exists~\mathrm{op}_{1},\dots,\mathrm{op}_{m}\text{ such that }\mathrm{op}_{1}\circ\dots\circ\mathrm{op}_{m}(\mathrm{S}_{i})=\emptyset\right\},\quad l=1,\dots,L,$
where each $\mathrm{op}_{k}$ is either $\operatorname{OP}_{(i,j),\text{CE}}$ or $\operatorname{OP}_{(i,j),\text{ER}}$ for some $(i,j)$ that can depend on $k$. Because the v-structures from two different partitions share no common vertex by construction and this property is maintained by both operations, we conclude that
$m_{\text{op}}(\mathrm{S}):=\sum_{l=1}^{L}m_{\text{op}}(\mathrm{S}_{l})$
lower bounds the number of steps taken to align the set of v-structures between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$.
Property 1 (Upper bound).
For CPDAGs $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$, we have
$d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq\min\left\{d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t}),d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t})\right\}.$
Property 2 (Lower bound).
For CPDAGs $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$, we have
$d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\geq|\operatorname{sk}(\mathcal{G}_{s})\triangle\operatorname{sk}(\mathcal{G}_{t})|+2m_{\text{op}}(\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t})).$
We prove Property 2 in Appendix F.1. Property 1 follows from recalling that the model-oriented distance is upper bounded by both the down-up and the up-down distance (both exist for CPDAGs since the poset has a least and greatest element), and appealing to Proposition 3.
4.3.2 Restriction to polytrees
We have established that the model-oriented distance for CPDAGs is inherently more complex than that for UGs or DAGs, reflecting both the increased structural complexity of the model-oriented poset and the richer graphical nature of CPDAGs themselves. We now study how the model-oriented distance may simplify if we restrict the class of CPDAGs we consider. Specifically, we focus on CPDAGs that are also polytrees, that is, CPDAGs with an acyclic skeleton. This class frequently arises in the causal inference literature (Rebane and Pearl, 1987; Sepehr and Materassi, 2019; Grüttemeier, Komusiewicz and Morawietz, 2021; Tramontano, Monod and Drton, 2022; Améndola et al., 2023)
Let $\mathfrak{G}_{\text{poly}}$ be the collection of completed partially directed graphs over the node set $V$ that are polytrees. Here, the model-oriented distance between two graphs $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}_{\text{poly}}$ is the number of neighbor steps needed to transform $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$, where each step is a move to a CPDAG that is also a polytree. Formally, let $\mathcal{L}_{\text{poly}}=(\mathfrak{G}_{\text{poly}},\preceq)$ be the poset underlying $\mathfrak{G}_{\text{poly}}$ using the same ordering relationship as the unrestricted setting (this is a subposet of $\mathcal{L}$). Then, the model-oriented distance $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ can be characterized as follows.
Theorem 3.
For any pair of graphs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ in $\mathfrak{G}_{\text{poly}}$, we have $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})=d_{\mathcal{L}_{\text{poly}},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$.
We prove this result in Appendix F.2. The proof proceeds by showing that $\mathcal{L}_{\text{poly}}$ is lower semimodular and then appeals to Theorem 2 to obtain the desired result.
4.3.3 Relationship to Separation Distance
There exist another distance for CPDAGs and probabilistic graphical model encoding only conditional independence statements more broadly, called separation distance (Wahl and Runge, 2025). It measures the distance between two graphs by counting the number of conditional independence statements encoded by only one of them. Similarly to our class of distances, the separation distance is a metric and model oriented, albeit on a different model interpretation. However, the separation distance framework implicitly assumes that any collection of conditional independence statements forms a valid conditional independence model. This is similar to how the SHD on CPDAGs implicitly assumes any simple graph with direct and undirected edges is a valid CPDAG (see Example 1). Conditional independence statements, however, may imply other such statements (Geiger, Verma and Pearl, 1989). In fact, determining which sets of triples correspond to a conditional independence model and which of these in turn correspond to a probabilistic graphical model is a challenging problem that has not yet been fully resolved (Sadeghi, 2017). There are two notable consequences of this idiosyncrasy. First, the separation distance between sparse graphs tends to be much larger than between dense graphs. Second, the separation distance between two graphs increases substantially when a disconnected node is added to both. We provide an example illustrating this and discuss the separation distance paradigm more thoroughly in Appendix H.2.
4.4 Causal MPDAGs
Suppose the graph class $\mathfrak{G}$ is the set of MPDAGs over vertex set $V$ with $\mathcal{M}:=\mathcal{M}_{\text{mpdag}}$ as the model map. The following result provides a graphical characterization for model containment.
Proposition 9.
Let $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ be two MPDAGs. Then, $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$ if and only if for every DAG $\mathcal{D}_{1}$ represented by $\mathcal{G}_{1}$, there exists a DAG $\mathcal{D}_{2}$ represented by $\mathcal{G}_{2}$ such that $\mathcal{D}_{1}$ is a directed subgraph of $\mathcal{D}_{2}$.
We prove Proposition 9 in Section˜A.2. Using this result, we can characterize the poset $\mathcal{L}=(P,\preceq)$ underlying MPDAGs. This poset consists of MPDAGs, ordered according to subgraph containment: graphs $\mathcal{G}_{1},\mathcal{G}_{2}\in L$ satisfy $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ if for every DAG $\mathcal{D}_{1}$ in $\mathcal{G}_{1}$, there exists a DAG $\mathcal{D}_{2}$ such that $\mathcal{D}_{1}$ is a subgraph of $\mathcal{D}_{2}$. The least element is the empty graph. The poset has a greatest element that is the fully connected undirected graph. Further, it is not meet or join semi-lattice nor is lower or upper semimodular; see Appendix E.3 for counter examples. Fig.˜3(d) shows the Hasse diagram associated with this poset. With this poset $d_{\mathcal{L}}$, we have a formal definition for model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ between two MPDAGs.
As with CPDAGs, the model‑oriented distance differs from the two variants of SHD. As a simple illustration, suppose $\mathcal{G}_{s}$ is an MPDAG among three nodes $\{v_{1},v_{2},v_{3}\}$ with two edges $v_{1}\to v_{2}$ and $v_{3}\to v_{2}$, and $\mathcal{G}_{t}$ is an MPDAG among the same node set containing two edges $v_{1}-v_{2}$ and $v_{3}-v_{2}$. Both variants of the structural Hamming distance give the value two. However, these measures implicitly assume that the graph with edge set $v_{1}-v_{2}$ and $v_{3}\to v_{2}$ is a valid MPDAG, which is incorrect — highlighting that the structural Hamming distance fails to account for model structure. In fact, the model-oriented distance between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ in this case is four.
Given the importance of gradedness for computing and deriving simplified characterizations of $d_{\mathcal{L}}$, we study whether the poset $\mathcal{L}$ is graded, that is, whether there exists a function $\mathrm{rank}(\cdot)$ assigning non-negative integers to graphs, such that each covering relation corresponds to a rank increase of exactly one. In previous graph classes, such a rank function was defined by the number of edges, so neighboring graphs always differed by exactly one edge. In the context of MPDAGs, however, a neighbor of a graph may contain many more edges. As an example, consider the graph $\mathcal{G}_{1}$ in Fig.˜6(a). Both graphs $\mathcal{G}_{2}$ and $\mathcal{G}_{3}$ in panels (b) and (c), respectively, satisfy $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$ and $\mathcal{G}_{1}\lessdot\mathcal{G}_{3}$. The neighbor $\mathcal{G}_{2}$ contains one additional edge compared to $\mathcal{G}_{1}$, whereas $\mathcal{G}_{3}$ contains n1n-1 additional edges. Note that $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$ by Proposition˜11. To see why $\mathcal{G}_{1}\lessdot\mathcal{G}_{3}$, note that any graph $\mathcal{G}$ satisfying $\mathcal{G}\preceq\mathcal{G}_{3}$, we must either direct undirected edges or drop edges. However, for $\mathcal{G}$ to also satisfy $\mathcal{G}_{1}\preceq\mathcal{G}$, none of the undirected edges in $\mathcal{G}_{3}$ can be modified in any way. Finally, dropping any subset of directed edges in $\mathcal{G}_{3}$ results in an invalid MPDAG (see Fig.˜2(b)).
This example rules out the number of edges as a valid rank function. However, it does not rule out the existence of some other valid rank function. The following proposition formally establishes that no such function exists, and thus the poset $\mathcal{L}$ is not graded.
Proposition 10.
There does not exists a function $\mathrm{rank}(\cdot)$ assigning non-negative integers to graphs in $\mathfrak{G}$, such that each covering relation corresponds to a rank increase of exactly one. Thus, the poset $\mathcal{L}$ is not graded.
We prove Proposition 10 in Appendix G.1. It constructs a pair of MPDAGs $\mathcal{G}_{s},\mathcal{G}_{t}$ with $\mathcal{G}_{s}\preceq\mathcal{G}_{t}$, where there are maximal chains from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ with different length. (Recall that a maximal chain in $\mathcal{L}$ is a sequence of graphs from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ in which every graph covers the proceeding one.) This difference in chain lengths contradicts the gradedness property, since every covering step increases the rank by exactly one, so any two maximal chains between the same endpoints would have to have equal length.
In addition to a lack of graded structure in $\mathcal{L}$, we do not have an explicit graphical characterization of neighbor structure. This makes it difficult to enumerate neighbors and compute $d_{\mathcal{L}}$ or the down-up and up-down distances $d_{\mathcal{L},\downarrow\uparrow}$ and $d_{\mathcal{L},\uparrow\downarrow}$. To address these challenges, we introduce a pseudo-rank function.
Definition 8 (Pseudo-rank).
For any MPDAG $\mathcal{G}\in\mathfrak{G}$, we define its pseudo-rank, denoted $\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G})$, by the the number of directed edges in $\mathcal{G}$ plus twice the number of undirected edges.
The pseudo-rank function serves as a surrogate for rank in a non-graded poset (see Proposition˜11) and reflects an intuitive measure of structural complexity (see Proposition˜12).
Proposition 11 (Sufficient condition for covering).
Consider $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ with $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$. If $\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G}_{1})+1=\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G}_{2})$, then, $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$.
Proof.
Suppose $\mathcal{G}\in\mathfrak{G}$ satisfies $\mathcal{G}_{1}\preceq\mathcal{G}\preceq\mathcal{G}_{2}$ and differs from both. By the pseudo-rank condition, the skeletons of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are either identical or differ by one edge. If their skeletons coincide, then $\mathcal{G}_{2}$ has exactly one undirected edge that is directed in $\mathcal{G}_{1}$, so the intermediate $\mathcal{G}$ would need to both orient an undirected edge of $\mathcal{G}_{1}$ and undirect an edge of $\mathcal{G}_{2}$, which is impossible. If instead $\mathcal{G}$ has the skeleton of $\mathcal{G}_{1}$, which differs from that of $\mathcal{G}_{2}$ by one edge, then $\mathcal{G}_{2}$ contains one directed edge that does not exist in $\mathcal{G}_{1}$, while all other edges are shared; however, $\mathcal{G}$ must contain at least one undirected edge that is directed in $\mathcal{G}_{1}$ and can therefore not satisfy $\mathcal{G}\preceq\mathcal{G}_{2}$. The symmetric case, where $\mathcal{G}$ has the skeleton of $\mathcal{G}_{2}$ which differs from that of $\mathcal{G}_{1}$, follows analogously.
Proposition 12.
(MPDAG model complexity) For any MPDAG $\mathcal{G}\in\mathfrak{G}$ there exists a sequence $\mathcal{G}=x_{1}\gtrdot\dots\gtrdot x_{k}=\mathcal{G}_{\hat{0}}$ of MPDAGs, where $G_{\hat{0}}$ is the empty MPDAG and $k=\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G})$.
We prove Proposition 12 in Appendix G.2. Using this pseudo-rank function, we define $P_{\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ as the set of paths $x_{0}:=\mathcal{G}_{s},x_{1},\dots,x_{r}:=\mathcal{G}_{t}$ in the Hasse diagram corresponding to $\mathcal{L}$, such that each step satisfies $|\mathrm{pseudo}\text{-}\mathrm{rank}(x_{i})-\mathrm{pseudo}\text{-}\mathrm{rank}(x_{i-1})|=1$ for all $i$. In other words, $P_{\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})\subseteq P(\mathcal{G}_{s},\mathcal{G}_{t})$, where relative to $P(\mathcal{G}_{s},\mathcal{G}_{t})$, $P_{\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ has the added constraint that the pseudo-rank changes by exactly one at each step in the path. This ensures that each transition $(x_{i-1},x_{i})$ between neighbor graphs corresponds to a minimal modification: either an undirected edge in one graph becomes directed in the other, or a directed edge in one graph is absent in the other. This constraint mimics graded posets, where neighbors differ by exactly one in rank, helping recover some useful structure in the non-graded setting of $\mathcal{L}$.
By Proposition 12, $P_{\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is a non-empty set. Since $P_{\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})\subseteq P(\mathcal{G}_{s},\mathcal{G}_{t})$, we can therefore define a distance based on $P_{\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ that upper bounds $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ formalized as follows.
Corollary 4.
Let $d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t}):=\min\{\allowbreak\mathrm{length}(p):p\in L_{\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})\}$, i.e., the length of the shortest path between $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$, among paths that lie inside $P_{\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$. Then, we have $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
Corollary 4 has practical implications: the pseudo-distance $d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is often significantly easier to compute than either the model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ or the down-up and up-down distances $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$ and $d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t})$. In particular, recall that computing the model-oriented distance or the down-up and up-down distances requires enumerating neighbor graphs. However, as demonstrated in the example in Fig.˜6, the structure in neighbor graphs can vary widely, and no simple graphical characterization of neighborhood is known. In contrast, computing $d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ only requires considering neighbors where the pseudo-rank changes by exactly one. This makes the enumeration process much simpler: for a graph $\mathcal{G}$, the relevant neighbors can be obtained by adding or removing any undirected edge, directing an undirected edge, or undirecting a directed edge — provided the resulting graph is a valid MPDAG in that it does not contain any induced subgraphs in Fig.˜2(b).
It may be tempting to claim $d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is equal to $\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})$, the second variant of SHD between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$. However, in the simple example among three nodes from earlier, $d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is four, while $\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})=2$ is two. In fact, $\operatorname{\mathrm{SHD}}_{2}$ and consequently $\operatorname{\mathrm{SHD}}_{1}$ both lower bound $d_{\mathcal{L},\text{pseudo}}$ as formalized next.
Proposition 13.
For any pair of MPDAGs $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, both variants of the SHD between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ is smaller than or equal to $d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{1},\mathcal{G}_{2})$; that is, $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{1},\mathcal{G}_{2})\leq\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{1},\mathcal{G}_{2})\leq d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{1},\mathcal{G}_{2})$.
We prove Proposition 13 in Appendix G.3. The bound in this proposition is used in the algorithm for computing $d_{\mathcal{L},\text{pseudo}}$; see Section 5.3. Note that while the SHD variants lower-bound $d_{\mathcal{L},\text{pseudo}}$, they do not always lower bound the model-oriented distance $d_{\mathcal{L}}$. For example, let $\mathcal{G}_{1}$ be the graph in Fig.˜6(a) and $\mathcal{G}_{3}$ be the graph in Fig.˜6(c). Then $d_{\mathcal{L}}(\mathcal{G}_{1},\mathcal{G}_{3})=1$, while both SHD variants return n1n-1.
Figure 6: An MPDAG $\mathcal{G}_{1}$ in (a) and two of its neighboring MPDAGs in (b) and (c). The neighbor in (c) has n1n-1 more edges than $\mathcal{G}_{1}$.
4.4.1 Restriction to polytrees for MPDAGs
Similar to the case of CPDAGs, we can restrict our attention to a subclass of MPDAGs that are polytrees. Let $\mathfrak{G}_{\text{poly}}$ be the collection of MPDAGs over node set $V$ that are polytrees. Define the subposet $\mathcal{L}_{\text{poly}}=(\mathfrak{G}_{\text{poly}},\preceq)$ as the restriction of $\mathcal{L}$ to this subclass and let $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ denote the corresponding model-oriented distance. Let $d_{\mathcal{L}_{\text{poly}},\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ be the analog of $d_{\mathcal{L},\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$, but defined on the subposet $\mathcal{L}_{\text{poly}}$. Using a similar argument as Theorem 4, we can conclude that $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq d_{\mathcal{L}_{\text{poly}},\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$. In fact, this inequality is tight: equality holds because the poset $\mathcal{L}_{\text{poly}}$ is graded with pseudo-rank as its ranking function. We formalize this next.
Theorem 4.
The poset $\mathcal{L}_{\text{poly}}$ is graded with rank function $\mathrm{pseudo}\text{-}\mathrm{rank}(\cdot)$. Thus, the model-oriented distance $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is equal to the pseudo-distance $d_{\mathcal{L}_{\text{poly}},\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
We prove Theorem 4 in Appendix G.4. This result confirms that in the polytree setting, the pseudo-rank function defines a proper grading on $\mathcal{L}_{\text{poly}}$, and the corresponding distance exactly matches the model-oriented distance. Theorem 4 provides concrete justification for using the pseudo-rank-based distance more broadly.
Recall that when restricting CPDAGs to polytrees, the corresponding poset is lower semimodular, allowing us to appeal to Theorem 2 to conclude that the model-oriented distance equals the down-up distance. However, we show in Appendix E.3 that the poset obtained by restricting MPDAGs to polytrees is not lower semimodular. Thus, a similar simplification is not possible and we need to rely on the $A^{*}$-algorithm to compute $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ instead.
5 Algorithms for computing model-oriented distances
In this section, we develop algorithms for either computing or approximating the model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ for a pair of graphs $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$ from different graphical model classes $\mathfrak{G}$, including CPDAGs and MPDAGs. In Section 5.1, we present a general-purpose A* procedure for finding $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ across different graph classes. In Sections 5.2 and 5.3, we specialize this algorithm to the class of CPDAGs and MPDAGs, respectively, and present numerical experiments comparing the model-oriented distance to the two SHD variants $\operatorname{\mathrm{SHD}}_{1}$ and $\operatorname{\mathrm{SHD}}_{2}$ while tracking the computational cost of our algorithms.
5.1 General A* procedure
To compute the model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ between two graphs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$, we use an A* search algorithm, detailed in Algorithm 1. The search is performed on the poset $\mathcal{L}$. The algorithm maintains a priority queue (openSet) initialized with the start graph $\mathcal{G}_{s}$. This queue is ordered by a costs function $f(\mathcal{G})=g(\mathcal{G})+h(\mathcal{G})$, where: (i) $g(\mathcal{G})$ tracks the exact costs (i.e., number of steps) from $\mathcal{G}_{s}$ to $\mathcal{G}$, and (ii) $h(\mathcal{G})$ is a heuristic function that estimates the remaining cost from $\mathcal{G}$ to $\mathcal{G}_{t}$. To ensure that A* returns the distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$, the heuristic $h$ must be admissible:
Definition 9 (Admissible heuristic).
A heuristic $h(\cdot)$ is admissible if $h(\mathcal{G})\leq d_{\mathcal{L}}(\mathcal{G},\mathcal{G}_{t})$ for all $\mathcal{G}\in\mathfrak{G}$, i.e., it never overestimates the true distance.
We provide some more detail about Algorithm 1. At each iteration, the graph $\mathcal{G}_{\text{current}}$ with the smallest $f$-score is removed from the queue and expanded. If $\mathcal{G}_{\text{current}}=\mathcal{G}_{t}$, the algorithm terminates and returns $g(\mathcal{G}_{t})$. Otherwise, the algorithm examines all neighbors of $\mathcal{G}_{\text{current}}$ in $\mathcal{L}$. Here, the function $\mathrm{n}_{\mathcal{L}}(\cdot)$ finds neighboring graphs in the poset $\mathcal{L}$. For each neighbor $\mathcal{G}_{\text{neighbor}}$, it computes a tentative distance via $\mathcal{G}_{\text{current}}$. If this tentative distance is shorter than the path previously recorded to $\mathcal{G}_{\text{neighbor}}$, the algorithm updates $g(\mathcal{G}_{\text{neighbor}})$ and adds it to the queue with priority $g(\mathcal{G}_{\text{neighbor}})+h(\mathcal{G}_{\text{neighbor}})$.
This A* framework is broadly applicable to any graph class $\mathfrak{G}$ that can be equipped with a connected poset $\mathcal{L}$. To use the algorithm in a given setting, one must provide: (i) an admissible heuristic that underestimates $d_{\mathcal{L}}(\mathcal{G}_{1},\mathcal{G}_{2})$ for all pairs $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, and (ii) a method for enumerating all neighbors of a given graph $\mathcal{G}\in\mathfrak{G}$, i.e., the graphs that cover or are covered by $\mathcal{G}$ in the poset.
input: Graph class $\mathfrak{G}$, pair of graphs $\mathcal{G}_{s},\mathcal{G}_{t}$, heuristic $h\colon\mathfrak{G}\to\mathbb{R}_{\geq 0}$, neighbor function $\mathrm{n}_{\mathcal{L}}\colon\mathfrak{G}\to 2^{\mathfrak{G}}$
output : Model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$
1 begin
 $openSet\coloneqq\{\mathcal{G}_{s}\}$;
 // priority queue ordered by $g+h$
2 $g(\mathcal{G}_{s})\coloneqq 0$;
3
4 while $openSet\neq\emptyset$ do
5    $\mathcal{G}_{\text{current}}\coloneqq$ graph in openSetopenSet minimizing $g(\mathcal{G}_{\text{current}})+h(\mathcal{G}_{\text{current}})$;
6    if $\mathcal{G}_{\text{current}}=\mathcal{G}_{t}$ then
7       return $g(\mathcal{G}_{t})$
8     remove $\mathcal{G}_{\text{current}}$ from openSetopenSet;
    // Neighbors according to poset $\mathcal{L}$
9    foreach $\mathcal{G}_{\text{neighbor}}\in\mathrm{n}_{\mathcal{L}}(\mathcal{G}_{\text{current}}$) do
10       $tentative\coloneqq g(\mathcal{G}_{\text{current}})+1$;
11       if $g(\mathcal{G}_{\text{neighbor}})$ not defined   or   $tentative<g(\mathcal{G}_{\text{neighbor}})$ then
12          $g(\mathcal{G}_{\text{neighbor}})\coloneqq tentative$;
13            add $\mathcal{G}_{\text{neighbor}}$ to $openSet$ with priority $g(\mathcal{G}_{\text{neighbor}})+h(\mathcal{G}_{\text{neighbor}})$;
Algorithm 1 An A* algorithm for computing the model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ .
5.2 Specialization to CPDAGs
Next, we consider the concrete problem of computing $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ for CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$. Here, $\mathcal{L}$ is the poset underlying all CPDAGs over a fixed vertex set. Using Algorithm 1 requires specifying a heuristic $h$ and the function $\mathrm{n}_{\mathcal{L}}$. As heuristic $h$, we use $|\mathrm{sk}(\mathcal{G})\setminus\mathrm{sk}(\mathcal{G}_{t})|+|\mathrm{sk}(\mathcal{G}_{t})\setminus\mathrm{sk}(\mathcal{G})|$ as a lower bound for the model-oriented distance between any $\mathcal{G}$ and $\mathcal{G}_{t}$; see Property 2 where we lower bound $m_{\text{op}}$ by zero as computing $m_{\text{op}}$ for every call of the heuristic can be costly. For the function $\mathrm{n}_{\mathcal{L}}$, we use characterization via the “Insert" and “Delete" operators introduced in Chickering (2002) (see Lemma˜2).
Computing $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ via Algorithm 1 can be expensive. We can compute the down-up and up-down distances (Property 1), to use them as upper bounds on $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$. We can also compute the lower bound from Property 2. These bounds are more efficient to compute, and when they coincide, they yield the exact value of $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$. We next describe how to compute the upper bounds based on the down-up and up-down distances, and present numerical experiments comparing the model-oriented distance to the two variants of SHD. The algorithmic details for computing the lower bound are provided in Appendix I.
5.2.1 Computing an upper bound based on down-up or up-down distances
Consider first the down-up distance $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$. One can adapt Algorithm 1 to search for the shortest down-up path between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$. This adaptation requires only a few conceptual changes: (i) The neighbor selection step is restricted to graphs that are covered by the current graph, ensuring that each move is downward. (ii) The search terminates early when a graph $\mathcal{G}^{\prime}\preceq\mathcal{G}_{t}$ is reached. At this point, the upward portion of the path is not explicitly traversed; instead, its length is computed directly as $\mathrm{rank}(\mathcal{G}_{t})-\mathrm{rank}(\mathcal{G}^{\prime})$. This uses the fact that in a graded poset, the shortest upward path from $\mathcal{G}_{1}$ to $\mathcal{G}_{2}$ with $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ has length exactly $\mathrm{rank}(\mathcal{G}_{2})-\mathrm{rank}(\mathcal{G}_{1})$; see Proposition 1. (iii) The down-up distance between any graph $\mathcal{G}^{\prime}$ and $\mathcal{G}_{t}$ in this poset is given by $\mathrm{rank}(\mathcal{G}_{t})+\mathrm{rank}(\mathcal{G}^{\prime})-2\max_{\mathcal{G}\preceq\mathcal{G}^{\prime},\mathcal{G}\preceq\mathcal{G}_{t}}\mathrm{rank}(\mathcal{G})$, which motivates the following heuristic $\mathrm{rank}(\mathcal{G}_{t})+\mathrm{rank}(\mathcal{G}^{\prime})-2|\mathrm{sk}(\mathcal{G}_{t})\cap\mathrm{sk}(\mathcal{G}^{\prime})|$. In Appendix F.3, we prove that for any two CPDAGs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, $\max_{\mathcal{G}\preceq\mathcal{G}_{1},\mathcal{G}\preceq\mathcal{G}_{2}}\mathrm{rank}(\mathcal{G})\leq|\mathrm{sk}(\mathcal{G}_{1})\cap\mathrm{sk}(\mathcal{G}_{2})|$, so the heuristic is a valid lower bound on the down-up distance and is therefore admissible.
Note that we check whether $\mathcal{G}^{\prime}\preceq\mathcal{G}_{t}$ using the characterization in Kocka, Bouckaert and Studený (2001), which provides necessary and sufficient conditions for conditional dependence inclusion between two CPDAGs. Since this check can be computationally costly, we consider a necessary criterion for $\mathcal{G}^{\prime}$ to be the inflection point in the shortest down-up path between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$. Appendix F.1 shows that the optimal inflection point is any maximizer of $\operatorname*{arg\,max}_{\mathcal{G}\preceq\mathcal{G}_{s},,\mathcal{G}\preceq\mathcal{G}_{t}}\mathrm{rank}(\mathcal{G})$, whose rank we denote by $\rho_{\text{meet}}(\mathcal{G}_{s},\mathcal{G}_{t})$. As previously discussed, this value is guaranteed to be less than or equal to $u=|\mathrm{sk}(\mathcal{G}_{s})\cap\mathrm{sk}(\mathcal{G}_{t})|$. Thus, before performing the check $\mathcal{G}^{\prime}\preceq\mathcal{G}_{t}$, we first verify whether $\mathrm{rank}(\mathcal{G}^{\prime})\leq\rho_{\text{meet}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq u$. The overall adapted A* algorithm for finding the shortest down-up distance is presented in Algorithm 2.
We present the algorithm for computing the up-down distance in Appendix I, which is the ‘dual’ of Algorithm 2. In particular, (i) the neighbor selection step is restricted to be upward; (ii) the search terminates when a graph $\mathcal{G}^{\prime}$ is found with $\mathcal{G}_{t}\preceq\mathcal{G}^{\prime}$; (iii) we use the admissible heuristic $2|\mathrm{sk}(\mathcal{G}_{t})\cup\mathrm{sk}(\mathcal{G}^{\prime})|-\mathrm{rank}(\mathcal{G}_{t})-\mathrm{rank}(\mathcal{G}^{\prime})$; and finally (iv) we verify whether $\mathrm{rank}(\mathcal{G}^{\prime})\geq|\mathrm{sk}(\mathcal{G}_{t})\cup\mathrm{sk}(\mathcal{G}^{\prime})|$ before checking if $\mathcal{G}_{t}\preceq\mathcal{G}^{\prime}$.
input: Graph class of CPDAGs $\mathfrak{G}$, pair of graphs $\mathcal{G}_{s},\mathcal{G}_{t}$, a heuristic $h\colon\mathfrak{G}\to\mathbb{R}_{\geq 0}$, downwards neighbor function $dn_{\mathcal{L}}\colon\mathfrak{G}\to 2^{\mathfrak{G}}$, upper bound $u\in\mathbb{R}_{\geq 0}$ on $\rho_{\text{meet}}(\mathcal{G}_{s},\mathcal{G}_{t})$
output : Down-up distance $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$
1 begin
 $openSet\coloneqq\{\mathcal{G}_{s}\}$;
 // priority queue ordered by $g+h$
2 $g(\mathcal{G}_{s})\coloneqq 0$;
3 while $openSet\neq\emptyset$ do
4    $\mathcal{G}_{\text{current}}\coloneqq$ graph in openSetopenSet minimizing $g(\mathcal{G}_{\text{current}})+h(\mathcal{G}_{\text{current}})$;
5    if $\mathrm{rank}(\mathcal{G}_{\text{current}})\leq u\,\,\text{ and }\,\,\mathcal{G}_{\text{current}}\preceq\mathcal{G}_{t}$ then
6       return $g(\mathcal{G}_{\text{current}})+\mathrm{rank}(\mathcal{G}_{t})-\mathrm{rank}(\mathcal{G}_{\text{current}})$
7     remove $\mathcal{G}_{\text{current}}$ from openSetopenSet;
8
9    foreach $\mathcal{G}_{\text{neighbor}}\in dn_{\mathcal{L}}(\mathcal{G}_{\text{current}})$ do
10       $tentative\coloneqq g(\mathcal{G}_{\text{current}})+1$;
11       if $g(\mathcal{G}_{\text{neighbor}})$ not defined   or   $tentative<g(\mathcal{G}_{\text{neighbor}})$ then
12          $g(\mathcal{G}_{\text{neighbor}})\coloneqq tentative$;
13            add $\mathcal{G}_{\text{neighbor}}$ to $openSet$ with priority $g(\mathcal{G}_{\text{neighbor}})+h(\mathcal{G}_{\text{neighbor}})$;
Algorithm 2 An A* algorithm for computing the down-up distance $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$.
5.2.2 Numerical experiments for CPDAGs
To study the characteristics of the model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ for CPDAGs as well as the computational cost of our algorithms, we perform a small simulation study. We consider two settings. First, all pairs of CPDAGs $(\mathcal{G}_{s},\mathcal{G}_{t})$ over four nodes (185 in total; Gillispie and Perlman, 2001) and second, all pairs over five nodes ($8782$ graphs). In both settings, We first compute lower and upper bounds on $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ using the procedures described earlier. When the bounds do not match, we apply Algorithm 1 to obtain the exact distance. We also compute both SHD variants (Definition 7) for comparison.
In the setting with four nodes we find the following results. For approximately $96\%$ of pairs, the bounds match exactly; for the remaining $\approx 4\%$, they differ by at most two. The average computation time is $\approx 0.00003$ seconds when bounds match, compared to $\approx 0.0001$ seconds otherwise, highlighting the efficiency of using bounds before invoking Algorithm 1. Fig.˜7(a) shows the histogram of the difference between SHD and the model-oriented distance. In $\approx 64\%$ of cases, the distances differ, and in $\approx 63\%$, the model-oriented distance is strictly smaller. Fig.˜8 presents three illustrative examples. In the first example in Figure 8(a), $\mathcal{G}_{s}\preceq\mathcal{G}_{t}$, and thus appealing to Proposition 1, $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=\mathrm{rank}(\mathcal{G}_{t})-\mathrm{rank}(\mathcal{G}_{s})=2$. The SHD variants give significantly different values for the distance: $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})=6$ and $\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})=8$. In the second example in Fig.˜8(b), we have $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=2$, where the shortest path between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ is an up-down path: moving a step upward from $\mathcal{G}_{s}$ to the fully connected CPDAG $\mathcal{G}$, and then moving a step downward from $\mathcal{G}$ to $\mathcal{G}_{t}$. Again, in this case, SHD variants are significantly different: $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})=6$ and $\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})=12$. In the final example in Fig.˜8(c), $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})=\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})=3$. In this case, unlike the other examples, the model-oriented distance is larger and equals $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=4$. Specifically, a shortest down-up path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ first removes the edges $v_{3}\rightarrow v_{4}$ and $v_{1}\rightarrow v_{3}$ in two downward steps to obtain a CPDAG $\mathcal{G}$ containing the edge $v_{2}-v_{4}$, and then adds the edges $v_{1}-v_{4}$ and $v_{3}-v_{4}$ in two upward steps.
In the setting with five nodes we find the following results. The upper and lower bounds on the model-oriented distance coincide in $89\%$ of cases, with an average computation time of $\approx 0.0001$ seconds. For the remaining cases, the average time is $\approx 0.002$ seconds. Figure 7(b) shows the histogram of differences between both SHD variants and the model-oriented distance. In $\approx 80\%$ of cases, the model-oriented distance differs from both SHD variants, and in $\approx 78\%$, it is strictly smaller. In about $3\%$ of cases, the shortest path follows a ‘zig-zag’ pattern rather than a down–up or up–down path. Fig.˜8(d) presents an example where the model-oriented distance and SHD differ significantly. Here, $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})=10$ and $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})=12$, whereas the model-oriented distance yields $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=2$. Since SHD is model-agnostic, it does not account for the fact that all conditional dependencies encoded by $\mathcal{G}_{s}$ are also encoded by $\mathcal{G}_{t}$, and that adding two edges to $\mathcal{G}_{s}$ – one between $v_{1}$ and $v_{2}$, and another between $v_{3}$ and $v_{4}$ – yields the CPDAG $\mathcal{G}_{t}$.
Figure 7: The difference between SHD variants and our model-oriented distance, computed over all CPDAG pairs with $p$ nodes. (a) $p=4$; (b) $p=5$.
Figure 8: Examples highlighting differences between model-oriented distances and SHD; see text in Sections 5.2.2 and 5.3.1.
5.3 Specialization to MPDAGs
Next, we consider computing $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ for MPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$, where $\mathcal{L}_{\text{poly}}$ is the poset underlying all MPDAGs that are polytrees over a fixed vertex set. In the setting of polytrees, recall from Theorem 4 that $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ simplifies to $d_{\mathcal{L}_{\text{poly}},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$. This simplification is important for multiple reasons: (i) appealing to Proposition 13, $\operatorname{\mathrm{SHD}}_{2}$ provides a lower bound for $d_{\mathcal{L}_{\text{poly}}}$ and thus serves as an admissible heuristic for Algorithm 1; and (ii) a graph’s neighbors in the poset $\mathcal{L}_{\text{poly}}$ are those that differ in pseudo-rank by one, and these neighbors can be computed efficiently (see Section 4.4). We supply the admissible heuristic and the neighbor function to Algorithm 1 to compute the model-oriented distance.
5.3.1 Numerical experiments for MPDAGs
We compute model-oriented distance $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ for each pair of MPDAGs $(\mathcal{G}_{s},\mathcal{G}_{t})$ over five nodes (6679 total). The average time to compute this distance is approximately $0.004$ seconds. In roughly $98\%$ of the instances, the model-oriented distance matches the distance obtained by $\operatorname{\mathrm{SHD}}_{2}$. Among the remaining approximately $2\%$: in about $98\%$ of these cases, the model-oriented distance is two units larger than $\operatorname{\mathrm{SHD}}_{2}$; in roughly $1.7\%$ of cases, it is four units larger; and in five instances, it is six units larger. Fig.˜8(d) shows an example of an instance where the difference between the two distances is six. Here, $\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})=4$, whereas the model-oriented distance yields $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})=10$. The discrepancy arises because SHD incorrectly assumes that directing any of the undirected edges toward $v_{5}$ results in a valid MPDAG. However, since $\mathcal{L}_{\text{poly}}$ only contains polytrees, the shortest path in the poset $\mathcal{L}_{\text{poly}}$ from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ involves first directing each undirected edge away from $v_{5}$ (toward the other node), then removing these directed edges, and finally adding directed edges from each node to $v_{5}$.
6 Discussion
This paper introduces a model-oriented framework for defining distances between graphs within a given class, where each graph is treated as a statistical model and the collection of graphs is organized as a poset based on model inclusion. The resulting metric, derived from the shortest path through neighboring models, provides a principled way to measure distance across various graph classes including causal DAGs, probabilistic CPDAGs, and causal MPDAGs.
Our work opens several avenues for future research. First, for efficient computation of the model-oriented distance for MPDAGs, it is important to develop an explicit graphical characterization of the neighborhood structure. Second, one can consider a weighted version of the shortest path, where a weight is assigned to each covering pair in the poset. One possible approach for selecting these weights is to base them on information-theoretic notions, as was done in the context of distances between two partitions (Meila, 2005). Moreover, computational efficiency in executing the A* algorithm can be substantially improved through better heuristics, parallelization, and memory management. The authors are currently pursuing a separate project to reduce the time it takes to compute the model-oriented distance. Finally, the proposed framework is applicable to other graph classes including local independence graphs for capturing dependence structure of events that occur in time (Didelez, 2008; Mogensen and Hansen, 2020), maximal ancestral graphs (Richardson and Spirtes, 2002) and partial ancestral graph (Spirtes, Glymour and Scheines, 2000) that account for model dependencies in the presence of latent variables.
{acks}
[Acknowledgments] The authors thank undergraduate students Arnav Mazumder, Jian Kang, and Carter Lembo at the University of Washington for helping with the implementations of our algorithms.
Appendix A Proofs: model maps and associated graphs
A.1 Injective model maps
We show that the models $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}},\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}},\mathcal{M}^{\mathrm{c}}_{\textsc{dag}},\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}$, as defined in Section˜2.2, are injective in the sense that no two different graphs represent the same model.
Assumption 1.
$\mathrm{V}$ is a finite vertex set. For every $v\in\mathrm{V}$, $\mathcal{X}_{v}$ is either infinite or finite with two or more elements.
Theorem 5.
Under ˜1, we have the following results.
  1. (i)
    $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}:\mathfrak{G}=\{\text{all UGs over vertex set $\mathrm{V}$}\}\rightarrow 2^{\mathcal{M}^{\mathrm{o}}_{+}}$ is injective.
  2. (ii)
    $\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}}:\mathfrak{G}=\{\text{all CPDAGs over vertex set $\mathrm{V}$}\}\rightarrow 2^{\mathcal{M}^{\mathrm{o}}}$ is injective.
  3. (iii)
    $\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}:\mathfrak{G}=\{\text{all DAGs over vertex set $\mathrm{V}$}\}\rightarrow 2^{\mathcal{M}^{\mathrm{c}}}$ is injective.
  4. (iv)
    $\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}:\mathfrak{G}=\{\text{all MPDAGs over vertex set $\mathrm{V}$}\}\rightarrow 2^{\mathcal{M}^{\mathrm{c}}}$ is injective.
Proof.
The proof for $|\mathrm{V}|=1$ is trivial. In what follows, we assume $|\mathrm{V}|\geq 2$.
  1. (i)
    For every $\mathcal{G}=(\mathrm{V},\mathrm{E})\in\mathfrak{G}$, $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}(\mathcal{G})$ can be characterized by the pairwise Markov property (Lauritzen, 1996, §3.2.1)
[…]
    Further, under ˜1, there exists $P\in\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}(\mathcal{G})$ that is faithful to $\mathcal{G}$, i.e.,
[…]
    This shows that the map must be injective.
  2. (ii)
    This directly follows from the characterization of Markov equivalent DAGs; see Andersson, Madigan and Perlman (1997).
  3. (iii)
    Fix $x\in\mathcal{X}$. For any $\mathcal{D}\in\mathfrak{G}$, by definition and $|\mathrm{V}|<\infty$, we have
    $P\left(X_{v}(x)=X_{v}(x_{\operatorname{Pa}_{\mathcal{D}}(v)}),\;\forall v\in\mathrm{V}\right)=1,\quad\forall P\in\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}(\mathcal{D}).$
    Under ˜1, there also exists $P^{\prime}\in\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}(\mathcal{D})$ such that
    $\mathrm{W}\neq\operatorname{Pa}_{\mathcal{D}}(v)\implies P^{\prime}\left(X_{v}(x)=X_{v}(x_{\mathrm{W}})\right)<1,\quad\forall v\in\mathrm{V},\;\mathrm{W}\subseteq\mathrm{V}.$
    Because $\mathcal{D}$’s edges can be determined from $\{\operatorname{Pa}_{\mathcal{D}}(v):v\in\mathrm{V}\}$, this shows that $\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}$ is injective.
  4. (iv)
    Consider $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, $\mathcal{G}_{1}\neq\mathcal{G}_{2}$. By the completeness of Meek’s rules (Meek, 1995), the sets of DAGs represented by $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ must be different. Without loss of generality, suppose there is a DAG $\mathcal{D}_{1}$ that is represented by $\mathcal{G}_{1}$ but not by $\mathcal{G}_{2}$. Fix $x\in\mathcal{X}$. By the proof of (iii), there exists $P^{\prime}\in\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}(\mathcal{D}_{1})\subseteq\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}(\mathcal{G}_{1})$ such that
    $\mathrm{W}\neq\operatorname{Pa}_{\mathcal{D}_{1}}(v)\implies P^{\prime}\left(X_{v}(x)=X_{v}(x_{\mathrm{W}})\right)<1,\quad\forall v\in\mathrm{V},\;\mathrm{W}\subseteq\mathrm{V}.$
    It then follows that $P^{\prime}\notin\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}(\mathcal{D}_{2})$ for every DAG $\mathcal{D}_{2}$ represented by $\mathcal{G}_{2}$, and hence $P^{\prime}\notin\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}(\mathcal{G}_{2})$. We have shown that $\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}(\mathcal{G}_{1})\neq\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}(\mathcal{G}_{2})$.
A.2 Graphical characterization of model containment
Proof of Proposition˜4.
The $\Leftarrow$’ direction directly follows from Eq.˜1. For the $\Rightarrow$’ direction, we prove by contradiction. Suppose $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}(\mathcal{G}_{1})\subseteq\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}(\mathcal{G}_{2})$ but $\mathcal{G}_{1}$ contains an edge $(u,v)$ that is not in $\mathcal{G}_{2}$. Then, there exists a distribution $P\in\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}(\mathcal{G}_{1})$ under which $X_{u}$ and $X_{v}$ are not conditionally independent given $X_{\mathrm{V}\setminus\{u,v\}}$. However, $P\in\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}(\mathcal{G}_{2})$ contradicts the conditional independence posed by $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}(\mathcal{G}_{2})$. ∎
Proof of Proposition˜6.
By definition of the causal model associated with a DAG (see Section˜2.2.2), if $\mathcal{G}_{1}$ is a subgraph of $\mathcal{G}_{2}$, then $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$. We now prove the other direction by contradiction. Suppose $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$ but $\mathcal{G}_{1}$ contains an edge $i\rightarrow j$ that is absent from $\mathcal{G}_{2}$. Recall that $\mathcal{X}=\prod_{v\in\mathrm{V}}\mathcal{X}_{v}$ is the sample space. Let us choose $\mathcal{X}_{i}=\mathbb{R}$ for every $i\in V$. We can choose a law $P\in\mathcal{M}(\mathcal{G}_{1})$ of $X(\cdot)$ such that $\operatorname{\mathbb{E}}[X_{j}(x)]=x_{i}$ holds for every $x\in\mathcal{X}$. However, under $\mathcal{M}(\mathcal{G}_{2})$, $\operatorname{\mathbb{E}}[X_{j}(x)]$ (when it exists) can only depend on $x_{\operatorname{Pa}_{\mathcal{G}_{2}}(j)}$ with $i\notin\operatorname{Pa}_{\mathcal{G}_{2}}(j)$, which is a contradiction. ∎
Proof of Proposition˜9.
Recall from Section˜2.2 that $\mathcal{M}(\mathcal{G}_{i}):=\cup_{\mathcal{D}\in\mathcal{G}_{i}}\mathcal{M}(\mathcal{D})$ for $i=1,2$, where $\mathcal{D}\in\mathcal{G}_{i}$ is any DAG represented by the MPDAG $\mathcal{G}_{i}$. It follows from definition of $\mathcal{M}(\mathcal{D})$ and Proposition˜6 that, if for every $\mathcal{D}_{1}\in\mathcal{G}_{1}$, there exists a certain $\mathcal{D}_{2}\in\mathcal{G}_{2}$ such that $\mathcal{D}_{1}$ is a subgraph of $\mathcal{D}_{2}$, then $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$. We now prove the other direction. Let us choose $\mathcal{X}_{i}=\mathbb{R}$ for every $i\in V$. Fix any $\mathcal{D}_{1}\in\mathcal{G}_{1}$. We want to show that $\mathcal{D}_{1}$ is a subgraph of some $\mathcal{D}\in\mathcal{G}_{2}$. We can choose a law $P\in\mathcal{M}(\mathcal{D}_{1})$ of $X(\cdot)$ such that for every $i\in V$, $x\in\mathcal{X}$, it holds that
$\operatorname{\mathbb{E}}[X_{i}(x)]=\begin{cases}\beta_{i}^{{\mathrm{\scriptscriptstyle T}}}x_{\operatorname{Pa}_{\mathcal{D}_{1}}(i)},&\quad\operatorname{Pa}_{\mathcal{D}_{1}}(i)\neq\emptyset\\ 0,&\quad\operatorname{Pa}_{\mathcal{D}_{1}}(i)=\emptyset\end{cases},$ (5)
where $\beta_{i}\in\mathbb{R}^{|\operatorname{Pa}_{\mathcal{D}_{1}}(i)|}$ is a fixed vector with non-zero coordinates. Since $P\in\mathcal{M}(\mathcal{D}_{1})\subseteq\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})=\cup_{\mathcal{D}\in\mathcal{G}_{2}}\mathcal{M}(\mathcal{D})$, there must exists $\mathcal{D}_{2}\in\mathcal{G}_{2}$ such that $P\in\mathcal{M}(\mathcal{D}_{2})$. By definition, under $\mathcal{M}(\mathcal{D}_{2})$, $\operatorname{\mathbb{E}}[X_{i}(x)]$ (when it exists) can only depend on $x_{\operatorname{Pa}_{\mathcal{D}_{2}}(i)}$. Hence, Eq.˜5 implies that $\operatorname{Pa}_{\mathcal{D}_{1}}(i)\subseteq\operatorname{Pa}_{\mathcal{D}_{2}}(i)$ for every $i\in V$. That is, $\mathcal{D}_{1}$ is a subgraph of $\mathcal{D}_{2}$. ∎
Appendix B Structural constraints of graph types
B.1 CPDAG
For a PDAG $\mathcal{G}$ to be a valid CPDAG, it must also satisfy the following conditions (Andersson, Madigan and Perlman, 1997, Theorem 4.1).
  1. (i)
    It contains no partially directed cycle.
  2. (ii)
    Each connected component induced by undirected edges is chordal;
  3. (iii)
    $v_{1}\rightarrow v_{2}-v_{3}$ does not occur as an induced subgraph of $\mathcal{G}$;
  4. (iv)
    Every directed edge $v_{1}\rightarrow v_{2}$ in $\mathcal{G}$ occurs in at least one of following four structures as an induced subgraph of $\mathcal{G}$:
B.2 MPDAG
A PDAG $\mathcal{G}$ is a valid MPDAG if the following two conditions (Wienöbst, Bannach and Liskiewicz, 2021) hold:
  1. (i)
    For each connected component $\mathrm{C}$ induced by undirected edges, the skeleton of $\mathcal{G}_{\mathrm{C}}$ is chordal;
  2. (ii)
    The four structures in Fig.˜2(b) do not occur as induced subgraphs of $\mathcal{G}$.
To construct an MPDAG from a CPDAG, one first orients certain undirected edges according to background knowledge, and then use the rules due to Meek (1995) to remove the forbidden structures.
Appendix C Proofs: poset structure enabling simplifications to $d_{\mathcal{L}}$
C.1 Simplification due to gradedness: proof of Proposition 1
Consider any path $p\in L(\mathcal{G}_{s},\mathcal{G}_{t})$ from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ where $\mathcal{G}_{s}\preceq\mathcal{G}_{t}$. Let $(x_{1}:=\mathcal{G}_{s},x_{2},\dots,x_{k}:=\mathcal{G}_{t})$ be the sequence of graphs specifying this path. By definition, for each $i=2,\dots,k$, either $x_{i}$ covers $x_{i-1}$ or $x_{i-1}$ covers $x_{i}$. Since the poset is graded, we know that for each step, the rank increases or decreases by one, i.e. $|\mathrm{rank}(x_{i})-\mathrm{rank}(x_{i-1})|=1$. Note, by telescoping sum, we have:
$\mathrm{rank}(\mathcal{G}_{t})-\mathrm{rank}(\mathcal{G}_{s})=\sum_{i=2}^{k}\mathrm{rank}(x_{i})-\mathrm{rank}(x_{i-1})$
A term in the decomposition above is equal to $1$ when the change from $x_{i-1}$ to $x_{i}$ is an upward move in the poset ($x_{i-1}\preceq x_{i}$), and is 1-1 when the change from $x_{i-1}$ to $x_{i}$ is downward. Hence,
[…]
We have thus shown that the length of any path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ is greater than or equal $\mathrm{rank}(\mathcal{G}_{t})-\mathrm{rank}(\mathcal{G}_{s})$. Note that the lower-bound can be achieved by considering an upward path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ with the sequence $(x_{1}:=\mathcal{G}_{s},x_{2},\dots,x_{k}:=\mathcal{G}_{t})$ where $x_{i}$ covers $x_{i-1}$ for every $i=2,3,\dots,k$. Here, the inequality above becomes an equality since # $\text{ downward moves}=0$.
Appendix D Proof of Theorem 2
In what follows, we prove that under Condition˜3 the following holds:
[…]
The case for the up-down distance follows from a dual argument.
Proof.
$\Rightarrow$’ direction.  We prove by induction. Consider $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$. When $\mathcal{G}_{s}=\mathcal{G}_{t}$, both distances are zero; when $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=1$, we must have either $\mathcal{G}_{s}\gtrdot\mathcal{G}_{t}$ or $\mathcal{G}_{s}\lessdot\mathcal{G}_{t}$ so $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})=1$. Now, suppose $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})=d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ holds for all $\mathcal{G}_{s},\mathcal{G}_{t}$ such that $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq k$ for a fixed $k\geq 1$. We argue by contradiction that the equality in distance must also hold for all $\mathcal{G}_{s},\mathcal{G}_{t}$ with $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=k+1$, if any. Suppose this is not true for some $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$. Then, the set of shortest paths between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ must not contain one of down-up shape. From the set, pick path $p=(\mathcal{G}_{s}=\mathcal{G}_{0},\mathcal{G}_{1},\dots,\mathcal{G}_{d}=\mathcal{G}_{t})$ that has the smallest number of up-down segments, and this number is at least one because $p$ cannot be down-up. Let $(\mathcal{G}_{a},\mathcal{G}_{b},\mathcal{G}_{c})$ be the maximal up-down segment that is first encountered when traversing from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ on $p$. That is, $p$ is of the form
$\mathcal{G}_{s}=\mathcal{G}_{0}\gtrdot\cdots\gtrdot\mathcal{G}_{a}\lessdot\cdots\lessdot\mathcal{G}_{b}\gtrdot\cdots\gtrdot\mathcal{G}_{c}\lessdot\cdots\mathcal{G}_{d}=\mathcal{G}_{t},$
where $0\leq a<b<c\leq d$. There are two cases.
  1. i.
    $a=0$ and $c=d$. Path $p$ is up-down (but not a chain). Using lower semimodularity, $p$ can be transformed into a down-up path $p^{\prime}$ with the same length (see Fig.˜4) so we have a contradiction.
  2. ii.
    $a>0$ or $c<d$. The subpath $p(\mathcal{G}_{a},\mathcal{G}_{c})$ has length smaller or equal to $k$. By our induction hypothesis, the subpath can be replaced by a down-up subpath without altering its length, but this leads to a shortest path between $\mathcal{G}_{s},\mathcal{G}_{t}$ with one less up-down segment. This contradicts the definition of $p$.
Hence, we must have a contradiction and the result is proven by induction.
$\Leftarrow$’ direction.  Consider $\mathcal{G}_{s},\mathcal{G}_{t},\mathcal{G}\in\mathfrak{G}$ such that $\mathcal{G}_{s}\neq\mathcal{G}_{t}$ and $\mathcal{G}\gtrdot\mathcal{G}_{s},\mathcal{G}_{t}$. By definition, we have $1\leq d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq 2$. Further, observe that $\mathcal{G}\gtrdot\mathcal{G}_{s},\mathcal{G}_{t}$ implies $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ are incomparable ($\mathcal{G}\gtrdot\mathcal{G}_{s}\succ\mathcal{G}_{t}$ $\implies$ $\mathcal{G}$ cannot cover $\mathcal{G}_{t}$) and hence $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})>1$. Therefore, we have $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=2$. By our hypothesis, there exists a down-up path of length 2 between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$. Using again the fact that $\mathcal{G}_{s},\mathcal{G}_{t}$ are incomparable, this path must be of the form $\mathcal{G}_{s}\gtrdot\mathcal{G}^{\prime}\lessdot\mathcal{G}_{t}$ for some $\mathcal{G}^{\prime}\in\mathfrak{G}$, which concludes the proof. ∎
Appendix E Properties of posets underlying different graph classes
E.1 DAGs
Proof of Proposition˜7.
  1. 1, 2 and 3.
    They directly follow from Proposition˜6 and the fact that given an arbitrary $\mathcal{G}\in\mathfrak{G}$, every subgraph of $\mathcal{G}$ is also in $\mathfrak{G}$.
  2. 4.
    For any $\mathcal{G}_{1}=(\mathrm{V},\mathrm{E}_{1})$ and $\mathcal{G}_{2}=(\mathrm{V},\mathrm{E}_{2})$ in $\mathfrak{G}$, we have $\mathcal{G}_{1}\wedge\mathcal{G}_{2}=(\mathrm{V},\mathrm{E}_{1}\cap\mathrm{E}_{2})$.
  3. 5.
    To see lower semimodularity, suppose $\mathcal{G}_{3}=(\mathrm{V},\mathrm{E}_{3})$ covers two different DAGs $\mathcal{G}_{1}=(\mathrm{V},\mathrm{E}_{1})$ and $\mathcal{G}_{2}=(\mathrm{V},\mathrm{E}_{2})$. By 3., we have $\mathrm{E}_{1}\not\subseteq\mathrm{E}_{2}$, $\mathrm{E}_{2}\not\subseteq\mathrm{E}_{1}$ and $|\mathrm{E}_{1}\triangle\mathrm{E}_{2}|=2$. It follows that $\mathcal{G}_{1},\mathcal{G}_{2}\gtrdot\mathcal{G}_{4}:=(\mathrm{V},\mathrm{E}_{1}\cap\mathrm{E}_{2})$.
It is worth mentioning that when $|\mathrm{V}|>1$, $\mathcal{L}$ is neither a join semi-lattice nor upper semimodular. To see this, consider two DAGs that have an edge in opposite directions. They have no common upper bound by Proposition˜6.
E.2 CPDAGs
Proof of Proposition˜8.
  1. 1.
    This directly follows from definition of $\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}}$.
  2. 2.
    We first show the $\Rightarrow$’ direction. Suppose $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$, which implies $\mathcal{G}_{1}\prec\mathcal{G}_{2}$. By 3. of Lemma˜2, there exists a sequence of DAGs $\mathcal{D}_{0},\mathcal{D}_{1},\dots,\mathcal{D}_{k}$ ($k\geq 1$) with $\mathcal{D}_{0}\in[\mathcal{G}_{1}]$, $\mathcal{D}_{k}\in[\mathcal{G}_{2}]$ and every adjacent pair corresponds to either operation (i) or (ii). Let the subsequence $\mathcal{D}_{j_{1}},\mathcal{D}_{j_{2}},\dots,\mathcal{D}_{j_{l}}$ $(1\leq j_{1}<\dots<j_{l}\leq k)$ be those result from operation (ii). First, observe that $l\geq 1$ because otherwise only operation (i) is involved and $\mathcal{D}_{0},\mathcal{D}_{k}$ must be Markov equivalent, contradicting $\mathcal{G}_{1}\prec\mathcal{G}_{2}$. Then, we argue that $l<2$. To see this, suppose $l\geq 2$. Let $\mathcal{G}_{3}$ be the CPDAG that represents the Markov equivalence class of $\mathcal{D}_{j_{1}}$. It then follows that $\mathcal{G}_{1}\prec\mathcal{G}_{3}\prec\mathcal{G}_{2}$, which contradicts $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$. Hence, we conclude that $l=1$. Because operation (i) stays in the Markov equivalence class, we have $\mathcal{D}_{j_{1}-1}\in[\mathcal{G}_{1}]$, $\mathcal{D}_{j_{1}}\in[\mathcal{G}_{2}]$ and $\mathcal{D}_{j_{1}-1}$ is a subgraph of $\mathcal{D}_{j_{1}}$ with exactly one edge fewer.
    We now show the $\Leftarrow$’ direction. For $\mathcal{G}_{s}=(\mathrm{V},\mathrm{E}_{s}^{-},\mathrm{E}_{s}^{\rightarrow})$, $\mathcal{G}_{t}=(\mathrm{V},\mathrm{E}_{t}^{-},\mathrm{E}_{t}^{\rightarrow})$, suppose $\mathcal{D}_{1}\in[\mathcal{G}_{s}]$, $\mathcal{D}_{2}\in[\mathcal{G}_{t}]$ and $\mathcal{D}_{1}$ is a subgraph of $\mathcal{D}_{2}$ with exactly one edge fewer, which implies
    $(|\mathrm{E}_{t}^{-}|+|\mathrm{E}_{t}^{\rightarrow}|)-(|\mathrm{E}_{s}^{-}|+|\mathrm{E}_{s}^{\rightarrow}|)=1.$
    We want to show $\mathcal{G}_{s}\lessdot\mathcal{G}_{t}$. By 3. of Lemma˜2, we know $\mathcal{G}_{s}\prec\mathcal{G}_{t}$. Suppose $\mathcal{G}_{s}$ is not covered by $\mathcal{G}_{t}$. Then, the chain can be ‘saturated’ as $\mathcal{G}_{s}=\mathcal{G}_{1}\lessdot\mathcal{G}_{2}\lessdot\dots\lessdot\mathcal{G}_{k}=\mathcal{G}_{t}$ with $k>2$. Applying the result in the $\Rightarrow$’ direction above to every adjacent pair in the chain, we have
    $(|\mathrm{E}_{t}^{-}|+|\mathrm{E}_{t}^{\rightarrow}|)-(|\mathrm{E}_{s}^{-}|+|\mathrm{E}_{s}^{\rightarrow}|)=(k-1),$
    contradicting the previous display. Hence, we have $\mathcal{G}_{s}\lessdot\mathcal{G}_{t}$.
  3. 3.
    By 2., every covering relation corresponds to increasing the number of edges by one. Further, by 1., the least element $\mathcal{G}_{\hat{0}}$ has no edge. This fulfills the definition of $\operatorname{\mathrm{rank}}(\cdot)$.
We now show that the model-oriented poset need not be a meet or join semi-lattice in general. Let $\mathcal{L}$ be the poset for CPDAGs over $\mathrm{V}=\{1,2,3\}$. Consider the graphs
$\mathcal{G}_{1}:1-2\phantom{-}3,\quad\mathcal{G}_{2}:1\phantom{-}2-3,\quad\mathcal{G}_{3}:1\rightarrow 2\leftarrow 3,\quad\mathcal{G}_{4}:1-2-3.$
By 2. of Proposition˜8, we have $\mathcal{G}_{1},\mathcal{G}_{2}\lessdot\mathcal{G}_{3},\mathcal{G}_{4}$. To see that $\mathcal{L}$ is not a join semi-lattice, observe that the upper bounds on $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ consists of $\mathcal{G}_{3},\mathcal{G}_{4}$ and the full graph $\mathcal{G}_{\hat{1}}$, where $\mathcal{G}_{3},\mathcal{G}_{4}\prec\mathcal{G}_{\hat{1}}$ but $\mathcal{G}_{3}$ and $\mathcal{G}_{4}$ are incomparable. Similarly, to see that $\mathcal{L}$ is not a meet semi-lattice, observe that the lower bounds on $\mathcal{G}_{3}$ and $\mathcal{G}_{4}$ consists of $\mathcal{G}_{1}$, $\mathcal{G}_{2}$ and the empty graph $\mathcal{G}_{\hat{0}}$. However, while $\mathcal{G}_{1},\mathcal{G}_{2}\succ\mathcal{G}_{\hat{0}}$, $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are incomparable.
Further, we show that the poset need not be upper or lower semimodular. Specifically, let $\mathcal{L}^{\prime}$ be the model-oriented poset for CPDAGs over $\mathrm{V}^{\prime}=\{1,2,3,4\}$. For upper semimodularity, consider graphs
$\mathcal{G}_{1}^{\prime}:\begin{array}[]{ccc}1&\leftarrow&2\\ \uparrow&&|\\ 4&-&3\end{array},\qquad\mathcal{G}_{2}^{\prime}:\begin{array}[]{ccc}1&\rightarrow&2\\ |&&\uparrow\\ 4&-&3\end{array},\qquad\mathcal{G}_{3}^{\prime}:\begin{array}[]{ccc}1&&2\\ |&&|\\ 4&-&3\end{array}.$
By 2. of Proposition˜8, we know $\mathcal{G}_{1}^{\prime},\mathcal{G}_{2}^{\prime}\gtrdot\mathcal{G}_{3}^{\prime}$. If $\mathcal{L}^{\prime}$ is upper semimodular, then there must exist a graph that covers both $\mathcal{G}_{1}^{\prime}$ and $\mathcal{G}_{2}^{\prime}$. Suppose $\mathcal{G}^{\prime}_{4}\gtrdot\mathcal{G}_{1}^{\prime},\mathcal{G}_{2}^{\prime}$. By the same proposition, the skeleton of $\mathcal{G}^{\prime}_{4}$ must be either that of $\mathcal{G}_{5}^{\prime}$ or that of $\mathcal{G}_{6}^{\prime}$:
$\mathcal{G}_{5}^{\prime}:\begin{array}[]{ccc}1&-&2\\ |&\diagdown&|\\ 4&-&3\end{array},\qquad\mathcal{G}_{6}^{\prime}:\begin{array}[]{ccc}1&-&2\\ |&\diagup&|\\ 4&-&3\end{array}.$
[…]
Now we argue that $\mathcal{L}^{\prime}$ is not lower semimodular. For this purpose, consider again $\mathcal{G}_{5}^{\prime}$ and $\mathcal{G}_{6}^{\prime}$ above. Proposition˜8 implies that the full graph covers $\mathcal{G}_{5}^{\prime}$ and $\mathcal{G}_{6}^{\prime}$. If $\mathcal{L}$ is lower semimodular, then there must exist $\mathcal{G}_{7}^{\prime}$ that is covered by both $\mathcal{G}_{5}^{\prime}$ and $\mathcal{G}_{6}^{\prime}$. By the same result, the skeleton of $\mathcal{G}_{7}^{\prime}$ is a 4-cycle and hence $\mathcal{G}_{7}^{\prime}$ must be either
[…]
up to graph isomorphism. One can check that neither can be covered by both $\mathcal{G}_{5}^{\prime}$ and $\mathcal{G}_{6}^{\prime}$.
E.3 MPDAGs
We first show that the model-oriented poset underlying MPDAGs is not a meet or join semi-lattice in general. Specifically, let $\mathcal{L}$ be the model-oriented poset underlying MPDAGs over $\mathrm{V}=\{1,2,3\}$. Consider the following four MPDAGs:
$\mathcal{G}_{1}:\begin{array}[]{ccc}1&\rightarrow&2\\ &&\uparrow\\ &&3\end{array},\qquad\mathcal{G}_{2}:\begin{array}[]{ccc}1&-&2\\ &&|\\ &&3\end{array},\qquad\mathcal{G}_{3}:\begin{array}[]{ccc}1&\rightarrow&2\\ &&\\ &&3\end{array}\qquad\mathcal{G}_{4}:\begin{array}[]{ccc}1&\phantom{-}&2\\ &&\uparrow\\ &&3\end{array},$
which satisfy $\mathcal{G}_{3},\mathcal{G}_{4}\prec\mathcal{G}_{1},\mathcal{G}_{2}$. The upper bounds of $\mathcal{G}_{3}$ and $\mathcal{G}_{4}$ consist of $\mathcal{G}_{1},\mathcal{G}_{2}$ and $\mathcal{G}_{\hat{1}}$, where $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are incomparable — this shows that $\mathcal{L}$ is not a join semi-lattice. Similarly, the lower bounds of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ consist of $\mathcal{G}_{3},\mathcal{G}_{4}$ and $\mathcal{G}_{\hat{0}}$, where $\mathcal{G}_{3},\mathcal{G}_{4}$ are incomparable — this shows that $\mathcal{L}$ is not a meet semi-lattice.
Then, we show that $\mathcal{L}$ is neither lower nor upper semimodular. Consider graphs
$\mathcal{G}_{1}^{\prime}:\begin{array}[]{ccc}1&\rightarrow&2\\ &&\downarrow\\ &&3\end{array},\qquad\mathcal{G}_{2}^{\prime}:\begin{array}[]{ccc}1&&2\\ &\nwarrow&\downarrow\\ &&3\end{array},\qquad\mathcal{G}_{3}^{\prime}:\begin{array}[]{ccc}1&\phantom{-}&2\\ &&\downarrow\\ &&3\end{array},$
which satisfy $\mathcal{G}_{3}^{\prime}\lessdot\mathcal{G}_{1}^{\prime},\mathcal{G}_{2}^{\prime}$. However, there is no graph that covers both $\mathcal{G}_{1}^{\prime}$ and $\mathcal{G}_{2}^{\prime}$ due to the acyclicity constraint, establishing that $\mathcal{L}$ is not upper semimodular. Further, consider the following graphs:
$\mathcal{G}_{1}^{\prime\prime}:\begin{array}[]{ccc}1&-&2\\ &&|\\ 4&&3\end{array},\qquad\mathcal{G}_{2}^{\prime\prime}:\begin{array}[]{ccc}1&{-}&2\\ \uparrow&\nearrow&\downarrow\\ 4&\rightarrow&3\end{array},\qquad\mathcal{G}_{3}^{\prime\prime}:\begin{array}[]{ccc}1&{-}&2\\ \uparrow&\nearrow&|\\ 4&\rightarrow&3\end{array}.$
By construction, $\mathcal{G}_{1}^{\prime\prime}$ and $\mathcal{G}_{2}^{\prime\prime}$ are both covered by $\mathcal{G}_{3}^{\prime\prime}$ (see text in Section 4.4). However, there exists no graph covered by both $\mathcal{G}_{1}^{\prime\prime}$ and $\mathcal{G}_{2}^{\prime\prime}$. To see why, the only two graphs that are covered by $\mathcal{G}_{1}^{\prime\prime}$ are $\mathcal{G}_{1}^{\prime\prime\prime}$ and $\mathcal{G}_{2}^{\prime\prime\prime}$ below:
$\mathcal{G}_{1}^{\prime\prime\prime}:\begin{array}[]{ccc}1&\leftarrow&2\\ &&|\\ 4&&3\end{array},\qquad\mathcal{G}_{2}^{\prime\prime\prime}:\begin{array}[]{ccc}1&{-}&2\\ &&\downarrow\\ 4&&3\end{array},\qquad\mathcal{G}:\begin{array}[]{ccc}1&-&2\\ \uparrow&\nearrow&\downarrow\\ 4&&3\end{array}.$
Note that $\mathcal{G}_{1}^{\prime\prime\prime}$ does not satisfy $\mathcal{G}_{1}^{\prime\prime\prime}\preceq\mathcal{G}_{2}^{\prime\prime}$. Moreover, there exists the graph $\mathcal{G}$ which is intermediate between $\mathcal{G}_{2}^{\prime\prime\prime}$ and $\mathcal{G}_{2}^{\prime\prime}$. Hence, $\mathcal{L}$ is not lower semimodular.
Appendix F Theoretical results for CPDAGs
F.1 Lower bound on $d_{\mathcal{L}}$: proof of Property 2
Proof of Property 2.
Any path from $\mathcal{G}_{s}$ from $\mathcal{G}_{t}$ must involve removing edges that are in the skeleton of $\mathcal{G}_{s}$ and not in the skeleton of $\mathcal{G}_{t}$ and adding edges that are in the skeleton of $\mathcal{G}_{t}$ but not in the skeleton of $\mathcal{G}_{s}$. Since every edge addition or deletion increases the path length by one, we have that $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\geq|\mathrm{sk}(\mathcal{G}_{s})\setminus\mathrm{sk}(\mathcal{G}_{t})|+|\mathrm{sk}(\mathcal{G}_{t})\setminus\mathrm{sk}(\mathcal{G}_{s})|$.
The lower bound above does not account for v-structures that are present in one but not in the other. Specifically, suppose $\mathcal{G}_{s}$ has a v-structure $i\rightarrow j\leftarrow k$ with $i$ and $k$ not connected, and $\mathcal{G}_{t}$ does not have this v-structure, but has an edge between $i$ and $j$, an edge between $j$ and $k$, and no edge between $i$ and $k$. Thus, any path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ must remove this v-structure, which can only be done if and only if at some step at least one of the following happens: $(i)$ the edge between $i$ and $j$ and/or between $j$ and $k$ is removed, and $(ii)$ there is a covering edge that is added between $i$ and $k$. For both $(i)$ and $(j)$, an extra operation afterwards is needed: for $(i)$, it is to add back the edge, and for $(ii)$, it is to remove the covered edge. Similar argument can be made if $\mathcal{G}_{t}$ has a v-structure that $\mathcal{G}_{s}$ does not.
It is natural to define the set:
$\begin{aligned} S:=\Big\{(a,b,c)\subseteq V\times V\times V:&(a,b),(b,c)\subseteq\operatorname{sk}(\mathcal{G}_{s})\cap\operatorname{sk}(\mathcal{G}_{t}),(a,c)\not\in\operatorname{sk}(\mathcal{G}_{s})\cup\operatorname{sk}(\mathcal{G}_{t}),\\ &\text{exactly one of }(a,b,c)\in\operatorname{v-str}(\mathcal{G}_{s})\text{ and }(a,b,c)\in\operatorname{v-str}(\mathcal{G}_{t})\Big\},\end{aligned}$
and say that associated with every element in $S$, we need at least two operations. That is, in addition to $|\mathrm{sk}(\mathcal{G}_{s})\setminus\mathrm{sk}(\mathcal{G}_{t})|+|\mathrm{sk}(\mathcal{G}_{t})\setminus\mathrm{sk}(\mathcal{G}_{s})|$, at least $2|S(\mathcal{G}_{s},\mathcal{G}_{t})|$ additional steps must be involved in any path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$. However, $2|S|$ may overcount the number of operations needed. For example, removing an edge between $i$ and $j$ may account for elements $(i,j,k)$ and $(i,j,m)$ in $S$, or adding a covered edge between $i$ and $k$ may account for elements $(i,j,k)$ and $(i,m,k)$ in $S$.
To avoid overcounting the number of operations, we consider the following strategy. For any pair $(i,j)$, we define two operations. The first, denoted $\operatorname{OP}_{(i,j),\text{CE}}(\cdot)$, takes a subset of $S$, and removes all the elements of the form $(i,k,j)$ or $(j,k,i)$ for any $k$; this operation is associated with adding a covered edge between $i$ and $j$. The second, denoted $\operatorname{OP}_{(i,j),\text{ER}}(\cdot)$, removes all elements of the form $(i,j,k)$, $(k,j,i)$, $(j,i,k),(k,i,j)$ for any $k$; this operation is associated with deleting an edge between $i$ and $j$. Let $\mathrm{OP}=\{\operatorname{OP}_{(i,j),\text{CE}},\operatorname{OP}_{(i,j),\text{ER}}\text{ for all }\text{ vertices }(i,j)\}$ be the collection of all such operations. Our goal is to find the smallest number of operations from $\mathrm{OP}$ such that, when applied consecutively to $S$, the resulting set is empty. That is, we seek a sequence of operations $(\mathrm{op}_{1},\dots,\mathrm{op}_{m})$ where each step $\mathrm{op}_{i}\in\mathrm{OP}$, such that $\mathrm{op}_{1}\odot\dots\odot\mathrm{op}_{m}(S)=\emptyset$, and $m$ is minimized. Let the minimum be $m_{\text{op}}$. Then, it follows that:
$d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\geq|\mathrm{sk}(\mathcal{G}_{s})\setminus\mathrm{sk}(\mathcal{G}_{t})|+|\mathrm{sk}(\mathcal{G}_{t})\setminus\mathrm{sk}(\mathcal{G}_{s})|+2m_{\text{op}}.$
F.2 Simplifications under polytree assumptions: proof of Theorem 3
The result is based on the following lemmas, which we prove first.
Lemma 3.
Let $\mathcal{G}_{a}$ and $\mathcal{G}_{b}$ be two graphs in $\mathfrak{G}$, the set of polytree CPDAGs over a common node set, with $\mathcal{G}_{a}\preceq\mathcal{G}_{b}$. Then, for every DAG $\mathcal{D}_{b}$ represented by $\mathcal{G}_{b}$, there exists a DAG $\mathcal{D}_{a}$ represented by $\mathcal{G}_{a}$, such that $\mathcal{D}_{a}$ is a directed subgraph of $\mathcal{D}_{b}$.
Proof of Lemma 3.
First, note that $\mathrm{sk}(\mathcal{G}_{a})\subseteq\mathrm{sk}(\mathcal{G}_{b})$; otherwise conditional dependencies of $\mathcal{G}_{a}$ are not a subset of those of $\mathcal{G}_{b}$. Further, if $\mathcal{G}_{a}$ has a v-structure $i\rightarrow j\leftarrow k$, then $\mathcal{G}_{b}$ must also have this v-structure. Suppose not, then, since $\mathcal{G}_{b}$ cannot have cycles and its skeleton is a superset of $\mathcal{G}_{a}$, then $\mathcal{G}_{b}$ must have an edge between $i$ to $j$ and $j$ to $k$ without an edge between $i$ to $k$. This leads to a contradiction, however, since $\mathcal{G}_{a}$ encodes the conditional dependencies $X_{i}\not\perp X_{j}|X_{S}$ for every set $S$ that contains $k$, while $\mathcal{G}_{b}$ not has have this dependency for a set $S$. Similarly, if $\mathcal{G}_{b}$ has a v-structure $i\rightarrow j\leftarrow k$, then either $\mathcal{G}_{a}$ has the same v-structure or the edge $i$ to $j$ or $j$ to $k$ is missing in $\mathcal{G}_{a}$. Suppose not, that is $i$ and $j$, and $j$ and $k$ are connected but not in a v-structure form. Then, $\mathcal{G}_{a}$ implies the conditional dependency $X_{i}\not\perp X_{j}|S$ for every $S$ that does not contain $k$. However, $\mathcal{G}_{b}$ implies the conditional independence $X_{i}\perp X_{j}|X_{S}$ for some $S$ that does not contain $k$.
Using the facts above, we conclude the desired result. Let $\mathcal{E}=\{(i,j)\}$ be collection of nodes that are connected in $\mathcal{G}_{b}$ but not in $\mathcal{G}_{a}$. By the analysis above, any v-structure in $\mathcal{G}_{a}$ is also present in $\mathcal{G}_{b}$. Furthermore, every v-structure $i\rightarrow k\leftarrow j$ in $\mathcal{G}_{b}$ is either present in $\mathcal{G}_{a}$ or at least one of the edges between pair of nodes $(i,k)$ and $(j,k)$ is missing. Consider any DAG $\mathcal{D}_{b}$ represented by $\mathcal{G}_{b}$. Let $\mathcal{D}_{a}$ be the DAG that is obtained by deleting the edges in $\mathcal{D}_{b}$ among pairs of nodes $\mathcal{E}$. By construction, $\mathcal{D}_{a}$ has the same skeleton as any DAG in $\mathcal{G}_{a}$. Since $\mathcal{D}_{a}$ is a polytree, every structure present in $\mathcal{D}_{a}$ or in $\mathcal{G}_{a}$ are also in $\mathcal{G}_{b}$. Suppose as a point of contradiction that there exists a v-structure $i\rightarrow k\leftarrow j$ that is in $\mathcal{D}_{a}$ and not in $\mathcal{G}_{a}$. Since this v-structure must also be present in $\mathcal{G}_{b}$, and from the previous statement, then one of the edges between pairs of nodes $(i,k)$ or $(j,k)$ must be missing in $\mathcal{G}_{a}$, which contradicts $\mathcal{D}_{a}$ and $\mathcal{G}_{a}$ have the same skeleton. A similar argument allows to conclude that it is not possible that $\mathcal{G}_{a}$ has a v-structure that is not present in $\mathcal{D}_{a}$. Thus, $\mathcal{D}_{a}$ and $\mathcal{G}_{a}$ have the same skeleton and v-structures. We conclude that $\mathcal{D}_{a}$ must be a DAG represented by $\mathcal{G}_{a}$. ∎
Lemma 4.
The poset $\mathcal{L}_{\text{poly}}$ underlying CPDAGs that are polytrees is graded with $\mathrm{rank}(\cdot)$ measuring the number of edges.
Proof.
Suppose as a point of contradiction that $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathcal{G}_{\text{poly}}$ and $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$ with $\mathrm{rank}(\mathcal{G}_{1})+1<\mathrm{rank}(\mathcal{G}_{2})$. By Lemma 3, for every DAG $\mathcal{D}_{2}=(\mathrm{V},E_{2})$ represented by $\mathcal{G}_{2}$, there exists a DAG $\mathcal{D}_{1}=(\mathrm{V},E_{1})$ represented by $\mathcal{G}_{1}$ such that $E_{1}\subseteq E_{2}$. Since $\mathrm{rank}(\mathcal{G}_{1})+1<\mathrm{rank}(\mathcal{G}_{2})$, there exists $E_{3}$, satisfying $E_{1}\subset E_{3}\subset E_{2}$. Let $\mathcal{G}_{3}$ be the CPDAG obtained by completing the DAG $\mathcal{D}_{3}=(\mathrm{V},E_{3})$. By construction, $\mathcal{G}_{1}\preceq\mathcal{G}_{3}\preceq\mathcal{G}_{2}$ with $\mathcal{G}_{3}$ being distinct from $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. This leads to a contradiction.
Proof of Theorem 3.
We will show that the poset is lower semimodular and then appeal to Theorem 2 to conclude the desired result.
Let $\mathcal{G}_{1},\mathcal{G}_{2},\mathcal{G}_{3}\in\mathcal{G}_{\text{poly}}$ be three CPDAGs that are polytrees. Suppose $\mathcal{G}_{1}\lessdot\mathcal{G}_{3}$ and $\mathcal{G}_{2}\lessdot\mathcal{G}_{3}$. Consider any DAG $\mathcal{D}_{3}=(V,E_{3})$ represented by $\mathcal{G}_{3}$. By Lemma 3, there exist DAGs $\mathcal{D}_{1}=(V,E_{1})$ and $\mathcal{D}_{2}=(V,E_{2})$, represented by $\mathcal{G}_{2}$ and $\mathcal{G}_{1}$, respectively, such that $E_{1}\subseteq E_{3}$ and $E_{2}\subseteq E_{3}$. Since $\mathcal{L}_{\text{poly}}$ is graded by Lemma 4, we have that $E_{3}\setminus E_{1}=e_{1}$ and $E_{3}\setminus{E}_{2}=e_{2}$, where $|e_{1}|=|e_{2}|=1$ and $e_{1}\neq e_{2}$. Let $\tilde{\mathcal{D}}=(\mathrm{V},E_{1}\cap E_{2})$, and define $\tilde{\mathcal{G}}$ to be the CPDAG obtained by completing $\tilde{\mathcal{D}}$. By construction, $\tilde{\mathcal{G}}$ is a polytree and satisfies $\tilde{\mathcal{G}}\preceq\mathcal{G}_{1}$ and $\tilde{\mathcal{G}}\preceq\mathcal{G}_{2}$. Since $\mathrm{rank}(\tilde{\mathcal{G}})=\mathrm{rank}(\mathcal{G}_{1})-1=\mathrm{rank}(\mathcal{G}_{2})-1$, we have that $\tilde{\mathcal{G}}\lessdot\mathcal{G}_{1}$ and $\tilde{\mathcal{G}}\lessdot\mathcal{G}_{2}$. ∎
F.3 Additional theoretical results
Lemma 5.
For the class of CPDAGs, and for any pair of CPDAGs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, we have:
$\max_{\mathcal{G}\preceq\mathcal{G}_{1},\mathcal{G}\preceq\mathcal{G}_{2}}\mathrm{rank}(\mathcal{G})\leq|\mathrm{sk}(\mathcal{G}_{1})\cap\mathrm{sk}(\mathcal{G}_{2})|.$
Proof of Lemma 5.
Notice that for any $\mathcal{G}\preceq\mathcal{G}_{1}$, we have that $\mathrm{sk}(\mathcal{G})\subseteq\mathrm{sk}(\mathcal{G}_{1})$; otherwise, conditional dependencies of $\mathcal{G}$ are not a subset of those of $\mathcal{G}_{1}$. Similarly, we have that for any $\mathcal{G}\preceq\mathcal{G}_{2}$, $\mathrm{sk}(\mathcal{G})\subseteq\mathrm{sk}(\mathcal{G}_{2})$. This allows us to conclude that for $\mathcal{G}$ with $\mathcal{G}\preceq\mathcal{G}_{1}$ and $\mathcal{G}\preceq\mathcal{G}_{2}$, $\mathrm{sk}(\mathcal{G})\subseteq\mathrm{sk}(\mathcal{G}_{1})\cap\mathrm{sk}(\mathcal{G}_{2})$. Since $\mathrm{rank}(\mathcal{G})=|\mathrm{sk}(\mathcal{G})|$, we have the desired result. ∎
Lemma 6.
For the class of CPDAGs, and for any pair of CPDAGs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, we have:
$\min_{\mathcal{G}_{1}\preceq\mathcal{G},\mathcal{G}_{2}\preceq\mathcal{G}}\mathrm{rank}(\mathcal{G})\geq|\mathrm{sk}(\mathcal{G}_{1})\cup\mathrm{sk}(\mathcal{G}_{2})|.$
Proof of Lemma 6.
Based on the same argument as proof of Lemma 5, we have that for $\mathcal{G}$ with $\mathcal{G}_{1}\preceq\mathcal{G}$ and $\mathcal{G}_{2}\preceq\mathcal{G}$, $\mathrm{sk}(\mathcal{G}_{1})\cup\mathrm{sk}(\mathcal{G}_{2})\subseteq\mathrm{sk}(\mathcal{G})$. Since $\mathrm{rank}(\mathcal{G})=|\mathrm{sk}(\mathcal{G})|$, we have the desired result. ∎
Appendix G Theoretical results for MPDAGs
G.1 Poset for MPDAGs is not graded: proof of Proposition 10
To prove that the poset $\mathcal{L}$ underlying MPDAGs is not graded, it suffices to find two maximal chains between a pair of graphs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ of different lengths. Fig.˜9 presents a collection of graphs that specify a maximal chain from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$. This chain has length $3$. Similarly, Fig.˜10 presents another collection of graphs that specify a maximal chain from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$. This chain, however, has length $5$. As the length of these two maximal chains are not equal, we conclude that $\mathcal{L}$ is not graded.
Figure 9: A maximal chain between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$, specified by the sequence $(\mathcal{G}_{1},\mathcal{G}_{2},\mathcal{G}_{3},\mathcal{G}_{4})$ where $\mathcal{G}_{i+1}$ covers $\mathcal{G}_{i+1}$ for $i=1,2,3$.
Figure 10: A maximal chain between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$, specified by the sequence $(\mathcal{G}_{1},\mathcal{G}_{2},\mathcal{G}_{3},\mathcal{G}_{4},\mathcal{G}_{5},\mathcal{G}_{6})$ where $\mathcal{G}_{i+1}$ covers $\mathcal{G}_{i+1}$ for $i=1,2,3,4,5$.
G.2 MPDAG model complexity: proof of Proposition 12
We construct an explicit sequence of MPDAG from $\mathcal{G}$ to the empty MPDAG by first directing one undirected at a time in $\mathcal{G}$ to obtain a sequence $\mathcal{G}=\mathcal{G}_{1},\dots,\mathcal{G}_{k}$ such that $\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G}_{i})-\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G}_{i+1})=1,\mathcal{G}_{i}\succeq\mathcal{G}_{i+1}$, $\mathcal{G}_{k}$ is a DAG and where $k$ is the number of undirected edges in $\mathcal{G}$. This is possible by repeatedly applying the result from Lemma 7. We append to this sequence another sequence $\mathcal{G}_{k}=\tilde{\mathcal{G}}_{1},\dots,\tilde{\mathcal{G}}_{l}$ such that $\mathrm{pseudo}\text{-}\mathrm{rank}(\tilde{\mathcal{G}}_{i})-\mathrm{pseudo}\text{-}\mathrm{rank}(\tilde{\mathcal{G}}_{i+1})=1,\tilde{G}_{i}\succeq\tilde{G}_{i+1}$, all $\tilde{\mathcal{G}}_{i}$ are DAGs, $\tilde{\mathcal{G}}_{l}$ is the empty graph and $l$ is the number of directed edges in $\mathcal{G}$. We can obtain such a sequence simply by deleting the edges in $G_{k}$ in any random order. By combining the two sequence we obtain a $\mathrm{pseudo}\text{-}\mathrm{rank}(G)$-length sequence of graphs from $G$ to the empty graph such that adjacent graphs on the sequence neighbors in $\mathcal{L}$ while having pseudo-ranks that differ by $1$.
Lemma 7.
Consider an MPDAG $\mathcal{G}$. Then there exists an undirected edge in $\mathcal{G}$ such that directing the edge results in an MPDAG $\tilde{\mathcal{G}}$ which is equal to $\mathcal{G}$ except for the one additional directed edge.
Figure 11: Forbidden induced subgraphs of an MPDAG (left) and induced subgraphs that may exist in an MPDAG and are one directed edge away from a forbidden subgraph (right).
Proof.
Consider $\mathcal{G}$ and suppose there exists a pair $(a,b)$, such that aba-b is in $\mathcal{G}$, $\mathrm{Sib}(a)\subseteq\mathrm{Sib}(b)\cup\mathrm{Ch}(b)$ and $b$ is a minimal element in the partial topological ordering of $\mathcal{G}$ among all nodes with an adjacent undirected edge. Here, $Sib(v)$ is the set of nodes $u$ such that uvu-v called the sibling of $v$, while $Ch(v)$ is the set of nodes $v\rightarrow u$ called the set of children of $v$. We first show that directing the edge to $a\rightarrow b$ results in an MPDAG $\tilde{\mathcal{G}}$ before showing that there must always exist such a pair. To show that $\tilde{\mathcal{G}}$ is an MPDAG it suffices to show that it does not contain any of the forbidden subgraphs shown on the left hand size of Fig. 11. In order for that to possibly occur $\mathcal{G}$ must contain a pre-forbidden structure as shown on the right hand size of Fig. 11. We now prove that if such a pre-forbidden structure exists we will never direct an edge that will result in a forbidden structure.
Consider the first pre-forbidden subgraph. The node in the middle always has at least one sibling that the two adjacent nodes do not have. Therefore, the middle node can never be selected as $b$ and we will never arrive at the forbidden structure. Consider the second pre-forbidden subgraph. The left most node is not minimal in the topological order among all nodes with an adjacent undirected edge and can therefore not be selected as $b$. The right most-node has one more sibling (the left most node) than the middle node. Therefore, we can never arrive at the forbidden subgraph. There is no pre-forbidden graph for the third structure. For the fourth pre-forbidden graph the node at the bottom-left is not minimal and can therefore never be selected as $B$.
It now remains to show that such a pair $(a,b)$ exists in any MPDAG that has at least one undirected edge. Consider the CPDAG $\mathcal{C}$ of $\mathcal{G}$. The connected undirected components of $\mathcal{C}$ are chordal and therefore each have a perfect numbering, i.e., an ordering $v_{1},\dots,v_{p}$ such that each $v_{i}$ is simplical in the induced subgraph of $\mathcal{C}$ on $v_{i},\dots,v_{p}$. Here a node $v$ is simplical if it’s neighborhood $\mathrm{Sib}(v)$ forms a clique. Furthermore, each member of a connected undirected component of $\mathcal{C}$ is equivalent in terms of the graphs topological order (Andersson, Madigan and Perlman, 1997). Therefore, the simplical node $s$ in the undirected component that is minimal in terms of the order satisfies i) minimality among all nodes with a sibling and ii) for any sibling $s^{\prime}$ of $s$ it holds that $\mathrm{Sib}(s)\subseteq\mathrm{Sib}(s^{\prime})$. In any MPDAG in which $s$ remains minimal among all nodes with siblings, any node $s^{\prime}$ topologically maximal among the siblings of $s$ satisfies $\mathrm{Sib}(s)\subseteq\mathrm{Sib}(s^{\prime})\cup\mathrm{Ch}(s^{\prime})$.
It thus only remains to consider the case that $s$ is not topologically minimal in $\mathcal{G}$. This implies that there exists an edge out of $s$ in $\mathcal{G}$ into a node $c$ that is a sibling of $s$ in $\mathcal{C}$. In order to avoid any forbidden structure of the first kind any undirected edge adjacent to $c$ in $\mathcal{C}$ that is towards a non-sibling of $s$ in $\mathcal{C}$ must also be directed in $\mathcal{G}$. Therefore, any sibling of $c$ in $\mathcal{G}$ must have been a sibling of $s$ in $\mathcal{G}$. As the siblings of $s$ in $\mathcal{G}$ form a clique it follows that there must exist a sibling $s^{\prime}$ of $c$ such that $\mathrm{Sib}(c)\subseteq\mathrm{Sib}(s^{\prime})\cup\mathrm{Ch}(s^{\prime})$. If $c$ is minimal it therefore satisfies the conditions we make on $b$. If not we repeat the argument until we arrive at such a $b$ through breadth first search.
If there are no undirected edges adjacent to $s$ in $\mathcal{G}$ we move to the next simplical node in the perfect ordering of the undirected subgraph of $\mathcal{C}$ and repeat the argument above.
G.3 Pseudo-distance upper-bounding SHD: proof of Proposition 13
Proof.
The proof is based on the following observation: for any pair of graphs $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ with $\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G}_{2})-\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G}_{1})=1$, the graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are identical except for a change in a single edge. This change can be one of the following: a missing edge becomes a directed edge, or a directed edge becomes an undirected edge.
Consider a pair of vertices $(v_{i},v_{j})$ such that the edge between them differs in $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$. Then, based on the observation above, for any path $p\in L_{\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$, there exists at least one move along the path between graphs $\mathcal{G}_{x}$ and $\mathcal{G}_{y}$ where only this edge changes. Note that if the edge is missing in one graph and undirected in the other, or if the edge is directed in opposite orientations in the two graphs, then at least two such moves exist in the path. This implies that $d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is lower bounded by the structural Hamming distance. ∎
G.4 Simplifications under polytree assumptions: proof of Theorem 4
The proof relies on the following lemma, which we prove first. Throughout, we note that an MPDAG $\mathcal{G}$ can be viewed as containing all DAGs that are members of a CPDAG $\mathcal{C}$ subject to a set of fixed directed edges.
Lemma 8.
Let $\mathcal{G}$ be a MPDAG that is a polytree and represents the collection of DAGs $\mathcal{D}_{1},\dots,\mathcal{D}_{k}$. Consider any pair of nodes $(i,j)$ that are connected in $\mathcal{G}$ and let $\tilde{\mathcal{D}}_{l}$ be the DAG obtained by removing the edge between nodes $(i,j)$ in $\mathcal{D}_{l}$. Then, the collection $\tilde{\mathcal{D}}_{1},\tilde{\mathcal{D}}_{2}\dots,\tilde{\mathcal{D}}_{k}$ forms a MPDAG $\tilde{\mathcal{G}}$ with $\tilde{\mathcal{G}}\preceq\mathcal{G}$.
Proof of Lemma˜8.
Since $\mathcal{G}$ is a valid MPDAG, all the DAGs it represents have the same skeleton and v-structures. By construction, all modified DAGs (after removing the edge) have the same skeleton. Further, they all have the same v-structure. In particular, each $\mathcal{D}_{i}$ is a polytree, and removing an edge does not uncover a v-structure. Further, if the edge removal removes a v-structure in one modified graph, it removes it for all the other modified graphs. We have thus shown that the collection of DAGs $\{\tilde{\mathcal{D}}_{1},\tilde{\mathcal{D}}_{2},\dots,\tilde{\mathcal{D}}_{k}\}$ are all Markov equivalent and members of a CPDAG $\tilde{\mathcal{C}}$.
It remains to check that the collection $\{\tilde{\mathcal{D}}_{1},\tilde{\mathcal{D}}_{2},\dots,\tilde{\mathcal{D}}_{k}\}$ forms a valid MPDAG. That is, the collection represents all DAGs that are members of $\tilde{\mathcal{C}}$, subject to a set of fixed directed edges $\tilde{\mathcal{E}}=\{i_{1}\to i_{2},i_{3}\to i_{4},\dots\}$. Let $\mathcal{C}$ be the CPDAG associated with the collection $\{\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{k}\}$. By definition, $\{\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{k}\}$ is all DAGs that are members of $\mathcal{C}$}, subject to a set of fixed directed edges ${\mathcal{E}}=\{j_{1}\to j_{2},j_{3}\to j_{4},\dots\}$. Let
$\tilde{\mathcal{E}}=\mathcal{E}\cup\{k\to j:(i,j,k)\in\operatorname{v-str}(\mathcal{C})\}\cup\{k\to i:(j,i,k)\in\operatorname{v-str}(\mathcal{C})\}.$
Let $\tilde{\mathcal{G}}$ be the MPDAG obtained by restricting $\tilde{\mathcal{C}}$ to the edge set $\tilde{\mathcal{E}}$. We will show that the collection of DAGs represented by $\tilde{\mathcal{\mathcal{G}}}$ is precisely $\{\tilde{\mathcal{D}}_{1},\tilde{\mathcal{D}}_{2},\dots,\tilde{\mathcal{D}}_{k}\}$. First, we show the collection of DAGs represented by $\tilde{\mathcal{\mathcal{G}}}$ is a subset of $\{\tilde{\mathcal{D}}_{1},\tilde{\mathcal{D}}_{2},\dots,\tilde{\mathcal{D}}_{k}\}$. Take a DAG $\tilde{\mathcal{D}}$ represented by $\tilde{\mathcal{G}}$. By construction, $\tilde{\mathcal{D}}$ is compatible with edge set $\mathcal{E}$ (since $\mathcal{E}\subseteq\tilde{\mathcal{E}}$). Add the edge between nodes $(i,j)$ in the same orientation as any DAG in $\mathcal{\mathcal{G}}$ to obtain a DAG $\bar{\mathcal{D}}$. Then, by construction, $\bar{\mathcal{D}}$ is a member of $\mathcal{C}$, and is compatible with the edge set $\mathcal{E}$. Thus, $\bar{\mathcal{D}}$ is among the collection $\{\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{k}\}$, which means that $\tilde{\mathcal{D}}$ is among the collection $\{\tilde{\mathcal{D}}_{1},\tilde{\mathcal{D}}_{2},\dots,\tilde{\mathcal{D}}_{k}\}$. Next, we show the collection of DAGs represented by $\tilde{\mathcal{\mathcal{G}}}$ is a superset of $\{\tilde{\mathcal{D}}_{1},\tilde{\mathcal{D}}_{2},\dots,\tilde{\mathcal{D}}_{k}\}$. Recall that $\tilde{\mathcal{D}}_{i}$ is a member of $\tilde{\mathcal{C}}$ for every $i=1,2,\dots,k$. Furthermore, by construction, each $\tilde{\mathcal{D}}_{i}$ is compatible with $\tilde{\mathcal{E}}$. This direction then follows.
Proof of Theorem 4.
Let $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ be two MPDAGs that are polytrees with $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$. Suppose as a point of contradiction that $\mathrm{rank}(\mathcal{G}_{1})+1<\mathrm{rank}(\mathcal{G}_{2})$ and there does not exist a distinct MPDAG that is a polytree in between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. We consider three cases and show that there is a contradiction in each case.
Case I: Two or more pairs of nodes are connected in $\mathcal{G}_{2}$ but not in $\mathcal{G}_{1}$. Let $(i,j)$ be one such pair. Let $\{\mathcal{D}_{1},\dots,\mathcal{D}_{k}\}$ be the collection of DAGs represented by $\mathcal{G}_{2}$. Let $\{\tilde{\mathcal{D}}_{1},\dots,\tilde{\mathcal{D}}_{k}\}$ to be the collection of DAGs obtained by removing the edge between pair $(i,j)$ from each $\mathcal{D}_{l},l=1,2,\dots,k$. By Lemma 8, the collection form an MPDAG $\tilde{\mathcal{G}}$ with $\tilde{\mathcal{\mathcal{G}}}\preceq\mathcal{G}_{2}$. Since $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$, we have that for every DAG $\mathcal{D}$ represented by $\mathcal{G}_{1}$, there exists a DAG ${\mathcal{D}}_{l}$ represented by $\mathcal{G}_{2}$ where $\mathcal{D}$ is a subgraph of ${\mathcal{D}}_{l}$. It then follows that $\mathcal{D}$ is also a subgraph of ${\mathcal{D}}_{l}$. So $\mathcal{G}_{1}\preceq\tilde{\mathcal{G}}$.
Case II: There is at least one pair of node that is connected via an undirected edge in $\mathcal{G}_{2}$ but is connected via a directed edge in $\mathcal{G}_{1}$, and there is one pair of nodes that are connected in $\mathcal{G}_{2}$ but not in $\mathcal{G}_{1}$. Suppose $\mathcal{G}_{2}$ is the MPDAG obtained by restricting the CPDAG $\mathcal{C}$ to edge set $\mathcal{E}$. Consider one of the undirected edges between nodes $(i,j)$ that is directed as $i\to j$ in $\mathcal{G}_{1}$. Let $\tilde{\mathcal{G}}$ be the MPDAG obtained by restricting $\mathcal{C}$ to the edge set $\mathcal{E}\cup\{i\to j\}$. Then, by construction, $\tilde{\mathcal{G}}\preceq\mathcal{G}_{2}$ with $\tilde{\mathcal{G}}\neq\mathcal{G}_{2}$. Furthermore, since $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$, for any DAG $\mathcal{D}_{1}$ represented by $\mathcal{G}_{1}$, there exists a DAG $\mathcal{D}_{2}$ represented by $\mathcal{G}_{2}$ such that $\mathcal{D}_{1}$ is a sub-graph of $\mathcal{D}_{2}$. Since $i\to j$ in $\mathcal{D}_{1}$, then $i\to j$ in $\mathcal{D}_{2}$. This implies that $\mathcal{D}_{2}$ is also represented by $\tilde{G}$, which consequently implies that $\mathcal{G}_{1}\preceq\tilde{\mathcal{G}}$. Since $\tilde{\mathcal{G}}$ has still at least one more edge than $\mathcal{G}_{1}$, we have that $\mathcal{G}_{1}\neq\tilde{\mathcal{G}}$.
Case III: There is at least two pair of node that are connected via an undirected edge in $\mathcal{G}_{2}$ but are connected via a directed edge in $\mathcal{G}_{1}$. Suppose $\mathcal{G}_{2}$ is the MPDAG obtained by restricting CPDAG $\mathcal{C}$ to edge set $\mathcal{E}$. Let $\tilde{\mathcal{G}}$ be the MPDAG obtained by restricting $\mathcal{C}$ to $\mathcal{E}\cup\{i\to j\}$. Following the argument as Case II, we have that $\tilde{\mathcal{G}}\preceq\mathcal{G}_{2}$ and $\mathcal{G}_{1}\preceq\tilde{\mathcal{G}}$. By construction, $\tilde{\mathcal{G}}\neq\mathcal{G}_{2}$. It remains to show that $\tilde{\mathcal{G}}\neq\mathcal{G}_{1}$.
Note that each pair of nodes that have a directed edge in $\mathcal{G}_{1}$ (but undirected in $\mathcal{G}_{2}$) cannot be part of a v-structure in $\mathcal{G}_{1}$. Otherwise, since $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are both polytrees, all the v-structures in $\mathcal{G}_{1}$ must also be present in $\mathcal{G}_{2}$, and so these edges would also need to be directed in $\mathcal{G}_{2}$.
We consider some sub-cases. In the first subcase, the two undirected edges are incident with one edge being between $(i,j)$ and the other between $(j,k)$. In this case, to avoid creating a v-structure, because of the polytree condition, $j\to k$ in $\mathcal{G}_{1}$. Suppose we let $\tilde{\mathcal{E}}=\mathcal{E}\cup\{j\to k\}$, and let $\tilde{\mathcal{G}}$ be the MPDAG obtained by restricting $\mathcal{C}$ to the edge set $\tilde{\mathcal{E}}$. Since the edge between $(i,j)$ is undirected in $\mathcal{G}_{2}$, we have that $\tilde{\mathcal{G}}$ has an undirected edge between node pair $(i,j)$. Thus $\tilde{\mathcal{G}}\neq\mathcal{G}_{1}$ in this subcase.
In the second subcase, the two undirected edges are incident with one edge being between $(i,j)$ and the other between $(i,k)$. Since the edge between $(i,k)$ is undirected in $\mathcal{G}_{2}$, we have that there must exist DAGs represented by $\mathcal{G}_{2}$ with the edge $k\to i$ and an edge $i\to j$ as well as DAGs with edge $i\to k$ and $i\to j$, as these do not create v-structures and there is no concern about violating acyclicity due to the polytree condition. We thus have that that $\tilde{\mathcal{G}}$ has an undirected edge between node pair $(i,k)$, and consequently $\tilde{\mathcal{G}}\neq\mathcal{G}_{1}$.
In the final subcase, the two undirected edges in $\mathcal{G}_{2}$ that are directed in $\mathcal{G}_{1}$ are not incident, i.e., they involve pairs $(i,j)$ and $(l,m)$. To be distinct from the previous subcases, together with the fact that $i\to j$ in $\mathcal{G}_{1}$ should not create a v-structure, means that all other edges involving $j$ must be of the form $j\to k$ in $\mathcal{G}_{1}$. This fact isolates the two edges between $(i,j)$ and $(l,m)$, meaning that taking among all DAGs represented by $\mathcal{G}_{2}$ that have the edge $i\to j$, there exists ones that have the edge $l\to m$ and ones that have the edge $m\to l$. Thus, the intermediate graph $\tilde{\mathcal{G}}$ is distinct from $\mathcal{G}_{1}$. ∎
Appendix H On the Idiosyncrasies of the structural interventional distance and separation distance
H.1 Structural intervention distance
The structural intervention distance (SID) measures the distance between a candidate graph $\mathcal{G}_{\mathrm{guess}}$ and a ground-truth graph $\mathcal{G}_{\mathrm{true}}$ by counting the number of interventional distributions of the form $f(y\mid do(X=x))$ that would be incorrectly inferred if parent adjustment were applied according to $\mathcal{G}_{\mathrm{guess}}$, while the true causal model conforms to $\mathcal{G}_{\mathrm{true}}$. SID was originally proposed as an alternative to SHD, aiming to account for the fact that the graphs being compared are causal rather than purely structural. In this sense, it represents an attempt to define a model-oriented distance specifically for causal DAGs.
However, as noted by Peters and Bühlmann (2015), SID is not a metric. The SID between a pair of DAGs is zero if and only if $\mathcal{G}_{\mathrm{guess}}$ is a subgraph of $\mathcal{G}_{\mathrm{true}}$. Consequently, SID does not satisfy the standard metric properties. It fails to ensure that $d_{\mathrm{SID}}(\mathcal{G}_{1},\mathcal{G}_{2})=0$ if and only if $\mathcal{G}_{1}=\mathcal{G}_{2}$, is not symmetric and does not obey the triangle inequality. It is, however, possible to symmetrize SID by averaging the two asymmetric distances for each pair of graphs. The resulting symmetrized SID satisfies the first two axioms of a metric, but as we will show with an example, it still fails to satisfy the triangle inequality.
Figure 12: Three DAGs for which symmetrized SID violates the triangle inequality as shown in Example 2.
Example 2 (Symmetrized SID violates triangle inequality).
Let $\mathcal{G}_{1}$, $\mathcal{G}_{2}$, and $\mathcal{G}_{3}$ be the three DAGs from Figure 12. Denote $d_{SID}$, the structural intervention distance for DAGs as defined by (Peters and Bühlmann, 2015). Using the open source implementation for computing SID by (Henckel, Würtzen and Weichwald, 2024) we obtain that $d_{SID}(\mathcal{G}_{1},\mathcal{G}_{2})=3,d_{SID}(\mathcal{G}_{2},\mathcal{G}_{1})=4,d_{SID}(\mathcal{G}_{1},\mathcal{G}_{3})=5,d_{SID}(\mathcal{G}_{2},\mathcal{G}_{3})=5,d_{SID}(\mathcal{G}_{2},\mathcal{G}_{3})=9$ and $d_{SID}(\mathcal{G}_{3},\mathcal{G}_{2})=9$. For the symmetrized version of SID, denoted $d_{\mathrm{sySID}}$, it therefore holds that
$d_{\mathrm{sySID}}(G_{1},G_{2})+d_{\mathrm{sySID}}(G_{2},G_{3})=\frac{3+4}{2}+\frac{5+5}{2}=8.5<9=\frac{9+9}{2}=d_{\mathrm{sySID}}(G_{1},G_{3}),$
which shows that it does not satisfy triangle inequality and therefore fails to be a metric.
H.2 Separation distance
We interpret a probabilistic graphical model as the collection of all random variables satisfying the conditional independence constraints encoded by the graph. A closely related and popular alternative is to interpret a probabilistic graphical model directly as the collection of conditional independence statements encoded by the graph or more formally the set of all triples $(X,Y,Z)$ corresponding to a conditional independence statement of the form $X\perp\!\!\!\!\perp Y\mid Z$ (Studenỳ, 1992). Based on this interpretation of a CPDAG as a collection of triples (Wahl and Runge, 2025) have proposed an alternative distance notion called the separation distance. It defines the distance between two CPDAGs as the size of the symmetric difference of the two corresponding sets of triples. Similarly to our class of distances, the separation distance is a metric and model oriented, albeit on a different model interpretation. It can also be generalized to any class of probabilistic graphical models that encode only conditional independence statements. In fact any separation distance can be viewed as a minimal path distance on a Hasse diagram: Consider the set of all triples $X,Y,Z\subseteq\{1,\dots,n\}$ and the poset structure induced by subset inclusion on this set. This poset forms a lattice, with the meet and join between two sets of triples defined as their intersection and union, respectively. As a result, the minimal path distance on the corresponding Hasse diagram equals the down-up distance, which in turn equals the size of the symmetric difference between the two sets.
However, the fact that the separation distance is implicitly defined on the set of all triples of subsets of the node set, while ignoring which of these correspond to a CPDAG or even valid conditional independence models for that matter has two serious consequences. First, the separation distance between sparse graphs tends to be much larger than between dense graphs, simply because the former encode more separation statements to begin with and as a result dropping an additional statement typically triggers a cascade of implications. Second, the separation distance between two graphs increases substantially when a disconnected node is added to both. We now illustrate these characteristics with an example.
Example 3.
Let $\mathcal{G}_{1}$ denote the empty CPDAG, $\mathcal{G}_{2}$ a CPDAG with a single (undirected) edge, $\mathcal{G}_{3}$ the fully connected CPDAG, and $\mathcal{G}_{4}$ the same graph with one (undirected) edge removed, all on $n$ nodes. Since all d-separation statements hold in $\mathcal{G}_{1}$, it can be represented by the set of all triples of subsets on the set of $n$ nodes. In $\mathcal{G}_{2}$, on the other hand, all statements hold except those of the form
$\mathrm{V}_{i}\perp_{\mathcal{G}_{2}}\mathrm{V}_{j}\mid Z,\quad Z\subseteq V\setminus\{\mathrm{V}_{i},\mathrm{V}_{j}\}.$
As a result, the separation distance between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ is the size of this latter set, that is, $2^{n-1}$. Furthermore, for every disconnected node we add to both $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, their distance doubles. The CPDAG $\mathcal{G}_{3}$, on the other hand, encodes no d-separation statements and therefore corresponds to the empty set, while $\mathcal{G}_{4}$ encodes the single statement $\mathrm{V}_{i}\perp\mathrm{V}_{j}\mid\mathrm{V}\setminus\{\mathrm{V}_{i},\mathrm{V}_{j}\}$. Hence, the separation distance between $\mathcal{G}_{3}$ and $\mathcal{G}_{4}$ equals $1$. Again adding a disconnected node to both graphs doubles this distance.
In contrast, $d_{\mathcal{L}}(\mathcal{G}_{1},\mathcal{G}_{2})=d_{\mathcal{L}}(\mathcal{G}_{3},\mathcal{G}_{4})=1$, as $\mathcal{G}_{2}$ is a covering element of $\mathcal{G}_{1}$ and $\mathcal{G}_{4}$ of $\mathcal{G}_{3}$ in $\mathcal{L}$. This reflects the fact that none of the triples intermediate between the triple representing $\mathcal{G}_{1}$ and that representing $\mathcal{G}_{2}$ correspond to a CPDAG, or even to a valid conditional independence model. Furthermore, these distances remain unchanged if we add a disconnected node to all four graphs.
Appendix I Algorithmic Aspects of Computing the $\mathcal{M}$-oriented distance
We present additional algorithms that are important for finding the $\mathcal{M}$-oriented distance in the context of CPDAGs.
Up-down distance: The algorithm for computing the up-down distance is the ‘dual’ of Algorithm 2. In particular, (i) the neighbor selection step is restricted to be upward; (ii) the search terminates when a graph $\mathcal{G}^{\prime}$ is found with $\mathcal{G}_{t}\preceq\mathcal{G}^{\prime}$; (iii) we use the admissible heuristic $2|\mathrm{sk}(\mathcal{G}_{t})\cup\mathrm{sk}(\mathcal{G}^{\prime})|-\mathrm{rank}(\mathcal{G}_{t})-\mathrm{rank}(\mathcal{G}^{\prime})$; and finally (iv) we verify whether $\mathrm{rank}(\mathcal{G}^{\prime})\geq u:=|\mathrm{sk}(\mathcal{G}_{t})\cup\mathrm{sk}(\mathcal{G}^{\prime})|$ before checking if $\mathcal{G}_{t}\preceq\mathcal{G}^{\prime}$. Here, $u$ serves as the upper bound for $\rho_{\mathrm{join}}(\mathcal{G}_{s},\mathcal{G}_{t}):=\min_{\mathcal{G}\in\mathfrak{G},\mathcal{G}_{s}\preceq\mathcal{G},\mathcal{G}_{t}\preceq\mathcal{G}}\mathrm{rank}(\mathcal{G})$; see Lemma 6 in Section F.3 for proof. Algorithm 3 presents the full description of the procedure for finding the up-down distance.
input : Graph class of CPDAGs $\mathfrak{G}$, pair of graphs $\mathcal{G}_{s},\mathcal{G}_{t}$, a heuristic $h\colon\mathfrak{G}\to\mathbb{R}_{\geq 0}$, upwards neighbor function $un_{\mathcal{L}}\colon\mathfrak{G}\to 2^{\mathfrak{G}}$, upper bound $u$ to $\rho_{\text{join}}(\mathcal{G}_{s},\mathcal{G}_{t})$
output : Up-down distance $d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t})$
1 begin
 $openSet\coloneqq\{\mathcal{G}_{s}\}$;
 // priority queue ordered by $g+h$
2 $g(\mathcal{G}_{s})\coloneqq 0$;
3 while $openSet\neq\emptyset$ do
4    $\mathcal{G}_{\text{current}}\coloneqq$ graph in openSetopenSet minimizing $g(\mathcal{G}_{\text{current}})+h(\mathcal{G}_{\text{current}})$;
5    if $u\leq\mathrm{rank}(\mathcal{G}_{\text{current}})\,\,\land\,\,\mathcal{G}_{t}\preceq\mathcal{G}_{\text{current}}$ then
6       return $g(\mathcal{G}_{\text{current}})+\mathrm{rank}(\mathcal{G}_{\text{current}})-\mathrm{rank}(\mathcal{G}_{t})$
7     remove $\mathcal{G}_{\text{current}}$ from openSetopenSet;
8
9    foreach $\mathcal{G}_{\text{neighbor}}\in$ un($\mathcal{G}_{\text{current}}$) do
10       $tentative\coloneqq g(\mathcal{G}_{\text{current}})+1$;
[…]
12          $g(\mathcal{G}_{\text{neighbor}})\coloneqq tentative$;
13            add $\mathcal{G}_{\text{neighbor}}$ to $openSet$ with priority $g(\mathcal{G}_{\text{neighbor}})+h(\mathcal{G}_{\text{neighbor}})$;
Algorithm 3 An A* algorithm for computing the up-down distance $d_{\L ,\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t})$
Finding a lower-bound for $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$: We next present the algorithm for computing the lower bound for $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ according to Property 2. Algorithm 4 presents the algorithm which uses Algorithm 5.
input : CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ among nodes $V$, functions getDisjointV and findMinOps
output : The lower bound for the $\mathcal{M}$-oriented distance between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$
1 begin
2 $\mathcal{G}_{t,\text{undir}}\coloneqq undirect(\mathcal{G}_{t})$
3 $\mathcal{G}_{s,\text{undir}}\coloneqq undirect(\mathcal{G}_{s})$
 $d\coloneq|\mathrm{sk}(\mathcal{G}_{t,\text{undir}})\triangle\mathrm{sk}(\mathcal{G}_{s,\text{undir}})|$ ;
 // skeleton difference
 // Get all v structures
 $vList\coloneqq[]$ ;
 // list of uncommon undirected v-structures in the skeletons of the two graphs
4 foreach node $i\in V$ do
5    foreach node $j\in V$ do
6       foreach node $k\in V$ do
7          if $i\neq j\land j\neq k\land i\neq k\land((i,j)\in E(\mathcal{G}_{t,\text{undir}})\land(j,k)\in E(\mathcal{G}_{t,\text{undir}}))\land((i,j)\in E(\mathcal{G}_{s,\text{undir}})\land(j,k)\in E(\mathcal{G}_{s,\text{undir}}))\land(i,k)\notin E(\mathcal{G}_{t,\text{undir}})\land(i,k)\notin E(\mathcal{G}_{s,\text{undir}})$ then
8             if $((i,j)\in E(\mathcal{G}_{t})\land(k,j)\in E(\mathcal{G}_{t})\land(j,i)\notin E(\mathcal{G}_{t})\land(j,k)\notin E(\mathcal{G}_{t})\land((i,j)\notin E(\mathcal{G}_{s}))\lor(k,j)\notin E(\mathcal{G}_{s})))\lor((i,j)\in E(\mathcal{G}_{s})\land(k,j)\in E(\mathcal{G}_{s})\land(j,i)\notin E(\mathcal{G}_{s})\land(j,k)\notin E(\mathcal{G}_{s})\land((i,j)\notin E(\mathcal{G}_{t}))\lor(k,j)\notin E(\mathcal{G}_{t})))$ then
9                  append v structure $(i,j,k)$ to vListvList
 $groups\coloneqq getDistjointVGroups(vList)$ // get v structure groups with no shared vertices
10 $h\coloneqq 0$
11 if $|vList|>0$ then
12    foreach $group\in groups$ do
13       $h\coloneq h+2(findMinOps(group))$
14 return h + d
Algorithm 4 Algorithm for computing the lower bound for the $\mathcal{M}$-oriented distance according to Property 2.
input : Set $S$ consisting of a collection of triplets of vertices, operations $\mathrm{OP}$
output : Returns the minimum number of operations to remove all elements in $S$
1 begin
 $queue\coloneqq[(S,0)]$ // tracks state and step count
 $visited\coloneqq\{\}$ // avoid repeat states
2
3 while $queue\neq\emptyset$ do
4    $current\coloneqq$ get first element of the queue
5      remove first element of queuequeue
6
7    $subS\coloneqq$ subset of S in currentcurrent
8    $steps\coloneqq$ steps in currentcurrent
9    if $subS$ is empty then
10       return steps
    // Get all pairs (i, j) from triples in S
    $pairs\coloneqq{}$ // set of unique pairs from triples
11    foreach $triple\in subS$ do
12         add all combinations of 2 vertices from the tripletriple of vertices to pairspairs
13    foreach $pair\in Lairs$ do
14       $i\coloneqq$ first node in pairpair
15       $j\coloneqq$ second node in pairpair
16
17       foreach $op\in\mathrm{OP}$ do
18          $newS\coloneqq op(subS,i,j)$
19          if $newS\notin visited$ then
20               add $newS$ to visitedvisited
21               add $(newS,steps+1)$ to queuequeue
 return NULL // No Solution
22
Algorithm 5 Algorithm findMinOps used in Algorithm 4.

### v2

@@
 While there are typically standard distances to adopt in a continuous space, this is not the case for combinatorial parameters such as graphs that represent statistical models.
-Existing proposals like the structural Hamming distance are defined on the graphs rather than the models they represent and can hence lead to undesirable behaviors.
-We propose a model-oriented framework for defining the distance between graphs that is applicable across many different graph classes.
+Defined on the graphs alone, existing proposals like the structural Hamming distance ignore the structure of the model space and can thus exhibit undesirable behaviors.
+We propose a model-oriented framework for defining the distance between graphs that is applicable across different graph classes.
 Our approach treats each graph as a statistical model and organizes the graphs in a partially ordered set based on model inclusion.
 This induces a neighborhood structure, from which we define the model-oriented distance as the length of a shortest path through neighbors, yielding a metric in the space of graphs.
-We apply this framework to both probabilistic graphical models (e.g., undirected graphs and completed partially directed acyclic graphs) and causal graphical models (e.g., directed acyclic graphs and maximally oriented partially directed acyclic graphs).
-We analyze the theoretical and empirical behaviors of model-oriented distances.
-Algorithmic tools are also developed for computing and bounding these distances.
+We apply this framework to probabilistic undirected graphs, causal directed acyclic graphs, probabilistic completed partially directed acyclic graphs, and causal maximally oriented partially directed acyclic graphs.
+We analyze theoretical and empirical behaviors of the model-oriented distance and draw comparison with existing distances.
+By exploiting the underlying poset structures, we develop algorithms for computing and bounding the proposed distance that scale to moderate-sized graphs.
 structural Hamming distance,
@@
 These graphs are used to determine if a causal effect can be identified from observational data (Huang and Valtorta, 2006; Shpitser and Pearl, 2006), and if so, to also inform the best way to estimate the effect (Rotnitzky and Smucler, 2020; Henckel, Perković and Maathuis, 2022; Guo, Perković and Rotnitzky, 2023).
-There is also a large literature on learning probabilistic or causal graphs from data, a problem known as structure learning or causal discovery; the reader is referred to Drton and Maathuis (2017); Glymour, Zhang and Spirtes (2019) for two recent surveys.
+There is also a large literature on learning probabilistic or causal graphs from data, a problem known as structure learning or causal discovery; the reader is referred to Drton and Maathuis (2017); Glymour, Zhang and Spirtes (2019) for two surveys.
 In whichever context these graphs are used, a key unresolved issue is how to define a suitable distance between them.
-Having a distance is essential for several tasks, including benchmarking causal discovery algorithms (Scutari, Graafland and Gutiérrez, 2019; Rios, Moffa and Kuipers, 2023), characterizing their statistical performance (Jin, Zhang and Zhang, 2012; Ke, Jin and Fan, 2012; Han et al., 2016) or finding a consensus graph among a collection of graphs through a Frechét mean (Ferguson and Meyer, 2023; Wang, Kolar and Drton, 2025).
-Other potential applications include constructing ball-like confidence regions for a graph and and conducting sensitivity analysis with respect to potential mistakes in the causal graph.
+Having a distance is essential for several tasks, including benchmarking causal discovery algorithms (Scutari, Graafland and Gutiérrez, 2019; Rios, Moffa and Kuipers, 2025), characterizing their statistical performance (Jin, Zhang and Zhang, 2012; Ke, Jin and Fan, 2012; Han et al., 2016) or finding a consensus graph among a collection of graphs through a Frechét mean (Ferguson and Meyer, 2023; Wang, Kolar and Drton, 2025).
+Other potential applications include constructing ball-like confidence regions for a graph and conducting sensitivity analyses with respect to potential mistakes in the causal graph.
 However, defining a suitable distance is challenging for two reasons.
@@
 It counts the number of edges that differ between two graphs with various proposals for how to weigh different types of edge mismatches (Perrier, Imoto and Miyano, 2008).
-Irrespective of the concrete weighing the SHD is computationally inexpensive, flexible with respect to the graph class, and guaranteed to define a metric.
+Irrespective of the concrete weighing the SHD is computationally cheap, flexible with respect to the graph class, and guaranteed to define a metric.
 It is, however, model-agnostic — it disregards the specific models the graphs represent and is solely defined on the graphs themselves.
@@
 Second, the SHD does not depend on the set of graphs in consideration — it is agnostic to constraints posed by the graph type and domain specific considerations (e.g., one may wish to consider acyclic graphs).
-Finally, by construction the SHD treats each edge individually.
-This ignores that not all edge configurations may give a valid graph in the graph class under consideration.
+Instead the SHD treats each edge individually, which ignores that not all edge configurations may give a valid graph in the graph class under consideration.
 We illustrate these issues with an example.
 Example 1.
-Consider the four CPDAGs in Fig.˜1(a-d).
-For two of the more prominent types of SHD (see Section˜4.3), we have $\mathrm{SHD}(\mathcal{G}_{1},\mathcal{G}_{2})=n+1$, respectively $n+2$.
-Similarly, $\mathrm{SHD}(\mathcal{G}_{1},\mathcal{G}_{3})=2\text{ or }4$, and for even $n$, $\mathrm{SHD}(\mathcal{G}_{1},\mathcal{G}_{4})=n/2+1\text{ or }n+2$.
-However, in the context of probabilistic models, $\mathcal{G}_{2}$ sits “just between” $\mathcal{G}_{1}$ and $\mathcal{G}_{3}$: $\mathcal{G}_{2}$ encodes fewer conditional dependencies than $\mathcal{G}_{1}$ and no graph encodes an intermediary collection of dependencies.
-The same is true for $\mathcal{G}_{3}$ with respect to $\mathcal{G}_{2}$.
-Logically, $\mathcal{G}_{3}$ should be closer to $\mathcal{G}_{2}$ than $\mathcal{G}_{1}$.
-The SHD overestimates the semantic distance between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, as it wrongly assumes that undirecting any edge in $\mathcal{G}_{2}$ leads to a valid CPDAG that is “in between” $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$.
-Similarly, since $\mathcal{G}_{4}$ removes far more dependencies from $\mathcal{G}_{1}$ than $\mathcal{G}_{2}$ does, we expect $\mathcal{G}_{1}$ to be closer to $\mathcal{G}_{2}$ than to $\mathcal{G}_{4}$.
-Our proposed model-oriented distance reflects the semantics of the graphs, and yields $d(\mathcal{G}_{1},\mathcal{G}_{2})=1$, $d(\mathcal{G}_{1},\mathcal{G}_{3})=2$, and $d(\mathcal{G}_{1},\mathcal{G}_{4})=n/2+1$.
-As a further illustration, we set $n=10$, and generate $100$ datasets of $1000$ samples each from a linear Gaussian structural equation model from graph $\mathcal{G}_{1}$, with unit edge weight and noise variance.
-We can fit a given graph $\mathcal{G}$ to the data and compute the Bayesian Information Criterion (BIC), which balances model fit and complexity.
-As shown in Fig.˜1(e-f), our proposed model-oriented distance better reflects fit to data: further away graphs tend to yield worse BIC scores.
-This indicates that using SHD to benchmark score-based causal discovery algorithms that rely on BIC, such as the well-known greedy equivalence search algorithm (Chickering, 2002), may be misleading.
-To address the limitation of model-agnostic distances, Peters and Bühlmann (2015) proposed an alternative, task-oriented approach, based on the idea that the distance between two graphs should be measured based on the specific task they serve.
-In the context of causal inference, the central task is to identify a causal effect of interest, i.e., to express the effect as a functional of the observed distribution that holds under every causal system that is compatible with a graph.
-Peters and Bühlmann (2015) introduced the structural intervention distance (SID) between two DAGs, which counts how many identifying formulae under one graph would be invalid if the truth lies within the other graph.
-This approach was also adapted and generalized by Henckel, Würtzen and Weichwald (2024) who establish that the SID is merely one of a class of closely related distances they call adjustment identification distances.
-The SID is, however, not a metric: it can be zero between two different graphs, is not symmetric and does not obey the triangle inequality.
-It is possible to define a symmetrized version of SID but we show in Section˜H.1 that this version still violates the triangle inequality.
-In general task-oriented distance are not constructed to be metrics, limiting their usefulness.
+Consider the four CPDAGs listed in Fig.˜1.
+Each graph represents a set of distributions over $n+1$ random variables that obey the probabilistic constraints it encodes, where a missing edge between two variables signifies the conditional independence between them given another set of variables.
+Based on the edges (ignoring the difference between directed and undirected edges for now), we can see $\mathcal{G}_{1}\succ\mathcal{G}_{2}\succ\mathcal{G}_{3}\succ\mathcal{G}_{4}$ in terms of the models they represent.
+Table˜1 shows the distance of graphs $\mathcal{G}_{2},\mathcal{G}_{3},\mathcal{G}_{4}$ relative to $\mathcal{G}_{1}$: the first two columns list the two versions of the SHD (see Definition˜7) and the third column is the distance we propose in this paper.
+Our distance reflects what we expect logically: relative to $\mathcal{G}_{1}$, the graph $\mathcal{G}_{2}$ is the closest, followed by $\mathcal{G}_{3}$, and lastly $\mathcal{G}_{4}$.
+In contrast, the SHD incorrectly deems $\mathcal{G}_{3}$ to be the closest (can be much closer than $\mathcal{G}_{2}$ and $\mathcal{G}_{4}$) and fails to capture the fact that $\mathcal{G}_{2}$ is strictly closer than $\mathcal{G}_{4}$.
+This is due to the fact that the SHD wrongly assumes that undirecting any edge in $\mathcal{G}_{2}$ leads to a valid CPDAG that lies in between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, thereby overestimating the distance between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$.
+Moreover, our distance agrees with the Bayesian Information Criterion (BIC) listed in the last column: we generate data under $\mathcal{G}_{1}$, specifically from a Gaussian structural equation model over $11$ variables (with unit edge weights and noise variances), and draw 100 datasets each with sample size 1000; we then fit models and estimate $\operatorname{\mathbb{E}}[\mathrm{BIC}(\hat{\mathcal{G}}_{i})-\mathrm{BIC}(\hat{\mathcal{G}}_{1})]$ (a lower score means a better model).
+Meanwhile, the discordance between BIC and SHD indicates that using SHD to benchmark score-based causal discovery algorithms (see Drton and Maathuis, 2017, §4) that rely on BIC or the like, such as the well-known greedy equivalence search algorithm (Chickering, 2002), can be misleading.
+Table 1: Distances and BIC ($\pm$ standard error) relative to $\mathcal{G}_{1}$ for graphs in Fig.˜1 (assuming $n$ is even)
+Graph $\operatorname{\mathrm{SHD}}_{1}$ $\operatorname{\mathrm{SHD}}_{2}$ Model-oriented distance BIC
+$\mathcal{G}_{2}$ $n+1$ $n+2$ 11 $690\pm 5$
+$\mathcal{G}_{3}$ 22 44 22 $1084\pm 6$
+$\mathcal{G}_{4}$ $n/2+1$ $n+2$ $n/2+1$ $10159\pm 20$
+To address the limitation of model-agnostic distances, Peters and Bühlmann (2015) proposed an alternative, task-oriented approach, where the distance between two graphs is measured according to a concrete task they serve.
+For causal graphs, the central task is effect identification, i.e., expressing a causal effect as a functional of the observed distribution valid under any system compatible with the graph.
+Motivated by this, Peters and Bühlmann (2015) introduced the Structural Intervention Distance (SID), which counts the number of causal effects identified in one graph that would be invalid if the true data generating process corresponded to the other.
+While Henckel, Würtzen and Weichwald (2024) generalized this approach into a broader class of adjustment identification distances, the SID is not a metric: it allows zero distance between distinct graphs, lacks symmetry, and violates the triangle inequality.
+We show in Section˜H.1 that even a symmetrized version of SID fails the triangle inequality, limiting the utility of such task-oriented measures.
 (a) $\mathcal{G}_{1}$
@@
 (d) $\mathcal{G}_{4}$
-(e) SHD distance vs BIC score
-(f) Proposed model-oriented distance vs BIC score
-Figure 1: (a–f) Four CPDAGs in Example˜1 and their associated BIC scores for synthetically generated data.
-1.1 Our contributions
+Figure 1: CPDAGs $\mathcal{G}_{1}\succ\mathcal{G}_{2}\succ\mathcal{G}_{3}\succ\mathcal{G}_{4}$ in terms of the models they represent.
+1.1 Main contributions and organization of the paper
 We introduce a model-oriented framework that defines distance between graphs through the models they represent.
 Given a class of graphs of the same type (e.g., the collection of CPDAGs over five vertices), our approach begins by interpreting a probabilistic or causal graph as a statistical model, namely a set of distributions over a common collection of random variables — factual or counterfactual — that obey the constraints encoded by the graph.
-This viewpoint naturally leads to the organization of graphs into a partially ordered set (poset), where the ordering is defined by model inclusion: for two graphs $\mathcal{G}$ and $\mathcal{G}^{\prime}$, we say $\mathcal{G}\preceq\mathcal{G}^{\prime}$ if and only if the model represented by $\mathcal{G}$ is contained by the model represented by $\mathcal{G}^{\prime}$.
-Posets are equipped with a natural neighborhood notion called covering elements.
-Exploiting this neighborhood notion, we define our distance as the length of a shortest path between two graphs when moving only through neighbors.
-By construction, our distance is a metric and applicable to many classes of graphs of varying type.
-We study in detail the model-oriented distance for four graph types, listed in order of increasing complexity of their associated posets: probabilistic UGs, causal DAGs, probabilistic CPDAGs, and causal MPDAGs.
-For each, we study algorithms for computing or bounding the model-oriented distance.
-Here, a key ingredient is a graphical characterization of model containment, and whenever possible, of neighborhood.
-When the latter proves difficult, we study the neighborhood for a restricted class of graphs where the characterization is tractable.
-Further, we show that poset-theoretic properties, such as gradedness and semimodularity, can facilitate computing the distances.
-We also show that, for CPDAGs and MPDAGs, the model-oriented distance can differ substantially from the structural Hamming distance.
-For CPDAGs and MPDAGs restricted to polytrees, we also prove that the model-oriented distance admits several simplifications that make it easier to compute and interpret.
-By establishing that many statistical graph classes are naturally equipped with partial orderings that can be studied using theoretical tools from order theory, we open a new avenue for investigating graphical models and their properties.
-Partial orders, for example, have been shown to be natural and useful structures for model selection over combinatorial objects, including graphs (Taeb, Bühlmann and Chandrasekaran, 2023).
-Similarly, neighborhood notions are central to score-based causal discovery methods where they used to decide how to explore the space of all graphs in a class while searching for a score-minimizing graph (Chickering, 2002).
+The notation and background for these models are described in Section˜2.
+We demonstrate in Section˜3 that this viewpoint naturally organizes graphs into a partially ordered set (poset) defined by model inclusion: for graphs $\mathcal{G}$ and $\mathcal{G}^{\prime}$, we say $\mathcal{G}\preceq\mathcal{G}^{\prime}$ if the model represented by $\mathcal{G}$ is contained by the model represented by $\mathcal{G}^{\prime}$.
+We leverage the poset’s intrinsic notion of neighborhood, called covering relations, to define the model-oriented distance as the length of a shortest path between two graphs when moving only through neighbors.
+By construction, this distance is a metric applicable to any class of graphs satisfying some mild assumptions, regardless of its edge types, constraints or semantics.
+We note that the notion of neighborhood is also central to score-based causal discovery, where an algorithm traverses neighborhoods to search for a score-minimizing model; see, e.g., Chickering (2002); Hauser and Bühlmann (2012); Ng et al.
+(2024).
+However, to the best of our knowledge, this is the first work to leverage the structural properties of the underlying model space to formally define and analyze a distance metric between graphs.
+To facilitate distance computation, in Section˜4 we analyze the structural properties of posets underlying four graph types, ordered by increasing complexity: probabilistic UGs, causal DAGs, probabilistic CPDAGs and causal MPDAGs.
+A key ingredient of this analysis is the graphical characterization of model containment and, where feasible, of neighborhood.
+In cases where characterizing the general neighborhood proves intractable, we restrict our study to the subclass of polytrees.
+We show that for CPDAGs and MPDAGs, this model-oriented distance can be substantially different from the structural Hamming distance.
+In Section˜5, we develop a generic A* algorithm to compute our distance.
+The algorithm is then specialized to CPDAGs and polytree MPDAGs, for which some numerical results are also presented.
+We conclude the paper with a brief discussion in Section˜6.
 2 Background
+2.1 Notation
+We generally use Roman capital letters (e.g., $\mathrm{V},\mathrm{B}$) to denote a set and lowercase letters (e.g., $v_{1},b_{2}$) to denote its elements.
+For two sets $\mathrm{A}$ and $\mathrm{B}$, we use $\mathrm{A}\triangle\mathrm{B}:=(\mathrm{A}\setminus\mathrm{B})\cup(\mathrm{B}\setminus\mathrm{A})$ to denote their symmetric difference.
+We use calligraphic letters (e.g., $\mathcal{G},\mathcal{D}$) for graphs and italic capital letters (e.g., $X,Y$) for random variables.
 2.2 Graphical preliminaries
-We consider graphs with nodes representing random variables and edges between them representing probabilistic or causal dependencies between these variables.
-Basic graphical concepts are briefly described in below.
-2.1.1 Node, edge, subgraph, cycle, PDAG
+We consider graphs with vertices representing random variables and edges between them representing probabilistic or causal dependencies between these variables.
+Basic graphical concepts are briefly described below.
+2.2.1 Vertex, edge, subgraph, cycle, PDAG
+We use letter $\mathrm{V}$ to denote a set of vertices, which we assume to be finite throughout; this implies that class over $\mathrm{V}$ is also finite.
+For convenience, $\mathrm{V}$ can be identified with $\{1,\dots,|\mathrm{V}|\}$.
 We write a graph over vertex set $\mathrm{V}$ with $k\geq 1$ types of edges (e.g., undirected ‘-’ and directed $\rightarrow$’) as $\mathcal{G}=(\mathrm{V},\mathrm{E}^{1},\dots,\mathrm{E}^{k})$.
@@
 A graph $\mathcal{G}_{1}=(\mathrm{V},\mathrm{E}_{1}^{1},\dots,\mathrm{E}_{1}^{k})$ is said to be a subgraph of $\mathcal{G}_{2}=(\mathrm{V},\mathrm{E}_{2}^{1},\dots,\mathrm{E}_{2}^{k})$ if $\mathrm{E}_{1}^{i}\subseteq\mathrm{E}_{2}^{i}$ for each edge type $i$.
-Given a graph $\mathcal{G}=(\mathrm{V},\mathrm{E}^{1},\dots,\mathrm{E}^{k})$ and a subset of vertices $\mathrm{V}^{\prime}\subseteq\mathrm{V}$, the induced graph of $\mathrm{V}^{\prime}$ is $\mathcal{G}_{\mathrm{V}^{\prime}}:=(\mathrm{V}^{\prime},{\mathrm{E}^{1}}^{\prime},\dots,{\mathrm{E}^{k}}^{\prime})$ where ${\mathrm{E}^{i}}^{\prime}:=\mathrm{E}^{i}\cap(\mathrm{V}^{\prime}\times\mathrm{V}^{\prime})$ for $i=1,\dots,k$.
+Given a graph $\mathcal{G}=(\mathrm{V},\mathrm{E}^{1},\dots,\mathrm{E}^{k})$ and a subset of vertices $\mathrm{V}^{\prime}\subseteq\mathrm{V}$, the induced subgraph of $\mathrm{V}^{\prime}$ is $\mathcal{G}_{\mathrm{V}^{\prime}}:=(\mathrm{V}^{\prime},{\mathrm{E}^{1}}^{\prime},\dots,{\mathrm{E}^{k}}^{\prime})$ where ${\mathrm{E}^{i}}^{\prime}:=\mathrm{E}^{i}\cap(\mathrm{V}^{\prime}\times\mathrm{V}^{\prime})$ for $i=1,\dots,k$.
 A graph is simple if between every two vertices there is at most one edge (of any type).
@@
 A directed cycle is a cycle of the form $v_{1}\rightarrow v_{2}\rightarrow\dots\rightarrow v_{k}$.
-In contrast, a partially directed cycle is a cycle where the edges can be directed or undirected such that (i) at least one edge is directed and (ii) all the directed edges in the cycle point to the same direction.
-A graph is chordal if every cycle of length 4 or more has a chord — an edge connecting two non-consecutive vertices.
+In contrast, a partially directed cycle is a cycle where the edges can be directed or undirected such that (i) at least one edge is directed and (ii) all the directed edges in the cycle point in the same direction.
+A graph is chordal if every cycle of length 4 or more has a chord, i.e., an edge connecting two non-consecutive vertices.
 A partially directed acyclic graph (PDAG) is a simple graph with two types of edges, directed and undirected, such that there is no directed cycle.
@@
 An undirected graph is a graph consisting of undirected edges.
-Suppose a random vector $X=(X_{v}:v\in\mathrm{V})$ is indexed by vertices of the undirected graph $\mathcal{G}=(\mathrm{V},\mathrm{E})$.
+Suppose a random vector $X=(X_{v}:v\in\mathrm{V})$ corresponds to the vertices of the undirected graph $\mathcal{G}=(\mathrm{V},\mathrm{E})$.
 An undirected graph $\mathcal{G}$ defines a probabilistic model for the distribution of $X$ through a set of semantics known as the Markov properties of UGs.
@@
 The exact choice is not essential for our approach.
-Yet, for simplicity, in Section˜2.2 we will assume regularity conditions so that these four notions are equivalent.
+Yet, for simplicity, in Section˜2.3 we will assume regularity conditions so that these four notions are equivalent.
 2.2.3 Directed acyclic graphs (DAGs)
@@
 For the purpose of this paper, we only consider DAGs as causal models since distinct DAGs (over the same set of vertices) may encode the same probabilistic model — these DAGs are called Markov equivalent — and we require the map from graph to model to be injective.
-For probabilistic modeling, we instead use a special type of graphs to represent the Markov equivalence classes of DAGs, which is described next.
+For probabilistic modeling, we instead use a special type of graph to represent the Markov equivalence classes of DAGs, which is described next.
 2.2.4 Completed partially directed acyclic graphs (CPDAGs)
 A CPDAG, also known as an essential graph, is a partially directed acyclic graph (PDAG) that represents a set of DAGs that are Markov equivalent.
-By a classical result (Verma and Pearl, 1990; Andersson, Madigan and Perlman, 1997), two DAGs are Markov equivalent if they share (1) the same skeletons (e.g., edges ignoring the directions) and (2) the same set of v-structures (i.e., unshielded colliders).
+By a classical result (Verma and Pearl, 1990; Andersson, Madigan and Perlman, 1997), two DAGs are Markov equivalent if they share (1) the same skeletons (i.e., edges ignoring the directions) and (2) the same set of v-structures (i.e., unshielded colliders).
 We have a v-structure at a vertex $v_{2}$ if there are two edges, e.g., $v_{1}\rightarrow v_{2}$ and $v_{2}\leftarrow v_{3}$, pointing to $v_{2}$, but there is no edge between $v_{1}$ and $v_{3}$.
@@
 In the causal context, suppose we also have the background knowledge that $v_{2}$ temporally precedes $v_{3}$.
-Then, we can narrow down to $v_{1}\rightarrow v_{2}\rightarrow v_{3}$ or $v_{1}\leftarrow v_{2}\rightarrow v_{3}$, which together can be represented by a PDAG $v_{1}-v_{2}\rightarrow v_{3}$.
+Then, the set of DAGS can be narrowed down to $v_{1}\rightarrow v_{2}\rightarrow v_{3}$ and $v_{1}\leftarrow v_{2}\rightarrow v_{3}$, which together are represented by a PDAG $v_{1}-v_{2}\rightarrow v_{3}$.
 PDAGs like this are MPDAGs, which represent the conjunction of several distinct causal models that are observationally Markov equivalent.
@@
 Therefore, in the rest of this paper, we will only focus on the probabilistic models represented by CPDAGs and the causal models represented by MPDAGs.
-Not every PDAG is valid MPDAG; see Section˜B.2 for structural constraints that an MPDAG must obey.
+Not every PDAG is a valid MPDAG; see Section˜B.2 for structural constraints that an MPDAG must obey.
 Figure 2: (a) A CPDAG representing three Markov equivalent DAGs. A subset of two DAGs can be represented by an MPDAG. (b) Structures that are forbidden from a valid MPDAG.
 2.3 Graphs as statistical models
-For the four types of graphs we introduced, we describe the models they represent.
-To ensure precision in deriving the corresponding model-oriented distances we do so rigorously, but the reader may set aside the technical details on first reading.
+For the four graph types we introduced, we describe the models they represent.
+To derive the corresponding model-oriented distances precisely, we do so rigorously, though the reader may set aside these technical details on first reading.
 2.3.1 Graphs as probabilistic models
 UGs and CPDAGs represent probabilistic models for a random vector $(X_{v}:v\in\mathrm{V})$.
-Suppose $X$ takes value in a product measurable space $\mathcal{X}=\prod_{v\in\mathrm{V}}\mathcal{X}_{v}$.
-Let $\mu=\prod_{v\in V}\mu_{v}$ be a product measure over $\mathcal{X}$.
+Suppose $X$ takes values in a product measurable space $\mathcal{X}=\prod_{v\in\mathrm{V}}\mathcal{X}_{v}$.
+Let $\mu=\prod_{v\in\mathrm{V}}\mu_{v}$ be a product measure over $\mathcal{X}$.
 Let $P$ be the law of $X$.
@@
 For $P\in\mathcal{M}^{\mathrm{o}}$, we use $p:=\mathop{}\!\mathrm{d}P/\mathop{}\!\mathrm{d}\mu$ to denote its density.
-For UGs, as explained in Section˜2.1.2, we pose additional regularity condition on the ground set such that four Markov properties all lead to the same model definition. Let the ground set be
+For UGs, as explained in Section˜2.2.2, we pose an additional regularity condition on the ground set such that all four Markov properties lead to the same model definition.
+Concretely, for UGs let the ground set be
 $\mathcal{M}^{\mathrm{o}}_{+}:=\left\{P\in\mathcal{M}^{\mathrm{o}}:p>0\text{ $\mu$-a.e.}\right\}.$
@@
 […]
-For a CPDAG $\mathcal{G}$, it represents the probabilistic model consisting of densities that factorize according any (and hence every) DAG $\mathcal{D}$ represented by $\mathcal{G}$ (i.e., $\mathcal{D}$ is in the Markov equivalence class represented by $\mathcal{G}$), given by
+A CPDAG $\mathcal{G}$ represents the probabilistic model consisting of densities that factorize according to any (and hence every) DAG $\mathcal{D}$ represented by $\mathcal{G}$ (i.e., $\mathcal{D}$ is in the Markov equivalence class represented by $\mathcal{G}$), given by
 $\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}}(\mathcal{G}):=\left\{P\in\mathcal{M}^{\mathrm{o}}:\,p(x)=\prod_{v\in{V}}p_{v}(x_{v}|x_{{\text{pa}_{\mathcal{D}}(v)}})\text{ for some DAG $\mathcal{D}$ represented by $\mathcal{G}$}\right\},$
@@
 We follow Zhao (2025) in defining causal models associated with DAGs and MPDAGs.
-Consider a causal system that generates observations of a random vector $X=(X_{v}:v\in\mathrm{V})$, which takes value in a product measurable space $\mathcal{X}\equiv\mathcal{X}_{\mathrm{V}}$.
-Here we use notation $\mathcal{X}_{I}:=\prod_{v\in\mathrm{I}}\mathcal{X}_{v}$ for every $\mathrm{I}\subseteq\mathrm{V}$.
+Consider a causal system that generates observations of a random vector $X=(X_{v}:v\in\mathrm{V})$, which takes values in a product space $\mathcal{X}\equiv\mathcal{X}_{\mathrm{V}}$ equipped a suitable product $\sigma$-algebra.
+Here we use notation $\mathcal{X}_{\mathrm{I}}:=\prod_{v\in\mathrm{I}}\mathcal{X}_{v}$ for every $\mathrm{I}\subseteq\mathrm{V}$.
 A causal model is a set of joint distributions of counterfactual variables as well as factual (i.e., observed) variables, both of which are part of the potential outcome schedule:
-$X(\cdot):=\left(X_{v}(x_{I}):v\in V,I\subseteq V,x_{I}\in\mathcal{X}_{I}\right),$
-where $X_{v}(x_{I})$ is the potential outcome of variable $v$ under the intervention that imposes value $x_{I}$ on $X_{I}$.
-In the special case when $I=\emptyset$, the margin $(X_{v}(\emptyset):v\in V)$ consists of all the factual variables, whose law is called the observed distribution.
-The potential outcome schedule is a stochastic process that takes value in $\mathcal{X}^{\prod_{I\subseteq V}\mathcal{X}_{I}}$, which can be equipped with a suitable $\sigma$-algebra to form a measurable space.
+$X(\cdot):=\left(X_{v}(x_{\mathrm{I}}):v\in\mathrm{V},\mathrm{I}\subseteq\mathrm{V},x_{\mathrm{I}}\in\mathcal{X}_{\mathrm{I}}\right),$
+where $X_{v}(x_{\mathrm{I}})$ is the potential outcome of variable $v$ under the intervention that imposes value $x_{\mathrm{I}}$ on $X_{\mathrm{I}}$.
+In the special case when $\mathrm{I}=\emptyset$, the margin $(X_{v}(\emptyset):v\in\mathrm{V})$ consists of all the factual variables, whose law is called the observed distribution.
+The potential outcome schedule is a stochastic process that takes values in $\mathcal{X}^{\prod_{\mathrm{I}\subseteq\mathrm{V}}\mathcal{X}_{\mathrm{I}}}$, which can be equipped with a suitable $\sigma$-algebra to form a measurable space.
 We use $P$ to denote the law of $X(\cdot)$ (i.e., a probability measure over the measurable space).
 Similar to the definition of $\mathcal{M}^{\mathrm{o}}$, we consider the ground causal model to be
-$\mathcal{M}^{\mathrm{c}}:=\big\{P:P\left(X(x_{I},x_{I^{\prime}})=X(x_{I})\mid X_{I^{\prime}}(x_{I})=x_{I^{\prime}}\right)=1,\\ \text{ for every disjoint $I,I^{\prime}\subseteq V$ and every $x\in\mathcal{X}$}\big\},$ (3)
+$\mathcal{M}^{\mathrm{c}}:=\big\{P:P\left(X(x_{\mathrm{I}},x_{\mathrm{I}^{\prime}})=X(x_{\mathrm{I}})\mid X_{\mathrm{I}^{\prime}}(x_{\mathrm{I}})=x_{\mathrm{I}^{\prime}}\right)=1,\\ \text{ for every disjoint $\mathrm{I},\mathrm{I}^{\prime}\subseteq\mathrm{V}$ and every $x\in\mathcal{X}$}\big\},$ (3)
 where $\mathrm{c}$’ stands for causal.
-In words, $\mathcal{M}^{\mathrm{c}}$ is the set of distributions of $X(\cdot)$ that maintain consistency through recursive substitution (see, e.g., Shpitser, Richardson and Robins, 2022, §41.2): in a world that sets $X_{I}$ to be $x_{I}$, if we observe $X_{I^{\prime}}$ to be $x_{I}^{\prime}$, then the observed $X$ equals $X(x_{I},x_{I^{\prime}})$.
+In words, $\mathcal{M}^{\mathrm{c}}$ is the set of distributions of $X(\cdot)$ that maintain consistency through recursive substitution (see, e.g., Shpitser, Richardson and Robins, 2022, §41.2): in a world that sets $X_{\mathrm{I}}$ to be $x_{\mathrm{I}}$, if we observe $X_{\mathrm{I}^{\prime}}$ to be $x_{\mathrm{I}^{\prime}}$, then the observed $X$ equals $X(x_{\mathrm{I}},x_{\mathrm{I}^{\prime}})$.
 Any causal model that we discuss next will be a subset of $\mathcal{M}^{\mathrm{c}}$.
@@
 A DAG $\mathcal{D}$ is represented by $\mathcal{G}$ if $\mathcal{D}$ can be obtained by orienting the undirected edges in $\mathcal{G}$ without introducing any directed cycles or v-structures.
-The causal model represented by MPDAG $\mathcal{G}$ is simply defined in terms of a disjunction:
+The causal model represented by the MPDAG $\mathcal{G}$ is simply defined in terms of a disjunction:
 $\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}(\mathcal{G})=\bigcup\big\{\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}(\mathcal{D}):\text{$\mathcal{D}$ is a DAG represented by $\mathcal{G}$}\big\}.$
 3 A general recipe for model-oriented distance using partially ordered sets
-Let $\mathfrak{G}$ be a finite collection of graphs over a common set of vertices $\mathrm{V}$.
+Let $\mathfrak{G}$ be a collection of graphs over a finite vertex set $\mathrm{V}$.
 Let $\mathcal{M}$ be an associated model map (i.e., semantics) that assigns a model to each graph in the class.
-Given $(\mathfrak{G},\mathcal{M})$, our goal is to define an $\mathcal{M}$-oriented distance $d(\mathcal{G}_{s},\mathcal{G}_{t})$ for every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$.
-For example, for $\mathrm{V}=\{1,2,3,4,5\}$, we can consider defining distances for $(\mathfrak{G}_{1},\mathcal{M}^{\mathrm{o}}_{\textsc{ug}})$ and $(\mathfrak{G}_{2},\mathcal{M}^{\mathrm{o}}_{\textsc{ug}})$ with
-[…]
+Given $(\mathfrak{G},\mathcal{M})$, our goal is to define a model-oriented distance $d(\mathcal{G}_{s},\mathcal{G}_{t})$ for every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$.
+By definition, $d$ depends on $\mathfrak{G}$ and $\mathcal{M}$.
+For example, for $\mathrm{V}=\{1,2,3,4,5\}$, we can define distances $d_{1},d_{2}$ respectively for $(\mathfrak{G}_{1},\mathcal{M}^{\mathrm{o}}_{\textsc{ug}})$, $(\mathfrak{G}_{2},\mathcal{M}^{\mathrm{o}}_{\textsc{ug}})$ with
+$\mathfrak{G}_{1}:=\{\text{all UGs over $\mathrm{V}$}\},\quad\mathfrak{G}_{2}:=\{\text{all acyclic UGs over $\mathrm{V}$}\}.$
 In Section 3.1, we use the model map $\mathcal{M}$ to define our model-oriented distance between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$.
@@
 By construction, neighborhood is a symmetric relation.
-Intuitively, $\mathcal{G}_{2}$ is a neighbor of $\mathcal{G}_{1}$ if it represents the simplest refinement or coarsening of $\mathcal{G}_{1}$’s model, with no graph lying strictly between them in the model hierarchy.
+Intuitively, $\mathcal{G}_{2}$ is a neighbor of $\mathcal{G}_{1}$ if it represents a finest refinement or coarsening of $\mathcal{G}_{1}$’s model, in the sense that there is no graph lying strictly between them in the model hierarchy.
 The model-oriented distance between two graphs counts the minimal number of such basic refinement or coarsening operations to transform one model into the other, which we define as follows.
@@
 The significance of the symmetry and anti-symmetry properties is straightforward.
-We illustrate the importance of triangle inequality from both statistical and algorithmic perspectives.
-From a statistical perspective, when proving consistency or convergence of an estimator, one often introduces an intermediate “ideal" object (e.g., the population version of the estimator) to bridge the gap between the data and the truth.
-The triangle inequality ensures that the distance between an estimated graph and the true graph can be controlled via the intermediate graph.
-From an algorithmic perspective, the triangle inequality helps in enumerating graphs within a small distance of a reference graph.
-For example, when the distance takes integer values, one can use the triangle inequality to enumerate all graphs within distance two of a reference graph by first finding all graphs at distance one, and then enumerating all graphs that are distance one from those.
-Without the triangle inequality, this decomposition is no longer possible, making enumeration much more expensive.
+We underscore the importance of the triangle inequality from both statistical and algorithmic perspectives.
+Statistically, the triangle inequality is essential for proving properties like the consistency of an estimator .
+It allows one to bound the total error between an estimate and the truth by introducing an intermediate ’population-level’ proxy, ensuring that convergence to the proxy implies convergence to the truth.
+Algorithmically, the triangle inequality facilitates enumerating graphs.
+For instance, to enumerate all graphs within distance $k$ of a given graph $\mathcal{G}$, it suffices to expand the neighborhood of all graphs at distance k1k-1 to $\mathcal{G}$.
+Without this property, the enumeration can be computationally prohibitive.
 For the model-oriented distance to be a metric, we require the following two conditions.
@@
 Condition 2 (Connectedness).
-For every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$, there exists a finite sequence of graphs $\mathcal{G}_{s}=\mathcal{G}_{1},\dots,\mathcal{G}_{k}=\mathcal{G}_{t}$ with $k\geq 1$ such that every $\mathcal{G}_{i},\mathcal{G}_{i+1}$ are neighbors.
+For every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$, there exists a sequence of graphs $\mathcal{G}_{s}=\mathcal{G}_{1},\dots,\mathcal{G}_{k}=\mathcal{G}_{t}$ with $k\geq 1$ such that every $\mathcal{G}_{i},\mathcal{G}_{i+1}$ are neighbors.
 Theorem 1 (Metric).
@@
 Proof.
-Let the model-oriented distance be denoted by $d$.
+Let $d$ denote the model-oriented distance.
 Under Condition˜1, function $d$ can be viewed as the shortest-path distance, which is a metric, defined on an undirected graph with vertex set $\mathfrak{G}$ and edges incident to neighbors.
@@
 In fact, in Section˜A.1, we show that all the four models $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}},\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}},\mathcal{M}^{\mathrm{c}}_{\textsc{dag}},\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}$ discussed in this paper satisfy Condition˜1.
-Typically, Condition˜2 can be ensured by the checking either of the two following stronger condition.
+Typically, Condition˜2 can be ensured by the checking either of the two following stronger conditions.
 Condition 3 (Least element).
@@
 When it exists, $\mathcal{G}_{\hat{0}}$ is the unique least element in $\mathfrak{G}$ and is usually taken to be the empty graph, which represents the smallest model.
-This condition has the following dual.
+Again, Condition 3 is satisfied by all four graph classes with $\mathcal{G}_{\hat{0}}$ being the empty graph.
+This condition also has the following dual.
 Condition 3 (Greatest element).
@@
 In both cases, the distance reflects how many basic, interpretable operations are required to transform one model into another.
-Here, we use the notion of neighbor to provide a parsimonious and model-oriented definition of what the basic operations are.
+Here, we use the notion of neighbor to provide a model-oriented definition of what the basic operations are.
 In what follows, we use posets to formalize our approach.
@@
 However, since the definition is implicit, it is not immediately clear how to compute the distance for any specific class.
-We now provide a constructive characterization of the model-oriented distance which we both use to study under what conditions it simplifies (see Sections 3.3 and 4) and to develop algorithmic tools for computing it (see Section 5).
+We use partially ordered sets to construct the model-oriented distance, which leads to structural properties that facilitate distance computation (see Sections˜3.3 and 4) and efficient algorithms (see Section˜5).
 Definition 3 (Poset, cover, connectedness, least element, comparability).
-A poset (partially ordered set) $\mathcal{L}:=(L,\preceq)$ is a collection $L$ of elements and a relation $\preceq$ that is reflexive ($x\preceq x,~\forall x\in L$), transitive ($x\preceq y,y\preceq z\Rightarrow x\preceq z,~\forall x,y,z\in L$), and anti-symmetric ($x\preceq y,y\preceq x\Rightarrow x=y,~\forall x,y\in L$).
+A poset (partially ordered set) $\mathcal{L}:=(\mathrm{L},\preceq)$ is a collection $\mathrm{L}$ of elements and a relation $\preceq$ that is reflexive ($x\preceq x,~\forall x\in\mathrm{L}$), transitive ($x\preceq y,y\preceq z\Rightarrow x\preceq z,~\forall x,y,z\in\mathrm{L}$), and anti-symmetric ($x\preceq y,y\preceq x\Rightarrow x=y,~\forall x,y\in\mathrm{L}$).
 We use notation $x\prec y$ if $x\preceq y$ and $x\neq y$.
-An element $y\in L$ covers $x\in L$, denoted as $x\lessdot y$, if $x\prec y$ and there is no $z\in L$ with $x\prec z\prec y$.
-The poset $\mathcal{L}$ is connected if for every $x,y\in L$, there exists a finite sequence of $k\in\mathbb{N}_{+}$ elements $x=x_{1},x_{2},\dots,x_{k}=y$ such that either $x_{i+1}\gtrdot x_{i}$ or $x_{i+1}\lessdot x_{i}$ for every $i=1,2,\dots,k-1$.
-An element $x\in L$ is called the least (greatest) element if $x\preceq y$ ($x\succeq y$) for all $y\in L$.
+An element $y\in\mathrm{L}$ covers $x\in\mathrm{L}$, denoted as $x\lessdot y$, if $x\prec y$ and there is no $z\in\mathrm{L}$ with $x\prec z\prec y$.
+The poset $\mathcal{L}$ is connected if for every $x,y\in\mathrm{L}$, there exists a sequence of $k$ elements $x=x_{1},x_{2},\dots,x_{k}=y$ such that either $x_{i+1}\gtrdot x_{i}$ or $x_{i+1}\lessdot x_{i}$ for every $i=1,2,\dots,k-1$.
+An element $x\in\mathrm{L}$ is a minimal (maximal) element if there is no $y\in\mathrm{L}$ such that $y\prec x$ ($y\succ x$).
+An element $x\in\mathrm{L}$ is called the least (greatest) element if $x\preceq y$ ($x\succeq y$) for all $y\in\mathrm{L}$.
 Elements $x,y\in\mathcal{L}$ are said to be comparable if either $x\preceq y$ or $y\succeq x$; otherwise, they are incomparable.
 Definition 4 (Model-oriented poset).
-Given a graph class $\mathfrak{G}$ and an associated model map $\mathcal{M}$ that satisfy Condition˜1, the corresponding model-oriented poset is $\mathcal{L}:=(\mathfrak{G},\preceq)$ with $\preceq$’ given by model containment: for $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, we say $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ if $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$.
+Given a graph class $\mathfrak{G}$ and an associated model map $\mathcal{M}$ that satisfies Condition˜1, the corresponding model-oriented poset is $\mathcal{L}:=(\mathfrak{G},\preceq)$ with $\preceq$’ given by model containment: for $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, we say $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ if $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$.
 Reflexivity and transitivity of $\preceq$’ hold by definition, while the anti-symmetry follows from the injectivity of $\mathcal{M}(\cdot)$ (Condition˜1).
 The poset $\mathcal{L}$ offers a natural way to organize the graphs $\mathcal{G}\in\mathfrak{G}$ based on the models they represent.
-The neighbors in Definition˜1 are exactly the covering relations in the poset: $\mathcal{G}_{2}$ is a neighbor of $\mathcal{G}_{1}$ if either $\mathcal{G}_{2}$ covers $\mathcal{G}_{1}$ or $\mathcal{G}_{1}$ covers $\mathcal{G}_{2}$.
+The neighbors in Definition˜1 are exactly given by the covering relations: $\mathcal{G}^{\prime}$ is a neighbor of $\mathcal{G}$ if either $\mathcal{G}$ covers $\mathcal{G}^{\prime}$ or $\mathcal{G}^{\prime}$ covers $\mathcal{G}$, i.e.,
+$\text{Neighbors}(\mathcal{G})=\{\mathcal{G}^{\prime}:\mathcal{G}^{\prime}\lessdot\mathcal{G}\}\cup\{\mathcal{G}^{\prime}:\mathcal{G}^{\prime}\gtrdot\mathcal{G}\},\quad\mathcal{G}\in\mathfrak{G}.$ (5)
 Note that Condition˜2 is equivalent to requiring that the poset $\mathcal{L}$ is connected, which is typically ensured by the existence of a least element $\mathcal{G}_{\hat{0}}$ (usually the empty graph) in $\mathfrak{G}$.
-Figure 3: Hasse diagrams for different statistical graphs over vertex set $\{1,2,3\}$: each box represents a graph, and each dashed arrow represents a covering relation, i.e., we draw $\mathcal{G}_{1}\dashrightarrow\mathcal{G}_{2}$ pointing upwards if $\mathcal{G}_{2}$ covers $\mathcal{G}_{1}$.
+Figure 3: Hasse diagrams for four different statistical graph classes over vertex set $\{1,2,3\}$: each box represents a graph, and each dashed arrow represents a covering relation, i.e., we draw $\mathcal{G}_{1}\dashrightarrow\mathcal{G}_{2}$ pointing upwards if $\mathcal{G}_{2}$ covers $\mathcal{G}_{1}$.
 We can determine whether $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ by checking whether there is a directed path from $\mathcal{G}_{1}$ to $\mathcal{G}_{2}$.
 In (b) and (d), only part of the poset is shown for simplicity.
-Each poset has a least element, but not all have a greatest element.
-The poset in panel (d) is not graded when the graphs have more than three vertices (see Proposition˜10).
-This means there is no consistent way to assign ranks to the graphs.
+Each poset has a least element; posets (a), (c) and (d) also have a greatest element.
+Posets (a), (b) and (c) and graded; (d) is not graded (see Proposition˜8).
 For example, consider $\mathfrak{G}$ to be all the UGs over the vertex set $\{1,2,3\}$ and let $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}$ be the model map.
 In this case, the partial order $\preceq$’ can be characterized graphically: for $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, one can show that $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ if and only if $\mathcal{G}_{1}$ is a subgraph of $\mathcal{G}_{2}$.
-Further, $\mathcal{G}_{2}$ covers $\mathcal{G}_{1}$ if and only if $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ and $\mathcal{G}_{2}$ has exactly one more edge than $\mathcal{G}_{1}$.
+Further, $\mathcal{G}_{2}$ covers $\mathcal{G}_{1}$ if and only if $\mathcal{G}_{2}$ has exactly one more edge than $\mathcal{G}_{1}$.
 The least element $\mathcal{G}_{\hat{0}}$ in the poset is the empty graph.
-This poset can be visualized with a Hasse diagram shown in Fig.˜3(a): each box is a a graph; if $\mathcal{G}_{2}$ covers $\mathcal{G}_{1}$, we then place $\mathcal{G}_{2}$ above $\mathcal{G}_{1}$ and draw an upward arrow $\mathcal{G}_{1}\dashrightarrow\mathcal{G}_{2}$.
+This poset can be visualized with a Hasse diagram shown in Fig.˜3(a): each box is a graph; if $\mathcal{G}_{2}$ covers $\mathcal{G}_{1}$, we then place $\mathcal{G}_{2}$ above $\mathcal{G}_{1}$ and draw an upward arrow $\mathcal{G}_{1}\dashrightarrow\mathcal{G}_{2}$.
 Partial order can be read off from the diagram: $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ if there exists a directed path (can be of length zero) from $\mathcal{G}_{1}$ to $\mathcal{G}_{2}$.
@@
 Finally, it is not hard to see that our model-oriented distance, denoted hereafter as $d_{\mathcal{L}}$ to highlight its dependence on $\mathcal{L}$ and hence $(\mathfrak{G},\mathcal{M})$, is given by
+$d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=\min\left\{\mathrm{length}(p):p\in\mathrm{P}_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\right\},\quad\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G},$
+where
 […]
-where the path can contain dashed arrows in both directions.
-We note that the ‘shortest path length’ is a standard way to measure distance in posets (Monjardet, 1981; Trotter, 1992; Brualdi, Graves and Lawrence, 1995), although its use has primarily been confined to the order theory literature.
-3.3 Poset structures that facilitate the computation of model-oriented distance
-Having established that the model-oriented distance corresponds to the shortest-path distance in the Hasse diagram of the model-oriented poset, we can compute it by using using the well-known $A^{*}$ algorithm.
-In general, however, this can be expensive, particularly because it requires repeatedly identifying all neighbors of a graph while traversing many candidate paths in the Hasse diagram.
-Often, however, we can exploit additional structure in the poset to simplify computation — either by deriving a closed-form expression for the distance (eliminating the need for $A^{*}$), by streamlining the neighborhood search, or by reducing the number of paths that must be considered.
-In the following, we study several such structural properties in general, as we will rely on them when analyzing the model-oriented distances corresponding to probabilistic UG and CPDAG models, as well as causal DAGs and MPDAG models (see Section 4).
-We also examine the conditions under which the model-oriented distance reduces to an edge-by-edge distance similarly to the SHD.
+$=\left\{(\mathcal{G}_{0},\dots,\mathcal{G}_{d}):\mathcal{G}_{0}=\mathcal{G}_{s},\mathcal{G}_{d}=\mathcal{G}_{t},\,\text{and for $1\leq i\leq d$, $\mathcal{G}_{i-1}\lessdot\mathcal{G}_{i}$ or $\mathcal{G}_{i-1}\gtrdot\mathcal{G}_{i}$}\right\}.$
+We note that the shortest path length is a standard way to measure distance in posets (Monjardet, 1981; Trotter, 1992; Brualdi, Graves and Lawrence, 1995), although its use has primarily been confined to the order theory literature.
+Remark 1.
+Although a poset always admits a dual poset by replacing $\preceq$’ with $\succeq$’, it is important to adopt and maintain a convention, since some properties are asymmetric.
+As will become clear in Section˜4, there is, for example, an asymmetry between the empty graph and the full graph, and likewise between lower and upper semimodularity (see Table˜2).
+This asymmetry is related to the fact that, starting from an arbitrary graph, we can more commonly obtain a graph in the same class by removing edges than adding edges.
+3.3 Poset structures that facilitate computing model-oriented distances
+Since the model-oriented distance corresponds to the shortest path in the poset’s Hasse diagram, it can, in principle, be computed using search algorithms like A*.
+However, a naive search is often computationally prohibitive, as it requires repeatedly generating all neighbors for numerous candidate paths.
+Hence, we investigate structural properties of the poset to save computation — by deriving a closed-form expression for distance, characterizing the neighbors, or pruning the search space.
+In this subsection, we analyze several such structural properties, laying the groundwork for our specific analysis of probabilistic UGs and CPDAGs, as well as causal DAGs and MPDAGs, in Section˜4.
 3.3.1 Graded poset
-A connected poset $\mathcal{L}$ is graded if there exists a function $\operatorname{\mathrm{rank}}:\mathfrak{G}\rightarrow\mathbb{N}$ such that $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$ implies $\operatorname{\mathrm{rank}}(\mathcal{G}_{2})=\operatorname{\mathrm{rank}}(\mathcal{G}_{1})+1$ and $\operatorname{\mathrm{rank}}(\mathcal{G})=0$ for a minimal $\mathcal{G}\in\mathfrak{G}$.
-It can be shown that if it exists, the rank function is unique.
-A graph $\mathcal{G}^{\prime}\in\mathfrak{G}$ is called a minimal element if there exists no $\mathcal{G}\in\mathfrak{G}$ such that $\mathcal{G}\prec\mathcal{G}^{\prime}$ ($\mathcal{G}\succ\mathcal{G}^{\prime}$).
+A connected poset $\mathcal{L}$ is graded if it admits a function $\operatorname{\mathrm{rank}}:\mathfrak{G}\rightarrow\mathbb{N}$ such that (i) $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$ implies $\operatorname{\mathrm{rank}}(\mathcal{G}_{2})=\operatorname{\mathrm{rank}}(\mathcal{G}_{1})+1$, and (ii) $\operatorname{\mathrm{rank}}(\mathcal{G})=0$ holds for some minimal element $\mathcal{G}$.
+It can be shown that if it exists, such a rank function is unique.
 If the model-oriented poset is graded, we can exploit this property to simplify computation in two ways.
 First, we obtain the following closed form characterization of the model-oriented distance between graphs comparable within the poset.
-Proposition 1.
-Suppose $\mathcal{L}=(\mathfrak{G},\preceq)$ is a graded poset with rank function $\mathrm{rank}(\cdot)$.
-For any $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$ with $\mathcal{G}_{s}\succeq\mathcal{G}_{t}$, every downward path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ has length $\mathrm{rank}(\mathcal{G}_{s})-\mathrm{rank}(\mathcal{G}_{t})$, which also equals $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
-This is a well-known result that follows from the property that in a graded poset every maximal chain that is bounded between two comparable elements are of the same length.
+Proposition 1 (Distance between comparable graphs in a graded poset).
+Suppose $\mathcal{L}=(\mathfrak{G},\preceq)$ is a graded poset with rank function $\operatorname{\mathrm{rank}}(\cdot)$.
+For any $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$ with $\mathcal{G}_{s}\succeq\mathcal{G}_{t}$, every downward path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ has length $\operatorname{\mathrm{rank}}(\mathcal{G}_{s})-\operatorname{\mathrm{rank}}(\mathcal{G}_{t})$, which also equals $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
+This is a well-known result that follows from the property that in a graded poset all maximal chains between two comparable elements have equal length.
 A subset of $\mathfrak{G}$ is a chain (i.e., totally ordered set) if every pair of elements are comparable.
@@
 For completeness, we provide a short proof in Section˜C.1.
-Second, gradedness makes it significantly easier to enumerate all neighbors of a given graph $\mathcal{G}$.
-Recall that $\mathcal{G}^{\prime}$ is a neighbor of $\mathcal{G}$ in $\mathcal{L}$ if either $\mathcal{G}^{\prime}\preceq\mathcal{G}$ and there is no distinct graph $\tilde{\mathcal{G}}$ with $\mathcal{G}^{\prime}\preceq\tilde{\mathcal{G}}\preceq\mathcal{G}$, or $\mathcal{G}\preceq\mathcal{G}^{\prime}$ and there is no distinct graph $\tilde{\mathcal{G}}$ with $\mathcal{G}\preceq\tilde{\mathcal{G}}\preceq\mathcal{G}^{\prime}$.
-In general, enumerating all neighbors can be cumbersome, since one must check for the absence of these intermediate elements.
-If $\mathcal{L}$ is graded, this process becomes much simpler since two graphs can only be neighbors if their ranks differ by exactly one.
-Combined with the requirement that one graph is either above or below the other in the poset, this condition becomes both necessary and sufficient.
-This characterization is particularly useful when the rank function has a simple form.
-For instance, we will show in Section 4.1 that the poset underlying undirected graphs (UGs) is graded with the rank of a graph equal to its number of edges.
-All neighbors of a UG $\mathcal{G}$ are therefore UGs $\mathcal{G}^{\prime}$ that differ by exactly one edge and are comparable to $\mathcal{G}$.
-In a non-graded poset, neighboring graphs may differ more substantially; for example, we show in Section 4.4 that for causal MPDAGs, a neighboring graph may have many more edges.
+Gradedness and the rank can greatly facilitate the enumeration of neighborhood, which is the basic routine to any search algorithm.
+Under gradedness, Eq.˜5 becomes
+$\text{Neighbors}(\mathcal{G})=\{\mathcal{G}^{\prime}:\mathcal{G}^{\prime}\preceq\mathcal{G},\,\operatorname{\mathrm{rank}}(\mathcal{G}^{\prime})=\operatorname{\mathrm{rank}}(\mathcal{G})-1\}\\ \cup\{\mathcal{G}^{\prime}:\mathcal{G}^{\prime}\succeq\mathcal{G},\,\operatorname{\mathrm{rank}}(\mathcal{G}^{\prime})=\operatorname{\mathrm{rank}}(\mathcal{G})+1\},\quad\mathcal{G}\in\mathfrak{G}.$
 3.3.2 Down-up and up-down distance, semi-lattice, semimodularity
 To reduce the number of paths we need to explore, it is natural to consider the shortest path among a restricted set of paths between two graphs in the Hasse diagram.
-Specifically, for $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$, we consider the set of down-up and up-down paths between them, denoted by $\mathrm{P}_{\downarrow\uparrow}{(\mathcal{G}_{s},\mathcal{G}_{t})}$ and $\mathrm{P}_{\uparrow\downarrow}{(\mathcal{G}_{s},\mathcal{G}_{t})}$ respectively.
+Specifically, for $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$, we consider the set of down-up and up-down paths between them, denoted by $\mathrm{P}_{\mathcal{L},\downarrow\uparrow}{(\mathcal{G}_{s},\mathcal{G}_{t})}$ and $\mathrm{P}_{\mathcal{L},\uparrow\downarrow}{(\mathcal{G}_{s},\mathcal{G}_{t})}$ respectively.
 A path is down-up if it takes the form of
 $\mathcal{G}_{s}=x_{0}\gtrdot\cdots\gtrdot x_{i^{\ast}}\lessdot\cdots\lessdot x_{d}=\mathcal{G}_{t},$
-where $i^{\ast}\in\{0,\dots,d\}$ is the inflection point; chains ($i^{\ast}=0$ or $i^{\ast}=d$) are special cases of down-up paths. Conversely, an up-down path takes the form of
+where $i^{\ast}\in\{0,\dots,d\}$ is the inflection point; downward and upward paths are special cases of down-up paths. Conversely, an up-down path takes the form of
 $\mathcal{G}_{s}=x_{0}\lessdot\cdots\lessdot x_{i^{\ast}}\gtrdot\cdots\gtrdot x_{d}=\mathcal{G}_{t}.$
 The next result characterizes when down-up (up-down) paths connect every pair of graphs.
-Proposition 2.
-The set of down-up paths $\mathrm{P}_{\downarrow\uparrow}{(\mathcal{G}_{1},\mathcal{G}_{2})}$ is non-empty for all $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ if and only if $\mathcal{L}$ has a least element (Condition˜3).
-Similarly, the set of up-down paths $\mathrm{P}_{\uparrow\downarrow}{(\mathcal{G}_{1},\mathcal{G}_{2})}$ is non-empty for all $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ if and only if $\mathcal{L}$ has a greatest element (Condition˜3).
+Proposition 2 (Existence of down-up and up-down paths).
+The set of down-up paths $\mathrm{P}_{\mathcal{L},\downarrow\uparrow}{(\mathcal{G}_{1},\mathcal{G}_{2})}$ is non-empty for all $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ if and only if $\mathcal{L}$ has a least element (Condition˜3).
+Similarly, the set of up-down paths $\mathrm{P}_{\mathcal{L},\uparrow\downarrow}{(\mathcal{G}_{1},\mathcal{G}_{2})}$ is non-empty for all $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ if and only if $\mathcal{L}$ has a greatest element (Condition˜3).
 Proof.
@@
 Further, except for DAGs, all of them also have a greatest element.
-See Table˜1 for a summary.
+See Table˜2 for a summary.
 The down-up and up-down paths lead to the definition of down-up and up-down distances.
 Definition 5 (Down-up and up-down distance).
-Under Condition˜3, the down-up distance is defined as $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t}):=\min\{\mathrm{length}(p):p\in\mathrm{P}_{\downarrow\uparrow}{(\mathcal{G}_{s},\mathcal{G}_{t})}\}$ for every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$.
-Under Condition˜3, the up-down distance is defined as $d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t}):=\min\{\mathrm{length}(p):p\in\mathrm{P}_{\uparrow\downarrow}{(\mathcal{G}_{s},\mathcal{G}_{t})}\}$ for every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$.
+Under Condition˜3, the down-up distance is defined as $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t}):=\min\{\mathrm{length}(p):p\in\mathrm{P}_{\mathcal{L},\downarrow\uparrow}{(\mathcal{G}_{s},\mathcal{G}_{t})}\}$ for every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$.
+Under Condition˜3, the up-down distance is defined as $d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t}):=\min\{\mathrm{length}(p):p\in\mathrm{P}_{\mathcal{L},\uparrow\downarrow}{(\mathcal{G}_{s},\mathcal{G}_{t})}\}$ for every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$.
 Unlike $d_{\mathcal{L}}$, the down-up and up-down distances may not be metrics as they are not guaranteed to obey the triangle inequality.
@@
 In particular, when $\mathcal{L}$ is graded, the down-up or up-down distances take a closed form.
-Proposition 3.
-Suppose $\mathcal{L}=(\mathfrak{G},\preceq)$ is graded with rank function $\mathrm{rank}(\cdot)$. Then, under Condition˜3, we have
+Proposition 3 (Down-up and up-down distances for graded posets).
+Suppose $\mathcal{L}=(\mathfrak{G},\preceq)$ is graded with rank function $\operatorname{\mathrm{rank}}(\cdot)$. Then, under Condition˜3, we have
 $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})=\operatorname{\mathrm{rank}}(\mathcal{G}_{s})+\operatorname{\mathrm{rank}}(\mathcal{G}_{t})-2\max_{\mathcal{G}\preceq\mathcal{G}_{s},\mathcal{G}_{t}}\operatorname{\mathrm{rank}}(\mathcal{G}),\quad\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}.$
@@
 Any down-up path between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ moves downward from $\mathcal{G}_{s}$, and then upward from an inflection point $\mathcal{G}$ to $\mathcal{G}_{t}$.
-Since $\mathcal{G}\preceq\mathcal{G}_{s}$ and $\mathcal{G}\preceq\mathcal{G}_{t}$, Proposition 1 implies that the length of the path is $\mathrm{rank}(\mathcal{G}_{s})+\mathrm{rank}(\mathcal{G}_{t})-2\mathrm{rank}(\mathcal{G})$.
-Therefore, the shortest down-up path corresponds to an inflection point $\mathcal{G}$ that maximizes $\mathrm{rank}(\mathcal{G})$ over all $\mathcal{G}\preceq\mathcal{G}_{s},\mathcal{G}\preceq\mathcal{G}_{t}$.
+Since $\mathcal{G}\preceq\mathcal{G}_{s}$ and $\mathcal{G}\preceq\mathcal{G}_{t}$, Proposition 1 implies that the length of the path is $\operatorname{\mathrm{rank}}(\mathcal{G}_{s})+\operatorname{\mathrm{rank}}(\mathcal{G}_{t})-2\operatorname{\mathrm{rank}}(\mathcal{G})$.
+Therefore, the shortest down-up path corresponds to an inflection point $\mathcal{G}$ that maximizes $\operatorname{\mathrm{rank}}(\mathcal{G})$ over all $\mathcal{G}\preceq\mathcal{G}_{s},\mathcal{G}\preceq\mathcal{G}_{t}$.
 The proof for the up-down distance follows similarly.
 ∎
-The following additional simplification requires notions of meet and join semi-lattice structures.
-In particular, graph $\mathcal{G}_{3}\in\mathfrak{G}$ is the least upper bound (or join) of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, denoted as $\mathcal{G}_{1}\vee\mathcal{G}_{2}$, if $\mathcal{G}_{3}\succeq\mathcal{G}_{1},\mathcal{G}_{2}$ and $\mathcal{G}_{3}\preceq\mathcal{G}^{\prime}$ for every $\mathcal{G}^{\prime}$ satisfying $\mathcal{G}^{\prime}\succeq\mathcal{G}_{1},\mathcal{G}_{2}$.
+For a further simplification, we require the notions of meet and join semi-lattice structures: graph $\mathcal{G}_{3}\in\mathfrak{G}$ is the least upper bound (or join) of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, denoted as $\mathcal{G}_{1}\vee\mathcal{G}_{2}$, if $\mathcal{G}_{3}\succeq\mathcal{G}_{1},\mathcal{G}_{2}$ and $\mathcal{G}_{3}\preceq\mathcal{G}^{\prime}$ for every $\mathcal{G}^{\prime}$ satisfying $\mathcal{G}^{\prime}\succeq\mathcal{G}_{1},\mathcal{G}_{2}$.
 The greatest lower bound (or meet) of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ is defined similarly and is denoted as $\mathcal{G}_{1}\wedge\mathcal{G}_{2}$.
 Poset $\mathcal{L}$ is a join (meet) semi-lattice if $\mathcal{G}\vee\mathcal{G}^{\prime}$ ($\mathcal{G}\wedge\mathcal{G}^{\prime}$) exists for every $\mathcal{G},\mathcal{G}^{\prime}\in\mathfrak{G}$.
-Corollary 1.
-Suppose $\mathcal{L}=(\mathfrak{G},\preceq)$ is graded with rank function $\mathrm{rank}(\cdot)$. Under Condition˜3, if $\mathcal{L}$ is also a meet semi-lattice, we have
+Corollary 1 (Simplification with semi-lattice structure).
+Suppose $\mathcal{L}=(\mathfrak{G},\preceq)$ is graded with rank function $\operatorname{\mathrm{rank}}(\cdot)$. Under Condition˜3, if $\mathcal{L}$ is a meet semi-lattice, we have
 $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})=\operatorname{\mathrm{rank}}(\mathcal{G}_{s})+\operatorname{\mathrm{rank}}(\mathcal{G}_{t})-2\operatorname{\mathrm{rank}}(\mathcal{G}_{s}\wedge\mathcal{G}_{t}),\quad\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}.$
@@
 […]
-Theorem 2.
+Theorem 2 (Simplification for semimodular posets).
 Given a graph class $\mathfrak{G}$ and a model map $\mathcal{M}$ that satisfy Condition˜1, let $\mathcal{L}=(\mathfrak{G},\preceq)$ be the model-oriented poset.
@@
 Fig.˜4 illustrates the basic idea: when $\mathcal{L}$ is semimodular, any shortest path can be transformed into a down-up or up-down path with the same length.
-We remark that, to our best knowledge, our characterization weakens the conditions for existing results in the literature and is well-suited for studying model-oriented graph distances.
-For example, Monjardet (1981, Theorem 8) additionally requires the poset to be graded, while Foldes and Radeleczki (2021, Proposition 2.5) instead requires the poset to be a semi-lattice.
-However, the model-oriented posets for graphs may not satisfy these conditions.
-In fact, in Section˜4.3 we will use Theorem˜2 to study polytree CPDAGs, where these earlier results are inapplicable.
+We remark that, to our best knowledge, Theorem˜2 weakens the conditions for existing results in the literature and is well-suited for studying model-oriented graph distances.
+For example, Monjardet (1981, Theorem 8) additionally requires the poset to be graded with a greatest element, while Foldes and Radeleczki (2021, Proposition 2.5) instead requires the poset to be a semi-lattice.
+We will use Theorem˜2 to study the model-oriented distance for polytree CPDAGs and MPDAGs in Sections˜4.3.2 and 4.4.1.
 3.3.3 Relation to edge-by-edge distances
-Recall that SHDs measure the distance between two graphs by counting how many edges are mismatched.
-They are therefore by construction edge-by-edge style distance, in the sense that they can be computed considering each edge at a time.
-Different types of mismatches are often assigned different weights (Perrier, Imoto and Miyano, 2008).
+Recall that the structural Hamming distance is given by the number of edges that are mismatched between the two graphs.
+Different types of mismatches can receive different weights (Perrier, Imoto and Miyano, 2008).
 For PDAGs, for example, there are two common variants of SHD (Tsamardinos, Brown and Aliferis, 2006) defined as follows.
@@
 Let PDAGs $\mathcal{G}_{s},\mathcal{G}_{t}$ over a vertex set be encoded by adjacency matrix $A^{s},A^{t}$ respectively. The two variants of SHD are given by
-$\begin{aligned} \operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})&:=\sum_{i<j}\mathbb{I}\left[A^{s}_{i,j}\neq A^{t}_{i,j}\text{ or }A^{s}_{j,i}\neq A^{t}_{j,i}\right],\\ \operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})&:=\sum_{i<j}\mathbb{I}\left[A^{s}_{i,j}\neq A^{t}_{i,j}\right]+\mathbb{I}\left[A^{s}_{j,i}\neq A^{t}_{j,i}\right].\end{aligned}$
-Here, $\mathbb{I}[\cdot]$ is the indicator function.
-In the following corollary to Theorem˜2, we study for general graph-types when the model-oriented distance also behaves edge-by-edge, and discuss the implications for certain PDAGs considered in this work.
-Lemma 1.
-Suppose we consider graphs $\mathfrak{G}$ with $k$ edge types (see Section 2.1.1) that define a poset $\mathcal{L}$, in which they are ordered based on subgraph containment; that is, $\mathcal{G}_{1}=(V,E^{1}_{1}\cup\dots\cup E_{1}^{k})\preceq\mathcal{G}_{2}=(V,E^{1}_{2}\cup\dots\cup E^{k}_{2})$ if $E^{i}_{1}\subseteq E^{i}_{2}$ for all $i$.
-Further, suppose $\mathcal{L}$ is graded with $\mathrm{rank}(\mathcal{G})=\sum_{i=1}^{k}|E_{i}|$ for any $\mathcal{G}\in\mathfrak{G}$, and that it is lower semimodular.
-Then, $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=\sum_{i=1}^{k}|E^{1}_{i}\triangle E^{2}_{i}|$, where $\triangle$ denotes the symmetric set difference.
-If the graph class contains only undirected edges, this is equal to $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})$.
-If the class contains only directed edges, this is equal to $\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})$.
+$\begin{aligned} \operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})&:=\sum_{i<j}\mathbb{I}\left[A^{s}_{i,j}\neq A^{t}_{i,j}\text{ or }A^{s}_{j,i}\neq A^{t}_{j,i}\right],\\ \operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})&:=\sum_{i,j}\mathbb{I}\left[A^{s}_{i,j}\neq A^{t}_{i,j}\right].\end{aligned}$
+Here $\mathbb{I}[\cdot]$ is the indicator function.
+Note that $\operatorname{\mathrm{SHD}}_{2}\geq\operatorname{\mathrm{SHD}}_{1}$ holds by definition.
+The following result outlines when our distance reduces to an edge-by-edge type.
+Lemma 1 (Edge-by-edge model oriented distances).
+Let $\mathfrak{G}$ be a class of graphs with $k\geq 1$ edge types (see Section˜2.2.1).
+Consider a poset $\mathcal{L}$ that orders graphs by subgraph containment, i.e., $\mathcal{G}_{s}=(\mathrm{V},\mathrm{E}^{1}_{s},\dots,\mathrm{E}_{s}^{k})\preceq\mathcal{G}_{t}=(\mathrm{V},\mathrm{E}^{1}_{t},\dots,\mathrm{E}^{k}_{t})$ if $\mathrm{E}^{i}_{s}\subseteq\mathrm{E}^{i}_{t}$ for $i=1,\dots,k$.
+Suppose $\mathcal{L}$ is lower semimodular and graded with $\operatorname{\mathrm{rank}}(\mathcal{G})=\sum_{i=1}^{k}|\mathrm{E}^{i}|$ for each $\mathcal{G}=(\mathrm{V},\mathrm{E}^{1},\dots,\mathrm{E}^{2})$.
+Then, we have $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=\sum_{i=1}^{k}|\mathrm{E}^{i}_{s}\triangle\mathrm{E}^{i}_{t}|$.
+Further, when $k=1$, $d_{\mathcal{L}}$ reduces to $\operatorname{\mathrm{SHD}}_{1}$ if $\mathfrak{G}$ is a class of undirected graphs, and to $\operatorname{\mathrm{SHD}}_{2}$ if $\mathfrak{G}$ is a class of directed graphs.
 Proof.
-Since the poset is lower semimodular, we have from Theorem 2 that $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$.
-By construction, $\mathcal{L}$ is a meet semi-lattice with $\mathcal{G}_{1}\wedge\mathcal{G}_{2}=(V,E_{1}^{1}\cap E_{2}^{1}\cup\dots\cup E_{1}^{k}\cap E_{2}^{k})$.
-Appealing to Corollary 1, we then have that: $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=\mathrm{rank}(\mathcal{G}_{s})+\mathrm{rank}(\mathcal{G}_{t})-2\mathrm{rank}(\mathcal{G}_{s}\wedge\mathcal{G}_{t})=\sum_{i=1}^{k}|E^{1}_{i}\triangle E^{2}_{i}|$.
+Since the poset is lower semimodular, we have from Theorem˜2 that $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$.
+By construction, $\mathcal{L}$ is a meet semi-lattice with $\mathcal{G}_{s}\wedge\mathcal{G}_{t}=(\mathrm{V},\mathrm{E}_{s}^{1}\cap\mathrm{E}_{t}^{1},\dots,\mathrm{E}_{s}^{k}\cap\mathrm{E}_{t}^{k})$.
+Appealing to Corollary˜1, we then have $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=\operatorname{\mathrm{rank}}(\mathcal{G}_{s})+\operatorname{\mathrm{rank}}(\mathcal{G}_{t})-2\operatorname{\mathrm{rank}}(\mathcal{G}_{s}\wedge\mathcal{G}_{t})=\sum_{i=1}^{k}\left(|\mathrm{E}^{i}_{s}|+|\mathrm{E}^{i}_{t}|-2|\mathrm{E}^{i}_{s}\cap\mathrm{E}^{i}_{t}|\right)=\sum_{i=1}^{k}|\mathrm{E}^{i}_{s}\triangle\mathrm{E}^{i}_{t}|$.
 ∎
-Effectively, this shows that the model-oriented distance is an edge-by-edge distance whenever the partial ordering can be determined by examining one edge at a time (subgraph containment) and the Hasse diagram is sufficiently dense (lower semimodularity).
-We will see that this typically holds for graph classes that impose few constraints, such as UGs and DAGs, but not for CPDAGs or MPDAGs.
-Which version of the SHD it coincides with then depends on whether the edges are directed or undirected.
+Effectively, this shows that the model-oriented distance is an edge-by-edge distance whenever the partial ordering can be determined by examining one edge at a time (subgraph containment) and the Hasse diagram is sufficiently dense (lower semimodular).
+This typically holds for graph classes that impose few constraints, such as UGs and DAGs, but not for CPDAGs or MPDAGs.
+The version of the SHD it coincides with depends on whether the edges are directed or undirected.
 In particular, as we will see in the next section, for UGs with model map $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}$, the model-oriented distance is equivalent to $\operatorname{\mathrm{SHD}}_{1}$.
-However, the model-oriented distance diverges from $\operatorname{\mathrm{SHD}}_{1}$ for all other graph types considered in this paper.
-4 Model-oriented distance $d_{\mathcal{L}}$ in probabilistic and causal graphs
-Next, we apply the poset framework described in Section˜3 to derive the model-oriented distance for UGs, DAGs, CPDAGs, and MPDAGs, presented in order of increasing complexity in their associated poset.
-Table˜1 highlights the different attributes for each associated poset, with the class of UGs being the most structured, and the class of MPDAGs being the least structured.
-We will formally show these properties in this section.
-For each graph class, we first study the corresponding model-oriented poset, then use it to derive the model-oriented distance, and finally examine simpler characterizations of the model-oriented distance, for example in terms of the SHD, or when this is not possible approximations to the model-oriented distance.
-Table 1: Structural attributes of the underlying model-oriented poset for different graph classes
+The model-oriented distance diverges from $\operatorname{\mathrm{SHD}}_{1}$ for all other graph types considered in this paper.
+We will compare our distance with SHD numerically in Section˜5.
+4 Model-oriented distance for probabilistic and causal graphs
+We are ready to apply the general poset framework to UGs, DAGs, CPDAGs, MPDAGs as well as simpler subclasses of CPDAGs and MPDAGs.
+Table˜2 summarizes the structural properties of the poset underlying each graph class, with UGs being the most structured and MPDAGs being the least structured.
+We will use this section to prove these properties and consider their implication on computing the model-oriented distance.
+Table 2: Structural properties of the underlying model-oriented poset for different graph classes
 Graph class Model Least element Greatest element Graded Semi-lattice Semimodular
@@
 MPDAG causal empty graph full graph
-Polytree MPDAG causal empty graph
+Polytree MPDAG causal empty graph lower
 4.1 Probabilistic UGs
 As a warm-up, consider $\mathfrak{G}$ the class of UGs over a finite vertex set $\mathrm{V}$ representing probabilistic models in the sense of $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}$ given by Eq.˜1.
-The next result states that the model inclusion becomes graph inclusion, which we prove in Section˜A.2.
-Proposition 4.
+The next result states that model inclusion becomes graph inclusion for UGs, which we prove in Section˜A.2.
+Proposition 4 (UG comparability).
 For UGs $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, we have $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}(\mathcal{G}_{1})\subseteq\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}(\mathcal{G}_{2})$ if and only if $\mathcal{G}_{1}$ is a subgraph of $\mathcal{G}_{2}$.
@@
 Fig.˜3(c) shows its Hasse diagram when $|\mathrm{V}|=3$.
-Proposition 5.
+Theorem 3 (UG poset).
 Let $\mathcal{L}$ be the model-oriented poset associated with $(\mathfrak{G},\mathcal{M}^{\mathrm{o}}_{\textsc{ug}})$, where $\mathfrak{G}$ is the class of UGs over a finite vertex set $\mathrm{V}$.
@@
     $\mathcal{L}$ is both lower and upper semimodular.
-From these properties, we obtain a closed form for the model-oriented distance.
+From Theorems˜3 and 1, we obtain a closed form for the model-oriented distance.
 Corollary 2.
@@
 $d_{\mathcal{L}}(\mathcal{G}_{1},\mathcal{G}_{2})=|\mathrm{E}_{1}\triangle\mathrm{E}_{2}|=\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{1},\mathcal{G}_{2}),\quad\mathcal{G}_{1}=(\mathrm{V},\mathrm{E}_{1}),\;\mathcal{G}_{2}=(\mathrm{V},\mathrm{E}_{2})\in\mathfrak{G}.$
-Proof.
-It follows from Proposition˜5 and Lemma˜1. ∎
 4.2 Causal DAGs
-Consider $\mathfrak{G}$ the class of DAGs over a finite vertex set representing causal models in the sense of $\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}$ given by Eq.˜4.
+Let $\mathfrak{G}$ be the class of DAGs over a finite vertex set representing causal models in the sense of $\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}$ given by Eq.˜4.
 The partial order is given by subgraph containment, stated as follows.
-Proposition 6.
+Proposition 5 (DAG comparability).
 For DAGs $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, we have $\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}(\mathcal{G}_{1})\subseteq\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}(\mathcal{G}_{2})$ if and only if $\mathcal{G}_{1}$ is a subgraph of $\mathcal{G}_{2}$.
-We prove Proposition˜6 in Section˜A.2.
-Using this result, the model-oriented poset can be characterized as follows; see Section˜E.1 for its proof.
+We prove Proposition˜5 in Section˜A.2.
+Using this result, the model-oriented poset is characterized by the next result, which is proved in Section˜E.1.
 Fig.˜3(b) shows the Hasse diagram of this poset when the vertex set is of size 3.
-Proposition 7.
+Theorem 4 (DAG poset).
 Let $\mathcal{L}$ be the model-oriented poset associated with $(\mathfrak{G},\mathcal{M}^{\mathrm{c}}_{\textsc{dag}})$, where $\mathfrak{G}$ is the class of DAGs over a finite vertex set $\mathrm{V}$.
@@
   1. 1.
-    $\mathcal{G}_{\hat{0}}$ is the empty graph. There is no $\mathcal{G}_{\hat{1}}$.
+    $\mathcal{G}_{\hat{0}}$ is the empty graph. There is no $\mathcal{G}_{\hat{1}}$ when $|\mathrm{V}|>1$.
   2. 2.
@@
     $\mathcal{L}$ is lower semimodular.
-From Propositions˜7 and 1, we obtain the model-oriented distance as follows.
+From Theorems˜4 and 1, we obtain a closed form for the model-oriented distance as follows.
 Corollary 3.
@@
 Let $\mathfrak{G}$ be the class of CPDAGs over a finite vertex set $\mathrm{V}$, which represent probabilistic models in the sense of $\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}}$ given by Eq.˜2.
-We use $\mathcal{G}]$ to denote the set of Markov equivalent DAGs represented by a CPDAG $\mathcal{G}$.
+For $\mathcal{G}\in\mathfrak{G}$, we use $\mathcal{G}]$ to denote the set of Markov equivalent DAGs it represents.
 The model containment is characterized by a basic result due to Chickering (2002), restated as follows.
-Lemma 2 (Chickering, 2002, Theorem 4).
+Lemma 2 (CPDAG comparability).
 For CPDAGs $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$, the following are equivalent:
@@
 For operation (i), $\mathcal{D}_{i-1}$ and $\mathcal{D}_{i}$ are Markov equivalent; for (ii), the d-separations in $\mathcal{D}_{i-1}$ contain those in $\mathcal{D}_{i}$.
-Here, an edge $u\rightarrow v$ is covered in DAG $\mathcal{D}$ if $\operatorname{Pa}_{\mathcal{D}}(v)=\operatorname{Pa}_{\mathcal{D}}(u)\cup\{u\}$.
-Using the result above, we can derive the following properties of CPDAG’s model-oriented poset.
+Here, a directed edge $u\rightarrow v$ is covered in DAG $\mathcal{D}$ if $\operatorname{Pa}_{\mathcal{D}}(v)=\operatorname{Pa}_{\mathcal{D}}(u)\cup\{u\}$.
+Using the result above, we can derive the properties of CPDAG’s model-oriented poset listed in Theorem˜5.
 Fig.˜3(a) shows the Hasse diagram of this poset when $|\mathrm{V}|=3$.
-Proposition 8.
+Remark 2.
+In light of the second result in Lemma˜2, it may be tempting to conjecture that for CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$, we have $\mathcal{G}_{s}\preceq\mathcal{G}_{t}$ if and only if there exist DAGs $\mathcal{D}_{s}\in[\mathcal{G}_{s}],\mathcal{D}_{t}\in[\mathcal{G}_{t}]$ such that $\mathcal{D}_{s}$ is a subgraph of $\mathcal{D}_{t}$.
+But the ‘only if’ part of this is not true; consider the following two graphs over $\{1,2,3,4\}$:
+$\mathcal{G}_{s}:\begin{array}[]{ccc}1&&2\\ &\searrow&\downarrow\\ 4&&3\end{array},\qquad\mathcal{G}_{t}:\begin{array}[]{ccc}1&\rightarrow&2\\ \uparrow&\nwarrow&\uparrow\\ 4&&3\end{array}.$
+Taeb, Bühlmann and Chandrasekaran (2024) highlighted some properties of the CPDAG poset, which we restate as follows.
+Theorem 5 (CPDAG poset).
 Let $\mathcal{L}$ be the model-oriented poset associated with $(\mathfrak{G},\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}})$, where $\mathfrak{G}$ is the class of CPDAGs over a finite vertex set $\mathrm{V}$.
@@
     $\mathcal{L}$ is graded with $\operatorname{\mathrm{rank}}(\cdot)$ equal to the number of edges.
-Additionally, $\mathcal{L}$ is neither a (join or meet) semi-lattice nor (upper or lower) semimodular in general. For proofs of these results, see Section˜E.2.
+Additionally, $\mathcal{L}$ is neither a (join or meet) semi-lattice nor (upper or lower) semimodular in general. For completeness, we prove these results in Section˜E.2.
 4.3.1 Upper and lower bounds on $d_{\mathcal{L}}$
-The lack of semi-lattice and semimodularity structures makes computing our model-oriented distance between CPDAGs much harder than UGs and DAGs.
+The lack of semi-lattice and semimodularity structures makes computing our model-oriented distance between CPDAGs harder than for UGs and DAGs.
 First, Lemma˜1 is no longer applicable and indeed, as we demonstrate in Example˜1, the SHD behaves very differently from our distance.
 Further, by Theorem˜2, because the poset is not semimodular, up-down and down-up distances can only provide upper bounds on the model-oriented distance.
-Indeed, consider the example in Fig.˜5, where the shortest path, taking a ‘zigzag’ shape, is of length 4, while both the shortest down-up and up-down paths have length 8.
-Figure 5: The model-oriented distance between probabilistic CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ is 4, as shown by the shortest path in red.
-Meanwhile, the up-down distance is 8 (blue path) and the down-up distance is also 8 (orange path).
-Due to these difficulties, in Section˜5, we propose using the $\text{A}^{\ast}$ algorithm to compute $d_{\mathcal{L}}$.
-As an informed search algorithm, $\text{A}^{\ast}$ relies on good, custom-made lower and upper bounds on the distance to speed up the search: the lower bound, often called the heuristic function in the literature, allows one to disregard some paths without discarding a potentially optimal path; the upper bound let one prune all the paths that are longer than the current best solution.
+Indeed, consider the example in Fig.˜5, where the shortest path, taking a zigzag shape, is of length 4, while both the shortest down-up and up-down paths have length 8.
+Figure 5: The model-oriented distance between probabilistic CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ is 4, as shown by the shortest path drawn in red.
+Meanwhile, the up-down (blue path) and the down-up (orange path) distances are both 8.
+Due to these difficulties, in Section˜5, we propose using the A* algorithm to compute $d_{\mathcal{L}}$.
+As an informed search algorithm, A* relies on good, custom-made lower and upper bounds on the distance to speed up the search: the lower bound, often called the heuristic function in the literature, allows one to disregard some paths without discarding a potentially optimal path; the upper bound enables early termination and further pruning of suboptimal paths.
 In this section, we investigate lower and upper bounds on $d_{\mathcal{L}}$ between CPDAGs.
-To formulate the lower bound we need some notation.
-For a PDAG $\mathcal{G}$, we use $\operatorname{sk}(\mathcal{G})$ to denote its skeleton, which is the set of edges after undirecting all the edges in $\mathcal{G}$.
-We also use $\operatorname{v-str}(\mathcal{G})$ to denote the set of triplets $(a,b,c)$ that form a v-structure at $b$ in $\mathcal{G}$ (see Section˜2.1.4); here we do not distinguish between $(a,b,c)$ and $(c,b,a)$.
-Let us fix two CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ over a vertex set $\mathrm{V}$.
+To formulate the lower bound, we need some notation.
+For a PDAG $\mathcal{G}$, we use $\operatorname{sk}(\mathcal{G})$ to denote its skeleton, which is the set of edges upon undirecting all the edges in $\mathcal{G}$.
+We also use $\operatorname{v-str}(\mathcal{G})$ to denote the set of triplets $(a,b,c)$ that form a v-structure at $b$ in $\mathcal{G}$ (see Section˜2.2.4 for its definition); here we do not distinguish between $(a,b,c)$ and $(c,b,a)$.
 Let us fix two CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ over a vertex set $\mathrm{V}$.
@@
 $\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t}):=\Big\{(a,b,c)\subseteq\mathrm{V}\times\mathrm{V}\times\mathrm{V}:(a,b),(b,c)\in\operatorname{sk}(\mathcal{G}_{s})\cap\operatorname{sk}(\mathcal{G}_{t}),\\ (a,c)\not\in\operatorname{sk}(\mathcal{G}_{s})\cup\operatorname{sk}(\mathcal{G}_{t}),\text{ and }(a,b,c)\in\operatorname{v-str}(\mathcal{G}_{s})\triangle\operatorname{v-str}(\mathcal{G}_{t})\Big\},$
-which consists of v-structures that appear in only one of the two graphs. The set above can always be partitioned as
+which consists of v-structures that appear in only one of the two graphs. We partition $\mathrm{S}$ as
 $\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t})=\mathrm{S}_{1}\cup\dots\cup\mathrm{S}_{L},\quad L\geq 1,$
-where two triplets $(a,b,c)$ and $(d,e,f)$ are placed in the same partition if they have any vertex in common (e.g., $a=d$ or $a=e$).
-For a pair of vertices $(i,j)$, we define operations $\operatorname{OP}_{(i,j),\text{CE}}$ (cover edge) and $\operatorname{OP}_{(i,j),\text{ER}}$ (edge removal) that map a subset of $\mathrm{S}$ to a subset of $\mathrm{S}$, given by
-$\begin{array}[]{ll}\operatorname{OP}_{(i,j),\text{CE}}(\mathrm{T})&:=\mathrm{T}\setminus\{(i,k,j):k\in\mathrm{V}\},\\ \operatorname{OP}_{(i,j),\text{ER}}(\mathrm{T})&:=\mathrm{T}\setminus\{(i,j,k),(j,i,k):k\in\mathrm{V}\}\end{array},\quad\mathrm{T}\subseteq\mathrm{S}.$
+where two triplets $(a,b,c)$ and $(d,e,f)$ are placed in the same partition if they share any common vertex (e.g., $a=d$ or $a=e$).
+For a pair of vertices $(i,j)$, we define operations $\operatorname{OP}_{(i,j),\text{CE}}$ (cover edge) and $\operatorname{OP}_{(i,j),\text{ER}}$ (edge removal) that map a subset of $\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t})$ to a subset of $\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t})$, given by
+$\begin{array}[]{ll}\operatorname{OP}_{(i,j),\text{CE}}(\mathrm{T})&:=\mathrm{T}\setminus\{(i,k,j):k\in\mathrm{V}\},\\ \operatorname{OP}_{(i,j),\text{ER}}(\mathrm{T})&:=\mathrm{T}\setminus\{(i,j,k),(j,i,k):k\in\mathrm{V}\}\end{array},\quad\mathrm{T}\subseteq\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t}).$
 Here, $\operatorname{OP}_{(i,j),\text{CE}}$ removes v-structures of the form $i\rightarrow k\leftarrow j$ by adding a covering edge (see Lemma˜2) between $i$ and $j$; $\operatorname{OP}_{(i,j),\text{ER}}$ removes v-structures of the form $i\rightarrow j\leftarrow k$ or $j\rightarrow i\leftarrow k$ by removing the edge between $i$ and $j$.
 For each partition $\mathrm{S}_{l}$, consider the minimum number of operations taken to eliminate all the v-structures in the partition, namely
-$m_{\text{op}}(\mathrm{S}_{l}):=\min\left\{m\in\mathbb{N}:\exists~\mathrm{op}_{1},\dots,\mathrm{op}_{m}\text{ such that }\mathrm{op}_{1}\circ\dots\circ\mathrm{op}_{m}(\mathrm{S}_{i})=\emptyset\right\},\quad l=1,\dots,L,$
-where each $\mathrm{op}_{k}$ is either $\operatorname{OP}_{(i,j),\text{CE}}$ or $\operatorname{OP}_{(i,j),\text{ER}}$ for some $(i,j)$ that can depend on $k$.
-Because the v-structures from two different partitions share no common vertex by construction and this property is maintained by both operations, we conclude that
-$m_{\text{op}}(\mathrm{S}):=\sum_{l=1}^{L}m_{\text{op}}(\mathrm{S}_{l})$
-lower bounds the number of steps taken to align the set of v-structures between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$.
-Property 1 (Upper bound).
-For CPDAGs $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$, we have
-$d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq\min\left\{d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t}),d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t})\right\}.$
-Property 2 (Lower bound).
-For CPDAGs $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$, we have
-$d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\geq|\operatorname{sk}(\mathcal{G}_{s})\triangle\operatorname{sk}(\mathcal{G}_{t})|+2m_{\text{op}}(\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t})).$
-We prove Property 2 in Appendix F.1.
-Property 1 follows from recalling that the model-oriented distance is upper bounded by both the down-up and the up-down distance (both exist for CPDAGs since the poset has a least and greatest element), and appealing to Proposition 3.
+$m_{\text{op}}(\mathrm{S}_{l}):=\min\left\{m\in\mathbb{N}:\exists~\mathrm{op}_{1},\dots,\mathrm{op}_{m},\;\mathrm{op}_{1}\circ\dots\circ\mathrm{op}_{m}(\mathrm{S}_{l})=\emptyset\right\},\quad l=1,\dots,L,$ (6)
+where each $\mathrm{op}_{k}$ ($1\leq k\leq m$) is either $\operatorname{OP}_{(i,j),\text{CE}}$ or $\operatorname{OP}_{(i,j),\text{ER}}$ for some $(i,j)$ that can depend on $k$.
+Because the v-structures from two different sets of the partition have no common vertex by construction and this property is maintained by both operations, we conclude that
+$m_{\text{op}}(\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t}))=\sum_{l=1}^{L}m_{\text{op}}(\mathrm{S}_{l}),$
+which lower bounds the number of steps taken to align the v-structures between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$.
+Proposition 6 (CPDAG $d_{\mathcal{L}}$ bounds).
+For CPDAGs $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$, we have upper bound
+$d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq\min\left\{d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t}),d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t})\right\},$
+where $d_{\mathcal{L},\downarrow\uparrow},d_{\mathcal{L},\uparrow\downarrow}$ can be computed using Proposition˜3. Meanwhile, we have lower bound
+$d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\geq|\operatorname{sk}(\mathcal{G}_{s})\triangle\operatorname{sk}(\mathcal{G}_{t})|+2\,m_{\mathrm{op}}(\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t})).$
+The upper bound follows from the definition of $d_{\mathcal{L}}$ and the fact that both the down-up and up-down distances are well-defined for CPDAGs, as guaranteed by the existence of the least and the greatest element in its poset.
+We prove the lower bound in Section˜F.1.
 4.3.2 Restriction to polytrees
-We have established that the model-oriented distance for CPDAGs is inherently more complex than that for UGs or DAGs, reflecting both the increased structural complexity of the model-oriented poset and the richer graphical nature of CPDAGs themselves.
-We now study how the model-oriented distance may simplify if we restrict the class of CPDAGs we consider.
-Specifically, we focus on CPDAGs that are also polytrees, that is, CPDAGs with an acyclic skeleton.
-This class frequently arises in the causal inference literature (Rebane and Pearl, 1987; Sepehr and Materassi, 2019; Grüttemeier, Komusiewicz and Morawietz, 2021; Tramontano, Monod and Drton, 2022; Améndola et al., 2023)
-Let $\mathfrak{G}_{\text{poly}}$ be the collection of completed partially directed graphs over the node set $V$ that are polytrees.
-Here, the model-oriented distance between two graphs $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}_{\text{poly}}$ is the number of neighbor steps needed to transform $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$, where each step is a move to a CPDAG that is also a polytree.
-Formally, let $\mathcal{L}_{\text{poly}}=(\mathfrak{G}_{\text{poly}},\preceq)$ be the poset underlying $\mathfrak{G}_{\text{poly}}$ using the same ordering relationship as the unrestricted setting (this is a subposet of $\mathcal{L}$).
-Then, the model-oriented distance $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ can be characterized as follows.
-Theorem 3.
-For any pair of graphs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ in $\mathfrak{G}_{\text{poly}}$, we have $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})=d_{\mathcal{L}_{\text{poly}},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$.
-We prove this result in Appendix F.2. The proof proceeds by showing that $\mathcal{L}_{\text{poly}}$ is lower semimodular and then appeals to Theorem 2 to obtain the desired result.
-4.3.3 Relationship to Separation Distance
-There exist another distance for CPDAGs and probabilistic graphical model encoding only conditional independence statements more broadly, called separation distance (Wahl and Runge, 2025).
-It measures the distance between two graphs by counting the number of conditional independence statements encoded by only one of them.
-Similarly to our class of distances, the separation distance is a metric and model oriented, albeit on a different model interpretation.
-However, the separation distance framework implicitly assumes that any collection of conditional independence statements forms a valid conditional independence model.
-This is similar to how the SHD on CPDAGs implicitly assumes any simple graph with direct and undirected edges is a valid CPDAG (see Example 1).
-Conditional independence statements, however, may imply other such statements (Geiger, Verma and Pearl, 1989).
-In fact, determining which sets of triples correspond to a conditional independence model and which of these in turn correspond to a probabilistic graphical model is a challenging problem that has not yet been fully resolved (Sadeghi, 2017).
-There are two notable consequences of this idiosyncrasy.
-First, the separation distance between sparse graphs tends to be much larger than between dense graphs.
-Second, the separation distance between two graphs increases substantially when a disconnected node is added to both.
-We provide an example illustrating this and discuss the separation distance paradigm more thoroughly in Appendix H.2.
+We have seen that the model-oriented distance for probabilistic CPDAGs does not admit a simple, closed-form expression as opposed to probabilistic UGs (Corollary˜2) and causal DAGs (Corollary˜3).
+This reflects the complexity of the poset underlying CPDAGs.
+We will use the A* algorithm to compute the distance in Section˜5.
+In the rest of this subsection, we show that the distance takes a much simpler form if we restrict the graph class to polytrees, namely those CPDAGs whose skeletons do not contain cycles.
+Due to its relative tractability, polytrees have been studied in the causal discovery literature; see, e.g., Rebane and Pearl (1987); Sepehr and Materassi (2019); Grüttemeier, Komusiewicz and Morawietz (2021); Tramontano, Monod and Drton (2022); Améndola et al.
+(2023).
+Formally, let $\mathfrak{G}_{\mathrm{poly}}$ be the class of CPDAGs over a vertex set $\mathrm{V}$ that are polytrees and let $\mathcal{L}_{\mathrm{poly}}$ be the model-oriented poset associated with $(\mathfrak{G}_{\mathrm{poly}},\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}})$.
+The poset $\mathcal{L}_{\mathrm{poly}}$ is called the subposet of $\mathcal{L}$ induced by $\mathfrak{G}_{\mathrm{poly}}$.
+This subposet has the following properties, which we prove in Section˜E.3.
+Theorem 6 (Polytree CPDAG poset).
+Let $\mathcal{L}_{\mathrm{poly}}$ be the model-oriented poset associated with $(\mathfrak{G}_{\mathrm{poly}},\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}})$, where $\mathfrak{G}_{\mathrm{poly}}$ is the class of polytree CPDAGs over a finite vertex set $\mathrm{V}$.
+The following results hold.
+  1. 1.
+    $\mathcal{G}_{\hat{0}}$ is the empty graph. There is no $\mathcal{G}_{\hat{1}}$ when $|\mathrm{V}|>2$.
+  2. 2.
+    For $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}_{\mathrm{poly}}$, we have $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ if and only if there exist DAGs $\mathcal{D}_{1}\in[\mathcal{G}_{1}]$, $\mathcal{D}_{2}\in[\mathcal{G}_{2}]$ such that $\mathcal{D}_{1}$ is a subgraph of $\mathcal{D}_{2}$.
+  3. 3.
+    $\mathcal{L}_{\mathrm{poly}}$ is graded with $\operatorname{\mathrm{rank}}(\cdot)$ equal to the number of edges.
+  4. 4.
+    $\mathcal{L}_{\mathrm{poly}}$ is lower semimodular.
+  5. 5.
+    $d_{\mathcal{L}_{\mathrm{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})=d_{\mathcal{L}_{\mathrm{poly}},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$ holds for every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}_{\mathrm{poly}}$.
+Importantly, the last result in this theorem says that we can restrict to down–up paths when finding the model-oriented distance, leading to substantial computational savings.
+We conclude this subsection on CPDAGs with the following remark.
+Remark 3 (Separation distance).
+[…]
 4.4 Causal MPDAGs
-Suppose the graph class $\mathfrak{G}$ is the set of MPDAGs over vertex set $V$ with $\mathcal{M}:=\mathcal{M}_{\text{mpdag}}$ as the model map.
-The following result provides a graphical characterization for model containment.
+In this subsection, we consider $\mathfrak{G}$ as the set of MPDAGs over a finite vertex set $\mathrm{V}$ equipped with the model map $\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}$.
+Recall from Section˜2.2.5 that an MPDAG $\mathcal{G}$ represents a subset of Markov equivalent (but causally distinct) DAGs that satisfy certain background knowledge; we use $\mathcal{G}]$ to denote the collection of DAGs that $\mathcal{G}$ represents.
+The partial order in this context is characterized graphically by the following result, which we prove in Section˜A.2; see also Lemma˜2 for comparison.
+Proposition 7 (MPDAG comparability).
+Consider two MPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$.
+We have $\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}(\mathcal{G}_{s})\subseteq\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}(\mathcal{G}_{t})$ if and only if for every DAG $\mathcal{D}_{s}\in[\mathcal{G}_{s}]$, there exists a DAG $\mathcal{D}_{t}\in[\mathcal{G}_{t}]$ such that $\mathcal{D}_{s}$ is a subgraph of $\mathcal{D}_{t}$.
+Proposition˜7 is central to studying the model-oriented poset $\mathcal{L}$ of MPDAGs.
+Fig.˜3(d) shows the corresponding Hasse diagram when $|\mathrm{V}|=3$.
+Table˜2 summarizes its structural properties.
+In particular, the least element $\mathcal{G}_{\hat{0}}$ is the empty graph and the greatest element $\mathcal{G}_{\hat{1}}$ is the fully connected undirected graph.
+However, unlike UGs, DAGs and CPDAGs, the MPDAG poset does not admit the total number of edges as the rank function, as illustrated by the next example.
+Example 2.
+Consider the MPDAGs in Fig.˜6 over $n$ vertices.
+One can use Proposition˜9 to show $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$.
+It also holds that $\mathcal{G}_{1}\lessdot\mathcal{G}_{3}$.
+To see this, first observe that $\mathcal{G}_{1}\prec\mathcal{G}_{3}$ by Proposition˜7; then, note that there cannot be a graph $\mathcal{G}$ that satisfies $\mathcal{G}_{1}\prec\mathcal{G}\prec\mathcal{G}_{3}$ because $\mathcal{G}$ must result from removing one or more directed edge from $\mathcal{G}_{3}$, which would introduce a forbidden structure to the MPDAG (see Fig.˜2).
+Hence, despite both being a neighbor of $\mathcal{G}_{1}$, the graph $\mathcal{G}_{2}$ has one more edge than $\mathcal{G}_{1}$, while the graph $\mathcal{G}_{3}$ has n1n-1 more edges than $\mathcal{G}_{1}$.
+Figure 6: MPDAGs $\mathcal{G}_{1},\mathcal{G}_{2},\mathcal{G}_{3}$ considered in Example˜2.
+Furthermore, the following result establishes that the poset does not admit any rank function; see Section˜E.4 for its proof.
+Proposition 8 (MPDAG non-gradedness).
+Let $\mathfrak{G}$ be the class of MPDAGs over a finite vertex set $\mathrm{V}$ with $|\mathrm{V}|\geq 4$.
+The model-oriented poset of $(\mathfrak{G},\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}})$ is not graded.
+In addition to the lack of gradedness in $\mathcal{L}$, we do not yet have an explicit graphical characterization of the neighborhood (i.e., the covering relation); indeed, as we have seen from Example˜2, such a characterization can be technically challenging because the structures of neighboring graphs can vary greatly.
+Hence, it remains very difficult to enumerate neighbors, which is an essential component of any search algorithm.
+Moreover, the down-up and up-down distances are less useful in the context of MPDAGs.
+First, $\mathcal{L}$ is neither upper nor lower semimodular (see Section˜E.4), so $d_{\mathcal{L},\downarrow\uparrow}$ and $d_{\mathcal{L},\uparrow\downarrow}$ are merely upper bounds of $d_{\mathcal{L}}$ (see Theorem˜2).
+Second, the lack of gradedness and semi-lattice structures (see Section˜E.4) prevents us from reducing these distances to a more computable form (see Propositions˜3 and 1).
+To remedy these challenges, we introduce the following notion of pseudo-rank.
+Definition 8 (MPDAG pseudo-rank).
+The pseudo-rank of an MPDAG $\mathcal{G}$ is defined as
+$\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}):=\text{number of directed edges}+\text{number of undirected edges}\times 2.$
+This pseudo-rank serves as a surrogate for the rank in a non-graded poset and reflects an intuitive measure of structural complexity, as shown by the next two results.
+Proposition 9 (MPDAG sufficient condition for covering).
+For MPDAGs $\mathcal{G}_{1},\mathcal{G}_{2}$ such that $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$, we have $\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{1})\leq\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{2})$, where the equality holds if and only if $\mathcal{G}_{1}=\mathcal{G}_{2}$.
+Additionally, if $\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{1})+1=\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{2})$, we have $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$.
+Proof.
+Given $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$, from Proposition˜7 it follows that every $v_{1}\rightarrow v_{2}$ present in $\mathcal{G}_{1}$ corresponds to either $v_{1}\rightarrow v_{2}$ or $v_{1}-v_{2}$ in $\mathcal{G}_{2}$, and every $v_{3}-v_{4}$ present in $\mathcal{G}_{1}$ corresponds to $v_{3}-v_{4}$ in $\mathcal{G}_{2}$.
+The first statement then follows from definition.
+Further, suppose $\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{1})+1=\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{2})$ but there exists $\mathcal{G}$ that satisfies $\mathcal{G}_{1}\prec\mathcal{G}\prec\mathcal{G}_{2}$.
+The skeletons of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are either identical or different by exactly one edge, which we discuss in three cases: (i) If their skeletons coincide, then $\mathcal{G}_{2}$ must have exactly one undirected edge that is directed in $\mathcal{G}_{1}$, so the intermediate $\mathcal{G}$ cannot exist.
+(ii) If $\operatorname{sk}(\mathcal{G}_{1})=\operatorname{sk}(\mathcal{G})$ has one fewer edge than $\operatorname{sk}(\mathcal{G}_{2})$, then the extra edge in $\mathcal{G}_{2}$ must be directed and all the other edges in $\mathcal{G}_{2}$ are identical to those in $\mathcal{G}_{1}$.
+Because $\mathcal{G}_{1}\prec\mathcal{G}$, $\mathcal{G}$ must have an undirected edge that corresponds to a directed edge in $\mathcal{G}_{1}$ and hence in $\mathcal{G}_{2}$, but this contradicts $\mathcal{G}\prec\mathcal{G}_{2}$.
+(iii) Similarly, we also have contradiction if $\operatorname{sk}(\mathcal{G})=\operatorname{sk}(\mathcal{G}_{2})$ has one more edge than $\operatorname{sk}(\mathcal{G}_{1})$.
+Therefore, by ruling out the intermediate $\mathcal{G}$, we conclude $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$.
+∎
 Proposition 10.
-Let $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ be two MPDAGs.
-Then, $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$ if and only if for every DAG $\mathcal{D}_{1}$ represented by $\mathcal{G}_{1}$, there exists a DAG $\mathcal{D}_{2}$ represented by $\mathcal{G}_{2}$ such that $\mathcal{D}_{1}$ is a directed subgraph of $\mathcal{D}_{2}$.
-We prove Proposition 9 in Section˜A.2.
-Using this result, we can characterize the poset $\mathcal{L}=(P,\preceq)$ underlying MPDAGs.
-This poset consists of MPDAGs, ordered according to subgraph containment: graphs $\mathcal{G}_{1},\mathcal{G}_{2}\in L$ satisfy $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ if for every DAG $\mathcal{D}_{1}$ in $\mathcal{G}_{1}$, there exists a DAG $\mathcal{D}_{2}$ such that $\mathcal{D}_{1}$ is a subgraph of $\mathcal{D}_{2}$.
-The least element is the empty graph.
-The poset has a greatest element that is the fully connected undirected graph.
-Further, it is not meet or join semi-lattice nor is lower or upper semimodular; see Appendix E.3 for counter examples.
-Fig.˜3(d) shows the Hasse diagram associated with this poset.
-With this poset $d_{\mathcal{L}}$, we have a formal definition for model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ between two MPDAGs.
-As with CPDAGs, the model‑oriented distance differs from the two variants of SHD.
-As a simple illustration, suppose $\mathcal{G}_{s}$ is an MPDAG among three nodes $\{v_{1},v_{2},v_{3}\}$ with two edges $v_{1}\to v_{2}$ and $v_{3}\to v_{2}$, and $\mathcal{G}_{t}$ is an MPDAG among the same node set containing two edges $v_{1}-v_{2}$ and $v_{3}-v_{2}$.
-Both variants of the structural Hamming distance give the value two.
-However, these measures implicitly assume that the graph with edge set $v_{1}-v_{2}$ and $v_{3}\to v_{2}$ is a valid MPDAG, which is incorrect — highlighting that the structural Hamming distance fails to account for model structure.
-In fact, the model-oriented distance between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ in this case is four.
-Given the importance of gradedness for computing and deriving simplified characterizations of $d_{\mathcal{L}}$, we study whether the poset $\mathcal{L}$ is graded, that is, whether there exists a function $\mathrm{rank}(\cdot)$ assigning non-negative integers to graphs, such that each covering relation corresponds to a rank increase of exactly one.
-In previous graph classes, such a rank function was defined by the number of edges, so neighboring graphs always differed by exactly one edge.
-In the context of MPDAGs, however, a neighbor of a graph may contain many more edges.
-As an example, consider the graph $\mathcal{G}_{1}$ in Fig.˜6(a).
-Both graphs $\mathcal{G}_{2}$ and $\mathcal{G}_{3}$ in panels (b) and (c), respectively, satisfy $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$ and $\mathcal{G}_{1}\lessdot\mathcal{G}_{3}$.
-The neighbor $\mathcal{G}_{2}$ contains one additional edge compared to $\mathcal{G}_{1}$, whereas $\mathcal{G}_{3}$ contains n1n-1 additional edges.
-Note that $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$ by Proposition˜11.
-To see why $\mathcal{G}_{1}\lessdot\mathcal{G}_{3}$, note that any graph $\mathcal{G}$ satisfying $\mathcal{G}\preceq\mathcal{G}_{3}$, we must either direct undirected edges or drop edges.
-However, for $\mathcal{G}$ to also satisfy $\mathcal{G}_{1}\preceq\mathcal{G}$, none of the undirected edges in $\mathcal{G}_{3}$ can be modified in any way.
-Finally, dropping any subset of directed edges in $\mathcal{G}_{3}$ results in an invalid MPDAG (see Fig.˜2(b)).
-This example rules out the number of edges as a valid rank function.
-However, it does not rule out the existence of some other valid rank function.
-The following proposition formally establishes that no such function exists, and thus the poset $\mathcal{L}$ is not graded.
-Proposition 10.
-There does not exists a function $\mathrm{rank}(\cdot)$ assigning non-negative integers to graphs in $\mathfrak{G}$, such that each covering relation corresponds to a rank increase of exactly one.
-Thus, the poset $\mathcal{L}$ is not graded.
-We prove Proposition 10 in Appendix G.1.
-It constructs a pair of MPDAGs $\mathcal{G}_{s},\mathcal{G}_{t}$ with $\mathcal{G}_{s}\preceq\mathcal{G}_{t}$, where there are maximal chains from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ with different length.
-(Recall that a maximal chain in $\mathcal{L}$ is a sequence of graphs from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ in which every graph covers the proceeding one.) This difference in chain lengths contradicts the gradedness property, since every covering step increases the rank by exactly one, so any two maximal chains between the same endpoints would have to have equal length.
-In addition to a lack of graded structure in $\mathcal{L}$, we do not have an explicit graphical characterization of neighbor structure.
-This makes it difficult to enumerate neighbors and compute $d_{\mathcal{L}}$ or the down-up and up-down distances $d_{\mathcal{L},\downarrow\uparrow}$ and $d_{\mathcal{L},\uparrow\downarrow}$.
-To address these challenges, we introduce a pseudo-rank function.
-Definition 8 (Pseudo-rank).
-For any MPDAG $\mathcal{G}\in\mathfrak{G}$, we define its pseudo-rank, denoted $\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G})$, by the the number of directed edges in $\mathcal{G}$ plus twice the number of undirected edges.
-The pseudo-rank function serves as a surrogate for rank in a non-graded poset (see Proposition˜11) and reflects an intuitive measure of structural complexity (see Proposition˜12).
-Proposition 11 (Sufficient condition for covering).
-Consider $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ with $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$.
-If $\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G}_{1})+1=\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G}_{2})$, then, $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$.
+(MPDAG model complexity) Between any MPDAG $\mathcal{G}$ and the empty graph $\mathcal{G}_{\hat{0}}$, there exists a maximal chain (i.e., of the form $\mathcal{G}_{\hat{0}}=:\mathcal{G}_{0}\lessdot\cdots\lessdot\mathcal{G}_{l}:=\mathcal{G}$) of length $\operatorname{\mathrm{pseudo-rank}}(\mathcal{G})$.
+This result is a consequence of the following lemma, which may be of independent interest.
+Lemma 3 (MPDAG single-edge orientation).
+Let $\mathcal{G}$ be any MPDAG with one or more undirected edges.
+Then, there exists an undirected edge in $\mathcal{G}$ such that by directing it and leaving all the other edges unchanged, we obtain another MPDAG $\tilde{\mathcal{G}}$.
+It follows that we can turn $\mathcal{G}$ into a directed graph by orienting the undirected edges one by one, and then into the empty graph by removing the directed edges one by one; see Section˜G.1 for a constructive proof of the previous two results.
+Based on Proposition˜10, we define $\mathrm{P}_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ as the subset of paths between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ such that every consecutive pair of graphs on the path differ by unit pseudo-rank, namely
+$\mathrm{P}_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t}):=\big\{(\mathcal{G}_{s}=:\mathcal{G}_{0},\dots,\mathcal{G}_{d}:=\mathcal{G}_{t})\in\mathrm{P}_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t}):\;\text{for $i=1,\dots,d$, }\\ |\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{i})-\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{i-1})|=1\big\}.$
+In other words, $\mathrm{P}_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ only consists of paths through restricted neighborhoods, i.e., those neighboring graphs that differ by exactly one edge between a certain pair of vertices $a,b$: the difference is either aba~~~b (no edge) versus $a\rightarrow b$, or $a\rightarrow b$ versus aba-b.
+We then define the pseudo-distance as
+$d_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t}):=\min\{\mathrm{length}(p):p\in\mathrm{P}_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t})\},\quad\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}.$ (7)
+Corollary 4.
+It holds that $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq d_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ for every $\mathcal{G}_{s},\mathcal{G}_{t}$.
 Proof.
-Suppose $\mathcal{G}\in\mathfrak{G}$ satisfies $\mathcal{G}_{1}\preceq\mathcal{G}\preceq\mathcal{G}_{2}$ and differs from both.
-By the pseudo-rank condition, the skeletons of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are either identical or differ by one edge.
-If their skeletons coincide, then $\mathcal{G}_{2}$ has exactly one undirected edge that is directed in $\mathcal{G}_{1}$, so the intermediate $\mathcal{G}$ would need to both orient an undirected edge of $\mathcal{G}_{1}$ and undirect an edge of $\mathcal{G}_{2}$, which is impossible.
-If instead $\mathcal{G}$ has the skeleton of $\mathcal{G}_{1}$, which differs from that of $\mathcal{G}_{2}$ by one edge, then $\mathcal{G}_{2}$ contains one directed edge that does not exist in $\mathcal{G}_{1}$, while all other edges are shared; however, $\mathcal{G}$ must contain at least one undirected edge that is directed in $\mathcal{G}_{1}$ and can therefore not satisfy $\mathcal{G}\preceq\mathcal{G}_{2}$.
-The symmetric case, where $\mathcal{G}$ has the skeleton of $\mathcal{G}_{2}$ which differs from that of $\mathcal{G}_{1}$, follows analogously.
-Proposition 12.
-(MPDAG model complexity) For any MPDAG $\mathcal{G}\in\mathfrak{G}$ there exists a sequence $\mathcal{G}=x_{1}\gtrdot\dots\gtrdot x_{k}=\mathcal{G}_{\hat{0}}$ of MPDAGs, where $G_{\hat{0}}$ is the empty MPDAG and $k=\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G})$.
-We prove Proposition 12 in Appendix G.2.
-Using this pseudo-rank function, we define $P_{\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ as the set of paths $x_{0}:=\mathcal{G}_{s},x_{1},\dots,x_{r}:=\mathcal{G}_{t}$ in the Hasse diagram corresponding to $\mathcal{L}$, such that each step satisfies $|\mathrm{pseudo}\text{-}\mathrm{rank}(x_{i})-\mathrm{pseudo}\text{-}\mathrm{rank}(x_{i-1})|=1$ for all $i$.
-In other words, $P_{\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})\subseteq P(\mathcal{G}_{s},\mathcal{G}_{t})$, where relative to $P(\mathcal{G}_{s},\mathcal{G}_{t})$, $P_{\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ has the added constraint that the pseudo-rank changes by exactly one at each step in the path.
-This ensures that each transition $(x_{i-1},x_{i})$ between neighbor graphs corresponds to a minimal modification: either an undirected edge in one graph becomes directed in the other, or a directed edge in one graph is absent in the other.
-This constraint mimics graded posets, where neighbors differ by exactly one in rank, helping recover some useful structure in the non-graded setting of $\mathcal{L}$.
-By Proposition 12, $P_{\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is a non-empty set.
-Since $P_{\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})\subseteq P(\mathcal{G}_{s},\mathcal{G}_{t})$, we can therefore define a distance based on $P_{\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ that upper bounds $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ formalized as follows.
-Corollary 4.
-Let $d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t}):=\min\{\allowbreak\mathrm{length}(p):p\in L_{\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})\}$, i.e., the length of the shortest path between $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$, among paths that lie inside $P_{\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
-Then, we have $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
-Corollary 4 has practical implications: the pseudo-distance $d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is often significantly easier to compute than either the model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ or the down-up and up-down distances $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$ and $d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t})$.
-In particular, recall that computing the model-oriented distance or the down-up and up-down distances requires enumerating neighbor graphs.
-However, as demonstrated in the example in Fig.˜6, the structure in neighbor graphs can vary widely, and no simple graphical characterization of neighborhood is known.
-In contrast, computing $d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ only requires considering neighbors where the pseudo-rank changes by exactly one.
-This makes the enumeration process much simpler: for a graph $\mathcal{G}$, the relevant neighbors can be obtained by adding or removing any undirected edge, directing an undirected edge, or undirecting a directed edge — provided the resulting graph is a valid MPDAG in that it does not contain any induced subgraphs in Fig.˜2(b).
-It may be tempting to claim $d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is equal to $\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})$, the second variant of SHD between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$.
-However, in the simple example among three nodes from earlier, $d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is four, while $\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})=2$ is two.
-In fact, $\operatorname{\mathrm{SHD}}_{2}$ and consequently $\operatorname{\mathrm{SHD}}_{1}$ both lower bound $d_{\mathcal{L},\text{pseudo}}$ as formalized next.
-Proposition 13.
-For any pair of MPDAGs $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, both variants of the SHD between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ is smaller than or equal to $d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{1},\mathcal{G}_{2})$; that is, $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{1},\mathcal{G}_{2})\leq\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{1},\mathcal{G}_{2})\leq d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{1},\mathcal{G}_{2})$.
-We prove Proposition 13 in Appendix G.3.
-The bound in this proposition is used in the algorithm for computing $d_{\mathcal{L},\text{pseudo}}$; see Section 5.3.
-Note that while the SHD variants lower-bound $d_{\mathcal{L},\text{pseudo}}$, they do not always lower bound the model-oriented distance $d_{\mathcal{L}}$.
-For example, let $\mathcal{G}_{1}$ be the graph in Fig.˜6(a) and $\mathcal{G}_{3}$ be the graph in Fig.˜6(c).
-Then $d_{\mathcal{L}}(\mathcal{G}_{1},\mathcal{G}_{3})=1$, while both SHD variants return n1n-1.
-Figure 6: An MPDAG $\mathcal{G}_{1}$ in (a) and two of its neighboring MPDAGs in (b) and (c). The neighbor in (c) has n1n-1 more edges than $\mathcal{G}_{1}$.
-4.4.1 Restriction to polytrees for MPDAGs
-Similar to the case of CPDAGs, we can restrict our attention to a subclass of MPDAGs that are polytrees.
-Let $\mathfrak{G}_{\text{poly}}$ be the collection of MPDAGs over node set $V$ that are polytrees.
-Define the subposet $\mathcal{L}_{\text{poly}}=(\mathfrak{G}_{\text{poly}},\preceq)$ as the restriction of $\mathcal{L}$ to this subclass and let $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ denote the corresponding model-oriented distance.
-Let $d_{\mathcal{L}_{\text{poly}},\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ be the analog of $d_{\mathcal{L},\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$, but defined on the subposet $\mathcal{L}_{\text{poly}}$.
-Using a similar argument as Theorem 4, we can conclude that $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq d_{\mathcal{L}_{\text{poly}},\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
-In fact, this inequality is tight: equality holds because the poset $\mathcal{L}_{\text{poly}}$ is graded with pseudo-rank as its ranking function.
-We formalize this next.
-Theorem 4.
-The poset $\mathcal{L}_{\text{poly}}$ is graded with rank function $\mathrm{pseudo}\text{-}\mathrm{rank}(\cdot)$.
-Thus, the model-oriented distance $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is equal to the pseudo-distance $d_{\mathcal{L}_{\text{poly}},\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
-We prove Theorem 4 in Appendix G.4.
-This result confirms that in the polytree setting, the pseudo-rank function defines a proper grading on $\mathcal{L}_{\text{poly}}$, and the corresponding distance exactly matches the model-oriented distance.
-Theorem 4 provides concrete justification for using the pseudo-rank-based distance more broadly.
-Recall that when restricting CPDAGs to polytrees, the corresponding poset is lower semimodular, allowing us to appeal to Theorem 2 to conclude that the model-oriented distance equals the down-up distance.
-However, we show in Appendix E.3 that the poset obtained by restricting MPDAGs to polytrees is not lower semimodular.
-Thus, a similar simplification is not possible and we need to rely on the $A^{*}$-algorithm to compute $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ instead.
+By Proposition˜10, given any $\mathcal{G}_{s},\mathcal{G}_{t}$, the set $\mathrm{P}_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is non-empty because it contains the path that concatenates two maximal chains: first from $\mathcal{G}_{s}$ to $\mathcal{G}_{\hat{0}}$, then from $\mathcal{G}_{\hat{0}}$ to $\mathcal{G}_{t}$.
+The result then follows by definition.
+∎
+Due to the ease of enumerating graphs in the restricted neighborhood, $d_{\mathcal{L},\operatorname{\mathrm{pseudo}}}$ is easier to compute than $_{\mathcal{L}}, d,d_{\mathcal{L},\downarrow\uparrow}$ or $d_{\mathcal{L},\uparrow\downarrow}$.
+We also compare it with the structural Hamming distances in the next result.
+Proposition 11 (MPDAG distance to SHD relationship).
+For any MPDAGs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, it holds that $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{1},\mathcal{G}_{2})\leq\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{1},\mathcal{G}_{2})\leq d_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{1},\mathcal{G}_{2})$.
+We prove this result in Section˜G.2.
+See also Fig.˜S14 in the Appendix that compares these distances computed for polytree MPDAGs over 5 vertices.
+In Section˜5.3, we will use $\operatorname{\mathrm{SHD}}_{2}$ as a lower bound in the A* algorithm for computing $d_{\mathcal{L},\operatorname{\mathrm{pseudo}}}$.
+Note that the structural Hamming distances do not lower bound the original $d_{\mathcal{L}}$.
+In the example of Fig.˜6, we have $d_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{1},\mathcal{G}_{3})=1$ but $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{1},\mathcal{G}_{3})=\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{1},\mathcal{G}_{3})=n-1$.
+4.4.1 Restriction to polytrees
+We consider again restricting the class of graphs to $\mathfrak{G}_{\mathrm{poly}}$, which in this context denotes all MPDAGs that are polytrees over a vertex set $\mathrm{V}$.
+This induces the subposet $\mathcal{L}_{\mathrm{poly}}=(\mathfrak{G}_{\mathrm{poly}},\preceq)$ and further the corresponding shortest-path distance $d_{\mathcal{L}_{\mathrm{poly}}}$ between every pair of graphs in $\mathfrak{G}_{\mathrm{poly}}$.
+Furthermore, let $d_{\mathcal{L}_{\mathrm{poly}},\operatorname{\mathrm{pseudo}}}$ be the analogue of $d_{\mathcal{L},\operatorname{\mathrm{pseudo}}}$ in Eq.˜7 but defined on the subposet $\mathcal{L}_{\mathrm{poly}}$ instead of $\mathcal{L}$.
+By an argument similar to the proof of Corollary˜4 (there still exists a path by concatenating two maximal chains of polytrees), we have $d_{\mathcal{L}_{\mathrm{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq d_{\mathcal{L}_{\mathrm{poly}},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ for every polytree $\mathcal{G}_{s},\mathcal{G}_{t}$.
+In fact, the result below shows that this inequality is an equality; see Section˜E.5 for its proof.
+Theorem 7 (Polytree MPDAG poset).
+Let $\mathcal{L}_{\mathrm{poly}}$ be the model-oriented poset associated with $(\mathfrak{G}_{\mathrm{poly}},\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}})$, where $\mathfrak{G}_{\mathrm{poly}}$ is the class of polytree MPDAGs over a finite vertex set $\mathrm{V}$.
+We have the following results.
+  1. 1.
+    $\mathcal{G}_{\hat{0}}$ is the empty graph. There is no $\mathcal{G}_{\hat{1}}$ when $|\mathrm{V}|>2$.
+  2. 2.
+    $\mathcal{L}_{\mathrm{poly}}$ is graded with rank function equal to $\operatorname{\mathrm{pseudo-rank}}(\cdot)$.
+  3. 3.
+    $\mathcal{L}_{\mathrm{poly}}$ is lower semimodular.
+  4. 4.
+    It holds that
+    $d_{\mathcal{L}_{\mathrm{poly}},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})=d_{\mathcal{L}_{\mathrm{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})=d_{\mathcal{L}_{\mathrm{poly}},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t}),\quad\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}_{\mathrm{poly}}.$
+  5. 5.
+    It holds that
+    $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})\leq\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})\leq d_{\mathcal{L}_{\mathrm{poly}},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t}),\quad\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}_{\mathrm{poly}}.$
 5 Algorithms for computing model-oriented distances
-In this section, we develop algorithms for either computing or approximating the model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ for a pair of graphs $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$ from different graphical model classes $\mathfrak{G}$, including CPDAGs and MPDAGs.
-In Section 5.1, we present a general-purpose A* procedure for finding $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ across different graph classes.
-In Sections 5.2 and 5.3, we specialize this algorithm to the class of CPDAGs and MPDAGs, respectively, and present numerical experiments comparing the model-oriented distance to the two SHD variants $\operatorname{\mathrm{SHD}}_{1}$ and $\operatorname{\mathrm{SHD}}_{2}$ while tracking the computational cost of our algorithms.
-5.1 General A* procedure
-To compute the model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ between two graphs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$, we use an A* search algorithm, detailed in Algorithm 1.
-The search is performed on the poset $\mathcal{L}$.
-The algorithm maintains a priority queue (openSet) initialized with the start graph $\mathcal{G}_{s}$.
-This queue is ordered by a costs function $f(\mathcal{G})=g(\mathcal{G})+h(\mathcal{G})$, where: (i) $g(\mathcal{G})$ tracks the exact costs (i.e., number of steps) from $\mathcal{G}_{s}$ to $\mathcal{G}$, and (ii) $h(\mathcal{G})$ is a heuristic function that estimates the remaining cost from $\mathcal{G}$ to $\mathcal{G}_{t}$.
-To ensure that A* returns the distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$, the heuristic $h$ must be admissible:
-Definition 9 (Admissible heuristic).
-A heuristic $h(\cdot)$ is admissible if $h(\mathcal{G})\leq d_{\mathcal{L}}(\mathcal{G},\mathcal{G}_{t})$ for all $\mathcal{G}\in\mathfrak{G}$, i.e., it never overestimates the true distance.
-We provide some more detail about Algorithm 1.
-At each iteration, the graph $\mathcal{G}_{\text{current}}$ with the smallest $f$-score is removed from the queue and expanded.
-If $\mathcal{G}_{\text{current}}=\mathcal{G}_{t}$, the algorithm terminates and returns $g(\mathcal{G}_{t})$.
-Otherwise, the algorithm examines all neighbors of $\mathcal{G}_{\text{current}}$ in $\mathcal{L}$.
-Here, the function $\mathrm{n}_{\mathcal{L}}(\cdot)$ finds neighboring graphs in the poset $\mathcal{L}$.
-For each neighbor $\mathcal{G}_{\text{neighbor}}$, it computes a tentative distance via $\mathcal{G}_{\text{current}}$.
-If this tentative distance is shorter than the path previously recorded to $\mathcal{G}_{\text{neighbor}}$, the algorithm updates $g(\mathcal{G}_{\text{neighbor}})$ and adds it to the queue with priority $g(\mathcal{G}_{\text{neighbor}})+h(\mathcal{G}_{\text{neighbor}})$.
-This A* framework is broadly applicable to any graph class $\mathfrak{G}$ that can be equipped with a connected poset $\mathcal{L}$.
-To use the algorithm in a given setting, one must provide: (i) an admissible heuristic that underestimates $d_{\mathcal{L}}(\mathcal{G}_{1},\mathcal{G}_{2})$ for all pairs $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, and (ii) a method for enumerating all neighbors of a given graph $\mathcal{G}\in\mathfrak{G}$, i.e., the graphs that cover or are covered by $\mathcal{G}$ in the poset.
-input: Graph class $\mathfrak{G}$, pair of graphs $\mathcal{G}_{s},\mathcal{G}_{t}$, heuristic $h\colon\mathfrak{G}\to\mathbb{R}_{\geq 0}$, neighbor function $\mathrm{n}_{\mathcal{L}}\colon\mathfrak{G}\to 2^{\mathfrak{G}}$
-output : Model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$
-1 begin
- $openSet\coloneqq\{\mathcal{G}_{s}\}$;
- // priority queue ordered by $g+h$
-2 $g(\mathcal{G}_{s})\coloneqq 0$;
+In this section, we develop algorithms for computing the model-based distance.
+We will first present a generic A* algorithm suited to any graph class.
+Then, we will specialize the algorithm to probabilistic CPDAGs and causal, polytree MPDAGs in a manner that exploits their respective poset properties to save computation.
+Numerical results will be presented to illustrate computation time and to provide a comparison with the structural Hamming distance.
+5.1 Generic A* algorithm with branch and bound
+A* (Hart, Nilsson and Raphael, 1968) is a widely-used search algorithm to find the shortest path between two vertices on a graph.
+By employing a heuristic function to guide the search, it is typically more efficient than naive breadth-first or depth-first search.
+Suppose we want to find the distance between $\mathcal{G}_{s}$ (source) and $\mathcal{G}_{t}$ (target).
+Starting from $\mathcal{G}_{s}$, the algorithm maintains a priority queue (known as the open set) for the graphs to be visited, sorted by the index $f(\mathcal{G}):=g(\mathcal{G})+h(\mathcal{G})$ for each graph $\mathcal{G}$ in the queue.
+Here, $g(\mathcal{G})$ is the distance from $\mathcal{G}_{s}$ to $\mathcal{G}$ and $h(\mathcal{G})$ is the heuristic that estimates from below the distance from $\mathcal{G}$ to $\mathcal{G}_{t}$.
+For each iteration, a graph with the smallest index is popped from the queue and its neighbors are pushed to the queue, upon which their $f$ and $g$ values are updated if smaller values are found.
+The algorithm terminates when $\mathcal{G}_{t}$ is popped from the queue and returns $g(\mathcal{G}_{t})$ as the distance between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$.
+When $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ are connected, A* is guaranteed to terminate.
+It can be shown that A* outputs the correct distance if the heuristic is admissible in the sense that $0\leq h(\mathcal{G})\leq d_{\mathcal{L}}(\mathcal{G},\mathcal{G}_{t})$ for every $\mathcal{G}\in\mathfrak{G}$.
+Depending on how the queue is implemented, A* reduces to breadth-first or depth-first search if a trivial heuristic $h\equiv 0$ is supplied.
+By using an admissible heuristic with smaller $d_{\mathcal{L}}(\mathcal{G},\mathcal{G}_{t})-h(\mathcal{G})$, A* becomes more efficient.
+To further save computation with branch and bound, we also employ an upper bound $u(\mathcal{G})$ on the distance from $\mathcal{G}$ to $\mathcal{G}_{t}$ ($u\equiv+\infty$ is a trivial upper bound).
+By using $u(\cdot)$ to maintain $u^{\ast}$, the smallest upper bound on the distance between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ witnessed so far, the algorithm can terminate early whenever $f(\mathcal{G})\geq u^{\ast}$ due to the relation $f(\mathcal{G})\leq d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq u^{\ast}$.
+Moreover, when expanding the neighborhood of $\mathcal{G}$, those graphs with $f$ no smaller than $u^{\ast}$ can be pruned immediately.
+Naturally, one can use $_{\mathcal{L},\uparrow\downarrow}, d,d_{\mathcal{L},\downarrow\uparrow}$ or their minimum as the upper bound.
+We present this generic A* in Algorithm˜1, which requires three subroutines: lower bound $h(\cdot)$, upper bound $u(\cdot)$ and $\textsc{EnumNeighbors}()$ that enumerates the neighborhood of a graph in the model-oriented poset.
+We establish the correctness of the algorithm in the next theorem, which we prove in Section˜I.1.
+Theorem 8 (Algorithm correctness).
+Suppose $\mathfrak{G}$ is a finite graph class and the associated model-oriented poset $\mathcal{L}=(\mathfrak{G},\preceq)$ is connected.
+For Algorithm˜1, suppose the subroutines $\textsc{EnumNeighbors}(\cdot)$, $u(\cdot)$ and $h(\cdot)$ satisfy the conditions specified at the beginning of the algorithm.
+Then, the algorithm will terminate and return $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
+Input: Graphs $\mathcal{G}_{s},\mathcal{G}_{t}$ in a finite graph class $\mathfrak{G}$ with model-oriented poset $\mathcal{L}=(\mathfrak{G},\preceq)$.
+Require : Subroutine $\textnormal{{EnumNeighbors}}(\mathcal{G})$ that outputs the neighbors of $\mathcal{G}$ in $\mathcal{L}$;
+subroutine $h(\mathcal{G})$ such that $0\leq h(\mathcal{G})\leq d_{\mathcal{L}}(\mathcal{G},\mathcal{G}_{t})$ for every $\mathcal{G}\in\mathfrak{G}$;
+subroutine $u(\mathcal{G})$ such that $d_{\mathcal{L}}(\mathcal{G},\mathcal{G}_{t})\leq u(\mathcal{G})\leq+\infty$ for every $\mathcal{G}\in\mathfrak{G}$.
+Output: Model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
+1
+[…]
+// Upper bound on $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$
+$g\leftarrow\textnormal{{Dict}}[]$
+// Length of the shortest path from $\mathcal{G}_{s}$ found so far
+2 $g[\mathcal{G}_{s}]\leftarrow 0$
+$\textnormal{{openSet}}\leftarrow\textnormal{{PriorityQueue}}\left((\mathcal{G}_{s},0),h(\mathcal{G}_{s})\right)$
+// Item $(\mathcal{G}_{s},g[\mathcal{G}_{s}])$ with index $g[\mathcal{G}_{s}]+h(\mathcal{G}_{s})$
 3
-4 while $openSet\neq\emptyset$ do
-5    $\mathcal{G}_{\text{current}}\coloneqq$ graph in openSetopenSet minimizing $g(\mathcal{G}_{\text{current}})+h(\mathcal{G}_{\text{current}})$;
-6    if $\mathcal{G}_{\text{current}}=\mathcal{G}_{t}$ then
-7       return $g(\mathcal{G}_{t})$
-8     remove $\mathcal{G}_{\text{current}}$ from openSetopenSet;
-    // Neighbors according to poset $\mathcal{L}$
-9    foreach $\mathcal{G}_{\text{neighbor}}\in\mathrm{n}_{\mathcal{L}}(\mathcal{G}_{\text{current}}$) do
-10       $tentative\coloneqq g(\mathcal{G}_{\text{current}})+1$;
-11       if $g(\mathcal{G}_{\text{neighbor}})$ not defined   or   $tentative<g(\mathcal{G}_{\text{neighbor}})$ then
-12          $g(\mathcal{G}_{\text{neighbor}})\coloneqq tentative$;
-13            add $\mathcal{G}_{\text{neighbor}}$ to $openSet$ with priority $g(\mathcal{G}_{\text{neighbor}})+h(\mathcal{G}_{\text{neighbor}})$;
-Algorithm 1 An A* algorithm for computing the model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ .
+4while $\textnormal{{openSet}}\neq\emptyset$ do
+ $(\mathcal{G},\textnormal{{g}}),\textnormal{{f}}\leftarrow\textnormal{{Pop}}(\textnormal{{openSet}})$
+ // Pop an item with the smallest index
+5
+6 if $\textnormal{{f}}\geq u^{\ast}$ then
+    return $u^{\ast}$
+    // Can terminate early because $\textnormal{{f}}\leq d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq u^{\ast}$
+7
+8
+9 if $\textnormal{{g}}>g[\mathcal{G}]$ then
+    Continue
+    // Skip a stale entry in the queue
+10
+11
+12 if $\mathcal{G}=\mathcal{G}_{t}$ then
+    return g
+    // Reached the target
+13
+14
+15 foreach $\mathcal{G}^{\prime}\in\textnormal{{EnumNeighbors}}(\mathcal{G})$ do
+    $\textnormal{{g}}^{\prime}\leftarrow\textnormal{{g}}+1$;
+    // Increment current distance by one
+16    $\textnormal{{u}}^{\prime}\leftarrow\textnormal{{g}}^{\prime}+u(\mathcal{G}^{\prime})$
+17    if $\textnormal{{u}}^{\prime}<u^{\ast}$ then
+       $u^{\ast}\leftarrow\textnormal{{u}}^{\prime}$
+       // Tighten the upper bound
+18
+19    if $\mathcal{G}^{\prime}\notin\textnormal{{Keys}}(g)$ or $\textnormal{{g}}^{\prime}<g[\mathcal{G}^{\prime}]$ then
+       $g[\mathcal{G}^{\prime}]\leftarrow\textnormal{{g}}^{\prime}$
+       // Path through $\mathcal{G}^{\prime}$ is created or shortened
+20       $\textnormal{{f}}^{\prime}\leftarrow\textnormal{{g}}^{\prime}+h(\mathcal{G}^{\prime})$
+21       if $\textnormal{{f}}^{\prime}<u^{\ast}$ then
+          $\textnormal{{Push}}(\textnormal{{openSet}},(\mathcal{G}^{\prime},\textnormal{{g}}^{\prime}),\textnormal{{f}}^{\prime})$
+          // Push a potentially better solution
+22
+23
+24
+25
+return 1-1
+// Error: poset is not connected
+Algorithm 1 Generic A* algorithm with branch and bound for computing the model-oriented distance
 5.2 Specialization to CPDAGs
-Next, we consider the concrete problem of computing $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ for CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$.
-Here, $\mathcal{L}$ is the poset underlying all CPDAGs over a fixed vertex set.
-Using Algorithm 1 requires specifying a heuristic $h$ and the function $\mathrm{n}_{\mathcal{L}}$.
-As heuristic $h$, we use $|\mathrm{sk}(\mathcal{G})\setminus\mathrm{sk}(\mathcal{G}_{t})|+|\mathrm{sk}(\mathcal{G}_{t})\setminus\mathrm{sk}(\mathcal{G})|$ as a lower bound for the model-oriented distance between any $\mathcal{G}$ and $\mathcal{G}_{t}$; see Property 2 where we lower bound $m_{\text{op}}$ by zero as computing $m_{\text{op}}$ for every call of the heuristic can be costly.
-For the function $\mathrm{n}_{\mathcal{L}}$, we use characterization via the “Insert" and “Delete" operators introduced in Chickering (2002) (see Lemma˜2).
-Computing $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ via Algorithm 1 can be expensive.
-We can compute the down-up and up-down distances (Property 1), to use them as upper bounds on $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
-We can also compute the lower bound from Property 2.
-These bounds are more efficient to compute, and when they coincide, they yield the exact value of $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
-We next describe how to compute the upper bounds based on the down-up and up-down distances, and present numerical experiments comparing the model-oriented distance to the two variants of SHD.
-The algorithmic details for computing the lower bound are provided in Appendix I.
-5.2.1 Computing an upper bound based on down-up or up-down distances
-Consider first the down-up distance $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$.
-One can adapt Algorithm 1 to search for the shortest down-up path between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$.
-This adaptation requires only a few conceptual changes: (i) The neighbor selection step is restricted to graphs that are covered by the current graph, ensuring that each move is downward.
-(ii) The search terminates early when a graph $\mathcal{G}^{\prime}\preceq\mathcal{G}_{t}$ is reached.
-At this point, the upward portion of the path is not explicitly traversed; instead, its length is computed directly as $\mathrm{rank}(\mathcal{G}_{t})-\mathrm{rank}(\mathcal{G}^{\prime})$.
-This uses the fact that in a graded poset, the shortest upward path from $\mathcal{G}_{1}$ to $\mathcal{G}_{2}$ with $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ has length exactly $\mathrm{rank}(\mathcal{G}_{2})-\mathrm{rank}(\mathcal{G}_{1})$; see Proposition 1.
-(iii) The down-up distance between any graph $\mathcal{G}^{\prime}$ and $\mathcal{G}_{t}$ in this poset is given by $\mathrm{rank}(\mathcal{G}_{t})+\mathrm{rank}(\mathcal{G}^{\prime})-2\max_{\mathcal{G}\preceq\mathcal{G}^{\prime},\mathcal{G}\preceq\mathcal{G}_{t}}\mathrm{rank}(\mathcal{G})$, which motivates the following heuristic $\mathrm{rank}(\mathcal{G}_{t})+\mathrm{rank}(\mathcal{G}^{\prime})-2|\mathrm{sk}(\mathcal{G}_{t})\cap\mathrm{sk}(\mathcal{G}^{\prime})|$.
-In Appendix F.3, we prove that for any two CPDAGs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, $\max_{\mathcal{G}\preceq\mathcal{G}_{1},\mathcal{G}\preceq\mathcal{G}_{2}}\mathrm{rank}(\mathcal{G})\leq|\mathrm{sk}(\mathcal{G}_{1})\cap\mathrm{sk}(\mathcal{G}_{2})|$, so the heuristic is a valid lower bound on the down-up distance and is therefore admissible.
-Note that we check whether $\mathcal{G}^{\prime}\preceq\mathcal{G}_{t}$ using the characterization in Kocka, Bouckaert and Studený (2001), which provides necessary and sufficient conditions for conditional dependence inclusion between two CPDAGs.
-Since this check can be computationally costly, we consider a necessary criterion for $\mathcal{G}^{\prime}$ to be the inflection point in the shortest down-up path between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$.
-Appendix F.1 shows that the optimal inflection point is any maximizer of $\operatorname*{arg\,max}_{\mathcal{G}\preceq\mathcal{G}_{s},,\mathcal{G}\preceq\mathcal{G}_{t}}\mathrm{rank}(\mathcal{G})$, whose rank we denote by $\rho_{\text{meet}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
-As previously discussed, this value is guaranteed to be less than or equal to $u=|\mathrm{sk}(\mathcal{G}_{s})\cap\mathrm{sk}(\mathcal{G}_{t})|$.
-Thus, before performing the check $\mathcal{G}^{\prime}\preceq\mathcal{G}_{t}$, we first verify whether $\mathrm{rank}(\mathcal{G}^{\prime})\leq\rho_{\text{meet}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq u$.
-The overall adapted A* algorithm for finding the shortest down-up distance is presented in Algorithm 2.
-We present the algorithm for computing the up-down distance in Appendix I, which is the ‘dual’ of Algorithm 2.
-In particular, (i) the neighbor selection step is restricted to be upward; (ii) the search terminates when a graph $\mathcal{G}^{\prime}$ is found with $\mathcal{G}_{t}\preceq\mathcal{G}^{\prime}$; (iii) we use the admissible heuristic $2|\mathrm{sk}(\mathcal{G}_{t})\cup\mathrm{sk}(\mathcal{G}^{\prime})|-\mathrm{rank}(\mathcal{G}_{t})-\mathrm{rank}(\mathcal{G}^{\prime})$; and finally (iv) we verify whether $\mathrm{rank}(\mathcal{G}^{\prime})\geq|\mathrm{sk}(\mathcal{G}_{t})\cup\mathrm{sk}(\mathcal{G}^{\prime})|$ before checking if $\mathcal{G}_{t}\preceq\mathcal{G}^{\prime}$.
-input: Graph class of CPDAGs $\mathfrak{G}$, pair of graphs $\mathcal{G}_{s},\mathcal{G}_{t}$, a heuristic $h\colon\mathfrak{G}\to\mathbb{R}_{\geq 0}$, downwards neighbor function $dn_{\mathcal{L}}\colon\mathfrak{G}\to 2^{\mathfrak{G}}$, upper bound $u\in\mathbb{R}_{\geq 0}$ on $\rho_{\text{meet}}(\mathcal{G}_{s},\mathcal{G}_{t})$
-output : Down-up distance $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$
-1 begin
- $openSet\coloneqq\{\mathcal{G}_{s}\}$;
- // priority queue ordered by $g+h$
-2 $g(\mathcal{G}_{s})\coloneqq 0$;
-3 while $openSet\neq\emptyset$ do
-4    $\mathcal{G}_{\text{current}}\coloneqq$ graph in openSetopenSet minimizing $g(\mathcal{G}_{\text{current}})+h(\mathcal{G}_{\text{current}})$;
-5    if $\mathrm{rank}(\mathcal{G}_{\text{current}})\leq u\,\,\text{ and }\,\,\mathcal{G}_{\text{current}}\preceq\mathcal{G}_{t}$ then
-6       return $g(\mathcal{G}_{\text{current}})+\mathrm{rank}(\mathcal{G}_{t})-\mathrm{rank}(\mathcal{G}_{\text{current}})$
-7     remove $\mathcal{G}_{\text{current}}$ from openSetopenSet;
-8
-9    foreach $\mathcal{G}_{\text{neighbor}}\in dn_{\mathcal{L}}(\mathcal{G}_{\text{current}})$ do
-10       $tentative\coloneqq g(\mathcal{G}_{\text{current}})+1$;
-11       if $g(\mathcal{G}_{\text{neighbor}})$ not defined   or   $tentative<g(\mathcal{G}_{\text{neighbor}})$ then
-12          $g(\mathcal{G}_{\text{neighbor}})\coloneqq tentative$;
-13            add $\mathcal{G}_{\text{neighbor}}$ to $openSet$ with priority $g(\mathcal{G}_{\text{neighbor}})+h(\mathcal{G}_{\text{neighbor}})$;
-Algorithm 2 An A* algorithm for computing the down-up distance $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$.
-5.2.2 Numerical experiments for CPDAGs
-To study the characteristics of the model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ for CPDAGs as well as the computational cost of our algorithms, we perform a small simulation study.
-We consider two settings.
-First, all pairs of CPDAGs $(\mathcal{G}_{s},\mathcal{G}_{t})$ over four nodes (185 in total; Gillispie and Perlman, 2001) and second, all pairs over five nodes ($8782$ graphs).
-In both settings, We first compute lower and upper bounds on $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ using the procedures described earlier.
-When the bounds do not match, we apply Algorithm 1 to obtain the exact distance.
-We also compute both SHD variants (Definition 7) for comparison.
-In the setting with four nodes we find the following results.
-For approximately $96\%$ of pairs, the bounds match exactly; for the remaining $\approx 4\%$, they differ by at most two.
-The average computation time is $\approx 0.00003$ seconds when bounds match, compared to $\approx 0.0001$ seconds otherwise, highlighting the efficiency of using bounds before invoking Algorithm 1.
-Fig.˜7(a) shows the histogram of the difference between SHD and the model-oriented distance.
-In $\approx 64\%$ of cases, the distances differ, and in $\approx 63\%$, the model-oriented distance is strictly smaller.
-Fig.˜8 presents three illustrative examples.
-In the first example in Figure 8(a), $\mathcal{G}_{s}\preceq\mathcal{G}_{t}$, and thus appealing to Proposition 1, $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=\mathrm{rank}(\mathcal{G}_{t})-\mathrm{rank}(\mathcal{G}_{s})=2$.
-The SHD variants give significantly different values for the distance: $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})=6$ and $\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})=8$.
-In the second example in Fig.˜8(b), we have $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=2$, where the shortest path between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ is an up-down path: moving a step upward from $\mathcal{G}_{s}$ to the fully connected CPDAG $\mathcal{G}$, and then moving a step downward from $\mathcal{G}$ to $\mathcal{G}_{t}$.
-Again, in this case, SHD variants are significantly different: $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})=6$ and $\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})=12$.
-In the final example in Fig.˜8(c), $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})=\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})=3$.
-In this case, unlike the other examples, the model-oriented distance is larger and equals $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=4$.
-Specifically, a shortest down-up path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ first removes the edges $v_{3}\rightarrow v_{4}$ and $v_{1}\rightarrow v_{3}$ in two downward steps to obtain a CPDAG $\mathcal{G}$ containing the edge $v_{2}-v_{4}$, and then adds the edges $v_{1}-v_{4}$ and $v_{3}-v_{4}$ in two upward steps.
-In the setting with five nodes we find the following results.
-The upper and lower bounds on the model-oriented distance coincide in $89\%$ of cases, with an average computation time of $\approx 0.0001$ seconds.
-For the remaining cases, the average time is $\approx 0.002$ seconds.
-Figure 7(b) shows the histogram of differences between both SHD variants and the model-oriented distance.
-In $\approx 80\%$ of cases, the model-oriented distance differs from both SHD variants, and in $\approx 78\%$, it is strictly smaller.
-In about $3\%$ of cases, the shortest path follows a ‘zig-zag’ pattern rather than a down–up or up–down path.
-Fig.˜8(d) presents an example where the model-oriented distance and SHD differ significantly.
-Here, $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})=10$ and $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})=12$, whereas the model-oriented distance yields $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=2$.
-Since SHD is model-agnostic, it does not account for the fact that all conditional dependencies encoded by $\mathcal{G}_{s}$ are also encoded by $\mathcal{G}_{t}$, and that adding two edges to $\mathcal{G}_{s}$ – one between $v_{1}$ and $v_{2}$, and another between $v_{3}$ and $v_{4}$ – yields the CPDAG $\mathcal{G}_{t}$.
-Figure 7: The difference between SHD variants and our model-oriented distance, computed over all CPDAG pairs with $p$ nodes. (a) $p=4$; (b) $p=5$.
-Figure 8: Examples highlighting differences between model-oriented distances and SHD; see text in Sections 5.2.2 and 5.3.1.
-5.3 Specialization to MPDAGs
-Next, we consider computing $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ for MPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$, where $\mathcal{L}_{\text{poly}}$ is the poset underlying all MPDAGs that are polytrees over a fixed vertex set.
-In the setting of polytrees, recall from Theorem 4 that $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ simplifies to $d_{\mathcal{L}_{\text{poly}},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
-This simplification is important for multiple reasons: (i) appealing to Proposition 13, $\operatorname{\mathrm{SHD}}_{2}$ provides a lower bound for $d_{\mathcal{L}_{\text{poly}}}$ and thus serves as an admissible heuristic for Algorithm 1; and (ii) a graph’s neighbors in the poset $\mathcal{L}_{\text{poly}}$ are those that differ in pseudo-rank by one, and these neighbors can be computed efficiently (see Section 4.4).
-We supply the admissible heuristic and the neighbor function to Algorithm 1 to compute the model-oriented distance.
-5.3.1 Numerical experiments for MPDAGs
-We compute model-oriented distance $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ for each pair of MPDAGs $(\mathcal{G}_{s},\mathcal{G}_{t})$ over five nodes (6679 total).
-The average time to compute this distance is approximately $0.004$ seconds.
-In roughly $98\%$ of the instances, the model-oriented distance matches the distance obtained by $\operatorname{\mathrm{SHD}}_{2}$.
-Among the remaining approximately $2\%$: in about $98\%$ of these cases, the model-oriented distance is two units larger than $\operatorname{\mathrm{SHD}}_{2}$; in roughly $1.7\%$ of cases, it is four units larger; and in five instances, it is six units larger.
-Fig.˜8(d) shows an example of an instance where the difference between the two distances is six.
-Here, $\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})=4$, whereas the model-oriented distance yields $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})=10$.
-The discrepancy arises because SHD incorrectly assumes that directing any of the undirected edges toward $v_{5}$ results in a valid MPDAG.
-However, since $\mathcal{L}_{\text{poly}}$ only contains polytrees, the shortest path in the poset $\mathcal{L}_{\text{poly}}$ from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ involves first directing each undirected edge away from $v_{5}$ (toward the other node), then removing these directed edges, and finally adding directed edges from each node to $v_{5}$.
+We specialize Algorithm˜1 to $\mathfrak{G}$, the class of CPDAGs over a vertex set $\mathrm{V}$, by specifying the subroutines required by Algorithm˜1.
+By Proposition˜6, an admissible heuristic can be constructed from
+$d_{\mathcal{L}}(\mathcal{G},\mathcal{G}_{t})\geq|\operatorname{sk}(\mathcal{G})\triangle\operatorname{sk}(\mathcal{G}_{t})|+2m_{\mathrm{op}}(\mathrm{S}(\mathcal{G},\mathcal{G}_{t}))\geq|\operatorname{sk}(\mathcal{G})\triangle\operatorname{sk}(\mathcal{G}_{t})|,\quad\mathcal{G}\in\mathfrak{G},$
+where $|\operatorname{sk}(\mathcal{G})\triangle\operatorname{sk}(\mathcal{G}_{t})|$ and $m_{\mathrm{op}}$ can be computed using Algorithms˜2 and 3 listed in the Appendix.
+By the same proposition, $u(\mathcal{G})$ can be constructed from the upper bound $d_{\mathcal{L}}(\mathcal{G},\mathcal{G}_{t})\leq\min\left\{d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G},\mathcal{G}_{t}),d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G},\mathcal{G}_{t})\right\}$ for every $\mathcal{G}\in\mathfrak{G}$.
+Next, we describe how to compute $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G},\mathcal{G}_{t})$ and the same logic also applies to $d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G},\mathcal{G}_{t})$.
+Because the poset for CPDAGs is graded, by Proposition˜3, we have
+$d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G},\mathcal{G}_{t})=\operatorname{\mathrm{rank}}(\mathcal{G})+\operatorname{\mathrm{rank}}(\mathcal{G}_{t})-2\operatorname{\mathrm{rank}}(\mathcal{G}^{\ast}),\quad\mathcal{G}\in\mathfrak{G},$
+where $\operatorname{\mathrm{rank}}(\cdot)$ is given by the total number of edges (i.e., $|\operatorname{sk}(\cdot)|$).
+Here, $\mathcal{G}^{\ast}$ is the inflection point, i.e., a graph with maximal rank among those that are below both $\mathcal{G}$ and $\mathcal{G}_{t}$.
+In fact, we can use Algorithm˜1 to search for $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G},\mathcal{G}_{t})$, only with the following adaptations: (i) $\textsc{EnumNeighbors}(\mathcal{G}^{\prime})$ is restricted to the neighbors below, i.e., those graphs covered by $\mathcal{G}^{\prime}$; (ii) the upper bound is set to
+$u(\mathcal{G}^{\prime}):=\begin{cases}\operatorname{\mathrm{rank}}(\mathcal{G}_{t})+\operatorname{\mathrm{rank}}(\mathcal{G}^{\prime}),&\quad\mathcal{G}^{\prime}\not\preceq\mathcal{G}_{t}\\ \operatorname{\mathrm{rank}}(\mathcal{G}_{t})-\operatorname{\mathrm{rank}}(\mathcal{G}^{\prime}),&\quad\mathcal{G}^{\prime}\preceq\mathcal{G}_{t}\end{cases},$
+where the first case corresponds to a path that first goes downward from $\mathcal{G}$ to $\mathcal{G}_{\hat{0}}$ and then upward to $\mathcal{G}_{t}$; (iii) the heuristic is set to
+$h(\mathcal{G}^{\prime}):=\operatorname{\mathrm{rank}}(\mathcal{G}_{t})+\operatorname{\mathrm{rank}}(\mathcal{G}^{\prime})-2|\mathrm{sk}(\mathcal{G}_{t})\cap\mathrm{sk}(\mathcal{G}^{\prime})|,$
+which is admissible as we show in Section˜F.2.
+Starting from $\mathcal{G}_{s}$, this algorithm searches downward until it finds the infection point $\mathcal{G}^{\ast}\preceq\mathcal{G}_{t}$, upon which we have $u(\mathcal{G}^{\ast})=h(\mathcal{G}^{\ast})=\operatorname{\mathrm{rank}}(\mathcal{G}_{t})-\operatorname{\mathrm{rank}}(\mathcal{G}^{\ast})$ and the algorithm terminates immediately by line 7 of Algorithm˜1.
+For checking $\mathcal{G}^{\prime}\preceq\mathcal{G}_{t}$, we use the following result due to Kocka, Bouckaert and Studený (2001).
+Lemma 4 (Characterization of CPDAG containment).
+For CPDAGs $\mathcal{G},\mathcal{G}^{\prime}$ with model map $\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}}$, we have $\mathcal{G}^{\prime}\preceq\mathcal{G}$ if and only if for every pair of distinct vertices $(u,v)$,
+[…]
+Figure 7: Joint distribution of our distance and the structural Hamming distance computed over all pairs of CPDAGs over 4 or 5 vertices, where the identity line is drawn as dashed.
+Finally, from Lemma˜2 it follows that for CPDAGs $\mathcal{G},\mathcal{G}^{\prime}$, we have $\mathcal{G}\lessdot\mathcal{G}^{\prime}$ if and only if there exist DAGs $\mathcal{D}\in[\mathcal{G}]$ and $\mathcal{D}^{\prime}\in[\mathcal{G}^{\prime}]$ such that $\mathcal{D}$ can be transformed to $\mathcal{D}^{\prime}$ through zero or more steps of covered edge reversal followed by a single edge addition.
+This can be turned into a subroutine that enumerates the neighbors of a CPDAG.
+In fact, the GES algorithm due to Chickering (2002) has already implemented this subroutine: the first phase of GES uses a set of valid Insert operations to enumerate the neighbors above (i.e., those that cover the current graph), and the second phase use a set of valid Delete operations to enumerate the neighbors below (i.e., those that are covered by the current graph).
+The reader is referred to Chickering (2002, §5) for a description of these operations.
+Numerical results
+To demonstrate our algorithm, we report results from a small numerical experiment.
+There are 185 CPDAGs over 4 vertices and 8782 CPDAGs over 5 vertices; see Gillispie and Perlman (2001).
+We use our algorithm to compute the model-oriented distance between every pair of CPDAGs over 4 or 5 vertices.
+Fig.˜7 compares our distance with the structural Hamming distance: both $\operatorname{\mathrm{SHD}}_{1}$ and $\operatorname{\mathrm{SHD}}_{2}$ are typically larger than the model-oriented distance.
+In particular, Fig.˜8 (a)-(d) highlight some pairs of graphs with large discrepancy between our distance and the $\operatorname{\mathrm{SHD}}$’s.
+We implement our algorithm in C++ and the results are obtained on an ARM64 machine with 4 CPU cores and 32 GB memory.
+The current version scales to a moderate number of vertices.
+For instance, for a pair of CPDAGs over 13 vertices randomly drawn from an Erdös–Rényi model with edge probability 0.2, the algorithm takes about 0.1 second on average to compute the distance.
+It is worth mentioning that in this case $\mathfrak{G}$ contains around $5\times 10^{30}$ graphs (OEIS Foundation Inc., 2025), so the computation is only feasible by algorithmically exploiting the structures of the poset.
+From our experiments, a significant amount of computation is saved by early termination upon attaining the upper bound (line 7 of Algorithm˜1).
+For CPDAGs over 4 or 5 vertices, around 96% or 89% of the pairs have $d_{\mathcal{L}}=\min(d_{\mathcal{L},\uparrow\downarrow},d_{\mathcal{L},\downarrow\uparrow})$.
+(a) $d_{\mathcal{L}}=2$, $\operatorname{\mathrm{SHD}}_{1}=6$, $\operatorname{\mathrm{SHD}}_{2}=8$
+(b) $d_{\mathcal{L}}=2$, $\operatorname{\mathrm{SHD}}_{1}=6$, $\operatorname{\mathrm{SHD}}_{2}=12$
+(c) $d_{\mathcal{L}}=4$, $\operatorname{\mathrm{SHD}}_{1}=3$, $\operatorname{\mathrm{SHD}}_{2}=3$
+(d) $d_{\mathcal{L}}=2$, $\operatorname{\mathrm{SHD}}_{1}=10$, $\operatorname{\mathrm{SHD}}_{2}=12$
+(e) $d_{\mathcal{L}}=10$, $\operatorname{\mathrm{SHD}}_{1}=4$, $\operatorname{\mathrm{SHD}}_{2}=4$
+Figure 8: Example graph pairs that exhibit large discrepancy between the model-oriented distance and the structural Hamming distance: (a-d) are CPDAGs, (e) is polytree MPDAGs.
+5.3 Specialization to polytree MPDAGs
+We consider using Algorithm˜1 to compute $d_{\mathcal{L}_{\mathrm{poly}}}$ for a pair of polytree MPDAGs, where $\mathcal{L}_{\mathrm{poly}}$ is the subposet induced by polytree MPDAGs over a fixed vertex set.
+This is feasible because we show in Theorem˜7 that $d_{\mathcal{L}_{\mathrm{poly}}}=d_{\mathcal{L}_{\mathrm{poly}},\downarrow\uparrow}$, so we can again adapt Algorithm˜1 to compute the down-up distance.
+By the same theorem, the neighborhood of a graph $\mathcal{G}$ in $\mathcal{L}_{\mathrm{poly}}$ precisely consists of those graphs that differs from $\mathcal{G}$’s by unit pseudo-rank, which leads to a straightforward EnumNeighbors subroutine.
+Finally, $h(\mathcal{G}):=\operatorname{\mathrm{SHD}}_{2}(\mathcal{G},\mathcal{G}_{t})$ is an admissible heuristic.
+Numerical results
+We compute the model-oriented distance for every pair of polytree MPDAGs over 5 vertices (6679 graphs in total), which takes 0.004 second per pair on average.
+Fig.˜S14 in the Appendix compares our distance to the structural Hamming distances: about 98% of the pairs have $d_{\mathcal{L}_{\mathrm{poly}}}=\operatorname{\mathrm{SHD}}_{2}$.
+Fig.˜8(d) shows an example where $d_{\mathcal{L}_{\mathrm{poly}}}=\operatorname{\mathrm{SHD}}_{2}+6$.
 6 Discussion
-This paper introduces a model-oriented framework for defining distances between graphs within a given class, where each graph is treated as a statistical model and the collection of graphs is organized as a poset based on model inclusion.
-The resulting metric, derived from the shortest path through neighboring models, provides a principled way to measure distance across various graph classes including causal DAGs, probabilistic CPDAGs, and causal MPDAGs.
-Our work opens several avenues for future research.
-First, for efficient computation of the model-oriented distance for MPDAGs, it is important to develop an explicit graphical characterization of the neighborhood structure.
-Second, one can consider a weighted version of the shortest path, where a weight is assigned to each covering pair in the poset.
-One possible approach for selecting these weights is to base them on information-theoretic notions, as was done in the context of distances between two partitions (Meila, 2005).
-Moreover, computational efficiency in executing the A* algorithm can be substantially improved through better heuristics, parallelization, and memory management.
-The authors are currently pursuing a separate project to reduce the time it takes to compute the model-oriented distance.
-Finally, the proposed framework is applicable to other graph classes including local independence graphs for capturing dependence structure of events that occur in time (Didelez, 2008; Mogensen and Hansen, 2020), maximal ancestral graphs (Richardson and Spirtes, 2002) and partial ancestral graph (Spirtes, Glymour and Scheines, 2000) that account for model dependencies in the presence of latent variables.
+This paper introduces a model-oriented framework for defining a distance between graphs relative to a class of graphs in consideration.
+By treating each graph as a statistical (i.e., probabilistic or causal) model, we organize the class of graphs as a poset ordered by model inclusion.
+The poset induces a notion of neighborhood, which then yields a shortest path distance through neighbors as a metric in the space of graphs.
+By studying and exploiting the poset structures that are summarized in Table˜2, we develop A* algorithms for computing the model-oriented distance that can scale up to a moderate number of vertices.
+By bridging graphical modeling with order theory, we open several avenues for future research.
+First, to compute the model-oriented distance between two MPDAGs, it remains an open problem to obtain an explicit graphical characterization of the neighborhood.
+Second, our distance can be extended to incorporate weights on the covering relations in the poset.
+In particular, one may introduce information-theoretic weights similar to those considered by Meila (2005) for partitions.
+Third, it is of interest to consider model selection in the spirit of Taeb, Bühlmann and Chandrasekaran (2024).
+Moreover, as an ongoing project, we are exploring ways to scale up the A* algorithm to larger graphs through better heuristic, upper bound, memory management and parallelization.
+Finally, our distance is in principle applicable to many other statistical graphs, such as maximal ancestral graphs (Richardson and Spirtes, 2002), partial ancestral graphs (Spirtes, Glymour and Scheines, 2000) and local independence graphs for event data (Didelez, 2008; Mogensen and Hansen, 2020).
 {acks}
-[Acknowledgments] The authors thank undergraduate students Arnav Mazumder, Jian Kang, and Carter Lembo at the University of Washington for helping with the implementations of our algorithms.
+[Acknowledgments] The authors thank undergraduate students Jian Kang, Carter Lembo and Arnav Mazumder at the University of Washington for helping with the implementation of our algorithms.
 Appendix A Proofs: model maps and associated graphs
 A.1 Injective model maps
-We show that the models $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}},\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}},\mathcal{M}^{\mathrm{c}}_{\textsc{dag}},\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}$, as defined in Section˜2.2, are injective in the sense that no two different graphs represent the same model.
-Assumption 1.
+We show that the models $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}},\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}},\mathcal{M}^{\mathrm{c}}_{\textsc{dag}},\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}$, as defined in Section˜2.3, are injective in the sense that no two different graphs represent the same model.
+Assumption S1.
 $\mathrm{V}$ is a finite vertex set. For every $v\in\mathrm{V}$, $\mathcal{X}_{v}$ is either infinite or finite with two or more elements.
-Theorem 5.
-Under ˜1, we have the following results.
+Theorem S9.
+Under ˜S1, we have the following results.
   1. (i)
@@
 […]
-    Further, under ˜1, there exists $P\in\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}(\mathcal{G})$ that is faithful to $\mathcal{G}$, i.e.,
+    Further, under ˜S1, there exists $P\in\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}(\mathcal{G})$ that is faithful to $\mathcal{G}$, i.e.,
 […]
@@
     $P\left(X_{v}(x)=X_{v}(x_{\operatorname{Pa}_{\mathcal{D}}(v)}),\;\forall v\in\mathrm{V}\right)=1,\quad\forall P\in\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}(\mathcal{D}).$
-    Under ˜1, there also exists $P^{\prime}\in\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}(\mathcal{D})$ such that
+    Under ˜S1, there also exists $P^{\prime}\in\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}(\mathcal{D})$ such that
     $\mathrm{W}\neq\operatorname{Pa}_{\mathcal{D}}(v)\implies P^{\prime}\left(X_{v}(x)=X_{v}(x_{\mathrm{W}})\right)<1,\quad\forall v\in\mathrm{V},\;\mathrm{W}\subseteq\mathrm{V}.$
@@
 ∎
-Proof of Proposition˜6.
-By definition of the causal model associated with a DAG (see Section˜2.2.2), if $\mathcal{G}_{1}$ is a subgraph of $\mathcal{G}_{2}$, then $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$.
+Proof of Proposition˜5.
+By definition of the causal model associated with a DAG (see Section˜2.3.2), if $\mathcal{G}_{1}$ is a subgraph of $\mathcal{G}_{2}$, then $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$.
 We now prove the other direction by contradiction.
@@
 ∎
-Proof of Proposition˜9.
-Recall from Section˜2.2 that $\mathcal{M}(\mathcal{G}_{i}):=\cup_{\mathcal{D}\in\mathcal{G}_{i}}\mathcal{M}(\mathcal{D})$ for $i=1,2$, where $\mathcal{D}\in\mathcal{G}_{i}$ is any DAG represented by the MPDAG $\mathcal{G}_{i}$.
-It follows from definition of $\mathcal{M}(\mathcal{D})$ and Proposition˜6 that, if for every $\mathcal{D}_{1}\in\mathcal{G}_{1}$, there exists a certain $\mathcal{D}_{2}\in\mathcal{G}_{2}$ such that $\mathcal{D}_{1}$ is a subgraph of $\mathcal{D}_{2}$, then $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$.
+Proof of Proposition˜7.
+Recall from Section˜2.3 that $\mathcal{M}(\mathcal{G}_{i}):=\cup_{\mathcal{D}\in\mathcal{G}_{i}}\mathcal{M}(\mathcal{D})$ for $i=1,2$, where $\mathcal{D}\in\mathcal{G}_{i}$ is any DAG represented by the MPDAG $\mathcal{G}_{i}$.
+It follows from definition of $\mathcal{M}(\mathcal{D})$ and Proposition˜5 that, if for every $\mathcal{D}_{1}\in\mathcal{G}_{1}$, there exists a certain $\mathcal{D}_{2}\in\mathcal{G}_{2}$ such that $\mathcal{D}_{1}$ is a subgraph of $\mathcal{D}_{2}$, then $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$.
 We now prove the other direction.
@@
 where $\beta_{i}\in\mathbb{R}^{|\operatorname{Pa}_{\mathcal{D}_{1}}(i)|}$ is a fixed vector with non-zero coordinates.
-Since $P\in\mathcal{M}(\mathcal{D}_{1})\subseteq\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})=\cup_{\mathcal{D}\in\mathcal{G}_{2}}\mathcal{M}(\mathcal{D})$, there must exists $\mathcal{D}_{2}\in\mathcal{G}_{2}$ such that $P\in\mathcal{M}(\mathcal{D}_{2})$.
+Since $P\in\mathcal{M}(\mathcal{D}_{1})\subseteq\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})=\cup_{\mathcal{D}\in\mathcal{G}_{2}}\mathcal{M}(\mathcal{D})$, there must exist $\mathcal{D}_{2}\in\mathcal{G}_{2}$ such that $P\in\mathcal{M}(\mathcal{D}_{2})$.
 By definition, under $\mathcal{M}(\mathcal{D}_{2})$, $\operatorname{\mathbb{E}}[X_{i}(x)]$ (when it exists) can only depend on $x_{\operatorname{Pa}_{\mathcal{D}_{2}}(i)}$.
-Hence, Eq.˜5 implies that $\operatorname{Pa}_{\mathcal{D}_{1}}(i)\subseteq\operatorname{Pa}_{\mathcal{D}_{2}}(i)$ for every $i\in V$.
+Hence, Eq.˜8 implies that $\operatorname{Pa}_{\mathcal{D}_{1}}(i)\subseteq\operatorname{Pa}_{\mathcal{D}_{2}}(i)$ for every $i\in V$.
 That is, $\mathcal{D}_{1}$ is a subgraph of $\mathcal{D}_{2}$.
@@
 Appendix B Structural constraints of graph types
+Recall that a partially directed acyclic graph (PDAG) is a simple graph with two types of edges, directed and undirected, such that there is no directed cycle.
 B.1 CPDAG
@@
   2. (ii)
-    Each connected component induced by undirected edges is chordal;
+    Each connected component spanned by undirected edges is chordal;
   3. (iii)
@@
   1. (i)
-    For each connected component $\mathrm{C}$ induced by undirected edges, the skeleton of $\mathcal{G}_{\mathrm{C}}$ is chordal;
+    For each connected component $\mathrm{C}$ spanned by undirected edges, the skeleton of $\mathcal{G}_{\mathrm{C}}$ ($G_{\mathrm{C}}$ is the subgraph induced by $\mathrm{C}$) is chordal;
   2. (ii)
@@
 By definition, for each $i=2,\dots,k$, either $x_{i}$ covers $x_{i-1}$ or $x_{i-1}$ covers $x_{i}$.
-Since the poset is graded, we know that for each step, the rank increases or decreases by one, i.e.
-$|\mathrm{rank}(x_{i})-\mathrm{rank}(x_{i-1})|=1$.
+Since the poset is graded, we know that for each step, the rank increases or decreases by one, i.e., $|\operatorname{\mathrm{rank}}(x_{i})-\operatorname{\mathrm{rank}}(x_{i-1})|=1$.
 Note, by telescoping sum, we have:
-$\mathrm{rank}(\mathcal{G}_{t})-\mathrm{rank}(\mathcal{G}_{s})=\sum_{i=2}^{k}\mathrm{rank}(x_{i})-\mathrm{rank}(x_{i-1})$
+$\operatorname{\mathrm{rank}}(\mathcal{G}_{t})-\operatorname{\mathrm{rank}}(\mathcal{G}_{s})=\sum_{i=2}^{k}\operatorname{\mathrm{rank}}(x_{i})-\operatorname{\mathrm{rank}}(x_{i-1})$
 A term in the decomposition above is equal to $1$ when the change from $x_{i-1}$ to $x_{i}$ is an upward move in the poset ($x_{i-1}\preceq x_{i}$), and is 1-1 when the change from $x_{i-1}$ to $x_{i}$ is downward.
@@
 […]
-We have thus shown that the length of any path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ is greater than or equal $\mathrm{rank}(\mathcal{G}_{t})-\mathrm{rank}(\mathcal{G}_{s})$.
+We have thus shown that the length of any path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ is greater than or equal $\operatorname{\mathrm{rank}}(\mathcal{G}_{t})-\operatorname{\mathrm{rank}}(\mathcal{G}_{s})$.
 Note that the lower-bound can be achieved by considering an upward path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ with the sequence $(x_{1}:=\mathcal{G}_{s},x_{2},\dots,x_{k}:=\mathcal{G}_{t})$ where $x_{i}$ covers $x_{i-1}$ for every $i=2,3,\dots,k$.
@@
 ∎
-Appendix E Properties of posets underlying different graph classes
+Appendix E Posets properties for different graph classes
 E.1 DAGs
-Proof of Proposition˜7.
+Proof of Theorem˜4.
   1. 1, 2 and 3.
-    They directly follow from Proposition˜6 and the fact that given an arbitrary $\mathcal{G}\in\mathfrak{G}$, every subgraph of $\mathcal{G}$ is also in $\mathfrak{G}$.
+    They directly follow from Proposition˜5 and the fact that given an arbitrary $\mathcal{G}\in\mathfrak{G}$, every subgraph of $\mathcal{G}$ is also in $\mathfrak{G}$.
   2. 4.
@@
 To see this, consider two DAGs that have an edge in opposite directions.
-They have no common upper bound by Proposition˜6.
+They have no common upper bound by Proposition˜5.
 E.2 CPDAGs
-Proof of Proposition˜8.
+Proof of Theorem˜5.
   1. 1.
@@
 By 2.
-of Proposition˜8, we have $\mathcal{G}_{1},\mathcal{G}_{2}\lessdot\mathcal{G}_{3},\mathcal{G}_{4}$.
+of Theorem˜5, we have $\mathcal{G}_{1},\mathcal{G}_{2}\lessdot\mathcal{G}_{3},\mathcal{G}_{4}$.
 To see that $\mathcal{L}$ is not a join semi-lattice, observe that the upper bounds on $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ consists of $\mathcal{G}_{3},\mathcal{G}_{4}$ and the full graph $\mathcal{G}_{\hat{1}}$, where $\mathcal{G}_{3},\mathcal{G}_{4}\prec\mathcal{G}_{\hat{1}}$ but $\mathcal{G}_{3}$ and $\mathcal{G}_{4}$ are incomparable.
@@
 By 2.
-of Proposition˜8, we know $\mathcal{G}_{1}^{\prime},\mathcal{G}_{2}^{\prime}\gtrdot\mathcal{G}_{3}^{\prime}$.
+of Theorem˜5, we know $\mathcal{G}_{1}^{\prime},\mathcal{G}_{2}^{\prime}\gtrdot\mathcal{G}_{3}^{\prime}$.
 If $\mathcal{L}^{\prime}$ is upper semimodular, then there must exist a graph that covers both $\mathcal{G}_{1}^{\prime}$ and $\mathcal{G}_{2}^{\prime}$.
@@
 For this purpose, consider again $\mathcal{G}_{5}^{\prime}$ and $\mathcal{G}_{6}^{\prime}$ above.
-Proposition˜8 implies that the full graph covers $\mathcal{G}_{5}^{\prime}$ and $\mathcal{G}_{6}^{\prime}$.
+Theorem˜5 implies that the full graph covers $\mathcal{G}_{5}^{\prime}$ and $\mathcal{G}_{6}^{\prime}$.
 If $\mathcal{L}$ is lower semimodular, then there must exist $\mathcal{G}_{7}^{\prime}$ that is covered by both $\mathcal{G}_{5}^{\prime}$ and $\mathcal{G}_{6}^{\prime}$.
@@
 up to graph isomorphism. One can check that neither can be covered by both $\mathcal{G}_{5}^{\prime}$ and $\mathcal{G}_{6}^{\prime}$.
-E.3 MPDAGs
+E.3 Polytree CPDAGs
+Proof of Theorem˜6.
+  1. 1.
+    This directly follows from definition.
+  2. 2.
+[…]
+    Using the facts above, we conclude the desired result.
+Let $\mathrm{F}=\{(i,j)\}$ be collection of vertices that are connected in $\mathcal{G}_{2}$ but not in $\mathcal{G}_{1}$.
+By the analysis above, any v-structure in $\mathcal{G}_{1}$ is also present in $\mathcal{G}_{2}$.
+Furthermore, every v-structure $i\rightarrow k\leftarrow j$ in $\mathcal{G}_{2}$ is either present in $\mathcal{G}_{1}$ or at least one of the edges between pair of vertices $(i,k)$ and $(j,k)$ is missing.
+Consider any DAG $\mathcal{D}_{2}$ represented by $\mathcal{G}_{2}$.
+Let $\mathcal{D}_{1}$ be the DAG that is obtained by deleting the edges in $\mathcal{D}_{2}$ among pairs of vertices $\mathrm{F}$.
+By construction, $\mathcal{D}_{1}$ has the same skeleton as any DAG in $\mathcal{G}_{1}$.
+Since $\mathcal{D}_{1}$ is a polytree, every structure present in $\mathcal{D}_{1}$ or in $\mathcal{G}_{1}$ are also in $\mathcal{G}_{2}$.
+Suppose as a point of contradiction that there exists a v-structure $i\rightarrow k\leftarrow j$ that is in $\mathcal{D}_{1}$ and not in $\mathcal{G}_{1}$.
+Since this v-structure must also be present in $\mathcal{G}_{2}$, and from the previous statement, then one of the edges between pairs of vertices $(i,k)$ or $(j,k)$ must be missing in $\mathcal{G}_{1}$, which contradicts $\mathcal{D}_{1}$ and $\mathcal{G}_{1}$ have the same skeleton.
+A similar argument allows us to conclude that it is not possible that $\mathcal{G}_{1}$ has a v-structure that is not present in $\mathcal{D}_{1}$.
+Thus, $\mathcal{D}_{1}$ and $\mathcal{G}_{1}$ have the same skeleton and v-structures.
+We conclude that $\mathcal{D}_{1}$ must be a DAG represented by $\mathcal{G}_{1}$.
+  3. 3.
+    Let $\operatorname{\mathrm{rank}}(\cdot)$ be the number of edges.
+It suffices to show that $\operatorname{\mathrm{rank}}(\mathcal{G})+1=\operatorname{\mathrm{rank}}(\mathcal{G}^{\prime})$ holds for every $\mathcal{G}\lessdot\mathcal{G}^{\prime}$ in $\mathcal{L}_{\mathrm{poly}}$.
+By the previous result, $\mathcal{G}\lessdot\mathcal{G}^{\prime}$ implies $\operatorname{\mathrm{rank}}(\mathcal{G}^{\prime})\geq\operatorname{\mathrm{rank}}(\mathcal{G})+1$.
+For contradiction, suppose there exists $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$ with $\operatorname{\mathrm{rank}}(\mathcal{G}_{1})+1<\operatorname{\mathrm{rank}}(\mathcal{G}_{2})$.
+By the previous result, for every DAG $\mathcal{D}_{2}=(\mathrm{V},\mathrm{E}_{2})$ represented by $\mathcal{G}_{2}$, there exists a DAG $\mathcal{D}_{1}=(\mathrm{V},\mathrm{E}_{1})$ represented by $\mathcal{G}_{1}$ such that $\mathrm{E}_{1}\subseteq\mathrm{E}_{2}$.
+Since $\operatorname{\mathrm{rank}}(\mathcal{G}_{1})+1<\operatorname{\mathrm{rank}}(\mathcal{G}_{2})$, there exists $\mathrm{E}_{3}$, satisfying $\mathrm{E}_{1}\subset\mathrm{E}_{3}\subset\mathrm{E}_{2}$.
+Let $\mathcal{G}_{3}$ be the CPDAG obtained by completing the DAG $\mathcal{D}_{3}=(\mathrm{V},\mathrm{E}_{3})$.
+By construction, we have $\mathcal{G}_{1}\preceq\mathcal{G}_{3}\preceq\mathcal{G}_{2}$ with $\mathcal{G}_{3}$ being distinct from $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, contradicting $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$.
+  4. 4.
+    Consider any $\mathcal{G}_{1},\mathcal{G}_{2},\mathcal{G}_{3}\in\mathfrak{G}_{\mathrm{poly}}$ with $\mathcal{G}_{1}\lessdot\mathcal{G}_{3}$ and $\mathcal{G}_{2}\lessdot\mathcal{G}_{3}$.
+We want to show that there is a graph that is covered by both $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$.
+Consider any DAG $\mathcal{D}_{3}=(\mathrm{V},\mathrm{E}_{3})\in[\mathcal{G}_{3}]$.
+By Result 2., there exist DAGs $\mathcal{D}_{1}=(\mathrm{V},\mathrm{E}_{1})\in[\mathcal{G}_{1}]$ and $\mathcal{D}_{2}=(\mathrm{V},\mathrm{E}_{2})\in[\mathcal{G}_{2}]$ with $\mathrm{E}_{1}\subseteq\mathrm{E}_{3}$ and $\mathrm{E}_{2}\subseteq\mathrm{E}_{3}$.
+Applying Result 3., we have $\mathrm{E}_{3}\setminus\mathrm{E}_{1}=e_{1}$ and $\mathrm{E}_{3}\setminus\mathrm{E}_{2}=e_{2}$ for two separate edges $e_{1},e_{2}$.
+Consider DAG $\tilde{\mathcal{D}}:=(\mathrm{V},\mathrm{E}_{1}\cap\mathrm{E}_{2})$ and let $\tilde{\mathcal{G}}$ be the CPDAG such that $\tilde{\mathcal{D}}\in[\tilde{\mathcal{G}}$].
+By construction, $\tilde{\mathcal{G}}$ is a polytree CPDAG that satisfies $\tilde{\mathcal{G}}\preceq\mathcal{G}_{1}$ and $\tilde{\mathcal{G}}\preceq\mathcal{G}_{2}$.
+Since $\operatorname{\mathrm{rank}}(\tilde{\mathcal{G}})=\operatorname{\mathrm{rank}}(\mathcal{G}_{1})-1=\operatorname{\mathrm{rank}}(\mathcal{G}_{2})-1$, we have $\tilde{\mathcal{G}}\lessdot\mathcal{G}_{1}$ and $\tilde{\mathcal{G}}\lessdot\mathcal{G}_{2}$.
+  5. 5.
+    This follows from the previous result and Theorem˜2.
+E.4 MPDAGs
 We first show that the model-oriented poset underlying MPDAGs is not a meet or join semi-lattice in general.
@@
 $\mathcal{G}_{1}^{\prime\prime}:\begin{array}[]{ccc}1&-&2\\ &&|\\ 4&&3\end{array},\qquad\mathcal{G}_{2}^{\prime\prime}:\begin{array}[]{ccc}1&{-}&2\\ \uparrow&\nearrow&\downarrow\\ 4&\rightarrow&3\end{array},\qquad\mathcal{G}_{3}^{\prime\prime}:\begin{array}[]{ccc}1&{-}&2\\ \uparrow&\nearrow&|\\ 4&\rightarrow&3\end{array}.$
-By construction, $\mathcal{G}_{1}^{\prime\prime}$ and $\mathcal{G}_{2}^{\prime\prime}$ are both covered by $\mathcal{G}_{3}^{\prime\prime}$ (see text in Section 4.4).
+By construction, $\mathcal{G}_{1}^{\prime\prime}$ and $\mathcal{G}_{2}^{\prime\prime}$ are both covered by $\mathcal{G}_{3}^{\prime\prime}$ (see Section˜4.4).
 However, there exists no graph covered by both $\mathcal{G}_{1}^{\prime\prime}$ and $\mathcal{G}_{2}^{\prime\prime}$.
@@
 Hence, $\mathcal{L}$ is not lower semimodular.
-Appendix F Theoretical results for CPDAGs
-F.1 Lower bound on $d_{\mathcal{L}}$: proof of Property 2
-Proof of Property 2.
+Finally, we show that the MPDAG poset need not be graded.
+Proof of Proposition˜8.
+To prove that the poset $\mathcal{L}$ underlying MPDAGs is not graded, it suffices to find two maximal chains between a pair of graphs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ of unequal lengths.
+We first show this under $|\mathrm{V}|=4$.
+Fig.˜S9 presents a sequence of graphs that specifies a maximal chain from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$.
+This chain has length $3$.
+Meanwhile, Fig.˜S10 presents another maximal chain from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ that has length 5.
+This shows that $\mathcal{L}$ is not graded.
+When $|\mathrm{V}|>4$, the same argument works by adding in the remaining vertices as isolated vertices with degree zero.
+∎
+Figure S9: A maximal chain between MPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$, where each graph covers its predecessor.
+To see $\mathcal{G}_{3}\lessdot\mathcal{G}_{4}$, observe that (i) $\mathcal{G}_{3}\prec\mathcal{G}_{4}$ and (ii) removing any directed edge from $\mathcal{G}_{4}$ introduces a structure that is forbidden from any MPDAG (see Fig.˜2), so no graph sits between $\mathcal{G}_{3}$ and $\mathcal{G}_{4}$.
+Figure S10: A maximal chain between MPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$, where each graph covers its predecessor.
+E.5 Polytree MPDAGs
+We first prove an auxiliary lemma.
+Lemma S5.
+Let $\mathcal{G}$ be a polytree MPDAG. By removing any edge from $\mathcal{G}$, we obtain an MPDAG $\mathcal{G}^{\prime}$ that satisfies $\mathcal{G}^{\prime}\prec\mathcal{G}$.
+Proof.
+Using the characterization of valid MPDAGs in Section˜B.2, $\mathcal{G}^{\prime}$ is still an MPDAG: it is a PDAG, and further (i) the chordality condition is satisfied, and (ii) no forbidden structure is introduced.
+It then follows from Proposition˜7 that $\mathcal{G}^{\prime}\prec\mathcal{G}$.
+∎
+Proof of Theorem˜7.
+  1. 1.
+    This directly follows from definition.
+  2. 2.
+    Let $\mathcal{G}_{1},\mathcal{G}_{2}$ be polytree MPDAGs with $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$.
+By Proposition˜9, we have $\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{1})+1\leq\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{2})$.
+For contradiction, suppose $\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{1})+1<\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{2})$, or equivalently, $\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{1})+2\leq\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{2})$.
+There are three cases to consider and we show that there is a contradiction in each case.
+    • Case I
+      Two or more pairs of vertices are connected in $\mathcal{G}_{2}$ but not in $\mathcal{G}_{1}$.
+Let $(u,v)$ be such a pair.
+Using Lemma˜S5, removing the edge between $u,v$ from $\mathcal{G}_{2}$ leads to an MPDAG $\tilde{\mathcal{G}}$ with $\tilde{\mathcal{G}}\prec\mathcal{G}_{2}$.
+It also follows from Lemma˜S5 that $\mathcal{G}_{1}\prec\tilde{\mathcal{G}}$.
+This contradicts $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$.
+    • Case II
+      There is at least one pair of vertices that is connected via an undirected edge in $\mathcal{G}_{2}$ but is connected via a directed edge in $\mathcal{G}_{1}$, and there is at least one pair of vertices, say $(u,v)$, that are connected in $\mathcal{G}_{2}$ but not in $\mathcal{G}_{1}$.
+By an argument similar to Case I, removing the edge between $u,v$ from $\mathcal{G}_{2}$ yields an MPDAG $\tilde{\mathcal{G}}$ satisfying $\mathcal{G}_{1}\prec\tilde{\mathcal{G}}\prec\mathcal{G}_{2}$ and hence a contradiction.
+    • Case III
+      There are at least two pairs of vertices, say $(i,j)$ and $(k,l)$, that are connected by an undirected edge in $\mathcal{G}_{2}$ but by a directed edge in $\mathcal{G}_{1}$.
+Suppose $i\to j$ and $k\to l$ appear in $\mathcal{G}_{1}$.
+Let $\mathcal{C}$ be the CPDAG that corresponds to $\mathcal{G}_{2}$.
+Because $\mathcal{G}_{2}$ is an MPDAG, it must result from imposing a set of edge orientations $\mathcal{E}$ (can be empty) on $\mathcal{C}$ and completing Meek’s rules.
+Let $\tilde{\mathcal{G}}$ be the MPDAG that represents $\mathcal{C}$ with orientations $\mathcal{E}\cup\{i\rightarrow j\}$.
+Following the argument from Case II, we have that $\mathcal{G}_{1}\preceq\tilde{\mathcal{G}}\preceq\mathcal{G}_{2}$.
+Due to $i\rightarrow j$ in $\tilde{\mathcal{G}}$ but iji-j in $\mathcal{G}_{2}$, we know $\tilde{\mathcal{G}}\neq\mathcal{G}_{2}$.
+To prove the result, we will show that $(i,j)$ and $(k,l)$ can be chosen in a way that ensures $\tilde{\mathcal{G}}\neq\mathcal{G}_{1}$.
+      More concretely, we show that $(i,j)$ and $(k,l)$ can be chosen such that klk-l instead of $k\rightarrow l$ appears in $\tilde{\mathcal{G}}$.
+To see this, note that in a polytree, the only rule in Meek (1995) that can be applied when directing an edge is the following: if there is a structure of the form $u\rightarrow v-w$, direct the edge vwv-w as $v\rightarrow w$.
+By construction of $\tilde{\mathcal{G}}$, klk-l becomes $k\rightarrow l$ in $\tilde{\mathcal{G}}$ only if it is compelled by this rule.
+Therefore, if $k\rightarrow l$ appears in $\tilde{\mathcal{G}}$, then there must exist an undirected path $i-j-\dots-k-l$ ($j$ may be identical to $k$) in $\mathcal{G}_{2}$ such that one ore more applications of the rule propagates $i\rightarrow j$ to $k\rightarrow l$ in $\tilde{\mathcal{G}}$.
+However, if we were to swap the choices of $(i,j)$ and $(k,l)$, then we would have $i\rightarrow j$ and klk-l in $\tilde{\mathcal{G}}$, which finishes the proof.
+  3. 3.
+    Let $\mathcal{G}_{1},\mathcal{G}_{2},\mathcal{G}_{3}$ be polytree MPDAGs with $\mathcal{G}_{2},\mathcal{G}_{3}\lessdot\mathcal{G}_{1}$ and $\mathcal{G}_{2}\neq\mathcal{G}_{3}$.
+By the previous result, we know
+    $\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{2})=\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{3})=\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{1})-1.$
+    Therefore, $\mathcal{G}_{2}$ and $\mathcal{G}_{3}$ differ from $\mathcal{G}_{1}$ by having either a single directed edge from $\mathcal{G}_{1}$ missing or having a single directed edge that is undirected in $\mathcal{G}_{1}$.
+Up to relabeling we therefore have three cases so let i) be the case where $\mathcal{G}_{2}$ and $\mathcal{G}_{3}$ both have an additional directed edge that is undirected in $\mathcal{G}_{1}$, ii) only $\mathcal{G}_{2}$ has such an edge and iii) neither do.
+    For the first case, note that when directing an undirected edge in an MPDAG that is a polytree, the resulting graph is guaranteed to be a polytree and therefore chordal.
+It is also an MPDAG unless the directed edge is now part of a forbidden structure of the form $\to v$- since the other three forbidden structure cannot exist in a polytree.
+Since $\mathcal{G}_{2}$ and $\mathcal{G}_{3}$ are MPDAGs by assumption, it follows that the edge directed in $\mathcal{G}_{2}$ (or $\mathcal{G}_{3}$) when compared to $\mathcal{G}_{1}$ does not point into another undirected edge in $\mathcal{G}_{1}$ or $\mathcal{G}_{2}$, respectively $\mathcal{G}_{3}$.
+Therefore, we can direct the edge directed in $\mathcal{G}_{3}$ also in $\mathcal{G}_{2}$ to obtain $\mathcal{G}_{4}$.
+By construction, $\mathcal{G}_{4}\lessdot\mathcal{G}_{2},\mathcal{G}_{3}$.
+For the second and third case, note that we can remove any directed edge from an MPDAG that is a polytree by Lemma S5.
+In both Case ii) and iii), we can therefore construct $\mathcal{G}_{4}$ by deleting the directed edge missing in $\mathcal{G}_{3}$ from $\mathcal{G}_{2}$.
+By construction, $\mathcal{G}_{4}\lessdot\mathcal{G}_{2},\mathcal{G}_{3}$.
+  4. 4.
+    From Result 2, we know $\mathrm{P}_{\mathcal{L}_{\mathrm{poly}},\operatorname{\mathrm{pseudo}}}=\mathrm{P}_{\mathcal{L}_{\mathrm{poly}}}$ and hence $d_{\mathcal{L}_{\mathrm{poly}},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t})=d_{\mathcal{L}_{\mathrm{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
+Further, by Result 3 and Theorem˜2, we also have $d_{\mathcal{L}_{\mathrm{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})=d_{\mathcal{L}_{\mathrm{poly}},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$.
+Appendix F Further results for CPDAGs
+F.1 Lower bound on $d_{\mathcal{L}}$: proof of Proposition˜6
+Proof of lower bound in Proposition˜6.
 Any path from $\mathcal{G}_{s}$ from $\mathcal{G}_{t}$ must involve removing edges that are in the skeleton of $\mathcal{G}_{s}$ and not in the skeleton of $\mathcal{G}_{t}$ and adding edges that are in the skeleton of $\mathcal{G}_{t}$ but not in the skeleton of $\mathcal{G}_{s}$.
-Since every edge addition or deletion increases the path length by one, we have that $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\geq|\mathrm{sk}(\mathcal{G}_{s})\setminus\mathrm{sk}(\mathcal{G}_{t})|+|\mathrm{sk}(\mathcal{G}_{t})\setminus\mathrm{sk}(\mathcal{G}_{s})|$.
+Since every edge addition or deletion increases the path length by one, we have that $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\geq|\operatorname{sk}(\mathcal{G}_{s})\triangle\operatorname{sk}(\mathcal{G}_{t})|$.
 The lower bound above does not account for v-structures that are present in one but not in the other.
 Specifically, suppose $\mathcal{G}_{s}$ has a v-structure $i\rightarrow j\leftarrow k$ with $i$ and $k$ not connected, and $\mathcal{G}_{t}$ does not have this v-structure, but has an edge between $i$ and $j$, an edge between $j$ and $k$, and no edge between $i$ and $k$.
-Thus, any path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ must remove this v-structure, which can only be done if and only if at some step at least one of the following happens: $(i)$ the edge between $i$ and $j$ and/or between $j$ and $k$ is removed, and $(ii)$ there is a covering edge that is added between $i$ and $k$.
-For both $(i)$ and $(j)$, an extra operation afterwards is needed: for $(i)$, it is to add back the edge, and for $(ii)$, it is to remove the covered edge.
+Thus, any path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ must remove this v-structure, which can only be done if and only if at some step at least one of the following happens: (i) the edge between $i$ and $j$ and/or between $j$ and $k$ is removed (this is the operation $\operatorname{OP}_{(i,j),\text{ER}}$); or (ii) there is a covering edge that is added between $i$ and $k$ (this is the operation $\operatorname{OP}_{(i,j),\text{CE}}$).
+For both (i) and (ii), an extra operation afterward is needed: for (i), it is to add back the edge; and for (ii), it is to remove the covered edge.
 Similar argument can be made if $\mathcal{G}_{t}$ has a v-structure that $\mathcal{G}_{s}$ does not.
-It is natural to define the set:
-$\begin{aligned} S:=\Big\{(a,b,c)\subseteq V\times V\times V:&(a,b),(b,c)\subseteq\operatorname{sk}(\mathcal{G}_{s})\cap\operatorname{sk}(\mathcal{G}_{t}),(a,c)\not\in\operatorname{sk}(\mathcal{G}_{s})\cup\operatorname{sk}(\mathcal{G}_{t}),\\ &\text{exactly one of }(a,b,c)\in\operatorname{v-str}(\mathcal{G}_{s})\text{ and }(a,b,c)\in\operatorname{v-str}(\mathcal{G}_{t})\Big\},\end{aligned}$
-and say that associated with every element in $S$, we need at least two operations.
-That is, in addition to $|\mathrm{sk}(\mathcal{G}_{s})\setminus\mathrm{sk}(\mathcal{G}_{t})|+|\mathrm{sk}(\mathcal{G}_{t})\setminus\mathrm{sk}(\mathcal{G}_{s})|$, at least $2|S(\mathcal{G}_{s},\mathcal{G}_{t})|$ additional steps must be involved in any path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$.
-However, $2|S|$ may overcount the number of operations needed.
-For example, removing an edge between $i$ and $j$ may account for elements $(i,j,k)$ and $(i,j,m)$ in $S$, or adding a covered edge between $i$ and $k$ may account for elements $(i,j,k)$ and $(i,m,k)$ in $S$.
-To avoid overcounting the number of operations, we consider the following strategy.
-For any pair $(i,j)$, we define two operations.
-The first, denoted $\operatorname{OP}_{(i,j),\text{CE}}(\cdot)$, takes a subset of $S$, and removes all the elements of the form $(i,k,j)$ or $(j,k,i)$ for any $k$; this operation is associated with adding a covered edge between $i$ and $j$.
-The second, denoted $\operatorname{OP}_{(i,j),\text{ER}}(\cdot)$, removes all elements of the form $(i,j,k)$, $(k,j,i)$, $(j,i,k),(k,i,j)$ for any $k$; this operation is associated with deleting an edge between $i$ and $j$.
-Let $\mathrm{OP}=\{\operatorname{OP}_{(i,j),\text{CE}},\operatorname{OP}_{(i,j),\text{ER}}\text{ for all }\text{ vertices }(i,j)\}$ be the collection of all such operations.
-Our goal is to find the smallest number of operations from $\mathrm{OP}$ such that, when applied consecutively to $S$, the resulting set is empty.
-That is, we seek a sequence of operations $(\mathrm{op}_{1},\dots,\mathrm{op}_{m})$ where each step $\mathrm{op}_{i}\in\mathrm{OP}$, such that $\mathrm{op}_{1}\odot\dots\odot\mathrm{op}_{m}(S)=\emptyset$, and $m$ is minimized.
-Let the minimum be $m_{\text{op}}$.
-Then, it follows that:
-$d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\geq|\mathrm{sk}(\mathcal{G}_{s})\setminus\mathrm{sk}(\mathcal{G}_{t})|+|\mathrm{sk}(\mathcal{G}_{t})\setminus\mathrm{sk}(\mathcal{G}_{s})|+2m_{\text{op}}.$
-F.2 Simplifications under polytree assumptions: proof of Theorem 3
-The result is based on the following lemmas, which we prove first.
-Lemma 3.
-Let $\mathcal{G}_{a}$ and $\mathcal{G}_{b}$ be two graphs in $\mathfrak{G}$, the set of polytree CPDAGs over a common node set, with $\mathcal{G}_{a}\preceq\mathcal{G}_{b}$.
-Then, for every DAG $\mathcal{D}_{b}$ represented by $\mathcal{G}_{b}$, there exists a DAG $\mathcal{D}_{a}$ represented by $\mathcal{G}_{a}$, such that $\mathcal{D}_{a}$ is a directed subgraph of $\mathcal{D}_{b}$.
-Proof of Lemma 3.
-First, note that $\mathrm{sk}(\mathcal{G}_{a})\subseteq\mathrm{sk}(\mathcal{G}_{b})$; otherwise conditional dependencies of $\mathcal{G}_{a}$ are not a subset of those of $\mathcal{G}_{b}$.
-Further, if $\mathcal{G}_{a}$ has a v-structure $i\rightarrow j\leftarrow k$, then $\mathcal{G}_{b}$ must also have this v-structure.
-Suppose not, then, since $\mathcal{G}_{b}$ cannot have cycles and its skeleton is a superset of $\mathcal{G}_{a}$, then $\mathcal{G}_{b}$ must have an edge between $i$ to $j$ and $j$ to $k$ without an edge between $i$ to $k$.
-This leads to a contradiction, however, since $\mathcal{G}_{a}$ encodes the conditional dependencies $X_{i}\not\perp X_{j}|X_{S}$ for every set $S$ that contains $k$, while $\mathcal{G}_{b}$ not has have this dependency for a set $S$.
-Similarly, if $\mathcal{G}_{b}$ has a v-structure $i\rightarrow j\leftarrow k$, then either $\mathcal{G}_{a}$ has the same v-structure or the edge $i$ to $j$ or $j$ to $k$ is missing in $\mathcal{G}_{a}$.
-Suppose not, that is $i$ and $j$, and $j$ and $k$ are connected but not in a v-structure form.
-Then, $\mathcal{G}_{a}$ implies the conditional dependency $X_{i}\not\perp X_{j}|S$ for every $S$ that does not contain $k$.
-However, $\mathcal{G}_{b}$ implies the conditional independence $X_{i}\perp X_{j}|X_{S}$ for some $S$ that does not contain $k$.
-Using the facts above, we conclude the desired result.
-Let $\mathcal{E}=\{(i,j)\}$ be collection of nodes that are connected in $\mathcal{G}_{b}$ but not in $\mathcal{G}_{a}$.
-By the analysis above, any v-structure in $\mathcal{G}_{a}$ is also present in $\mathcal{G}_{b}$.
-Furthermore, every v-structure $i\rightarrow k\leftarrow j$ in $\mathcal{G}_{b}$ is either present in $\mathcal{G}_{a}$ or at least one of the edges between pair of nodes $(i,k)$ and $(j,k)$ is missing.
-Consider any DAG $\mathcal{D}_{b}$ represented by $\mathcal{G}_{b}$.
-Let $\mathcal{D}_{a}$ be the DAG that is obtained by deleting the edges in $\mathcal{D}_{b}$ among pairs of nodes $\mathcal{E}$.
-By construction, $\mathcal{D}_{a}$ has the same skeleton as any DAG in $\mathcal{G}_{a}$.
-Since $\mathcal{D}_{a}$ is a polytree, every structure present in $\mathcal{D}_{a}$ or in $\mathcal{G}_{a}$ are also in $\mathcal{G}_{b}$.
-Suppose as a point of contradiction that there exists a v-structure $i\rightarrow k\leftarrow j$ that is in $\mathcal{D}_{a}$ and not in $\mathcal{G}_{a}$.
-Since this v-structure must also be present in $\mathcal{G}_{b}$, and from the previous statement, then one of the edges between pairs of nodes $(i,k)$ or $(j,k)$ must be missing in $\mathcal{G}_{a}$, which contradicts $\mathcal{D}_{a}$ and $\mathcal{G}_{a}$ have the same skeleton.
-A similar argument allows to conclude that it is not possible that $\mathcal{G}_{a}$ has a v-structure that is not present in $\mathcal{D}_{a}$.
-Thus, $\mathcal{D}_{a}$ and $\mathcal{G}_{a}$ have the same skeleton and v-structures.
-We conclude that $\mathcal{D}_{a}$ must be a DAG represented by $\mathcal{G}_{a}$.
-∎
-Lemma 4.
-The poset $\mathcal{L}_{\text{poly}}$ underlying CPDAGs that are polytrees is graded with $\mathrm{rank}(\cdot)$ measuring the number of edges.
-Proof.
-Suppose as a point of contradiction that $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathcal{G}_{\text{poly}}$ and $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$ with $\mathrm{rank}(\mathcal{G}_{1})+1<\mathrm{rank}(\mathcal{G}_{2})$.
-By Lemma 3, for every DAG $\mathcal{D}_{2}=(\mathrm{V},E_{2})$ represented by $\mathcal{G}_{2}$, there exists a DAG $\mathcal{D}_{1}=(\mathrm{V},E_{1})$ represented by $\mathcal{G}_{1}$ such that $E_{1}\subseteq E_{2}$.
-Since $\mathrm{rank}(\mathcal{G}_{1})+1<\mathrm{rank}(\mathcal{G}_{2})$, there exists $E_{3}$, satisfying $E_{1}\subset E_{3}\subset E_{2}$.
-Let $\mathcal{G}_{3}$ be the CPDAG obtained by completing the DAG $\mathcal{D}_{3}=(\mathrm{V},E_{3})$.
-By construction, $\mathcal{G}_{1}\preceq\mathcal{G}_{3}\preceq\mathcal{G}_{2}$ with $\mathcal{G}_{3}$ being distinct from $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$.
-This leads to a contradiction.
-Proof of Theorem 3.
-We will show that the poset is lower semimodular and then appeal to Theorem 2 to conclude the desired result.
-Let $\mathcal{G}_{1},\mathcal{G}_{2},\mathcal{G}_{3}\in\mathcal{G}_{\text{poly}}$ be three CPDAGs that are polytrees.
-Suppose $\mathcal{G}_{1}\lessdot\mathcal{G}_{3}$ and $\mathcal{G}_{2}\lessdot\mathcal{G}_{3}$.
-Consider any DAG $\mathcal{D}_{3}=(V,E_{3})$ represented by $\mathcal{G}_{3}$.
-By Lemma 3, there exist DAGs $\mathcal{D}_{1}=(V,E_{1})$ and $\mathcal{D}_{2}=(V,E_{2})$, represented by $\mathcal{G}_{2}$ and $\mathcal{G}_{1}$, respectively, such that $E_{1}\subseteq E_{3}$ and $E_{2}\subseteq E_{3}$.
-Since $\mathcal{L}_{\text{poly}}$ is graded by Lemma 4, we have that $E_{3}\setminus E_{1}=e_{1}$ and $E_{3}\setminus{E}_{2}=e_{2}$, where $|e_{1}|=|e_{2}|=1$ and $e_{1}\neq e_{2}$.
-Let $\tilde{\mathcal{D}}=(\mathrm{V},E_{1}\cap E_{2})$, and define $\tilde{\mathcal{G}}$ to be the CPDAG obtained by completing $\tilde{\mathcal{D}}$.
-By construction, $\tilde{\mathcal{G}}$ is a polytree and satisfies $\tilde{\mathcal{G}}\preceq\mathcal{G}_{1}$ and $\tilde{\mathcal{G}}\preceq\mathcal{G}_{2}$.
-Since $\mathrm{rank}(\tilde{\mathcal{G}})=\mathrm{rank}(\mathcal{G}_{1})-1=\mathrm{rank}(\mathcal{G}_{2})-1$, we have that $\tilde{\mathcal{G}}\lessdot\mathcal{G}_{1}$ and $\tilde{\mathcal{G}}\lessdot\mathcal{G}_{2}$.
-∎
-F.3 Additional theoretical results
-Lemma 5.
+The set $\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t})$ precisely defines the set of relevant triplets described in the previous paragraph.
+Let $\mathrm{OP}:=\{\operatorname{OP}_{(i,j),\text{CE}},\operatorname{OP}_{(i,j),\text{ER}}\forall i,j\in\mathrm{V}\}$.
+Our goal then becomes finding the smallest number of operations, from the set of operations $\mathrm{OP}$ such that, when applied consecutively to $S(\mathcal{G}_{s},\mathcal{G}_{t})$, the resulting set is empty.
+In other words, we obtain that:
+$d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\geq|\mathrm{sk}(\mathcal{G}_{s})\setminus\mathrm{sk}(\mathcal{G}_{t})|+|\mathrm{sk}(\mathcal{G}_{t})\setminus\mathrm{sk}(\mathcal{G}_{s})|+2m_{\text{opt}}(\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t})).$
+where
+$m_{\mathrm{op}}(\mathrm{S}):=\min\{m\in\mathbb{N}:\exists\mathrm{op}_{1},\dots,\mathrm{op}_{m}\text{ such that }\mathrm{op}_{1}\circ\dots\circ\mathrm{op}_{m}(\mathrm{S})=\emptyset\}.$
+Consider the partition $\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t})=\mathrm{S}_{1}\cup\dots\mathrm{S}_{L}$. Since triplets in distinct partitions do not share any common vertices, we have that:
+$m_{\mathrm{op}}(\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t}))=\sum_{\ell=1}^{L}m_{\mathrm{opt}}(\mathrm{S}_{\ell}).$
+F.2 Additional results
+Lemma S6.
 For the class of CPDAGs, and for any pair of CPDAGs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, we have:
-$\max_{\mathcal{G}\preceq\mathcal{G}_{1},\mathcal{G}\preceq\mathcal{G}_{2}}\mathrm{rank}(\mathcal{G})\leq|\mathrm{sk}(\mathcal{G}_{1})\cap\mathrm{sk}(\mathcal{G}_{2})|.$
-Proof of Lemma 5.
+$\max_{\mathcal{G}\preceq\mathcal{G}_{1},\mathcal{G}\preceq\mathcal{G}_{2}}\operatorname{\mathrm{rank}}(\mathcal{G})\leq|\mathrm{sk}(\mathcal{G}_{1})\cap\mathrm{sk}(\mathcal{G}_{2})|.$
+Proof of Lemma S6.
 Notice that for any $\mathcal{G}\preceq\mathcal{G}_{1}$, we have that $\mathrm{sk}(\mathcal{G})\subseteq\mathrm{sk}(\mathcal{G}_{1})$; otherwise, conditional dependencies of $\mathcal{G}$ are not a subset of those of $\mathcal{G}_{1}$.
@@
 This allows us to conclude that for $\mathcal{G}$ with $\mathcal{G}\preceq\mathcal{G}_{1}$ and $\mathcal{G}\preceq\mathcal{G}_{2}$, $\mathrm{sk}(\mathcal{G})\subseteq\mathrm{sk}(\mathcal{G}_{1})\cap\mathrm{sk}(\mathcal{G}_{2})$.
-Since $\mathrm{rank}(\mathcal{G})=|\mathrm{sk}(\mathcal{G})|$, we have the desired result.
+Since $\operatorname{\mathrm{rank}}(\mathcal{G})=|\mathrm{sk}(\mathcal{G})|$, we have the desired result.
 ∎
-Lemma 6.
+Lemma S7.
 For the class of CPDAGs, and for any pair of CPDAGs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, we have:
-$\min_{\mathcal{G}_{1}\preceq\mathcal{G},\mathcal{G}_{2}\preceq\mathcal{G}}\mathrm{rank}(\mathcal{G})\geq|\mathrm{sk}(\mathcal{G}_{1})\cup\mathrm{sk}(\mathcal{G}_{2})|.$
-Proof of Lemma 6.
-Based on the same argument as proof of Lemma 5, we have that for $\mathcal{G}$ with $\mathcal{G}_{1}\preceq\mathcal{G}$ and $\mathcal{G}_{2}\preceq\mathcal{G}$, $\mathrm{sk}(\mathcal{G}_{1})\cup\mathrm{sk}(\mathcal{G}_{2})\subseteq\mathrm{sk}(\mathcal{G})$.
-Since $\mathrm{rank}(\mathcal{G})=|\mathrm{sk}(\mathcal{G})|$, we have the desired result.
+$\min_{\mathcal{G}_{1}\preceq\mathcal{G},\mathcal{G}_{2}\preceq\mathcal{G}}\operatorname{\mathrm{rank}}(\mathcal{G})\geq|\mathrm{sk}(\mathcal{G}_{1})\cup\mathrm{sk}(\mathcal{G}_{2})|.$
+Proof of Lemma S7.
+Based on the same argument as proof of Lemma S6, we have that for $\mathcal{G}$ with $\mathcal{G}_{1}\preceq\mathcal{G}$ and $\mathcal{G}_{2}\preceq\mathcal{G}$, $\mathrm{sk}(\mathcal{G}_{1})\cup\mathrm{sk}(\mathcal{G}_{2})\subseteq\mathrm{sk}(\mathcal{G})$.
+Since $\operatorname{\mathrm{rank}}(\mathcal{G})=|\mathrm{sk}(\mathcal{G})|$, we have the desired result.
 ∎
-Appendix G Theoretical results for MPDAGs
-G.1 Poset for MPDAGs is not graded: proof of Proposition 10
-To prove that the poset $\mathcal{L}$ underlying MPDAGs is not graded, it suffices to find two maximal chains between a pair of graphs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ of different lengths.
-Fig.˜9 presents a collection of graphs that specify a maximal chain from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$.
-This chain has length $3$.
-Similarly, Fig.˜10 presents another collection of graphs that specify a maximal chain from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$.
-This chain, however, has length $5$.
-As the length of these two maximal chains are not equal, we conclude that $\mathcal{L}$ is not graded.
-Figure 9: A maximal chain between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$, specified by the sequence $(\mathcal{G}_{1},\mathcal{G}_{2},\mathcal{G}_{3},\mathcal{G}_{4})$ where $\mathcal{G}_{i+1}$ covers $\mathcal{G}_{i+1}$ for $i=1,2,3$.
-Figure 10: A maximal chain between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$, specified by the sequence $(\mathcal{G}_{1},\mathcal{G}_{2},\mathcal{G}_{3},\mathcal{G}_{4},\mathcal{G}_{5},\mathcal{G}_{6})$ where $\mathcal{G}_{i+1}$ covers $\mathcal{G}_{i+1}$ for $i=1,2,3,4,5$.
-G.2 MPDAG model complexity: proof of Proposition 12
-We construct an explicit sequence of MPDAG from $\mathcal{G}$ to the empty MPDAG by first directing one undirected at a time in $\mathcal{G}$ to obtain a sequence $\mathcal{G}=\mathcal{G}_{1},\dots,\mathcal{G}_{k}$ such that $\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G}_{i})-\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G}_{i+1})=1,\mathcal{G}_{i}\succeq\mathcal{G}_{i+1}$, $\mathcal{G}_{k}$ is a DAG and where $k$ is the number of undirected edges in $\mathcal{G}$.
-This is possible by repeatedly applying the result from Lemma 7.
-We append to this sequence another sequence $\mathcal{G}_{k}=\tilde{\mathcal{G}}_{1},\dots,\tilde{\mathcal{G}}_{l}$ such that $\mathrm{pseudo}\text{-}\mathrm{rank}(\tilde{\mathcal{G}}_{i})-\mathrm{pseudo}\text{-}\mathrm{rank}(\tilde{\mathcal{G}}_{i+1})=1,\tilde{G}_{i}\succeq\tilde{G}_{i+1}$, all $\tilde{\mathcal{G}}_{i}$ are DAGs, $\tilde{\mathcal{G}}_{l}$ is the empty graph and $l$ is the number of directed edges in $\mathcal{G}$.
-We can obtain such a sequence simply by deleting the edges in $G_{k}$ in any random order.
-By combining the two sequence we obtain a $\mathrm{pseudo}\text{-}\mathrm{rank}(G)$-length sequence of graphs from $G$ to the empty graph such that adjacent graphs on the sequence neighbors in $\mathcal{L}$ while having pseudo-ranks that differ by $1$.
-Lemma 7.
-Consider an MPDAG $\mathcal{G}$.
-Then there exists an undirected edge in $\mathcal{G}$ such that directing the edge results in an MPDAG $\tilde{\mathcal{G}}$ which is equal to $\mathcal{G}$ except for the one additional directed edge.
-Figure 11: Forbidden induced subgraphs of an MPDAG (left) and induced subgraphs that may exist in an MPDAG and are one directed edge away from a forbidden subgraph (right).
+Appendix G Further results for MPDAGs
+G.1 MPDAG model complexity: proof of Proposition 10
+We will use the following notation and terminology in this subsection.
+The siblings of a vertex $v$ in graph $\mathcal{G}$, denoted by $\operatorname{Sib}_{\mathcal{G}}(v)$, are those vertices that are adjacent to $v$ via an undirected edge.
+We use $\operatorname{Pa}_{\mathcal{G}}(v)$ and $\operatorname{Ch}_{\mathcal{G}}(v)$ to denote the parents and the children of $v$ in $\mathcal{G}$ respectively.
+The neighbors of $v$ in $\mathcal{G}$, denoted by $\operatorname{N}_{\mathcal{G}}(v)$, is the set of vertices that are adjacent to $v$.
+By definition, it holds that $\operatorname{N}_{\mathcal{G}}(v)=\operatorname{Sib}_{\mathcal{G}}(v)\cup\operatorname{Pa}_{\mathcal{G}}(v)\cup\operatorname{Ch}_{\mathcal{G}}(v)$.
+The graph in the subscripts will be suppressed when it is clear from the context.
+A vertex set is called a clique if every pair of vertices in the set are adjacent.
+A vertex $v$ in $\mathcal{G}$ is called simplicial if $\operatorname{N}_{\mathcal{G}}(v)$ is a clique.
+The topological order is a partial order on the set of vertices: for two vertices $u$ and $v$, we say $u\preceq v$ if $u\rightarrow\dots\rightarrow v$.
+Given a vertex set $\mathrm{S}$ that contains a vertex $u$, we say $u$ is topologically minimal in $\mathrm{S}$ if there exists no vertex $v$ in $\mathrm{S}$ such that $v\prec u$.
+We first prove a supporting lemma.
+Lemma S8.
+Let $\mathcal{H}$ be an MPDAG where every pair of vertices are connected by an undirected path.
+Suppose $u$ is a simplicial vertex and $u$ has a child $v$ with usvu-s-v in $\mathcal{H}$.
+Then $v$ is also simplicial.
+[…]
+Figure S11: Graphs that appear in the proof of Lemma˜S8.
 Proof.
-Consider $\mathcal{G}$ and suppose there exists a pair $(a,b)$, such that aba-b is in $\mathcal{G}$, $\mathrm{Sib}(a)\subseteq\mathrm{Sib}(b)\cup\mathrm{Ch}(b)$ and $b$ is a minimal element in the partial topological ordering of $\mathcal{G}$ among all nodes with an adjacent undirected edge.
-Here, $Sib(v)$ is the set of nodes $u$ such that uvu-v called the sibling of $v$, while $Ch(v)$ is the set of nodes $v\rightarrow u$ called the set of children of $v$.
-We first show that directing the edge to $a\rightarrow b$ results in an MPDAG $\tilde{\mathcal{G}}$ before showing that there must always exist such a pair.
-To show that $\tilde{\mathcal{G}}$ is an MPDAG it suffices to show that it does not contain any of the forbidden subgraphs shown on the left hand size of Fig.
-11.
-In order for that to possibly occur $\mathcal{G}$ must contain a pre-forbidden structure as shown on the right hand size of Fig.
-11.
-We now prove that if such a pre-forbidden structure exists we will never direct an edge that will result in a forbidden structure.
-Consider the first pre-forbidden subgraph.
-The node in the middle always has at least one sibling that the two adjacent nodes do not have.
-Therefore, the middle node can never be selected as $b$ and we will never arrive at the forbidden structure.
-Consider the second pre-forbidden subgraph.
-The left most node is not minimal in the topological order among all nodes with an adjacent undirected edge and can therefore not be selected as $b$.
-The right most-node has one more sibling (the left most node) than the middle node.
-Therefore, we can never arrive at the forbidden subgraph.
-There is no pre-forbidden graph for the third structure.
-For the fourth pre-forbidden graph the node at the bottom-left is not minimal and can therefore never be selected as $B$.
-It now remains to show that such a pair $(a,b)$ exists in any MPDAG that has at least one undirected edge.
-Consider the CPDAG $\mathcal{C}$ of $\mathcal{G}$.
-The connected undirected components of $\mathcal{C}$ are chordal and therefore each have a perfect numbering, i.e., an ordering $v_{1},\dots,v_{p}$ such that each $v_{i}$ is simplical in the induced subgraph of $\mathcal{C}$ on $v_{i},\dots,v_{p}$.
-Here a node $v$ is simplical if it’s neighborhood $\mathrm{Sib}(v)$ forms a clique.
-Furthermore, each member of a connected undirected component of $\mathcal{C}$ is equivalent in terms of the graphs topological order (Andersson, Madigan and Perlman, 1997).
-Therefore, the simplical node $s$ in the undirected component that is minimal in terms of the order satisfies i) minimality among all nodes with a sibling and ii) for any sibling $s^{\prime}$ of $s$ it holds that $\mathrm{Sib}(s)\subseteq\mathrm{Sib}(s^{\prime})$.
-In any MPDAG in which $s$ remains minimal among all nodes with siblings, any node $s^{\prime}$ topologically maximal among the siblings of $s$ satisfies $\mathrm{Sib}(s)\subseteq\mathrm{Sib}(s^{\prime})\cup\mathrm{Ch}(s^{\prime})$.
-It thus only remains to consider the case that $s$ is not topologically minimal in $\mathcal{G}$.
-This implies that there exists an edge out of $s$ in $\mathcal{G}$ into a node $c$ that is a sibling of $s$ in $\mathcal{C}$.
-In order to avoid any forbidden structure of the first kind any undirected edge adjacent to $c$ in $\mathcal{C}$ that is towards a non-sibling of $s$ in $\mathcal{C}$ must also be directed in $\mathcal{G}$.
-Therefore, any sibling of $c$ in $\mathcal{G}$ must have been a sibling of $s$ in $\mathcal{G}$.
-As the siblings of $s$ in $\mathcal{G}$ form a clique it follows that there must exist a sibling $s^{\prime}$ of $c$ such that $\mathrm{Sib}(c)\subseteq\mathrm{Sib}(s^{\prime})\cup\mathrm{Ch}(s^{\prime})$.
-If $c$ is minimal it therefore satisfies the conditions we make on $b$.
-If not we repeat the argument until we arrive at such a $b$ through breadth first search.
-If there are no undirected edges adjacent to $s$ in $\mathcal{G}$ we move to the next simplical node in the perfect ordering of the undirected subgraph of $\mathcal{C}$ and repeat the argument above.
-G.3 Pseudo-distance upper-bounding SHD: proof of Proposition 13
+Because $u$ is simplicial, it suffices to show that $w\in\{u\}\cup\operatorname{N}(u)$ holds for every $w\in\operatorname{N}(v)$.
+For each $w\in\operatorname{N}(v)$, there are three cases as illustrated in Fig.˜S11.
+For each case, the argument is based on the four induced subgraphs listed in Fig.˜2(b) that are forbidden in any MPDAG.
+  1. (I)
+    $w\in\operatorname{Sib}(v)$.
+The result holds trivially if $w=s$.
+For a contradiction, suppose $w\neq s$ and $w\notin\operatorname{N}(u)$.
+Now observe that $u\rightarrow v-w$, as an induced subgraph, is forbidden.
+  2. (II)
+    $w\in\operatorname{Pa}(v)$.
+Suppose $u,w$ are not adjacent.
+To avoid forbidden structure $w\rightarrow v-s$, we have an edge between $s$ and $w$.
+Further, the edge cannot be undirected because otherwise the graph induced by $\{u,v,w,s\}$ is a forbidden structure.
+Therefore, the edge must be directed, in particular, as $w\rightarrow s$ due to the undirected edge svs-v.
+However, this induces another forbidden structure $w\rightarrow s-u$.
+  3. (III)
+    $w\in\operatorname{Ch}(v)$.
+Because every pair of vertices is connected by an undirected path in $\mathcal{H}$, let $p$ be a shortest undirected path that connects $w$ and $\{u,s,v\}$.
+We prove $w\in\operatorname{N}(u)$, in particular, $u\rightarrow w$, by induction on the length of $p$.
+    1. (i)
+      Base case: $p$ is of length 1.
+Observe that in this case we must have an undirected edge between $s$ and $w$.
+Without an edge between $u$ and $w$, the graph induced by $\{s,u,v,w\}$ is a forbidden structure.
+Hence, we have an edge between $u$ and $v$, in particular, $u\rightarrow w$ by acyclicity.
+    2. (ii)
+      Suppose the result holds when $p$ is of length $l$ or shorter for $l\geq 1$.
+Now we show that the result still holds when $p$ is of length $l+1$.
+We write the path as $p=\langle w\equiv w_{0},w_{1},\dots,w_{l+1}\rangle$ for $w_{l+1}\in\{s,u,v\}$.
+To avoid the forbidden structure $v\rightarrow w-w_{1}$, we know $v$ and $w_{1}$ are adjacent.
+There are two further cases.
+      1. (a)
+        $v\rightarrow w_{1}$.
+Observe that $w_{1}$ plays the role of $w$ and $w_{1}$ is connected to $\{s,u,v\}$ by an undirected path of length $l$.
+By our induction hypothesis, we have $u\rightarrow w_{1}$.
+To avoid the forbidden structure $u\rightarrow w_{1}-w$, there must be an edge between $u$ and $w$.
+In particular, we have $u\rightarrow w$.
+      2. (b)
+        $v-w_{1}$.
+By case (I), we know $u$ and $w_{1}$ are adjacent.
+To prove by contradiction, suppose there is no edge betweeb $u$ and $w$.
+Observe that the edge between $u$ and $w_{1}$ cannot be undirected since otherwise $\{u,v,w,w_{1}\}$ induces a forbidden structure.
+Further, due to $w_{1}-v$, we cannot have $w_{1}\rightarrow u$.
+Hence, we must have $u\rightarrow w_{1}$.
+However, we see a contradiction because this leads to a forbidden structure $u\rightarrow w_{1}-w$.
+Hence, there must be an edge between $u$ and $w$, in particular, $u\rightarrow w$.
+Figure S12: Induced subgraphs that are forbidden in an MPDAG (left) and their corresponding pre-structures (right).
+Each pre-structure can exist as an induced subgraph in an MPDAG, but upon directing aba-b to $a\rightarrow b$, it becomes a forbidden structure.
+We are ready to prove Lemma˜3.
+Proof of Lemma˜3.
+The proof consists of two parts.
+First, we show that by choosing an undirected edge aba-b that meets certain criteria (assuming such aba-b exists in $\mathcal{G}$) and directing it to $a\rightarrow b$ and leaving everything else intact, we obtain an MPDAG $\tilde{\mathcal{G}}$.
+Second, we argue that an edge aba-b that meets our selection criteria always exists in $\mathcal{G}$.
+We show the first part. We select an edge aba-b in $\mathcal{G}$ that meets the following criteria:
+  1. (i)
+    $\operatorname{Sib}_{\mathcal{G}}(b)\subseteq\{a\}\cup\operatorname{Sib}_{\mathcal{G}}(a)\cup\operatorname{Ch}_{\mathcal{G}}(a)$, and
+  2. (ii)
+    there is no vertex $c$ such that $b\rightarrow c$ and bscb-s-c exist in $\mathcal{G}$.
+Let $\tilde{\mathcal{G}}$ be a copy of $\mathcal{G}$ except that aba-b is directed to $a\rightarrow b$.
+We argue that $\tilde{\mathcal{G}}$ is an MPDAG by showing that it satisfies the graphical constraints listed in Section˜B.2.
+The chordality constraint is satisfied because $\mathcal{G}$ is already an MPDAG and we have only directed an undirected edge.
+Thus, it remains to show that the four forbidden structures listed in Fig.˜2(b) are not introduced.
+By inspecting these forbidden structures, we list in Fig.˜S12 the only possible ways that a forbidden structure can be introduced upon directing a single edge: forbidden structure III cannot be introduced; for structures I, II and IV to be introduced, a corresponding pre-forbidden structure must exist in $\mathcal{G}$ and aba-b must be chosen as specified in the figure.
+To finish proving the first part, we argue that our selection criteria prevent aba-b from being chosen in such ways:
+  • Structure I.
+    This violates (i);
+  • Structure II.
+    The first case violates (i) and the second case violates (ii);
+  • Structure IV.
+    This violates (ii).
+We now prove the second part, namely the existence of aba-b in $\mathcal{G}$ that meets our selection criteria.
+By our assumption, $\mathcal{G}$ has one or more undirected edge.
+Accordingly, let $\mathrm{C}$ be a (maximal) connected component spanned by undirected edges that has two or more vertices and let $\mathcal{H}:=\mathcal{G}_{\mathrm{C}}$.
+We first choose $b$ in $\mathcal{H}$ that satisfies criterion (ii).
+By Section˜B.2, $\mathcal{H}$ is an MPDAG and the skeleton of $\mathcal{H}$ is chordal.
+By a classical result on chordal graphs (Lauritzen, 1996, Lemma 2.9), there exists a simplicial vertex $b^{\prime}$ in $\mathcal{H}$ with $\operatorname{Sib}_{\mathcal{H}}(b)\neq\emptyset$.
+If $b^{\prime}$ satisfies (ii), let $b:=b^{\prime}$.
+Otherwise, whenever $b^{\prime}\rightarrow c$ and $b^{\prime}-s-c$ holds in $\mathcal{G}$ (and hence in $\mathcal{H}$ by definition), replace $b^{\prime}$ with $c$.
+We repeat this process until (ii) is satisfied and let $b$ be the final $b^{\prime}$.
+Because $\mathcal{H}$ has no directed cycle, the process must terminate in a finite number of steps.
+Further, by Lemma˜S8, the process maintains the property that $b^{\prime}$ is simplicial.
+Meanwhile, the fact that $\operatorname{Sib}_{\mathcal{H}}(b^{\prime})\neq\emptyset$ is also maintained.
+Hence, the resulting $b$ is simplicial in $\mathcal{H}$ with $\operatorname{Sib}_{\mathcal{H}}(b)\neq\emptyset$.
+Finally, we choose $a\in\operatorname{Sib}_{\mathcal{H}}(b)$ such that $a$ is topologically minimal in $\operatorname{Sib}_{\mathcal{H}}(b)$.
+It only remains to show that criterion (i) holds.
+Since $b$ is simplicial in $\mathcal{H}$, for each $u\in\operatorname{Sib}_{\mathcal{H}}(b)$, we have $u=a$ or $u\in\mathcal{N}_{\mathcal{H}}(a)$.
+Further, by the way that $a$ is chosen, we know $u\notin\operatorname{Pa}_{\mathcal{H}}(a)$.
+Hence, we have $u\in\{a\}\cup\operatorname{Sib}_{\mathcal{H}}(a)\cup\operatorname{Ch}_{\mathcal{H}}(a)$, i.e.,
+$\operatorname{Sib}_{\mathcal{H}}(b)\subseteq\{a\}\cup\operatorname{Sib}_{\mathcal{H}}(a)\cup\operatorname{Ch}_{\mathcal{H}}(a).$
+By the fact that $\operatorname{Sib}_{\mathcal{G}}(b)=\operatorname{Sib}_{\mathcal{H}}(b)$, $\operatorname{Sib}_{\mathcal{G}}(a)=\operatorname{Sib}_{\mathcal{H}}(a)$ and $\operatorname{Ch}_{\mathcal{H}}(a)\subseteq\operatorname{Ch}_{\mathcal{G}}(a)$, we obtain
+$\operatorname{Sib}_{\mathcal{G}}(b)\subseteq\{a\}\cup\operatorname{Sib}_{\mathcal{G}}(a)\cup\operatorname{Ch}_{\mathcal{G}}(a),$
+i.e., criterion (i). ∎
+Proof of Proposition˜10.
+Suppose $\mathcal{G}$ has $l$ undirected edges and $k$ directed edges, so we have $\operatorname{\mathrm{pseudo-rank}}(\mathcal{G})=2l+k$. The maximal chain is given by
+$\mathcal{G}=:\mathcal{G}_{2l+k}\gtrdot\mathcal{G}_{2(l-1)+k+1}\gtrdot\dots\gtrdot\mathcal{G}_{l+k}\gtrdot\mathcal{G}_{l+k-1}\gtrdot\dots\gtrdot\mathcal{G}_{0}:=\mathcal{G}_{\hat{0}}$,
+which consists of two segments: (i) starting from $\mathcal{G}$, we repeatedly apply Lemma˜3 to convert one undirected edge to a directed edge until we reach $\mathcal{G}_{l+k}$ that only contains directed edges; (ii) then, we remove directed edges iteratively in any order till we reach the empty graph.
+∎
+G.2 Pseudo-distance upper-bounding SHD: proof of Proposition 11
 Proof.
-The proof is based on the following observation: for any pair of graphs $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ with $\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G}_{2})-\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G}_{1})=1$, the graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are identical except for a change in a single edge.
-This change can be one of the following: a missing edge becomes a directed edge, or a directed edge becomes an undirected edge.
-Consider a pair of vertices $(v_{i},v_{j})$ such that the edge between them differs in $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$.
-Then, based on the observation above, for any path $p\in L_{\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$, there exists at least one move along the path between graphs $\mathcal{G}_{x}$ and $\mathcal{G}_{y}$ where only this edge changes.
-Note that if the edge is missing in one graph and undirected in the other, or if the edge is directed in opposite orientations in the two graphs, then at least two such moves exist in the path.
-This implies that $d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is lower bounded by the structural Hamming distance.
+The proof is based on the following observation: for any pair of neighboring graphs $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$ with $\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{2})-\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{1})=1$, the graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are identical except for a single edge: either a missing edge versus a directed edge, or a directed edge versus an undirected edge.
+Consider a pair of vertices such that the edge between them differs in $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$.
+Then, based on the observation above, for any path $p\in\mathrm{P}_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t})$, then $p$ must go through some neighboring graphs $\mathcal{G}_{i},\mathcal{G}_{i+1}$ such that they differ exactly at this edge.
+If the edge is missing in one graph and undirected in the other, or if the edge is directed in opposite orientations in the two graphs, then at least two such moves exist in the path.
+This implies that $d_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is lower bounded by $\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})$.
 ∎
-G.4 Simplifications under polytree assumptions: proof of Theorem 4
-The proof relies on the following lemma, which we prove first.
-Throughout, we note that an MPDAG $\mathcal{G}$ can be viewed as containing all DAGs that are members of a CPDAG $\mathcal{C}$ subject to a set of fixed directed edges.
-Lemma 8.
-Let $\mathcal{G}$ be a MPDAG that is a polytree and represents the collection of DAGs $\mathcal{D}_{1},\dots,\mathcal{D}_{k}$.
-Consider any pair of nodes $(i,j)$ that are connected in $\mathcal{G}$ and let $\tilde{\mathcal{D}}_{l}$ be the DAG obtained by removing the edge between nodes $(i,j)$ in $\mathcal{D}_{l}$.
-Then, the collection $\tilde{\mathcal{D}}_{1},\tilde{\mathcal{D}}_{2}\dots,\tilde{\mathcal{D}}_{k}$ forms a MPDAG $\tilde{\mathcal{G}}$ with $\tilde{\mathcal{G}}\preceq\mathcal{G}$.
-Proof of Lemma˜8.
-Since $\mathcal{G}$ is a valid MPDAG, all the DAGs it represents have the same skeleton and v-structures.
-By construction, all modified DAGs (after removing the edge) have the same skeleton.
-Further, they all have the same v-structure.
-In particular, each $\mathcal{D}_{i}$ is a polytree, and removing an edge does not uncover a v-structure.
-Further, if the edge removal removes a v-structure in one modified graph, it removes it for all the other modified graphs.
-We have thus shown that the collection of DAGs $\{\tilde{\mathcal{D}}_{1},\tilde{\mathcal{D}}_{2},\dots,\tilde{\mathcal{D}}_{k}\}$ are all Markov equivalent and members of a CPDAG $\tilde{\mathcal{C}}$.
-It remains to check that the collection $\{\tilde{\mathcal{D}}_{1},\tilde{\mathcal{D}}_{2},\dots,\tilde{\mathcal{D}}_{k}\}$ forms a valid MPDAG.
-That is, the collection represents all DAGs that are members of $\tilde{\mathcal{C}}$, subject to a set of fixed directed edges $\tilde{\mathcal{E}}=\{i_{1}\to i_{2},i_{3}\to i_{4},\dots\}$.
-Let $\mathcal{C}$ be the CPDAG associated with the collection $\{\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{k}\}$.
-By definition, $\{\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{k}\}$ is all DAGs that are members of $\mathcal{C}$}, subject to a set of fixed directed edges ${\mathcal{E}}=\{j_{1}\to j_{2},j_{3}\to j_{4},\dots\}$.
-Let
-$\tilde{\mathcal{E}}=\mathcal{E}\cup\{k\to j:(i,j,k)\in\operatorname{v-str}(\mathcal{C})\}\cup\{k\to i:(j,i,k)\in\operatorname{v-str}(\mathcal{C})\}.$
-Let $\tilde{\mathcal{G}}$ be the MPDAG obtained by restricting $\tilde{\mathcal{C}}$ to the edge set $\tilde{\mathcal{E}}$.
-We will show that the collection of DAGs represented by $\tilde{\mathcal{\mathcal{G}}}$ is precisely $\{\tilde{\mathcal{D}}_{1},\tilde{\mathcal{D}}_{2},\dots,\tilde{\mathcal{D}}_{k}\}$.
-First, we show the collection of DAGs represented by $\tilde{\mathcal{\mathcal{G}}}$ is a subset of $\{\tilde{\mathcal{D}}_{1},\tilde{\mathcal{D}}_{2},\dots,\tilde{\mathcal{D}}_{k}\}$.
-Take a DAG $\tilde{\mathcal{D}}$ represented by $\tilde{\mathcal{G}}$.
-By construction, $\tilde{\mathcal{D}}$ is compatible with edge set $\mathcal{E}$ (since $\mathcal{E}\subseteq\tilde{\mathcal{E}}$).
-Add the edge between nodes $(i,j)$ in the same orientation as any DAG in $\mathcal{\mathcal{G}}$ to obtain a DAG $\bar{\mathcal{D}}$.
-Then, by construction, $\bar{\mathcal{D}}$ is a member of $\mathcal{C}$, and is compatible with the edge set $\mathcal{E}$.
-Thus, $\bar{\mathcal{D}}$ is among the collection $\{\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{k}\}$, which means that $\tilde{\mathcal{D}}$ is among the collection $\{\tilde{\mathcal{D}}_{1},\tilde{\mathcal{D}}_{2},\dots,\tilde{\mathcal{D}}_{k}\}$.
-Next, we show the collection of DAGs represented by $\tilde{\mathcal{\mathcal{G}}}$ is a superset of $\{\tilde{\mathcal{D}}_{1},\tilde{\mathcal{D}}_{2},\dots,\tilde{\mathcal{D}}_{k}\}$.
-Recall that $\tilde{\mathcal{D}}_{i}$ is a member of $\tilde{\mathcal{C}}$ for every $i=1,2,\dots,k$.
-Furthermore, by construction, each $\tilde{\mathcal{D}}_{i}$ is compatible with $\tilde{\mathcal{E}}$.
-This direction then follows.
-Proof of Theorem 4.
-Let $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ be two MPDAGs that are polytrees with $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$.
-Suppose as a point of contradiction that $\mathrm{rank}(\mathcal{G}_{1})+1<\mathrm{rank}(\mathcal{G}_{2})$ and there does not exist a distinct MPDAG that is a polytree in between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$.
-We consider three cases and show that there is a contradiction in each case.
-Case I: Two or more pairs of nodes are connected in $\mathcal{G}_{2}$ but not in $\mathcal{G}_{1}$.
-Let $(i,j)$ be one such pair.
-Let $\{\mathcal{D}_{1},\dots,\mathcal{D}_{k}\}$ be the collection of DAGs represented by $\mathcal{G}_{2}$.
-Let $\{\tilde{\mathcal{D}}_{1},\dots,\tilde{\mathcal{D}}_{k}\}$ to be the collection of DAGs obtained by removing the edge between pair $(i,j)$ from each $\mathcal{D}_{l},l=1,2,\dots,k$.
-By Lemma 8, the collection form an MPDAG $\tilde{\mathcal{G}}$ with $\tilde{\mathcal{\mathcal{G}}}\preceq\mathcal{G}_{2}$.
-Since $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$, we have that for every DAG $\mathcal{D}$ represented by $\mathcal{G}_{1}$, there exists a DAG ${\mathcal{D}}_{l}$ represented by $\mathcal{G}_{2}$ where $\mathcal{D}$ is a subgraph of ${\mathcal{D}}_{l}$.
-It then follows that $\mathcal{D}$ is also a subgraph of ${\mathcal{D}}_{l}$.
-So $\mathcal{G}_{1}\preceq\tilde{\mathcal{G}}$.
-Case II: There is at least one pair of node that is connected via an undirected edge in $\mathcal{G}_{2}$ but is connected via a directed edge in $\mathcal{G}_{1}$, and there is one pair of nodes that are connected in $\mathcal{G}_{2}$ but not in $\mathcal{G}_{1}$.
-Suppose $\mathcal{G}_{2}$ is the MPDAG obtained by restricting the CPDAG $\mathcal{C}$ to edge set $\mathcal{E}$.
-Consider one of the undirected edges between nodes $(i,j)$ that is directed as $i\to j$ in $\mathcal{G}_{1}$.
-Let $\tilde{\mathcal{G}}$ be the MPDAG obtained by restricting $\mathcal{C}$ to the edge set $\mathcal{E}\cup\{i\to j\}$.
-Then, by construction, $\tilde{\mathcal{G}}\preceq\mathcal{G}_{2}$ with $\tilde{\mathcal{G}}\neq\mathcal{G}_{2}$.
-Furthermore, since $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$, for any DAG $\mathcal{D}_{1}$ represented by $\mathcal{G}_{1}$, there exists a DAG $\mathcal{D}_{2}$ represented by $\mathcal{G}_{2}$ such that $\mathcal{D}_{1}$ is a sub-graph of $\mathcal{D}_{2}$.
-Since $i\to j$ in $\mathcal{D}_{1}$, then $i\to j$ in $\mathcal{D}_{2}$.
-This implies that $\mathcal{D}_{2}$ is also represented by $\tilde{G}$, which consequently implies that $\mathcal{G}_{1}\preceq\tilde{\mathcal{G}}$.
-Since $\tilde{\mathcal{G}}$ has still at least one more edge than $\mathcal{G}_{1}$, we have that $\mathcal{G}_{1}\neq\tilde{\mathcal{G}}$.
-Case III: There is at least two pair of node that are connected via an undirected edge in $\mathcal{G}_{2}$ but are connected via a directed edge in $\mathcal{G}_{1}$.
-Suppose $\mathcal{G}_{2}$ is the MPDAG obtained by restricting CPDAG $\mathcal{C}$ to edge set $\mathcal{E}$.
-Let $\tilde{\mathcal{G}}$ be the MPDAG obtained by restricting $\mathcal{C}$ to $\mathcal{E}\cup\{i\to j\}$.
-Following the argument as Case II, we have that $\tilde{\mathcal{G}}\preceq\mathcal{G}_{2}$ and $\mathcal{G}_{1}\preceq\tilde{\mathcal{G}}$.
-By construction, $\tilde{\mathcal{G}}\neq\mathcal{G}_{2}$.
-It remains to show that $\tilde{\mathcal{G}}\neq\mathcal{G}_{1}$.
-Note that each pair of nodes that have a directed edge in $\mathcal{G}_{1}$ (but undirected in $\mathcal{G}_{2}$) cannot be part of a v-structure in $\mathcal{G}_{1}$.
-Otherwise, since $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are both polytrees, all the v-structures in $\mathcal{G}_{1}$ must also be present in $\mathcal{G}_{2}$, and so these edges would also need to be directed in $\mathcal{G}_{2}$.
-We consider some sub-cases.
-In the first subcase, the two undirected edges are incident with one edge being between $(i,j)$ and the other between $(j,k)$.
-In this case, to avoid creating a v-structure, because of the polytree condition, $j\to k$ in $\mathcal{G}_{1}$.
-Suppose we let $\tilde{\mathcal{E}}=\mathcal{E}\cup\{j\to k\}$, and let $\tilde{\mathcal{G}}$ be the MPDAG obtained by restricting $\mathcal{C}$ to the edge set $\tilde{\mathcal{E}}$.
-Since the edge between $(i,j)$ is undirected in $\mathcal{G}_{2}$, we have that $\tilde{\mathcal{G}}$ has an undirected edge between node pair $(i,j)$.
-Thus $\tilde{\mathcal{G}}\neq\mathcal{G}_{1}$ in this subcase.
-In the second subcase, the two undirected edges are incident with one edge being between $(i,j)$ and the other between $(i,k)$.
-Since the edge between $(i,k)$ is undirected in $\mathcal{G}_{2}$, we have that there must exist DAGs represented by $\mathcal{G}_{2}$ with the edge $k\to i$ and an edge $i\to j$ as well as DAGs with edge $i\to k$ and $i\to j$, as these do not create v-structures and there is no concern about violating acyclicity due to the polytree condition.
-We thus have that that $\tilde{\mathcal{G}}$ has an undirected edge between node pair $(i,k)$, and consequently $\tilde{\mathcal{G}}\neq\mathcal{G}_{1}$.
-In the final subcase, the two undirected edges in $\mathcal{G}_{2}$ that are directed in $\mathcal{G}_{1}$ are not incident, i.e., they involve pairs $(i,j)$ and $(l,m)$.
-To be distinct from the previous subcases, together with the fact that $i\to j$ in $\mathcal{G}_{1}$ should not create a v-structure, means that all other edges involving $j$ must be of the form $j\to k$ in $\mathcal{G}_{1}$.
-This fact isolates the two edges between $(i,j)$ and $(l,m)$, meaning that taking among all DAGs represented by $\mathcal{G}_{2}$ that have the edge $i\to j$, there exists ones that have the edge $l\to m$ and ones that have the edge $m\to l$.
-Thus, the intermediate graph $\tilde{\mathcal{G}}$ is distinct from $\mathcal{G}_{1}$.
-∎
-Appendix H On the Idiosyncrasies of the structural interventional distance and separation distance
+Appendix H Idiosyncrasies of the structural interventional distance and separation distance
 H.1 Structural intervention distance
@@
 The resulting symmetrized SID satisfies the first two axioms of a metric, but as we will show with an example, it still fails to satisfy the triangle inequality.
-Figure 12: Three DAGs for which symmetrized SID violates the triangle inequality as shown in Example 2.
-Example 2 (Symmetrized SID violates triangle inequality).
-Let $\mathcal{G}_{1}$, $\mathcal{G}_{2}$, and $\mathcal{G}_{3}$ be the three DAGs from Figure 12.
+Figure S13: Three DAGs for which symmetrized SID violates the triangle inequality as shown in Example S3.
+Example S3 (Symmetrized SID violates triangle inequality).
+Let $\mathcal{G}_{1}$, $\mathcal{G}_{2}$, and $\mathcal{G}_{3}$ be the three DAGs from Figure S13.
 Denote $d_{SID}$, the structural intervention distance for DAGs as defined by (Peters and Bühlmann, 2015).
@@
 H.2 Separation distance
-We interpret a probabilistic graphical model as the collection of all random variables satisfying the conditional independence constraints encoded by the graph.
-A closely related and popular alternative is to interpret a probabilistic graphical model directly as the collection of conditional independence statements encoded by the graph or more formally the set of all triples $(X,Y,Z)$ corresponding to a conditional independence statement of the form $X\perp\!\!\!\!\perp Y\mid Z$ (Studenỳ, 1992).
-Based on this interpretation of a CPDAG as a collection of triples (Wahl and Runge, 2025) have proposed an alternative distance notion called the separation distance.
-It defines the distance between two CPDAGs as the size of the symmetric difference of the two corresponding sets of triples.
-Similarly to our class of distances, the separation distance is a metric and model oriented, albeit on a different model interpretation.
-It can also be generalized to any class of probabilistic graphical models that encode only conditional independence statements.
-In fact any separation distance can be viewed as a minimal path distance on a Hasse diagram: Consider the set of all triples $X,Y,Z\subseteq\{1,\dots,n\}$ and the poset structure induced by subset inclusion on this set.
-This poset forms a lattice, with the meet and join between two sets of triples defined as their intersection and union, respectively.
-As a result, the minimal path distance on the corresponding Hasse diagram equals the down-up distance, which in turn equals the size of the symmetric difference between the two sets.
+[…]
 However, the fact that the separation distance is implicitly defined on the set of all triples of subsets of the node set, while ignoring which of these correspond to a CPDAG or even valid conditional independence models for that matter has two serious consequences.
@@
 We now illustrate these characteristics with an example.
-Example 3.
-Let $\mathcal{G}_{1}$ denote the empty CPDAG, $\mathcal{G}_{2}$ a CPDAG with a single (undirected) edge, $\mathcal{G}_{3}$ the fully connected CPDAG, and $\mathcal{G}_{4}$ the same graph with one (undirected) edge removed, all on $n$ nodes.
-Since all d-separation statements hold in $\mathcal{G}_{1}$, it can be represented by the set of all triples of subsets on the set of $n$ nodes.
+Example S4.
+Let $\mathcal{G}_{1}$ denote the empty CPDAG, $\mathcal{G}_{2}$ a CPDAG with a single (undirected) edge, $\mathcal{G}_{3}$ the fully connected CPDAG, and $\mathcal{G}_{4}$ the same graph with one (undirected) edge removed, all on $n$ vertices.
+Since all d-separation statements hold in $\mathcal{G}_{1}$, it can be represented by the set of all triples of subsets on the set of $n$ vertices.
 In $\mathcal{G}_{2}$, on the other hand, all statements hold except those of the form
-$\mathrm{V}_{i}\perp_{\mathcal{G}_{2}}\mathrm{V}_{j}\mid Z,\quad Z\subseteq V\setminus\{\mathrm{V}_{i},\mathrm{V}_{j}\}.$
-As a result, the separation distance between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ is the size of this latter set, that is, $2^{n-1}$.
-Furthermore, for every disconnected node we add to both $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, their distance doubles.
-The CPDAG $\mathcal{G}_{3}$, on the other hand, encodes no d-separation statements and therefore corresponds to the empty set, while $\mathcal{G}_{4}$ encodes the single statement $\mathrm{V}_{i}\perp\mathrm{V}_{j}\mid\mathrm{V}\setminus\{\mathrm{V}_{i},\mathrm{V}_{j}\}$.
-Hence, the separation distance between $\mathcal{G}_{3}$ and $\mathcal{G}_{4}$ equals $1$.
-Again adding a disconnected node to both graphs doubles this distance.
+[…]
+[…]
 In contrast, $d_{\mathcal{L}}(\mathcal{G}_{1},\mathcal{G}_{2})=d_{\mathcal{L}}(\mathcal{G}_{3},\mathcal{G}_{4})=1$, as $\mathcal{G}_{2}$ is a covering element of $\mathcal{G}_{1}$ and $\mathcal{G}_{4}$ of $\mathcal{G}_{3}$ in $\mathcal{L}$.
@@
 Furthermore, these distances remain unchanged if we add a disconnected node to all four graphs.
-Appendix I Algorithmic Aspects of Computing the $\mathcal{M}$-oriented distance
-We present additional algorithms that are important for finding the $\mathcal{M}$-oriented distance in the context of CPDAGs.
-Up-down distance: The algorithm for computing the up-down distance is the ‘dual’ of Algorithm 2.
-In particular, (i) the neighbor selection step is restricted to be upward; (ii) the search terminates when a graph $\mathcal{G}^{\prime}$ is found with $\mathcal{G}_{t}\preceq\mathcal{G}^{\prime}$; (iii) we use the admissible heuristic $2|\mathrm{sk}(\mathcal{G}_{t})\cup\mathrm{sk}(\mathcal{G}^{\prime})|-\mathrm{rank}(\mathcal{G}_{t})-\mathrm{rank}(\mathcal{G}^{\prime})$; and finally (iv) we verify whether $\mathrm{rank}(\mathcal{G}^{\prime})\geq u:=|\mathrm{sk}(\mathcal{G}_{t})\cup\mathrm{sk}(\mathcal{G}^{\prime})|$ before checking if $\mathcal{G}_{t}\preceq\mathcal{G}^{\prime}$.
-Here, $u$ serves as the upper bound for $\rho_{\mathrm{join}}(\mathcal{G}_{s},\mathcal{G}_{t}):=\min_{\mathcal{G}\in\mathfrak{G},\mathcal{G}_{s}\preceq\mathcal{G},\mathcal{G}_{t}\preceq\mathcal{G}}\mathrm{rank}(\mathcal{G})$; see Lemma 6 in Section F.3 for proof.
-Algorithm 3 presents the full description of the procedure for finding the up-down distance.
-input : Graph class of CPDAGs $\mathfrak{G}$, pair of graphs $\mathcal{G}_{s},\mathcal{G}_{t}$, a heuristic $h\colon\mathfrak{G}\to\mathbb{R}_{\geq 0}$, upwards neighbor function $un_{\mathcal{L}}\colon\mathfrak{G}\to 2^{\mathfrak{G}}$, upper bound $u$ to $\rho_{\text{join}}(\mathcal{G}_{s},\mathcal{G}_{t})$
-output : Up-down distance $d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t})$
-1 begin
- $openSet\coloneqq\{\mathcal{G}_{s}\}$;
- // priority queue ordered by $g+h$
-2 $g(\mathcal{G}_{s})\coloneqq 0$;
-3 while $openSet\neq\emptyset$ do
-4    $\mathcal{G}_{\text{current}}\coloneqq$ graph in openSetopenSet minimizing $g(\mathcal{G}_{\text{current}})+h(\mathcal{G}_{\text{current}})$;
-5    if $u\leq\mathrm{rank}(\mathcal{G}_{\text{current}})\,\,\land\,\,\mathcal{G}_{t}\preceq\mathcal{G}_{\text{current}}$ then
-6       return $g(\mathcal{G}_{\text{current}})+\mathrm{rank}(\mathcal{G}_{\text{current}})-\mathrm{rank}(\mathcal{G}_{t})$
-7     remove $\mathcal{G}_{\text{current}}$ from openSetopenSet;
-8
-9    foreach $\mathcal{G}_{\text{neighbor}}\in$ un($\mathcal{G}_{\text{current}}$) do
-10       $tentative\coloneqq g(\mathcal{G}_{\text{current}})+1$;
+Appendix I Algorithms and additional numerical results
+I.1 Proof of Theorem˜8
+Proof.
+Because $\mathcal{L}$ is connected and $\mathfrak{G}$ is finite, $d_{\mathcal{L}}$ is well-defined for every pair of graphs in $\mathfrak{G}$.
+Suppose the algorithm does not terminate.
+Then, it must be due to graphs continuing to be pushed to openSet in line 22.
+However, this means line 19 of the algorithm is executed infinitely many times, where after each time $g[\mathcal{G}^{\prime}]$ is either created or strictly decreased.
+We have a contradiction because this implies the condition in line 18 must fail at some point.
+In what follows, we show that the algorithm returns the correct distance.
+Observe that the while loop cannot exit without returning either at line 8 or 12, because if so $\mathcal{G}_{t}$ is never visited and that contradicts the connectedness of $\mathcal{L}$.
+We first show that if the algorithm returns at line 8, $u^{\ast}$ upon return equals the desired distance.
+To see this, we argue that there must exist a shortest path that is routed through a graph $\mathcal{G}^{\prime}$ in openSet for which it holds that $f(\mathcal{G}^{\prime})\leq d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
+Specifically, consider any shortest path $p^{\ast}$ from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$.
+Because $\mathcal{G}_{t}$ is not yet reached, we have $p^{\ast}\cap\textsf{openSet}\neq\emptyset$ and we choose $\mathcal{G}^{\prime}$ to be the graph in this intersection that is pushed to openSet most recently.
+Indeed, for this graph on the frontier, we have $g[\mathcal{G}^{\prime}]=d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}^{\prime})$ and hence $f(\mathcal{G}^{\prime})=g[\mathcal{G}^{\prime}]+h(\mathcal{G}^{\prime})\leq d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}^{\prime})+d_{\mathcal{L}}(\mathcal{G}^{\prime},\mathcal{G}_{t})=d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ by the admissibility of $h$.
+Then, by the property of the priority queue, it follows that $\textsf{f}\leq f(\mathcal{G}^{\prime})\leq d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq u^{\ast}$.
+Meanwhile, we also have $\textsf{f}\geq u^{\ast}$ by line 7 so we conclude $\textsf{f}=u^{\ast}=d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
+Now we argue that g returned at line 12 equals $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
+Suppose this does not hold.
+Observe that $\textsf{g}=g[\mathcal{G}_{t}]$ when $\mathcal{G}_{t}$ was pushed into openSet and by construction $g[\mathcal{G}_{t}]\geq d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ holds, so we must have $\textsf{g}>d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
+Let $p^{\ast}$ be any shortest path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$.
+Similar to the argument in the preceding paragraph, let $\mathcal{G}^{\ast}$ be the graph on $p^{\ast}$ that was last pushed to openSet right before $\mathcal{G}_{t}$ was popped at line 6.
+By construction of the algorithm and the fact that $p^{\ast}$ is a shortest path, we have $g[\mathcal{G}^{\ast}]=d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}^{\ast})$.
+Further, using admissibility of $h$, we know
+$f(\mathcal{G}^{\ast})=g[\mathcal{G}^{\ast}]+h(\mathcal{G}^{\ast})\leq d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}^{\ast})+d_{\mathcal{L}}(\mathcal{G}^{\ast},\mathcal{G}_{t})=d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t}).$
+However, because the priority queue pops the item with the smallest index, we also have $f(\mathcal{G}^{\ast})\geq f(\mathcal{G}_{t})=\textsf{g}+0>d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ and hence a contradiction.
+∎
+I.2 Subroutines for computing the distance between CPDAGs
+Algorithm˜2 computes the lower bound on the model-oriented distance between two CPDAGs given by Proposition˜6. Algorithm˜3 computes $m_{\text{op}}(\cdot)$ for a set of v-structure triplets.
+Input: CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ over a vertex set $\mathrm{V}$
+Require : Subroutines $\textnormal{{FindMinOps}}(\cdot)$ and $\textnormal{{DisjointTriplets}}(\cdot)$
+Output: The lower bound on $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ given by Proposition˜6
+1 $\mathcal{G}_{s,\operatorname{sk}}\leftarrow\operatorname{sk}(\mathcal{G}_{s})$
+2 $\mathcal{G}_{t,\operatorname{sk}}\leftarrow\operatorname{sk}(\mathcal{G}_{t})$
+$d\leftarrow|\mathcal{G}_{s,\operatorname{sk}}\triangle\mathcal{G}_{s,\operatorname{sk}}|$
+// skeleton difference
+$\mathrm{S}^{\ast}\leftarrow\{\}$
+// $\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t})$
+3 foreach $a\in\mathrm{V}$ do
+4 foreach $b\in\mathrm{V}$ do
+5    foreach $c\in\mathrm{V}$ do
+6       if $a=b$ or $a=c$ or $b=c$ then
+7          Continue
+8       if $(a,b)\in\mathcal{G}_{s,\operatorname{sk}}\cap\mathcal{G}_{t,\operatorname{sk}}$ and $(b,c)\in\mathcal{G}_{s,\operatorname{sk}}\cap\mathcal{G}_{t,\operatorname{sk}}$ and $(a,c)\notin\mathcal{G}_{s,\operatorname{sk}}\cup\mathcal{G}_{t,\operatorname{sk}}$ then
+9          if $(a\rightarrow b\leftarrow c$ in $\mathcal{G}_{s})$ and $(a\rightarrow b\leftarrow c$ not in $\mathcal{G}_{t})$ then
+10             $\textnormal{{Push}}(\mathrm{S}^{\ast},(a,b,c))$
+11          else if $(a\rightarrow b\leftarrow c$ in $\mathcal{G}_{t})$ and $(a\rightarrow b\leftarrow c$ not in $\mathcal{G}_{s})$ then
+12             $\textnormal{{Push}}(\mathrm{S}^{\ast},(a,b,c))$
+13
+14
+15
+16
 […]
-12          $g(\mathcal{G}_{\text{neighbor}})\coloneqq tentative$;
-13            add $\mathcal{G}_{\text{neighbor}}$ to $openSet$ with priority $g(\mathcal{G}_{\text{neighbor}})+h(\mathcal{G}_{\text{neighbor}})$;
-Algorithm 3 An A* algorithm for computing the up-down distance $d_{\L ,\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t})$
-Finding a lower-bound for $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$: We next present the algorithm for computing the lower bound for $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ according to Property 2.
-Algorithm 4 presents the algorithm which uses Algorithm 5.
-input : CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ among nodes $V$, functions getDisjointV and findMinOps
-output : The lower bound for the $\mathcal{M}$-oriented distance between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$
-1 begin
-2 $\mathcal{G}_{t,\text{undir}}\coloneqq undirect(\mathcal{G}_{t})$
-3 $\mathcal{G}_{s,\text{undir}}\coloneqq undirect(\mathcal{G}_{s})$
- $d\coloneq|\mathrm{sk}(\mathcal{G}_{t,\text{undir}})\triangle\mathrm{sk}(\mathcal{G}_{s,\text{undir}})|$ ;
- // skeleton difference
- // Get all v structures
- $vList\coloneqq[]$ ;
- // list of uncommon undirected v-structures in the skeletons of the two graphs
-4 foreach node $i\in V$ do
-5    foreach node $j\in V$ do
-6       foreach node $k\in V$ do
-7          if $i\neq j\land j\neq k\land i\neq k\land((i,j)\in E(\mathcal{G}_{t,\text{undir}})\land(j,k)\in E(\mathcal{G}_{t,\text{undir}}))\land((i,j)\in E(\mathcal{G}_{s,\text{undir}})\land(j,k)\in E(\mathcal{G}_{s,\text{undir}}))\land(i,k)\notin E(\mathcal{G}_{t,\text{undir}})\land(i,k)\notin E(\mathcal{G}_{s,\text{undir}})$ then
-8             if $((i,j)\in E(\mathcal{G}_{t})\land(k,j)\in E(\mathcal{G}_{t})\land(j,i)\notin E(\mathcal{G}_{t})\land(j,k)\notin E(\mathcal{G}_{t})\land((i,j)\notin E(\mathcal{G}_{s}))\lor(k,j)\notin E(\mathcal{G}_{s})))\lor((i,j)\in E(\mathcal{G}_{s})\land(k,j)\in E(\mathcal{G}_{s})\land(j,i)\notin E(\mathcal{G}_{s})\land(j,k)\notin E(\mathcal{G}_{s})\land((i,j)\notin E(\mathcal{G}_{t}))\lor(k,j)\notin E(\mathcal{G}_{t})))$ then
-9                  append v structure $(i,j,k)$ to vListvList
- $groups\coloneqq getDistjointVGroups(vList)$ // get v structure groups with no shared vertices
-10 $h\coloneqq 0$
-11 if $|vList|>0$ then
-12    foreach $group\in groups$ do
-13       $h\coloneq h+2(findMinOps(group))$
-14 return h + d
-Algorithm 4 Algorithm for computing the lower bound for the $\mathcal{M}$-oriented distance according to Property 2.
-input : Set $S$ consisting of a collection of triplets of vertices, operations $\mathrm{OP}$
-output : Returns the minimum number of operations to remove all elements in $S$
-1 begin
- $queue\coloneqq[(S,0)]$ // tracks state and step count
- $visited\coloneqq\{\}$ // avoid repeat states
+18 foreach $\mathrm{S}\in\textnormal{{DisjointTriplets}}(\mathrm{S}^{\ast})$ do
+19 $h\leftarrow h+2\cdot\textnormal{{FindMinOps}}(\mathrm{S})$
+return $h+d$
+Algorithm 2 Computing the lower bound in Proposition˜6 on $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
+Input: Set $\mathrm{S}$ consisting of a collection of vertex triplets
+Output: $m_{\text{op}}(\mathrm{S})$
+$\textnormal{{Q}}\leftarrow\textnormal{{Dict}}()$
+// tracks state and step count
+1 $\textnormal{{Q}}[S]\leftarrow 0$
 2
-3 while $queue\neq\emptyset$ do
-4    $current\coloneqq$ get first element of the queue
-5      remove first element of queuequeue
-6
-7    $subS\coloneqq$ subset of S in currentcurrent
-8    $steps\coloneqq$ steps in currentcurrent
-9    if $subS$ is empty then
-10       return steps
-    // Get all pairs (i, j) from triples in S
-    $pairs\coloneqq{}$ // set of unique pairs from triples
-11    foreach $triple\in subS$ do
-12         add all combinations of 2 vertices from the tripletriple of vertices to pairspairs
-13    foreach $pair\in Lairs$ do
-14       $i\coloneqq$ first node in pairpair
-15       $j\coloneqq$ second node in pairpair
+3while Q is not empty do
+4 $(S^{\prime},k)\leftarrow\textnormal{{Pop}}(Q)$
+5 if $S^{\prime}=\emptyset$ then
+6    return $k$
+7 $\textnormal{{pairs}}\leftarrow\{\}$
+8 foreach $(a,b,c)\in S^{\prime}$ do
+9    $\textnormal{{Push}}(\textnormal{{pairs}},(a,b))$; $\textnormal{{Push}}(\textnormal{{pairs}},(a,c))$; $\textnormal{{Push}}(\textnormal{{pairs}},(b,c))$
+10
+11 foreach $(i,j)\in\textnormal{{pairs}}$ do
+12    foreach $\mathrm{op}\in\left\{\mathrm{OP}_{(i,j),\mathrm{CE}},\mathrm{OP}_{(i,j),\mathrm{ER}}\right\}$ do
+13       $S^{\prime\prime}\leftarrow\mathrm{op}(S^{\prime})$
+14       $\textnormal{{Q}}[S^{\prime\prime}]\leftarrow k+1$
+15
 16
-17       foreach $op\in\mathrm{OP}$ do
-18          $newS\coloneqq op(subS,i,j)$
-19          if $newS\notin visited$ then
-20               add $newS$ to visitedvisited
-21               add $(newS,steps+1)$ to queuequeue
- return NULL // No Solution
-22
-Algorithm 5 Algorithm findMinOps used in Algorithm 4.
+return 1-1
+// No Solution
+Algorithm 3 Subroutine FindMinOps for computing $m_{\text{op}}(\cdot)$ in Eq.˜6.
+I.3 Additional numerical results
+Fig.˜S14 compares the $d_{\mathcal{L},\operatorname{\mathrm{pseudo}}}$ and the structural Hamming distances for polytree MPDAGs over 5 vertices.
+Figure S14: The joint distribution of $d_{\mathcal{L},\operatorname{\mathrm{pseudo}}}$ (7) and the structural Hamming distance computed for all pairs of polytree MPDAGs over 5 vertices, where the identity line is drawn as dashed.